### v1

Abstract
Experiments deliver credible treatment effects estimates, but are often localized to specific sites, populations, or mechanisms. When such estimates are insufficient to extrapolate effects for broader policy questions as for external validity and general-equilibrium (GE) effects, researchers combine trials with external evidence from reduced-form or structural observational estimates, or prior experiments. We develop a unified framework for designing experiments in this setting: the researcher selects which parameters to identify experimentally from a feasible set (which treatment arms and/or individuals to include in the experiment), allocates sample size, and specifies how to weight experimental and observational estimators. Because observational inputs may be biased in ways unknown ex ante, we adopt a minimax proportional-regret objective that evaluates any candidate design relative to an oracle that knows the bias and jointly chooses the design and estimator. This yields a transparent bias-variance trade-off that requires no prespecified bias bound and depends only on information about the precision of the estimators and the estimand’s sensitivity to underlying parameters. We illustrate the framework by (i) designing small-scale cash-transfer experiments aimed at estimating GE effects and (ii) optimizing site selection for microfinance interventions.
1 Introduction
Randomized controlled trials have transformed empirical economics by delivering credible, internally valid estimates. However, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism. These effects, although useful, are often not sufficient to answer broader questions about external validity, generalizability, or equilibrium effects. In response, a growing literature in development economics (e.g., Meghir et al., 2022; Gechter et al., 2024; Bassi et al., 2022; Attanasio et al., 2012), political economy (e.g., de Albuquerque et al., 2025), education (e.g., Allende et al., 2019; Larroucau et al., 2024), and labor economics (e.g., Chetty et al., 2016) complements localized experiments with external evidence from reduced-form or structural observational estimates, and/or trials in other contexts, with the goal of estimating complex counterfactuals that no single experiment can fully identify. For example, in our review of AEA journals (in 2015-2025), over 30% of experimental papers report, in addition to the experiment, at least one observational estimate – either reduced-form (e.g., via matching or instrumental variable regression), or structural (see Figure 12). This raises a design question: given constraints on what experiments can identify, which experiments should be run (and how) when they will be combined with external evidence to estimate those counterfactuals?
For an illustrative example, consider a government piloting a cash-transfer program in a small set of districts to increase children’s school attendance. The trial delivers a local average effect, but policy decisions often require counterfactuals at scale, allowing prices and wages to adjust (e.g. Todd and Wolpin, 2006; Egger et al., 2022). Researchers therefore combine experimental evidence with a supply-demand model that leverages observational data to map local impacts into economy-wide outcomes. The design question is twofold: (i) which experiment to run (i.e., which parameters/effects are the most valuable to learn experimentally) given feasibility constraints and (ii) how to allocate sample across sites and arms so that, once combined with external inputs, we precisely estimate the effect at scale.
To answer these questions, we develop a framework for designing experiments to be used alongside external evidence. By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments conducted in other times or places. Our main contribution is a joint procedure that selects which experiments to run subject to a budget (i.e., which treatment arms and/or sub-populations to include in the study), sets their precision via sample allocation, and prescribes how to combine experimental estimates with external evidence. A central consideration in our design is that external inputs may be biased in ways not known ex ante.
We consider a setting where the object of interest is a known function $\tau(\theta)$ with $\theta$ a vector of unknown parameters; $\tau$ may be univariate or multivariate and typically encodes a policy-relevant target. Researchers have access to external estimates of $\theta$ that may be biased—henceforth, observational estimates. They then run one or more experiments in the population of interest to learn a subset of parameters without bias, choosing a sample allocation across sites and arms subject to a budget. We parameterize the design so that some, but not necessarily all, components of $\theta$ can be learned experimentally. After the experiment, each experimental estimate is combined with its observational counterpart via parameter-specific shrinkage weights (chosen jointly with the design), while parameters not learned experimentally rely on the observational estimates alone.
Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in a region receive the subsidy. The parameter vector $\theta$ bundles (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium, and (iii) wage/price adjustments that shift the returns to schooling. A large-scale trial that identifies GE effects is often infeasible; instead, one can run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$. For instance, researchers may choose between two (small-scale) treatment arms such as a CCT to estimate the direct effect or an unconditional-cash arm to measure the income elasticity, and extrapolate the GE effects using price effects from other experimental or observational studies. When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ may target the average effect across sites.
A natural starting point is to minimize the mean-squared error (MSE) for estimating $\tau(\theta)$, perhaps in a worst-case sense over bias in the observational estimates. In practice, however, the size and even the direction of bias are unknown (before the experiment is run). A worst-case approach would therefore be overly-conservative and disregard information about the variance. We instead adapt the definition of proportional (adaptation) regret previously studied when combining estimators generated by a fixed design (Armstrong et al., 2024; Tsybakov, 1998) and generalize it to the experimental design problem. Here, the regret is the ratio of the worst-case researcher’s mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the (shrinkage) estimator. Taking the supremum of this ratio over any value of the bias yields a procedure that is robust to any level of misspecification without requiring priors about the bias.
Our main result gives an explicit, easy-to-compute characterization of proportional regret. For any design and shrinkage weights, the regret is the maximum of two normalized components. The first is a variance component: the estimator’s sampling variance under the chosen design and weights, divided by the smallest variance attainable by any feasible design. The second is a bias component: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias. The variance component depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the shrinkage weights, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters. This gradient is evaluated, under mild conditions, at the observational estimates. [^4] For this result to hold, we assume that $\tau(\theta)$ is a smooth and differentiable function in $\theta$. For non-linear $\tau(\theta)$ in $\theta$, we also require that the observational estimates bias is local to zero in the spirit of Andrews et al. (2020); Bonhomme and Weidner (2022), but consider a more general framework where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error. Section 4.1 provides details. By contrast, the bias component depends only on the sensitivity vector and the shrinkage weights, both known ex-ante.
This characterization makes the bias-variance trade-off transparent. At the optimum, the two normalized components tend to be equalized: when the bias component dominates, weight shifts toward experimental evidence and the design prioritizes learning high-sensitivity coordinates; when variance dominates, the procedure invests sample where it most reduces variance and relies more (via shrinkage) on the most precise observational inputs. This yields a nested procedure: we (i) find (regret-optimal) shrinkage weights for each candidate design to balance the two ratios; (ii) allocate sample, subject to the budget, to reduce the variance; and (iii) select the experiment set, trading off designs that improve precision the most against those targeting parameters to which the policy estimand is the most sensitive. The resulting design and estimator only require knowledge of the expected variance-covariance matrix (standard in experimental design problems, Gerber and Green, 2012) and of the observational estimates. It can then be reported in a pre-analysis plan.
We illustrate the framework in the cash‐transfer application. As an illustrative example, we consider a researcher evaluating a conditional cash transfer (CCT) to increase schooling in rural Kenya. Budget and feasibility constraints permit only small-scale (partial-equilibrium) randomization, while the goal is to predict general-equilibrium effects. As preliminary (external) inputs, the researcher uses evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006). We are concerned that these preliminary estimates may lack external validity in Kenya. We compare three designs: (i) a CCT arm to estimate the direct schooling effect; (ii) an unconditional cash transfer (UCT) arm to estimate income effects; and (iii) a two-arm design that allocates sample across CCT and UCT under a common budget. Running a single arm captures settings with fixed costs per arm; allowing both arms with budgeted allocation captures cases with non-binding fixed costs but binding variable costs. When both arms are feasible, the regret-optimal allocation puts most participants in the CCT and a small fraction in the UCT, reflecting the target’s greater sensitivity to the direct effect despite the CCT’s higher variance. If only one arm can be run, the choice hinges on sample size: for very small $n$, the lower-variance of the UCT arm makes this preferred over a (too noisy) CCT arm; once $n\geq 500$, the more informative but noisier CCT yields lower regret. Relative to an oracle that knows the bias, the optimal design closely tracks oracle performance, and regret approaches one as sample size grows.
As a second application, we design where to run a microfinance experiment in rural India, integrating evidence from earlier nonrandomized microfinance introductions. We start from the observational estimates in Banerjee et al. (2024) and design an experiment that selects one or more areas for randomization. Because Banerjee et al. (2024) also implement a separate randomized expansion, we can calibrate performance of our design (that is agnostic of the bias) by comparing each candidate design’s MSE to that of an oracle that knows the bias, using the experimental estimates to calibrate the bias. The optimal design concentrates sample in high-sensitivity, high-variance areas and then spreads as the budget grows; compared to a benchmark that randomly chooses areas and splits sample equally, it reduces MSE by roughly 20% under the calibrated bias.
In summary, this paper links the experimental-design literature—which typically focuses on settings where all parameters are identified within the experiment and leaves aside questions of data combination—to recent work that integrates experimental and (reduced-form or structural) observational evidence to extrapolate effects in complex scenarios.
Recent advances for experimental design include balancing and variance-minimizing allocations (Tabord-Meehan, 2018; Bai, 2019; Cytrynbaum, 2021; Kallus, 2018; Bertsimas et al., 2015), adaptive designs for policy choice (Kasy and Sautmann, 2019; Russo et al., 2018; Kato et al., 2024; Cesa-Bianchi et al., 2025), and experimental design under correct model specification (Silvey, 2013; Chaudhuri and Mykland, 1993; Chaloner and Verdinelli, 1995; Higbee, 2024; Kiefer and Wolfowitz, 1959; Reeves et al., 2024; Viviano, 2020; Kasy, 2016). Traditional work on experimental design for robust-model estimation either focused on testing competing models (Atkinson and Fedorov, 1975; López-Fidalgo et al., 2007), or on using a-priori knowledge of (worst-case) bias for the design of an experiment (Box and Draper, 1959; Wiens, 1998; Tsirpitzi et al., 2023; Sacks and Ylvisaker, 1984). We complement this literature by introducing and studying the question of which experiment to design when combined with observational evidence (when not all parameters can be learned experimentally).
Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design. References include Manski and Tetenov (2016), Banerjee et al. (2020), Manski (2004), Dominitz and F. Manski (2017), Olea et al. (2024), Hu et al. (2024), Breza et al. (2025) among others. These focus on settings where researchers have only access to experimental variation. Here instead we study the problem of choosing which experiment to run (and how) when combined with observational evidence. We also complement literature that leverages correctly-specified models for site selection (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
A related line of work analyzes estimation under misspecification (Armstrong and Kolesár, 2018; Armstrong et al., 2024; Andrews et al., 2017, 2020; Bonhomme and Weidner, 2022; Christensen and Connault, 2023; James et al., 1961), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; Bhattacharya, 2013; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024). We adapt the worst-case proportional regret idea from Armstrong et al. (2024) and Tsybakov (1998) here studied in the context of experimental design. Our definition of sensitivity of the estimand to each parameter directly links to sensitivity analysis in Andrews et al. (2020). Unlike both strands of this literature where the design is fixed and researchers optimize over the choice of the estimator only, here we optimize the design itself. Consequently, the objective changes: regret is computed against the best design-estimator combination, and the resulting estimator is learned jointly with the design.
Finally, we connect to a growing empirical literature that combines observational (or model-based) evidence with experimental variation to learn economic quantities – spanning general equilibrium effects (Egger et al., 2022; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023), external-validity counterfactuals (Gechter et al., 2024; Banerjee et al., 2024), large-scale information campaigns (de Albuquerque et al., 2025; Larroucau et al., 2024), and market-system experiments (Bergquist and Dinerstein, 2020; Allende et al., 2019). We provide a unifying framework to decide what to learn experimentally, how to split the sample across experiments, and how to integrate experimental and observational evidence.
2 Problem description
Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$. The estimand may arise from a reduced-form (e.g., treatment effect in a given area) or a structural model. For simplicity, we assume that the estimand is univariate, while Section 4.3 shows how our framework directly extends to multivariate $\tau(\theta)\in\mathbb{R}^{q}$ for $q>1$.
The goal is to construct an estimator $\hat{\tau}$ that accurately approximates $\tau(\theta)$ by combining existing evidence (e.g., observational studies) with experimental variation designed by the researcher. Our question is how to design such experiments under feasibility constraints.
2.1 Setup
We assume that researchers have access to estimators (and their variance) of $\theta$ denoted as $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$. We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model. However, $\tilde{\theta}^{\mathrm{obs}}$, that for exposition we define observational estimates, have unknown biases collected in a vector $b\in\mathbb{R}^{p}$. Examples of $\tilde{\theta}^{\mathrm{obs}}$ include a structural model estimate with confounding, an estimate from an instrumental variable regression that may fail the exclusion restriction, or an estimate from a country different from the one of interest that may lack external validity.
Researchers may also collect evidence for a subset of parameters that they know is unbiased. For instance, researchers may generate exogenous variation via an experiment (or more generally acquiring more data) that identifies one or some of the parameters of interest. We refer to these parameters as experimental estimates, adopting here the definition of experimental estimates as parameters that the researcher knows are unbiased. [^5] In some applications researchers may be confident that a few observational estimates have no bias. It is possible to extend our framework to these settings with a change of notation where those observational estimates assumed to be unbiased are defined as experimental estimates, and such experimental estimates are always contained in the set $\mathcal{E}$ defined below.
Setting 1 (Main setup).
For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a known strictly positive-definite matrix $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})\subset\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$ with uniformly bounded entries, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
$\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{obs}}\right]-\theta=b,\qquad\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\right]-\theta_{\mathcal{E}}=0,$
for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance as
$\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E}).$ (1)
Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators
$\hat{\tau}_{\gamma}\equiv\tau(\hat{\theta}(\gamma)),\qquad\hat{\theta}_{j}(\gamma)=\begin{cases}\gamma_{j}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
where $\gamma=(\gamma_{j})_{j\in\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$.
Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design (which parameters are learned and with what precision). When clear, we omit the subscript $\Sigma$ on $\tilde{\theta}^{\mathrm{exp}}$.
We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012) (in practice, any consistent estimator, e.g., from pilot studies would suffice). In addition, we assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite. In addition $\Sigma$ is strictly positive definite, therefore assuming that the variance is bounded away from zero. [^6] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance. See Section 4.1 for details.
We let $\mathcal{E}\in\mathcal{S},\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$ for some constraint sets $(\mathcal{S},\mathcal{G}(\mathcal{E}))$ encoding feasibility or budget constraints. In our framework, restrictions on the experiments researchers can run $\mathcal{E}\in\mathcal{S}$ (and on their precision $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$) can take any desired form. In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or standard lower bound constraints on power for experimental estimates, which are common in empirical practice (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011)
We write compactly $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible experiments (e.g., sample size, etc.), with $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ for a choice of experiments $\mathcal{E}$.
Our goal is to optimize both over the design and the weights $\gamma$. Before introducing our main research question, we impose the following assumption.
Assumption 1 (First-order estimation error).
For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume
$\tau(\theta)-\tau\!\big(\hat{\theta}(\gamma)\big)=\sum_{j=1}^{p}\omega_{j}\,\big(\theta_{j}-\hat{\theta}_{j}(\gamma)\big),$ (2)
for known weights $\omega\in\mathbb{R}^{p}$, with $|\omega_{j}|>0$ for all $j$.
Assumption 1 holds exactly for linear $\tau(\cdot)$ and serves as a first-order approximation for smooth nonlinear $\tau(\cdot)$.
Remark 1 (Non-linear $\tau(\theta)$).
Section 4.1 (and Example 3 below) extends our framework to non-linear estimands and shows how Assumption 1 serves as an approximation via a Taylor expansion within a local asymptotic framework, where the bias is local to zero, but at a possibly faster, equal or even slower rate than the estimator’s standard error. For non-linear $\tau$, $\omega$ is replaced by the gradient of $\tau$ evaluated at the baseline observational estimates (i.e., $\omega=\frac{\partial\tau(\theta)}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}+o_{p}(1)$), which is typically observed before the experiment is conducted.
Remark 2 (When researchers can only identify linear combinations of the parameters).
In some applications, researchers may be able to identify via an experiment a linear combination of the parameters of interest (up-to a first order approximation), defined $\phi\equiv W^{\top}\theta$, where $W\in\mathbb{R}^{p\times s}$ is a known matrix. Suppose that $\widetilde{\phi}$ is an unbiased estimator of $\phi$, and the researcher can access a subset of its entries via an experiment.
Our framework extends to this setting as follows. Define the coefficient of the linear projection of $\omega$ onto the column space of $W$: $\pi=\left(W^{\top}W\right)^{-1}W^{\top}\omega$ (known ex-ante). Define the residual $v=\omega-W\pi$. Then,
$\tau=\left[W\pi+v\right]^{\top}\theta=\pi^{\top}\phi+v^{\top}\theta\implies\tau=\check{\omega}^{\top}\check{\theta},$
where $\check{\omega}\equiv(\beta^{\top},v^{\top})^{\top}$ and $\check{\theta}=(\phi^{\top},\theta^{\top})^{\top}$. This corresponds to our original setting where the researcher’s feasible set only allows access to unbiased estimates of a subset of the parameters, corresponding to the entries of $\phi$. ∎
2.2 Research questions
Our design problem can be described in three steps:
  1. 1.
    Preliminary step: weights. For each $(\mathcal{E},\Sigma(\mathcal{E}))\in\mathcal{D}$, choose $\gamma$ to combine experimental and observational estimates on the selected coordinates.
  2. 2.
    Middle step: precision. For each $\mathcal{E}$, choose $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$, where $\mathcal{G}(\mathcal{E})$ denotes an arbitrary set of positive definite covariances under feasibility constraints.
  3. 3.
    Outer step: experiment choice. Choose $\mathcal{E}\in\mathcal{S}$, the feasible set of parameters to learn experimentally for given arbitrary constraint set $\mathcal{S}$.
Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025). We study this question here in combination with the experimental design problem, step 2 and 3, which has not been studied by the references above; this, as we show, will change the underlying optimization problem (also for $\gamma$). The central challenge is that performance depends on the unknown bias $b$ in the observational estimates.
2.3 Mapping to a simple two parameters model
While we derive our results in general form, we will build intuition using a stylized description with a two-parameter model.
Setting 2 (Illustration with two-parameter model).
Consider two parameters $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates
$\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\qquad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\qquad j=1,2.$
We may run one experiment: $\mathcal{S}=\{\{1\},\{2\}\}$. If we pick $j$, we estimate
$\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\qquad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$
and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$
Example 1 (Choosing the site for an experiment for external validity).
Gechter et al. (2024) study where to run an experiment. For illustration, consider two sites $j\in\{1,2\}$ with site-specific ATEs $\theta_{j}$ and target the cross-site average
$\tau(\theta)=\tfrac{1}{2}(\theta_{1}+\theta_{2})=\omega^{\top}\theta,\qquad\omega=\tfrac{1}{2}(1,1)^{\top}.$
Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al. (2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so $\mathcal{S}=\{\{1\},\{2\}\}$. If site $j$ is chosen, we obtain an unbiased $\tilde{\theta}^{\mathrm{exp}}_{j}$; the other site $k\neq j$ remains observational. Given $\mathcal{E}\in\mathcal{S}$,
$\widehat{\tau}_{\gamma}=\tfrac{1}{2}\!\left[\gamma_{j}\,\tilde{\theta}^{\mathrm{exp}}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j}+\tilde{\theta}^{\mathrm{obs}}_{k}\right],\qquad k\neq j.$
Our question is how to choose the set $\{j\}$ where to conduct the experiment (jointly with $\gamma$).
Example 2 (Choosing which survey to conduct).
Egger et al. (2022) study the efficacy of cash-transfer programs on the marginal propensity to consume (MPC). Measuring the MPC requires capturing both short- and long-run effects. Because survey rounds are limited, the authors complement experimental data that lacks short-run effects with auxiliary information from prior studies that use short-run surveys (collected in other regions; see Egger et al. (2022)). This raises the question of which survey to conduct (and with which frequency).
In stylized form, suppose researchers can observe, for $s\in\{1,2\}$, potential outcomes $Y_{s}(t)$ denoting consumption in period $s$ when measured $t$ periods after the intervention. The authors have auxiliary estimates from previous studies,
$\alpha\;=\;\mathbb{E}\!\big[Y_{1}(t=1)-Y_{1}(t=\infty)\big],\qquad\beta\;=\;\mathbb{E}\!\big[Y_{2}(t=1)-Y_{2}(t=\infty)\big],$
and wish to estimate the total effect $\tau=\alpha+\beta$. Researchers consider two survey designs:
  • (i)
    Early survey design to estimate $\alpha$ precisely;
  • (ii)
    Later survey to estimate $\beta$ precisely.
Our goal is to study which survey design to implement. (More complex designs with additional time periods or mixed precision across rounds are also possible here.)
Example 3 (Supply or demand experiment with non linear target).
(Bergquist and Dinerstein, 2020) conduct demand and supply experiments in food markets in Kenya. Suppose here we are interested in similar applications in Uganda. For exposition, consider a basic linear demand and supply
$Q^{D}=a-\beta_{D}P+u_{D},\qquad Q^{S}=c+\beta_{S}P+u_{S},\qquad\theta=(\beta_{D},\beta_{S})^{\top},$
with $\beta_{D}\neq-\beta_{S}$. Several estimands may be of interest; one of such estimands is the effect of a tariff $t$ on prices
$\tau\equiv\frac{\beta_{S}}{\beta_{D}+\beta_{S}}\,t.$
Let $(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top}$ denote baseline estimates from Kenya, that may lack external validity.
Due to cost constraints, $\mathcal{S}=\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms). Let
$\big(\tilde{\theta}^{\mathrm{obs}}-\theta\big)\ \sim\ \mathcal{N}\big(b,\Sigma\big),\quad\tilde{\theta}^{\mathrm{obs}}\equiv\sqrt{n}(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top},\quad\theta\equiv\sqrt{n}(\beta_{D},\beta_{S})^{\top}$
with $\theta$ and $\tilde{\theta}$ denoting the parameter and estimator here rescaled by the square-root of the sample size, and $b$ capturing bias. Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$. For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda, under local asymptotics and a first-order Taylor expansion described in Section 4.1, we can write [^7] Here, we are using a local asymptotic framework where the bias of the estimates in Kenya is non-negligible but small relative to $n$, i.e., $b/\sqrt{n}=o(1)$, so that $\omega$ can be consistently estimated using observational data.
$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\check{\omega}(\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}}))^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{b}{n^{1/2}}\right)}_{\text{small higher order effects}},\qquad\check{\omega}(\beta)=\frac{t}{(\beta_{D}+\beta_{S})^{2}}\begin{pmatrix}-\beta_{S}\\ \beta_{D}\end{pmatrix}$,
where we can replace $\check{\omega}$ with its consistent counterpart $\check{\omega}(\tilde{\beta}^{\mathrm{obs}})$. Our goal is to choose between a supply or demand experiment accounting for first-order bias (and variance) effect.
3 Robust experimental design
In this section we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$, a common measure of precision. In Section 4.2 we extend the framework to minimize the length of the confidence intervals.
3.1 Adaptation regret for experimental design
For a given design $(\mathcal{E},\Sigma(\mathcal{E}))$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[(\widehat{\tau}_{\gamma}-\tau)^{2}\right],$
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$. The MSE is a measure of precision of $\hat{\tau}$ and ideally, one would minimize $\mathrm{MSE}_{b}$ over $(\mathcal{E},\Sigma(\mathcal{E}),\gamma)\in\mathcal{D}$.
However, because $b$ is unknown, we first consider an uncertainty set given by an $\ell_{\infty}$-ball,
$\mathcal{B}(B)\equiv\{b:\ \|b\|_{\infty}\leq B\},$
with $B\geq 0$ an upper bound on the largest coordinate-wise bias. The $\ell_{\infty}$ choice has a desirable property relative to $\ell_{1}$ or $\ell_{2}$: it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters. The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
If an oracle knew $B$, a natural choice would be to pick the minimax design
$\mathrm{MSE}^{*}(B)\;\equiv\;\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\ \sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma).$
However, having to specify $B$ can pose a large burden on the researchers and make the choice of the experiment sensitive to $B$. Nevertheless, we could seek designs that perform as close as possible to the oracle that knows $B$, uniformly over the values of $B$.
This idea of performing as close as possible to the oracle is based on an extensive literature on regret minimization (Manski, 2004; Manski and Tetenov, 2007; Montiel Olea et al., 2023; Kitagawa and Tetenov, 2018; Manski and Tetenov, 2016). We minimize the worst-case proportional increase in MSE relative to the oracle across all bias radii:
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;\equiv\;\sup_{B\geq 0}\ \frac{\displaystyle\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\displaystyle\mathrm{MSE}^{*}(B)},$
defined adaptation regret by Armstrong et al. (2024) (building in turn on Tsybakov, 1998).
Whereas the above references consider the choice of an estimator for fixed design, here, the regret is relative to an oracle that chooses both the estimator and the design. Our key contribution – different from the references above – is the study of the optimal experimental design informed by biased (observational) estimates. This leads to a different definition of regret optimization, both because we optimize over the class of designs and $\Sigma$, but also in $\gamma$ due to the different definition of oracle. The optimal solution consists of a pair of design and estimator that can be pre-specified.
3.2 Optimal experimental design: main result
Before stating the result, we introduce compact notation. For a given $(\mathcal{E},\gamma)$, with $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ define
$\tilde{\omega}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega_{j},&j\notin\mathcal{E},\\ \omega_{j}(1-\gamma_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}_{\text{exp}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ\gamma$
where $\circ$ denotes the component-wise product and $\omega_{\mathcal{E}}$ is the subvector of $\omega$ corresponding to the entries indexed by $\mathcal{E}$. Let
$\tilde{\omega}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$ (3)
The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},$ (4)
with $|1-\gamma|$ indicating the absolute value of each entry of the vector $1-\gamma$.
Define the best achievable variance and worst-case bias divided by $B^{2}$ as
$\alpha^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\alpha(\mathcal{E},\Sigma,\gamma),\qquad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\beta(\mathcal{E},\gamma)=\min_{\mathcal{E}\in\mathcal{S}}\Big(||\omega||_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2}.$ (5)
Here $\alpha^{\star}$ is the variance of the most precise feasible design, while $\beta^{\star}$ is the squared bias of the least biased design. [^8] The expression for $\beta^{\star}$ corresponds to the following experiment choice: choose $\mathcal{E}$ with the largest $\omega_{\mathcal{E}}||_{1}$ and set $\gamma_{\mathcal{E}}\equiv 1$ at those coordinates. Both quantities depend on the class of designs $\mathcal{S},\mathcal{G}(\mathcal{E})$ and are easy to compute: $\alpha^{\star}$ is a constrained quadratic minimization; $\beta^{\star}$ requires enumerating $\mathcal{E}\in\mathcal{S}$.
Ideally, we would like a design with variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$. Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other.
Instead, in our main theorem below we show that the regret-optimal pair of design and estimator trades-off the bias and variance and depends on how far each of these two components are from their smallest attainable value.
Theorem 1.
Consider Setting 1 and let Assumption 1 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}.$
Proof.
See Appendix B.1. ∎
Theorem 1 provides an explicit expression of the regret (which does not require researchers to specify $B$). To our knowledge, it is the first result to study adaptation regret for experimental design. Its derivation requires us to find precise binding patterns as a function of $B$, which accounts for its effect on optimizing the oracle solution over the design and estimator.
The adaptation regret balances the variance of a given design relative to the smallest possible variance $\alpha^{\star}$ against the worst-case bias $\beta$ relative to the smallest possible worst-case bias $\beta^{\star}$.
To gain insight into the optimal solution, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), we may expect the minimizer $(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})\in\arg\min\mathcal{R}(\mathcal{E},\Sigma,\gamma)$ to equalize
$\frac{\alpha(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})}{\alpha^{\star}}\;=\;\frac{\beta(\mathcal{E}^{\star},\gamma^{\star})}{\beta^{\star}}.$
That is, at the optimum the design equalizes the estimator’s variance $\mathbb{V}(\hat{\tau}_{\gamma})$ (over the design choice) divided by the smallest feasible variance $\alpha^{\star}$ to the worst-case squared bias (equal to $\beta B^{2}$) divided by the smallest achievable worst-case squared bias $\beta^{\star}B^{2}$.
[…]
Figure 1: Each feasible design $(\mathcal{E},\Sigma,\gamma)$ maps to a point $\big(\alpha/\alpha^{\star},\ \beta/\beta^{\star}\big)$: the xx–axis is the variance ratio and the yy–axis is the worst–case bias ratio. The blue curve depicts the attainable frontier as we vary shrinkage $\gamma$ and precision $\Sigma$. Level sets of the objective $\mathcal{R}=\max\{\alpha/\alpha^{\star},\ \beta/\beta^{\star}\}$ are axis–aligned squares (the dotted inverted “L” shows the smallest such square touching the frontier). The minimizer outside boundary solutions is where the frontier meets the $45^{\circ}$ line (red dot), i.e., where the two normalized components are equalized: $\alpha(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})/\alpha^{\star}=\beta(\mathcal{E}^{\star},\gamma^{\star})/\beta^{\star}$.
Theorem 1 provides us an immediate and simple to compute solution to the optimal design problem via backward induction. Specifically, we compute the estimator, the optimal variance and the experimental choice as follows:
$\gamma^{\star}(\mathcal{E},\Sigma)$ $\in\mathrm{arg}\min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}$
$\Sigma^{\star}(\mathcal{E})$ $\in\mathrm{arg}\min_{\Sigma\in\mathcal{G}(\mathcal{E})}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\mathcal{E},\Sigma))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma))}{\beta^{\star}}\right\}$
$\mathcal{E}^{\star}$ $\in\mathrm{arg}\min_{\mathcal{E}\in\mathcal{S}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma^{\star}(\mathcal{E}),\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\beta^{\star}}\right\}.$
Optimization proceeds as follows:
  • we first choose $\gamma^{\star}$ for each design $\mathcal{E},\Sigma(\mathcal{E})$. Importantly, the choice of $\gamma^{\star}$ depends on the class of designs available to researchers ($\mathcal{D}$), because $\mathcal{D}$ determines the oracle values $\alpha^{\star},\beta^{\star}$ in the denominators of the objective function (a larger class of designs improves the oracle solution);
  • the choice of the experimental variance $\Sigma$ (how precise each experiment is under budget constraints) also depends on the bias component $\beta$ because $\beta$ affects how we select the shrinkage weights $\gamma^{\star}$ in the previous step;
  • the optimal design $\mathcal{E}^{\star}$ then accounts for both choices of $\gamma^{\star}$ and $\Sigma^{\star}$.
3.3 Example with two parameters
Consider the two-parameters model of Setting 2 where for simplicity we fix the choice of the variances $\sigma^{2},v^{2}>0$. The smallest variance and worst–case squared bias (per unit $B^{2}$) are
$\alpha^{\star}=\min_{k\in\{1,2\}}\underbrace{\left\{\ \omega_{-k}^{2}\sigma_{-k}^{2}\;+\;\omega_{k}^{2}\frac{\sigma_{k}^{2}\,v_{k}^{2}}{\sigma_{k}^{2}+v_{k}^{2}}\ \right\}}_{\text{Variance with variance-optimal }\gamma},\qquad\beta^{\star}=\min_{k\in\{1,2\}}\underbrace{\omega_{k}^{2}}_{\text{Bias}^{2}/B^{2}\text{ with bias-optimal }\gamma}.$
Here, $\alpha^{\star}$ follows from the variance-only optimal shrinkage $\gamma_{k}=\sigma_{k}^{2}/(\sigma_{k}^{2}+v_{k}^{2})$ (i.e., the oracle choice when $B=0$); $\beta^{\star}$ follows from the fact that the oracle that minimizes the worst-case bias chooses the experiment $k$ with the largest $\omega_{k}$, sets $\gamma_{k}=1$ and incurs a bias proportional to $|\omega_{-k}|$, since no experiment is conducted for k-k.
The variance and worst-case squared bias (divided by $B^{2}$) incurred by an analyst choosing experiment $j$ and $\gamma_{j}$ are, respectively
$\alpha(j,\gamma_{j})$ $\equiv\underbrace{\omega_{-j}^{2}\sigma_{-j}^{2}}_{\text{var obs estimate}}+\underbrace{\omega_{j}^{2}\!\big[(1-\gamma_{j})^{2}\sigma_{j}^{2}+\gamma_{j}^{2}v_{j}^{2}\big]}_{\text{var exp and obs estimate}},\quad\beta(j,\gamma_{j})$ $\equiv\big(\underbrace{|\omega_{-j}|}_{\begin{subarray}{c}\text{bias obs}\\ \text{estimate}/B\end{subarray}}+\underbrace{(1-\gamma_{j})|\omega_{j}|}_{\begin{subarray}{c}\text{bias obs minus}\\ \text{ bias exp}/B\end{subarray}}\big)^{2}.$
By Theorem 1, the worst-case adaptation regret for choosing experiment $\{j\}$ and $\gamma_{j}$, equals
$\mathcal{R}(\{j\},\gamma_{j})=\max\Big\{\underbrace{\alpha(j,\gamma_{j})/\alpha^{\star}}_{\text{Variance/}\alpha^{\star}},\ \underbrace{\beta(j,\gamma_{j})/\beta^{\star}}_{\text{Bias}^{2}/(B^{2}\beta^{\star})}\Big\}.$
Both maximand components are evaluated with the same choice of $\gamma$.
3.3.1 Choosing $\gamma_{j}$ for given design $\{j\}$
The first step is to choose the shrinkage weight $\gamma$, noting that regret is computed relative to an oracle that optimizes both $\gamma$ and the experiment design $\{j\}$. Minimizing $\mathcal{R}$ over $\gamma_{j}$ yields a rule that either equalizes the two components or selects a boundary solution.
Corollary 1.
Consider Setting 2 and let Assumption 1 hold. Then
[…]
Proof.
See Appendix B.2. ∎
At the boundaries, if the relative variance $\alpha/\alpha^{\star}$ is always smaller than the relative bias $\beta/\beta^{\star}$, the optimal weight is $\gamma_{j}^{\star}=1$ (i.e., use only the experimental estimate). If instead the variance term always dominates the bias term, the optimal choice is the variance–minimizing weight $\gamma_{j}^{\star}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$. In most applications, the solution is interior: we choose $\gamma_{j}^{\star}$ strictly between the boundaries to equalize the two contributions, $\frac{\alpha(j,\gamma_{j}^{\star})}{\alpha^{\star}}=\frac{\beta(j,\gamma_{j}^{\star})}{\beta^{\star}},$ so that the variance ratio matches the worst–case squared bias ratio.
The optimal solution depends not only on the features of the experiment design $j$ but also on the features of the other feasible design j-j ($v_{-j}^{2},\sigma_{-j}^{2}$ and $\omega_{-j}$) through $\alpha^{\star}$ and $\beta^{\star}$. This differs from choosing the optimal $\gamma$ for a single fixed design, since the choice of $\gamma$ will then interact with the choice of the class of design the researcher (and oracle) can choose.
Illustration in Figure 2
Figure 2 illustrates three cases for $j=2$. In the top row, we plot $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. In the bottom row, we plot the corresponding $\gamma_{2}^{\star}$.
Observation 1: small $\omega_{2}$ pushes $\gamma_{2}^{\star}$ toward one. The first (left) column sets $v_{1}=v_{2}/2$ (the first experiment is more precise) and $\omega_{1}=1$ (with $\sigma_{1}=\sigma_{2}=1$). For small $\omega_{2}$ (low sensitivity of the second coordinate to bias), the optimal choice is $\gamma_{2}^{\star}=1$. The reason is twofold. First, the bias ratio $\beta(2,\gamma)/\beta^{\star}=\big(|\omega_{1}|+|1-\gamma||\omega_{2}|\big)^{2}/\beta^{\star}$ is normalized by $\beta^{\star}=\min\{|\omega_{1}|,|\omega_{2}|\}^{2}=\omega_{2}^{2}$ when $\omega_{2}<\omega_{1}$; thus even a small observational component $|1-\gamma|>0$ makes the ratio increase substantially. Second, the variance ratio $\alpha(2,\gamma)/\alpha^{\star}$ is multiplied by $\omega_{2}^{2}$ in its $j=2$ contribution, so its dependence on $\gamma$ becomes negligible as $\omega_{2}\downarrow 0$. Together these forces make the bias ratio dominant and push $\gamma_{2}^{\star}$ to the boundary at $1$. The implication is that for parameters with low $\omega_{2}$, we typically shrink $\hat{\theta}_{2}$ more toward the experimental estimate $\tilde{\theta}_{2}^{\mathrm{exp}}$.
Observation 2: when $\omega_{2}$ is of the same order as $\omega_{1}$, an interior solution emerges. As $\omega_{2}$ increases toward $\omega_{1}$ (roughly $0.6\,\omega_{1}$ and above), the normalization ceases to penalize $j=2$ as harshly, and an interior solution appears: $\gamma_{2}^{\star}$ moves down from $1$ toward the variance–only weight, trading off bias and variance. The first kink in the bottom-left panel reflects this regime change, where $\gamma_{2}^{\star}$ is selected to equalize bias and variance.
Observation 3: for $\omega_{2}>1$, $\gamma_{2}^{\star}$ increases again. At $\omega_{2}=\omega_{1}$ we observe a second kink, since the worst–case bias normalizer switches to $\beta^{\star}=\omega_{1}^{2}$. This raises the bias cost of putting weight on the observational estimate; as $\omega_{2}$ increases above $1$, $\gamma_{2}^{\star}$ rises again. This pattern is driven by the oracle benchmark, which optimizes both the estimator and the design.
Observation 4: for $\omega_{2}\gtrsim 1.25$, $\gamma_{2}^{\star}$ increases slowly and stays interior. Around $\omega_{2}\approx 1.25$ we observe a third kink: the variance normalizer $\alpha^{\star}$ switches branches (from favoring $k=1$ to favoring $k=2$ in the variance–only comparison). The optimal $\gamma_{2}^{\star}$ remains interior but grows more slowly, because increases in $\gamma_{2}^{\star}$ now have a larger relative impact on $\alpha/\alpha^{\star}$. [^9] A further kink could occur if the solution transitioned into a pure variance–dominant regime (where the boundary weight is chosen), which does not arise for the range of $\omega_{2}$ shown here.
Observation 5: $\gamma_{2}^{\star}$ increases with experimental precision and decreases with observational precision The second and third columns vary $v_{2}$ and $\sigma_{2}$, respectively. In both, we fix $v_{1}=1$ and take $\omega_{1}=0.9\,\omega_{2}$ (so here $\omega_{1}=0.9$, $\omega_{2}=1$), with $\sigma_{1}=\sigma_{2}=1$ unless varied. Across $v_{2},\sigma_{2}\in[0.5,2]$, the solution is interior. As $v_{2}\downarrow 0.5$ or $\sigma_{2}\uparrow 2$, the optimal weight $\gamma_{2}^{\star}$ increases toward $1$, placing nearly all weight on the experimental estimate. Conversely, as $v_{2}$ becomes large or $\sigma_{2}$ becomes small, $\gamma_{2}^{\star}$ approaches its lower bound, the variance–minimizing weight $\sigma_{2}^{2}/(\sigma_{2}^{2}+v_{2}^{2})$. This is a typical (and desired) behavior of shrinkage estimators.
Table 1: Regimes for $\gamma_{2}^{\star}$ as a function of $\omega_{2}$ for example in Figure 2 with $\sigma_{1}^{2}=\sigma_{2}^{2}=1,v_{1}=1,v_{2}=0.5,\omega_{1}=1$. Each row summarizes each observation 1-4 for $\gamma_{2}^{\star}$.
Regime $\beta^{\star}$ $\alpha^{\star}$ Solution regime ($\gamma_{2}^{\star}$) Trend of $\gamma_{2}^{\star}$
$\boldsymbol{\omega_{2}\ll\omega_{1}}$ $|\omega_{2}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Bias–dominant Constant $\gamma_{2}^{\star}=1$
$\boldsymbol{\omega_{2}<\omega_{1}}$ $|\omega_{2}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Decrease in $\omega_{2}$
$\boldsymbol{\omega_{2}>\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Increase in $\omega_{2}$
$\boldsymbol{\omega_{2}\gg\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{1}^{2}\sigma_{1}^{2}\;+\;\omega_{2}^{2}\dfrac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ Interior Slow increase in $\omega_{2}$
Figure 2: Illustration for experiment $j=2$. Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$. Bottom row: the optimal weight $\gamma_{2}^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
3.3.2 Choosing the design $\{j\}$
After computing $\gamma_{j}^{\star}$, we then optimize over the design index $j$, by choosing
$j^{\star}\in\arg\min_{j\in\{1,2\}}\max\Big\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\ \beta(j,\gamma_{j}^{\star})/\beta^{\star}\Big\}.$ (6)
To build intuition, suppose $\gamma_{j}^{\star}$ is interior for both $j\in\{1,2\}$. Then $\alpha(j,\gamma_{j}^{\star})/\alpha^{\star}=\beta(j,\gamma_{j}^{\star})/\beta^{\star}$, which implies
$j^{\star}\in\arg\min_{j\in\{1,2\}}\underbrace{\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}}_{\text{overall variance at }\gamma_{j}^{\star}}=\arg\min_{j\in\{1,2\}}\underbrace{\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}}_{\text{worst--case bias}/|B|\text{ at }\gamma_{j}^{\star}}.$
In this case, optimal $\gamma_{j}^{\star}$ balances variance and worst–case bias. Because $\alpha^{\star}$ and $\beta^{\star}$ do not depend on $j$, $j^{\star}$ minimizes either (and therefore both) criteria $\alpha(j,\gamma_{j}^{\star})$ and $\beta(j,\gamma_{j}^{\star})$.
A second case arises when one experiment (say $j=1$) is bias–dominant (i.e., $\gamma_{1}^{\star}=1$) or variance–dominant (i.e., $\gamma_{1}^{\star}=\sigma_{1}^{2}/(\sigma_{1}^{2}+v_{1}^{2})$), while the other admits an interior solution $\gamma_{2}^{\star}$ for which $\beta(2,\gamma_{2}^{\star})/\beta^{\star}=\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$. In this case, the optimal design selects the experiment that yields the smaller value of the dominating criterion (worst–case bias in the bias–dominant scenario, variance in the variance–dominant scenario).
Although we view these as the leading scenarios, cases at the extreme boundaries can occur. The first is when the relative variance $\alpha/\alpha^{\star}$ or the worst–case bias $\beta/\beta^{\star}$ uniformly dominates the other for both experiments. Then the optimal design minimizes the dominating term; alternatively, if for one experiment $j$ the variance ratio $\alpha/\alpha^{\star}$ is dominant while for the other experiment j-j the worst–case bias ratio $\beta/\beta^{\star}$ is dominant, the minimax regret is the smaller of the two dominating values. These are desiderable properties as the optimal design always prioritizes the dominating term. Table 2 summarizes the cases.
Table 2: Optimal shrinkage and experiment choice across different regimes. The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases). The other cases correspond to boundary solutions.
Case (condition) Optimal $\gamma_{j}^{\star}$ Optimal $j^{\star}$ solves Solution $(j^{\star},\gamma_{j}^{\star})$
Leading cases
Case 1 no bias/ variance dominance $\gamma_{j}^{\star}$ in the interior $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star}$ minimizes bias and variance
($\alpha(j,\gamma_{j})/\alpha^{\star}-\beta(j,\gamma_{j})/\beta^{\star}$ can flip sign) for $j\in\{1,2\}$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $\gamma_{j}^{\star}$ balances bias/variance
Case 2 $j=1$ is variance dominant $\gamma_{1}^{\star}=\frac{\sigma_{1}^{2}}{v_{1}^{2}+\sigma_{1}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes variance
$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
Case 3 $j=1$ is bias dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes bias
$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
Other (boundary) solutions
Case 4 Both $j\in\{1,2\}$ are variance-dominant $\gamma_{j}^{\star}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\frac{\sigma_{j}^{2}v_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}\right\}$ Minimizes variance
($\alpha/\alpha^{\star}>\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
Case 5 Both $j\in\{1,2\}$ are bias-dominant $\gamma_{j}^{\star}=1$ $\mathrm{arg}\max_{j}|\omega_{j}|$ Minimizes bias
($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
Case 6 bias dominant vs. variance dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}$ Minimizes bias $\beta/\beta^{\star}$ of $j=1$
($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for $j=1$ and vice versa for $j=2$) $\gamma_{2}^{\star}=\frac{\sigma_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ $\left\{\frac{|\omega_{2}|^{2}}{|\omega_{1}|^{2}}1\{j=1\}+(\omega_{1}^{2}\sigma_{1}^{2}+\omega_{2}^{2}\frac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}})1\{j=2\}\right\}$ vs. variance $\alpha/\alpha^{\star}$ of $j=2$
Illustration in Figure 3
For illustration, Figure 3 compares the maximum regret of choosing experiment $j=1$ versus $j=2$ in three scenarios. The vertical dashed line marks the value where the two curves intersect and the designer is indifferent between $j=1$ and $j=2$. In each column, we keep the same parameterizations as in Figure 2.
Observation 1: The regret for $j=2$ decreases rapidly in $\omega_{2}$ and then decreases more slowly. In the first column, we vary $\omega_{2}$ with the first experiment $j=1$ having a smaller experimental variance, $v_{1}=v_{2}/2$. The panel shows that a small $\omega_{2}$ makes the regret of choosing experiment $2$ larger than that of choosing experiment $1$. This is because a small $\omega_{2}$ corresponds to a small bias from not choosing $j=2$. As $\omega_{2}$ increases, the max–regret curve for $j=2$ declines until it reaches $\omega_{2}\approx 1.25$; after this point, the regret curve for $j=2$ rises slowly, since a further increase in $\omega_{2}$ is associated with an increase in estimator variance.
Observation 2: The regret for $j=1$ decreases slowly and then increases rapidly in $\omega_{2}$. The regret curve for choosing $j=1$ as we vary $\omega_{2}$ (fixing $\omega_{1}=1$) first decreases slowly and then increases. The reason is that for $\omega_{2}<\omega_{1}$, $\gamma_{1}^{\star}$ is an interior solution and the oracle squared bias is $\beta^{\star}=\omega_{2}^{2}$. Therefore, an increase in $\omega_{2}^{2}$ raises the oracle’s bias. When $\omega_{2}>\omega_{1}$, $\gamma_{1}^{\star}$ becomes a boundary solution and $\beta^{\star}=\omega_{1}^{2}$. In this case, a larger $\omega_{2}$ increases the variance of the estimator while the oracle bias $\beta^{\star}$ remains constant in $\omega_{2}$. As a result, it becomes more attractive for the analyst to run the experiment with $j=2$.
Observation 3: The regret for $j=2$ is monotonically increasing in $v_{2}^{2}$, and the opposite holds for $j=1$. The second plot shows that the regret from choosing $j=2$ increases monotonically with the experimental variance $v_{2}^{2}$, while the regret from choosing $j=1$ decreases correspondingly. The kinks in the curves are driven by regime shifts in the oracle solutions $(\beta^{\star},\alpha^{\star})$.
Observation 4: The regret for $j=2$ is monotonically decreasing in $\sigma_{2}^{2}$. The third plot shows that the regret for $j=2$ decreases monotonically with the observational variance $\sigma_{2}^{2}$. This is expected, since a larger $\sigma_{2}^{2}$ makes choosing $j=2$ more appealing relative to relying on observation.
Observation 5: The regret for $j=1$ first decreases slowly and then increases rapidly in $\sigma_{2}^{2}$. As we vary $\sigma_{2}^{2}$, the regret for $j=1$ is initially (very) slowly decreasing. The reason is that a larger $\sigma_{2}^{2}$ increases the oracle variance $\alpha^{\star}$ faster than the variance of the estimator for $j=1$ (with $\gamma_{1}^{\star}\approx 1$). However, this behavior does not affect the optimal solution: for smaller values of $\sigma_{2}^{2}$, choosing $j=1$ remains preferable to choosing $j=2$. When $\sigma_{2}^{2}$ is larger than a tipping point, however, the regret of choosing $j=1$ increases rapidly with $\sigma_{2}^{2}$, making the first experiment no longer preferable to the second. This aligns with the intuition that, as observational variance grows, we should favor conducting the experiment for $j=2$.
In sum, these patterns show how our framework disentangles the competing forces across signal strength and observational/experimental noise, guiding the analyst to transparent design choices even in complex regimes. Table 3 summarizes the discussion.
Table 3: Qualitative behavior of maximum regret as $\omega_{2}$, $v_{2}$, or $\sigma_{2}$ increase (rows) in Figure 3. Entries summarize the direction and relative speed of change in the maximum regret for choosing experiment $j=1$ or $j=2$.
Small regime ($\omega_{2}<\omega_{1}$) Large regime ($\omega_{2}>\omega_{1}$)
Trend regret $j=1$ regret $j=2$ regret $j=1$ regret $j=2$
$\uparrow~\omega_{2}$ slow decrease fast decrease fast increase slow increase
$\uparrow~v_{2}$ decrease increase decrease increase
$\uparrow~\sigma_{2}$ slow decrease decrease fast increase decrease
Figure 3: Regret comparisons. Top row: $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Bottom row: optimal weights $\gamma_{j}^{\star}$ for $j=1,2$, which explain how sensitivity and precision interact to drive the design switch at $x^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
4 Extensions
4.1 Nonlinear estimand $\tau(\theta)$
This subsection extends the framework to smooth, potentially nonlinear $\tau(\theta)$, and consider a local asymptotic framework in the spirit of Andrews et al. (2020); Bonhomme and Weidner (2022).
Setup.
Take any given $(\mathcal{E},\Sigma)$ and $\gamma$. Let $\tilde{\theta}^{\mathrm{obs}}$ and $\tilde{\theta}^{\mathrm{exp}}$ denote the corresponding observational and experimental estimators of $\theta\in\mathbb{R}^{p}$. Assume a common $\sqrt{n}$ rate and the following local asymptotics:
$\sqrt{n}\!\left(\tilde{\theta}^{\mathrm{exp}}-\theta\right)$ $\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,\Sigma^{\mathrm{exp}}),$
$\sqrt{n}\!\left(\tilde{\theta}^{\mathrm{obs}}-\theta-b_{n}\right)$ $\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,\Sigma^{\mathrm{obs}}),\qquad b_{n}\to 0,$
where $b_{n}\in\mathbb{R}^{p}$ is an unknown drift capturing local misspecification, and $\Sigma^{\mathrm{exp}},\Sigma^{\mathrm{obs}}$ are finite positive–definite matrices. We take the experimental and observational estimators to be asymptotically independent with common rates; this is not necessary but simplifies exposition. Thus $\tilde{\theta}^{\mathrm{obs}}\stackrel{{\scriptstyle p}}{{\to}}\theta$ (consistency), while $\sqrt{n}\,b_{n}$ may be bounded or even diverge. To justify first order linearization when $\sqrt{n}\,b_{n}$ may grow, we impose $\sqrt{n}\,\|b_{n}\|^{2}\to 0$ (e.g., $\|b_{n}\|=n^{-\alpha}$ with $\alpha\in(1/4,1/2]$). This implies that even if $b_{n}$ converges to zero, the asymptotic bias is possibly comparable to, smaller or larger than the asymptotic variance, therefore imposing weak restrictions on the magnitude of $\sqrt{n}b_{n}$ relative to the asymptotic variance.
Linearization of $\tau$.
Let $\tau:\mathbb{R}^{p}\to\mathbb{R}$ be twice continuously differentiable in a neighborhood of $\theta$, with Lipschitz gradient and bounded Hessian. A Taylor expansion around $\theta$ yields
$\tau(\hat{\theta}_{\gamma})-\tau(\theta)=\omega(\theta)^{\top}(\hat{\theta}_{\gamma}-\theta)\;+\;\tfrac{1}{2}(\hat{\theta}_{\gamma}-\theta)^{\top}H_{\tau}(\bar{\theta})\,(\hat{\theta}_{\gamma}-\theta),$
[…]
$\sqrt{n}\bigl(\tau(\hat{\theta}_{\gamma})-\tau(\theta)\bigr)=\omega(\theta)^{\top}\sqrt{n}\,(\hat{\theta}_{\gamma}-\theta)\;+\;o_{p}(1).$ (7)
Using the fact that $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$, we can consistently estimate the asymptotic MSE $n\mathbb{E}\left[(\tau(\hat{\theta}_{\gamma})-\tau(\theta))^{2}\right]$ by replacing $\omega(\theta)$ with $\omega(\tilde{\theta}^{\mathrm{obs}})\rightarrow_{p}\omega(\theta)$ and ignoring the $o_{p}(1)$ on the right-hand side. As a result all our results continue to hold under such linearization up-to a negligible $o_{p}(1)$.
4.2 Confidence interval length as a regret criterion
In this subsection we consider an alternative loss based on the length of confidence intervals (CIs). For a candidate design $(\mathcal{E},\Sigma(\mathcal{E}))$, shrinkage weights $\gamma$, and a bias vector $b\in\mathbb{R}^{p}$, let $\tilde{\omega}(\mathcal{E},\gamma)$ be as in (3). Define the two–sided, $(1-\eta)$ bias–aware lower and upper bounds for $\tau(\theta)$ by
$\ell_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;-\;z_{1-\eta/2}\,\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},$
$u_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;+\;z_{1-\eta/2}\,\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},$
where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile.
An audience, indexed by a worst–case bias radius $B\geq 0$, forms the worst–case CI
$L_{B}(\mathcal{E},\Sigma,\gamma)~\equiv~\left[\inf_{\|b\|_{\infty}\leq B}\ell_{b}(\mathcal{E},\Sigma,\gamma),\;\sup_{\|b\|_{\infty}\leq B}u_{b}(\mathcal{E},\Sigma,\gamma)\right],$
where we write $|L_{B}(\mathcal{E},\Sigma,\gamma)|$ for its total length. The proportional regret of $(\mathcal{E},\Sigma,\gamma)$ is
$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{B\geq 0}\;\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{\inf_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}\!.$
Theorem 2.
Consider Setting 1 and let Assumption 1 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
[…]
with $\alpha$ and $\beta$ as defined in Equation (4) and $\alpha^{\star},\beta^{\star}$ as defined in Equation (5).
Proof.
See Appendix C.1
Theorem 2 shows that the objective function is the same as the MSE-objective function up to a monotonic (square-root) transformation which does not affect the choice of the minimizer.
4.3 Vector valued estimands
Next, we extend our framework to vector value estimands of the form $\tau(\theta)\in\mathbb{R}^{q}$ for $q\geq 1$.
Setting 3 (Vector-valued estimands).
Consider a vector-valued target estimand with entries $\tau^{\ell}(\theta)\equiv(\omega^{\ell})^{\top}\theta$ for $\ell=1,\dots L$, and $\omega^{\ell},\theta\in\mathbb{R}^{k}$.
For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a positive-definite matrix $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})\subset\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
$\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{obs}}\right]-\theta=b,\qquad\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\right]-\theta_{\mathcal{E}}=0,$
for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance as
$\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E})$
assumed to be known, with uniformly bounded entries and strictly positive definite.
Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators to allow the optimal shrinkage weights to depend on the specific entry of $\tau$.
$\hat{\tau}^{\ell}_{\gamma}\equiv\tau^{\ell}(\hat{\theta}^{\ell}(\gamma)),\qquad\hat{\theta}^{\ell}_{j}(\gamma)=\begin{cases}\gamma_{j}^{\ell}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j}^{\ell})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
where $\gamma=(\gamma^{\ell}_{j})_{\ell\in[L],j\in\mathcal{E}}\in\mathbb{R}^{L\times|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$. Let $\hat{\theta}(\gamma)\equiv(\hat{\theta}^{1}(\gamma),\dots,\hat{\theta}^{L}(\gamma))\in\mathbb{R}^{p\times L}$ and $\hat{\tau}_{\gamma}\equiv(\hat{\tau}^{1}_{\gamma}(\hat{\theta}^{1}(\gamma)),\dots,\hat{\tau}^{L}_{\gamma}(\hat{\theta}^{L}(\gamma)))^{\top}\in\mathbb{R}^{L}$.
Setting 3 is a flexible generalization of Setting 1, allowing each entry of $\tau$ to have its own vector of shrinkage weights.
Similar to the scalar case, we make the following assumption regarding the first-order estimation error.
Assumption 2 (First-order estimation error).
For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume for all $\ell=1,\dots,L$
$\tau^{\ell}(\theta)-\tau^{\ell}\!\big(\hat{\theta}^{\ell}(\gamma)\big)=(\omega^{\ell})^{\top}\,\big(\theta-\hat{\theta}^{\ell}(\gamma)\big),$ (8)
for a known weighting vector $\omega^{\ell}\in\mathbb{R}^{p}$, with $|\omega^{\ell}_{i}|>0$, for all $i,\ell$.
We begin by generalizing the mean squared error to higher-dimensional parameters. For a given design $(\mathcal{E},\Sigma)$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[\|\widehat{\tau}_{\gamma}-\tau\|_{2}^{2}\right],$ (9)
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$, and $\|\cdot\|_{2}$ denotes the $L^{2}$ Euclidean norm.
As before, we introduce more compact notation before proceeding to the main result. For a given $(\mathcal{E},\gamma)$ define
$\tilde{\omega}^{\ell}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega^{\ell}_{j},&j\notin\mathcal{E},\\ \omega^{\ell}_{j}(1-\gamma^{\ell}_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}^{\ell}_{\text{exp},j}(\mathcal{E},\gamma)=\begin{cases}0,&j\notin\mathcal{E},\\ \omega^{\ell}_{j}\gamma^{\ell}_{j},&j\in\mathcal{E},\end{cases}$
and let
$\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}^{\ell}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}^{\ell}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$
Using this notation, we can write $\hat{\tau}^{\ell}_{\gamma}=\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}\,\hat{\theta}$, where $\hat{\theta}=(\tilde{\theta}^{\text{obs}\top},\tilde{\theta}^{\text{exp}\top})^{\top}\in\mathbb{R}^{p+|\mathcal{E}|}$. Stacking across the $\ell$ index gives the estimator for the full vector $\tau$: $\hat{\tau}_{\gamma}=\tilde{\omega}(\mathcal{E},\gamma)\,\hat{\theta}$, where the $\ell^{th}$ row of $\tilde{\omega}(\mathcal{E},\gamma)\in\mathbb{R}^{L\times(p+|\mathcal{E}|)}$ is $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}$.
The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ $\equiv\text{Trace}\left(\tilde{\omega}(\mathcal{E},\gamma)\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\right),$ (10)
$\check{\beta}(\mathcal{E},\gamma)$ $\equiv\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}\bar{v}(\gamma)-(\omega^{\ell}_{\,\mathcal{E}})^{\top}\bar{v}(\gamma)_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(\bar{v}(\gamma)_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
where $\omega^{\ell}_{\mathcal{E}}$ is the subvector of $\omega^{\ell}$ corresponding to the entries indexed by the experiment $\mathcal{E}$, and
$\bar{v}(\gamma)\in\operatorname*{arg\,max}_{\{v\in\mathbb{R}^{p}:|v_{j}|\leq 1\forall j\}}\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\,\mathcal{E}})^{\top}v_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(v_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
where the dependence of $\bar{v}(\gamma)$ on $\mathcal{E}$ is omitted for conciseness.
The smallest worst-case variance and bias are defined as
$\check{\alpha}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\beta(\mathcal{E},\gamma).$
As before, we are interesting in minimizing
$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\equiv\sup_{B\geq 0}\frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}.$
The result below shows how our results directly extend to this case.
Theorem 3.
Consider Setting 3 and let Assumption 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\check{\alpha}(\mathcal{E},\Sigma,\gamma)}{\check{\alpha}^{\star}},\ \ \frac{\check{\beta}(\mathcal{E},\gamma)}{\check{\beta}^{\star}}\right\}.$
Proof.
See Appendix C.2
5 Empirical applications
In this section we provide two empirical applications. In the first application we study the problem of choosing where to conduct an experiment using observational data from Banerjee et al. (2024). In the second application we illustrate the method for choosing treatment arms when estimating a structural model, calibrating our estimation to the PROGRESA experiment (Todd and Wolpin, 2006; Attanasio et al., 2012). We solve the optimization program by solving for each choice of experiments, both over $\gamma$ and the allocation of sample size via numerical optimizationand then enumerate the solutions across all possible choices.
5.1 Choosing experimental participants
As a first exercise, we show how observational evidence can inform whom to recruit into an experiment. We consider the setting in Banerjee et al. (2024), who study how the expansion of microfinance affects village social networks in Karnataka, India.
In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance. Because program rollout was not randomized across villages, they estimate effects using difference-in-differences (DiD). While DiD is informative, the lack of randomized assignment may bias estimates in the presence of selection and lack of parallel-trends (e.g., Marx et al., 2024; Ghanem et al., 2022). Motivated by these limitations, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods. We only use observational variation for choosing the experimental design and estimator. We then use experimental variation generated by Banerjee et al. (2024) to validate our procedure.
Using preliminary observational estimates from Banerjee et al. (2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled? Concretely, we consider designs that randomize the introduction of microfinance in $n$ villages within one or more selected areas and study how the choice of areas and the allocation of $n$ vary with design constraints. Our objective is to design an experiment that yields an externally valid estimate of the effect of microfinance for rural Karnataka. We focus on the first outcome reported by Banerjee et al. (2024) corresponding to the density of the network, i.e., the percentage of connections a random household in a village has relative to the village size.
In practice, implementation costs often depend on how geographically dispersed the study sites are. Coordinating with microfinance partners, training and supervising field teams, and conducting surveys become more expensive when sites are far apart or span multiple administrative units. We therefore consider operational constraints that limit the dispersion of enrolled villages and number of villages enrolled in the area.
Figure 4: Observational villages and area-level DiD effects in Karnataka. The background heat map partitions the state into the four contiguous areas used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value. Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey). Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
Observational study
We use observational evidence from Banerjee et al. (2024) to construct area-level difference-in-differences (DiD) estimates for Karnataka. Specifically, we partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period. Figure 4 maps these four areas and reports the corresponding DiD point estimates. Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas.
Table 4 summarizes, for each area, the number of treated and untreated villages, the pre-treatment sample variances, and the DiD point estimates together with their variances, computed assuming independence across villages. [^10] Alternative variance estimators (e.g., allowing within-area correlation) can be incorporated without changing the design logic.
These area-specific DiD estimates and their variances serve as our primary observational inputs to the experimental design problem. [^11] We use DiD estimates to mimic the estimator used by the authors. An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity.
Table 4: Area-level observational inputs (pre-prior) for Karnataka villages. The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village). For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$.
Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$
11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3
22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140
33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187
44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130
Experimental design
We consider a family of designs that (i) select $E\in\{1,\ldots,4\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$). Here, $E=1$ corresponds to recruiting from a single area, while $E=4$ recruits from all four areas. We examine a grid of sample sizes that varies the number of participants from $10$ up to $104$ ($52$ treated units), with the latter corresponding to the size of the experiment in Banerjee et al. (2024). For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 4. We assume that the variance of a single treated–control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
Results
We begin by studying which areas are selected and how sample is allocated when the total experimental sample is large ($n_{1}=52$) and the number of admissible areas may or may not be constrained. Figure 5 visualizes the resulting allocation across Karnataka’s four observational areas for $E\in\{1,2,3,4\}$. With $E=1$, the algorithm concentrates recruitment in the Tumakuru area—the location with the highest uncertainty in the observational estimates. As $E$ relaxes to $2$ and $3$, additional areas enter and the total sample splits across them in roughly—but not exactly—equal shares, reflecting a trade-off between (i) exploiting heterogeneity in the observational variance and (ii) keeping experimental variance low. When $E=4$, all areas are eligible and the allocation smooths further across space. In the rest of our discussion, we discuss properties when not all areas may be selected (E<4E<4). [^12] The solution for $E=4$ always favors minimizing the bias first, since $\beta^{\star}=0$.
Figure 6 tracks how area-level allocations evolve with the total number of villages. When the experiment is small (e.g., $10$ villages) and $E=1$, the algorithm favors an area with relatively low experimental variance—even if its observational estimate is less uncertain—because, at small $n$, experimental noise is first-order. As the total sample increases, the weight shifts toward areas with more uncertain observational estimates. Consistently, the area with the largest experimental variance (Area 2) is typically excluded unless all four areas can be chosen.
Figure 7 reports the optimal shrinkage $\gamma_{j}^{\star}$ by area as total villages vary. The solution is interior for all $E<4$: $\gamma_{j}^{\star}$ rises with sample size, approaching one from below as experimental noise diminishes. For example, at a moderate sample (around 15 villages), $\gamma^{\star}$ is about $0.65$ when $E\in\{2,3\}$; with $E=1$, concentrating the sample within one area lowers experimental variance and therefore yields a larger $\gamma^{\star}$.
Main MSE comparison
We compare three strategies: (i) our chosen design which chooses the area, sample size and $\gamma^{\star}$ as described in Section 3, (ii) a standard benchmark that selects areas uniformly at random, allocates villages evenly across the chosen areas, and sets $\gamma_{j}=1$ for experimental estimates, and (iii) an oracle that knows the bias vector $b$ and optimizes both design (areas and allocation) and $\gamma$.
Because $b$ is unknown in practice, for this exercise, we calibrate it using the difference between the Hyderabad RCT ATE obtained from the follow-up experiment of Banerjee et al. (2024) and the Karnataka DiD average effect across sites, treating $b$ as common across areas.
Figure 8 plots the ratio $\text{MSE}/\text{MSE}^{\star}$, where $\text{MSE}^{\star}$ is the oracle’s MSE. As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone. Across $E\in\{1,2,3\}$, the random benchmark remains well above the oracle—by roughly 3–5 percentage points (pp) for $E=1$, about 10 pp for $E=2$, and around 20 pp for $E=3$ at the top of the grid—whereas our design tracks the oracle closely (within a few percentage points) even without knowledge of $b$. The gap is the largest when a larger number of areas $E$ can be included, because our procedure can better emulate the oracle by selecting (and differentially weighting) favorable areas. On the other hand, the random benchmark ignores information about observational and experimental variance heterogeneity. As sample size increases, the gap between the two stabilizes: with large $n_{1}$, optimal shrinkage approaches $\gamma^{\star}\!\approx\!1$ and an even allocation across the selected areas is nearly optimal once the area set is well chosen.
This empirical calibration highlights the benefit of the procedure that is able to track the oracle even in the absence of knowledge of the bias $b$, while significantly outperforming standard alternative benchmarks such as a random allocation strategy.
Figure 5: Optimal area selection and village allocation for a large experiment ($n_{1}=52$) under constraints on the number of areas $E$. Each panel corresponds to a value of $E$; shading indicates the total number of recruited villages in each area. With $E=1$ the design concentrates in the area with the noisiest observational estimate; as $E$ increases, allocation spreads across areas, reflecting a trade-off between observational uncertainty and experimental variance.
Figure 6: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for each constraint on the number of eligible areas $E$. Lines show the number of villages assigned to each area.
Figure 7: Optimal shrinkage $\gamma_{j}^{\star}$ by area as a function of total treated villages $n_{1}$ (from ten to fifty-two villages), under different constraints $E$. For $E<4$, the solution is interior ($0<\gamma_{j}^{\star}<1$) and increases with $n_{1}$, approaching one as experimental noise falls.
Figure 8: Relative MSE across designs under an hypothetical scenario where the bias equal the difference between the experimental and observational ATE estimate from Banerjee et al. (2024) as $n_{1}$ varies (from ten to fifty-two villages). The figure reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$ for (i) the proposed design (area targeting and optimal $\gamma^{\star}$), (ii) a random benchmark (uniform area selection and equal allocation), where $\mathrm{MSE}^{*}$ denotes the MSE of the oracle that knows the bias, as $n_{1}$ varies and for each $E$. The proposed design closely tracks the oracle across $n_{1}$ and $E$, while the random benchmark remains substantially above, especially when $E$ is small.
5.2 Choosing treatment arms for model estimation
Next, we illustrate how our framework selects among treatment arms when the goal is to estimate and deploy a structural model. Combining experimental or pre-program observational data with structural models is increasingly used for program evaluation (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022). Because large-scale experimentation may be infeasible due to cost or fairness constraints (Muralidharan and Niehaus, 2017), researchers are often interested in designing small-scale experiments whose information, combined with a model, is informative about general-equilibrium (GE) effects.
We consider a researcher evaluating a conditional cash-transfer (CCT) for sending children to school in rural Kenya where such a program is not in place. Due to budget and feasibility constraints, the researcher can only randomize at small scale (partial equilibrium), but ultimately wishes to predict large-scale GE effects. As preliminary observational estimates, we use information from the Mexican PROGRESA experiment.
5.2.1 Model
Individual choice model
Following Bonhomme and Weidner (2022) (and in turn Todd and Wolpin (2006)), let $S\in\{0,1\}$ denote school attendance, $C$ consumption, $Y$ (pre-transfer) household income, $W$ the child’s potential wage, and $t$ the stipend when enrolled. Abstracting from covariates for exposition (we will introduce covariates in the estimation), utility is
$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{3}-\xi_{1}-\xi_{2})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (11)
with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS. The parametrization $(\xi_{3}-\xi_{1}-\xi_{2})$ is without loss and simplifies expressions below. Enrollment satisfies
$S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$
Letting $Z(Y,W,t)\equiv\xi_{1}W-\xi_{2}Y-\xi_{3}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$.
General equilibrium effects
We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya. GE feedback is allowed through income and wages: for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, we assume
$Y_{i}(t)=y(t)+\varepsilon_{Yi},\qquad W_{i}(t)=w(t)+\varepsilon_{Wi}.$
Let $y_{0}\equiv\partial_{t}y(t)|_{t=0}$ and $w_{0}\equiv\partial_{t}w(t)|_{t=0}$. Define $\phi_{0}\equiv\mathbb{E}\!\left[\phi\!\big(-Z\big(Y(0),W(0),0\big)\big)\right]$. Our estimand of interest is the marginal effect of an increase in a small transfer $t$ to all eligible households
$\left.\frac{\partial\,\mathbb{E}\!\left[\Pr(S{=}1\mid Y(t),W(t),t)\right]}{\partial t}\right|_{t=0}=\;\phi_{0}\cdot\big(\,\xi_{3}+\xi_{2}y_{0}-\xi_{1}w_{0}\,\big),$ (12)
which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{2}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$.
Income and wage effects
We estimate $y_{0}=1.5$ using a local fiscal/income multiplier $M_{\text{tot}}\approx 2.5$ from Egger et al. (2022), which in turn implies $y_{0}=M_{\text{tot}}-1$. [^13] Egger et al. (2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$. We write $w_{0}$ as a function of the parameters: Under simple market clearing for labor supply and demand, we can show [^14] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$. Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working. From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{3}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{1}$ at $t=0$. Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{3}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{1}H$. Dividing the equation by $H$ we obtain the desired result.
$w_{0}=\;\frac{\phi_{0}\,\xi_{3}}{\phi_{0}\,\xi_{1}-d},$ (13)
where $d$ denotes the slope of the demand curve divided by total hours of work. We calibrate $d$ to $0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained form the labor demand elasticity in Espey and Thilmany (2000). For expositional convenience, we treat $y_{0},d$ as constant, while $\phi_{0}\xi_{3}$ and $\phi_{0}\xi_{1}$ are parameters of the model and $w_{0}$ is a (non-linear) function of such parameters (it is possible to treat also $y_{0},d$ as additional possibly misspecified parameters, omitted for brevity). [^15] We treat $d$ as a fixed constant given the extensive literature on estimating demand elasiticities, and the fact that $S_{0}$, $W_{0}$ can be calibrated to pre-experimental data (for simplicity here using the PROGRESA pre-experimental data), and $y_{0}$ because computed experimentally in Kenya. When we are concerned with misspecification of those, researchers can augment $\theta$ to also incorporate $y_{0},d$ without having to estimate those estimated in the experiment.
Map to the design parameters
We can now define the parameters of interest as
$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{3},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{1},$ (14)
with
$\tau(\theta)\;=\;\theta_{2}\;+\;y_{0}\,\theta_{1}\;+\;w_{0}(\theta)\,\theta_{3},\qquad w_{0}(\theta)\;=\;\frac{\theta_{2}}{-\theta_{3}-d},$ (15)
where the explicit form of $w_{0}(\theta)$ follows from (13). Given observational estimates $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2},\tilde{\theta}^{\mathrm{obs}}_{3})$ described below, and following Section 4.1, we use $\omega=\frac{\partial\tau}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}$. [^16] Because $w_{0}$ is a function of $\theta$, we have $\frac{\partial\tau}{\partial\theta_{1}}=y_{0},\frac{\partial\tau}{\partial\theta_{2}}=\frac{d}{\theta_{3}+d},\frac{\partial\tau}{\partial\theta_{3}}=-\,\frac{\theta_{2}\,d}{(-\theta_{3}-d)^{2}}.$
5.2.2 Observational study and design of the experiment
Using experimental data from Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ via probit with standard controls (age, distance to school, eligibility, year, highest grade attained). Estimation is conducted separately by gender, focusing here on the effects on female students. We pool treated and control observations to estimate the schooling effect $\xi_{4}$, controlling for the individual-level subsidy. In turn, we obtain observational estimates $\big(\tilde{\theta}_{1}^{\mathrm{obs}},\tilde{\theta}_{2}^{\mathrm{obs}},\tilde{\theta}_{3}^{\mathrm{obs}}\big)$ that map to the model parametrization. Standard errors are constructed using the Delta method with clustering at the village level. Point estimates, estimated sensitivity $\omega$, and standard errors are reported below.
Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sigma$
[…]
Observational estimates for female students, corresponding sensitivity parameter $\omega$, and standard errors.
$\Sigma^{\mathrm{obs}}\;=\;\begin{bmatrix}4.31&-11.31&5.57\\ -11.31&973.11&-126.16\\ 5.57&-126.16&1038.56\end{bmatrix}\times e^{-9}$
Variance–covariance matrix of observational estimates for female students.
We observe heterogeneity in both the sensitivity weights $\omega$ and (more significantly) the precision of the observational estimates. Our goal is to trade off these features, together with prospective experimental variation, to design the experiment in Kenya. Because the setting differs in time and country, PROGRESA-based estimates provide informative but possibly biased baselines.
Candidate small-scale treatment-arms in Kenya (partial equilibrium)
We consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, as they may not be able to randomize treatments across the full population of interest. We examine three possible experiments for a given sample size $n$:
  • $j=\{1\}$: Unconditional transfer (income shock). The researcher randomizes a small income shock to a small fraction of households (implying no general-equilibrium effects). Under a first-order (Taylor) approximation, the experiment identifies
    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,0)]}{\partial Y}\right|_{t=0}=\theta_{1},$
    with precision $v_{1}^{2}/n$. The researcher optimizes over $\gamma_{1}$, the weight used to combine the experimental and observational estimates of $\theta_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
  • $j=\{2\}$: Conditional cash transfer (stipend). The researcher randomizes a small stipend $t$, conditional on attending school, to a small fraction of households (with prices held fixed). Under a first-order approximation, this identifies
    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,t)]}{\partial t}\right|_{t=0}=\theta_{2},$
    with precision $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, the weight used to combine the experimental and observational estimates of $\theta_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
  • $j=\{1,2\}$: Two-arm design. The researcher runs (j=1j{=}1) and (j=2j{=}2) on independent samples, identifying $(\theta_{1},\theta_{2})$ with precisions $(v_{1}^{2}/n_{1},\ v_{2}^{2}/n_{2})$ where $n_{1}+n_{2}=n$. They optimize over both $\gamma_{1}$ and $\gamma_{2}$, while $\theta_{3}$ is calibrated to the PROGRESA study.
We consider two scenarios. In the first, the researcher (and oracle) can choose one or both experiments. In the second, they can choose only one of the two (either $j=\{1\}$ or $j=\{2\}$). Regret is defined accordingly, conditional on the set of feasible designs.
We calibrate $v_{1}^{2}$ as $\sigma_{1}^{2}\times n^{\mathrm{obs}}$, where $\sigma_{1}^{2}$ is the variance of $\tilde{\theta}_{1}^{\mathrm{obs}}$ and $n^{\mathrm{obs}}$ is the observational-sample size ($n^{\mathrm{obs}}=1089$ for the female sample using data from eligible students in Todd and Wolpin (2006)). We calibrate $v_{2}^{2}$ analogously as $\sigma_{2}^{2}\times n^{\mathrm{obs}}$.
Experiment type Identified parameter Choice variables Experiment standard error Description
$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $6.56e^{-5}/\sqrt{n}$ Unconditional transfer (income shock, UCT). Randomize an income shock to a small fraction of households.
$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $9.86e^{-4}/\sqrt{n}$ Conditional cash transfer (education shock, CCT). Randomize a conditional transfer to a small fraction of households.
$j=\{1,2\}$ $(\theta_{1},\,\theta_{2})$ $\gamma_{1},\ \gamma_{2},\ n_{1}$ $6.56e^{-5}/\sqrt{n_{1}}$ (arm 1), $9.86e^{-4}/\sqrt{n-n_{1}}$ (arm 2) Two-arm design. Run (j=1j{=}1) and (j=2j{=}2) on independent samples with sample sizes respectively $n_{1},n-n_{1}$.
Table 5: Design options, identified parameters, choice variables, and per-unit variances.
5.2.3 Results
Figure 9: Regret–optimal sample allocation across treatment arms when both the unconditional cash transfer (UCT; income shock) and the conditional cash transfer (CCT; stipend) arms are available. The optimal solution is interior: the vast majority of participants (over 90%) are assigned to CCT, with a small but nonzero fraction assigned to UCT to hedge misspecification. This pattern reflects the higher payoff to learning about CCT in this application relative to UCT.
Figure 10: Optimal shrinkage weights and design choice as functions of the total sample size $n$. Left panel: When both arms are run, the weight on the UCT experimental estimator is essentially one across $n$, while the weight on the CCT experimental estimator, $\gamma^{\star}_{\text{CCT}}$, rises with precision (from about 0.5 at small $n$ to roughly 0.9 by $n=1000$), reflecting the increasing value of experimental evidence. Right panel: When restricted to a single arm, the optimal choice switches from UCT at small $n$ (variance dominates under CCT) to CCT once $n$ crosses a threshold (around $n\approx 500$), where bias considerations dominate and favor CCT. Notes: $\gamma_{j}^{\star}$ is the optimal weight on the experimental (vs. observational) estimator for parameter $j$; we set $\gamma_{j}^{\star}=0$ to indicate that arm $j$ is not chosen.
Figure 11: Worst–case regret by design. Left panel: Allowing both arms with an interior allocation yields regret that approaches one as $n$ increases, indicating near–oracle performance under our normalization. In contrast, single–arm designs exhibit nonvanishing regret because bias remains first–order even as variance shrinks. Right panel: Focusing on the single–arm problem, UCT dominates only at small $n$; beyond the same threshold ($\approx 500$), CCT delivers uniformly lower regret as variance becomes second–order relative to bias. Notes: Regret is the worst–case objective normalized so that a value of one corresponds to the oracle benchmark; UCT = unconditional cash transfer, CCT = conditional cash transfer.
We now summarize results as we vary both the set of experiments the researcher can run and the total sample size $n$. Specifically, Figure 9 reports the sample allocation implied by the regret-optimal design when both treatment arms are available. The optimal allocation is interior: most participants (over roughly $90\%$) are assigned to the CCT arm (education), with a small but nonzero fraction assigned to the UCT arm (income shock). Two forces rationalize this pattern. First, the CCT parameter is more misspecification-sensitive ($\omega_{2}\approx 1.98$ vs. $\omega_{1}\approx 1.5$), so generating experimental evidence on CCT has a larger payoff in bias reduction. Second, the income effect $\theta_{1}$’s variance is about an order of magnitude smaller than the stipend effect $\theta_{2}$’s variance in our PROGRESA calibration, lowering the marginal returns to learn about UCT relative to CCT. Hence, while the designer keeps some allocation on UCT, most of the sample is placed on CCT; our framework makes this bias–variance trade-off explicit.
Alongside the sample allocation, our method optimizes the shrinkage weights on the parameters of interest. Figure 10 (left panel) shows that when both arms can be run, we place essentially full weight on the experimental estimator for UCT ($\gamma^{\star}_{\text{UCT}}\approx 1$ across $n$), while for CCT we gradually shift toward the experimental estimator as precision improves: $\gamma^{\star}_{\text{CCT}}$ rises from about $0.5$ at small $n$ to approximately $0.9$ by $n=1000$. Intuitively, higher $n$ lowers experimental variance, making it optimal to rely more heavily on experimental evidence for the CCT parameter.
It is also informative to compare optimal choices when the researcher is restricted to one arm. Figure 10 (right panel) displays the selected arm as a function of $n$ (together with the corresponding $\gamma_{j}^{\star}$). We encode “not chosen” as $\gamma_{j}^{\star}=0$. For small samples ($n\lesssim 500$), the optimal choice is UCT: the experimental CCT estimator would be too noisy, so the resulting variance under CCT dominates, while the bias advantage of CCT over UCT is limited because $\omega_{1}$ and $\omega_{2}$ are of similar magnitude. Once $n$ crosses a critical threshold (around $n\approx 500$), CCT becomes dominant: by combining observational and experimental estimates for CCT, we can achieve comparable variance, and the worst-case bias is lower than under UCT.
Finally, Figure 11 reports worst-case regret for each design choice. The left panel shows that the regret of our method (green line, running both arms with an interior allocation) declines toward one as $n$ increases, reflecting near-oracle performance under our normalization. In contrast, single-arm designs exhibit nonvanishing regret because bias remains first order even as variance shrinks. The right panel focuses on the single-arm problem: both the oracle and the researcher can only choose a single arm. In this case, UCT is preferable only at small $n$; beyond the same threshold ($\approx 500$), CCT yields uniformly lower regret as variance becomes second order relative to bias. The regret vanishes relative to the oracle that can only choose a single arm.
These results illustrate how our approach systematically balances precision and misspecification risk to balance experimental variation with structural models.
6 Implications for practice
This paper studies experimental design in the presence of complementary observational evidence. Practitioners often have access to observational inputs—e.g., estimates from models trained on existing data (which may be misspecified), results from prior experiments that may not transport to the target context, or estimates from observational designs that may suffer from confounding. Our goal is to combine such evidence with new experimental data, which deliver unbiased estimators for specific parameters of interest. We ask which experiment to run, and how to run it, under budget constraints that limit the number of arms and/or impose per-unit (variable) costs on effective sample size.
A key challenge is that the bias of the observational estimators is unknown in practice. We adopt a minimax proportional regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias. This reveals a fundamental trade-off between precision and robustness. The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap). We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
In practice, the workflow is:
  • Define the estimand(s) of interest. Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$. For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations. Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify.
  • Assemble informative observational evidence. Collect observational estimates $\tilde{\theta}^{\rm obs}$ and their covariance $\tilde{\Sigma}^{\rm obs}$. These serve as informative (but potentially biased) baselines. Such evidence may come from observational designs, structural estimates with pre-experimental data, or prior experiments conducted in different contexts or periods.
  • Compute the sensitivity parameters. Using the observational baseline, compute the sensitivity weights $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\big|_{\tilde{\theta}^{\rm obs}}$, which quantify how bias in each coordinate of $\theta$ propagates to $\tau(\theta)$. Large $|\omega_{j}|$ indicates greater payoff to learning the $j$th component.
  • Specify feasibility constraints and calibrate experimental variance. Enumerate the admissible design set $\mathcal{S}$. That is, $\mathcal{S}$ is the set containing combinations of parameters that can be learned jointly via experiments (which arms, sites, or mechanisms can be randomized). Define, for each combination of parameters $\mathcal{E}\in\mathcal{S}$, the corresponding precision set $\mathcal{G}(\mathcal{E})$ as the feasible set of experimental variances for a given experiment $\mathcal{E}$. This requires specifying how the sample size $n_{j}$ allocated to each arm maps to the variance of each experimental estimate. Calibrate per-unit experimental variances using pilot studies or historical data.
  • Run the method and decide. For each feasible set of experiments $\mathcal{E}\in\mathcal{S}$ and precisions $\Sigma\in\mathcal{G}(\mathcal{E})$, we first optimize over the shrinkage vector $\gamma$ (how to weight observational vs. experimental estimates). We then optimize over the precision/allocation $\Sigma$ (e.g., the sample size $n_{j}$ assigned to each arm). Finally, we optimize over the experiment choice $\mathcal{E}\in\mathcal{S}$. In practice, this involves enumerating feasible $\mathcal{E}$ and simple numerical optimization for $(\gamma,\Sigma)$. The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$).
  • Reporting and diagnostics. In a pre-analysis plan, we recommend reporting (i) the estimands of interest; (ii) the chosen arm(s) and the final sample size allocation, (iii) the shrinkage weights $\gamma^{\star}$ by parameter between the observational and experimental study, and (iv) the two normalized components $\alpha/\alpha^{\star}$ (variance-gap) and $\beta/\beta^{\star}$ (bias-gap), that can help assess the quality of the allocation. Before committing to a single experiment choice, researchers may also want to explore sensitivity to a range of values for the experimental variances.
The applicability of our framework spans a wide range of settings in economics and beyond. Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024). In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis. Beyond economics, medical applications include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025). In all cases, the same recipe identifies which parameters to target, how much to learn about each, and how to blend experimental and observational evidence.
Several open questions remain for future research. These include settings where researchers have well-specified priors about bias (e.g., from meta-analysis), sequential or adaptive experimental choices, and environments where the exogenous source of variation yields only partial identification of some parameters of interest.
Appendix A Additional figures
Figure 12: Share of experimental papers published in AEA journals also presenting experimental results in combination with observational estimates (either from a reduced form, or from a structural model or from both).
Appendix B Proofs of main results
B.1 Proof of Theorem 1
Preliminary notation
It suffices to prove the theorem for any given feasible design and weights $(\mathcal{E},\Sigma(\mathcal{E}),\gamma)$. Throughout, we denote by $\mathcal{E}^{c}\equiv\{1,\dots,p\}\setminus\mathcal{E}$ the complement of $\mathcal{E}$.
Let $\tilde{\omega}(\mathcal{E},\gamma)=[\,\tilde{\omega}^{\mathrm{obs}}(\mathcal{E},\gamma)^{\top},\ \tilde{\omega}^{\mathrm{exp}}(\mathcal{E},\gamma)^{\top}\,]^{\top}$ as in (3), with
$\tilde{\omega}^{\mathrm{obs}}_{\mathcal{E}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ(1-\gamma),\quad\tilde{\omega}^{\mathrm{obs}}_{\mathcal{E}^{c}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}^{c}},\quad\tilde{\omega}^{\mathrm{exp}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ\gamma,$
so that $\tilde{\omega}^{\mathrm{obs}}\in\mathbb{R}^{p}$ and $\tilde{\omega}^{\mathrm{exp}}\in\mathbb{R}^{|\mathcal{E}|}$. We let $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ and denote $|1-\gamma|$ a vector of dimension $|\mathcal{E}|$ with entry $j$ equal to $|1-\gamma_{j}|$.
For an observational bias $b\in\mathbb{R}^{p}$ (i.e., $E[\tilde{\theta}^{\mathrm{obs}}]-\theta=b$), the mean–squared error is
$\mathrm{MSE}_{\mathcal{E},\Sigma,b}\!\left[\hat{\tau}_{\gamma}\right]~=~\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)\;+\;\bigl(\tilde{\omega}^{\mathrm{obs}}(\mathcal{E},\gamma)^{\top}b\bigr)^{2}.$
For an oracle that knows $\|b\|_{\infty}\leq\bar{B}$,
$R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{\|b\|_{\infty}\leq\bar{B}}\mathrm{MSE}_{\mathcal{E},\Sigma,b}\!\left[\hat{\tau}_{\gamma}\right]~=~\bar{B}^{2}\!\left(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|_{\mathcal{E}}^{\top}|\omega_{\mathcal{E}}|\right)^{\!2}+\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),$
since $\sup_{\|b\|_{\infty}\leq\bar{B}}\tilde{\omega}^{\mathrm{obs}\,\top}b=\bar{B}\,\|\tilde{\omega}^{\mathrm{obs}}\|_{1}=\bar{B}\big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|_{\mathcal{E}}^{\top}|\omega_{\mathcal{E}}|\big)$.
B.1.1 The oracle problem
Given $(\mathcal{E},\Sigma)$, the oracle first chooses the shrinkage vector
$\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)~\in~\arg\min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}R_{\bar{B}}(\mathcal{E},\Sigma,\gamma),$
and then (by backward induction) the design
$\big(\mathcal{E}^{\star}_{o}(\bar{B}),\ \Sigma^{\star}_{o}(\bar{B})\big)~\in~\arg\min_{\mathcal{E},\,\Sigma}R_{\bar{B}}\big(\mathcal{E},\Sigma,\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)\big).$
Let $R^{\star}_{o}(\bar{B})\equiv\min_{\mathcal{E},\Sigma,\gamma}R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)$ denote the oracle’s minimized worst–case MSE.
Lemma 1 (Convexity and strict convexity).
For fixed $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$, the map $\gamma\mapsto R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)$ is convex on $\gamma$. Moreover, if $|\omega_{j}|>0$ for all $j\in\mathcal{E}$, then $\gamma\mapsto R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)$ is strictly convex on $\gamma$.
Proof.
Write $R_{\bar{B}}=\bar{B}^{2}g+h$ with
$g(\gamma)=\Big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\Big)^{2},\qquad h(\gamma)=\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma).$
(i) Convexity of $g$. The map $\gamma\mapsto|1-\gamma|^{\top}|\omega_{\mathcal{E}}|$ is convex (sum of absolute values of affine functions). Adding the constant $\|\omega_{\mathcal{E}^{c}}\|_{1}$ preserves convexity, and squaring preserves convexity.
(ii) Strict convexity of $h$. Let $D\equiv\mathrm{diag}(\omega_{\mathcal{E}})$. Without loss, reorder the entries of $\theta$ so that we have $\theta=(\theta_{\mathcal{E}^{c}},\theta_{\mathcal{E}})$. Then
$\tilde{\omega}(\mathcal{E},\gamma)=\underbrace{\begin{bmatrix}\omega_{\mathcal{E}^{c}}\\ D\mathbf{1}\\ 0\end{bmatrix}}_{=:d}\;+\;\underbrace{\begin{bmatrix}0\\ -D\\ D\end{bmatrix}}_{=:A}\gamma,\quad\text{so}\quad h(\gamma)=(A\gamma+d)^{\top}\Sigma(\mathcal{E})\,(A\gamma+d).$
Hence $h(\gamma)=\gamma^{\top}(A^{\top}\Sigma A)\gamma+2\gamma^{\top}A^{\top}\Sigma d+d^{\top}\Sigma d$, and the Hessian is
$\nabla^{2}h(\gamma)=2\,A^{\top}\Sigma(\mathcal{E})\,A.$
Because $\Sigma(\mathcal{E})\succ 0$ and, when $|\omega_{j}|>0$ for all $j\in\mathcal{E}$, the columns of A=[ 0;D;D]A=[\,0;\,-D;\,D\,] are linearly independent, we have for any nonzero $u$,
$u^{\top}A^{\top}\Sigma A\,u=(Au)^{\top}\Sigma(Au)>0.$
Thus $A^{\top}\Sigma A\succ 0$ and $h$ is strictly convex (or weakly convex if $\omega_{j}=0$ for some $j$).
(iii) Conclusion. A sum of a convex function ($\bar{B}^{2}g$) and a strictly convex function ($h$) is strictly convex; when some $\omega_{j}=0$, the sum remains convex. ∎
Next, we show that as the worst-case bias becomes large enough, the oracle will put zero weight on the observational estimators.
Lemma 2.
Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$ and write $A\equiv\|\omega_{\mathcal{E}^{c}}\|_{1}$, with $|\omega_{j}|>0$ for all $j$. Let
$R_{\bar{B}}(\gamma)=\bar{B}^{2}\Big(A+t(\gamma)\Big)^{2}+\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad t(\gamma)\equiv|1-\gamma|^{\top}|\omega_{\mathcal{E}}|,$
for $\gamma\in\mathbb{R}^{|\mathcal{E}|}$. Then:
(i) If $A>0$ (equivalently, $\|\omega_{\mathcal{E}}\|_{1}<\|\omega\|_{1}$), there exists
$\bar{B}^{\max}(\mathcal{E},\Sigma,\omega)\;\leq\;\sqrt{\,2\,\|\Sigma(\mathcal{E})\|_{\infty}\cdot\max\!\Big\{2,\ \frac{\|\omega\|_{1}}{A}\Big\}\,}$
such that every minimizer satisfies $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)=\mathbf{1}$ for all $\bar{B}\geq\bar{B}^{\max}$.
(ii) If $A=0$ (equivalently, $\|\omega_{\mathcal{E}}\|_{1}=\|\omega\|_{1}$), then for any sequence $\bar{B}\to\infty$ every sequence of minimizers obeys $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)\to\mathbf{1}$.
Proof.
Let $x(\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)$ and $x(1)\equiv\tilde{\omega}(\mathcal{E},\mathbf{1})$. Using $a^{\top}\Sigma a-b^{\top}\Sigma b=(a-b)^{\top}\Sigma(a+b)$ and the Holder’s bound $|u^{\top}\Sigma v|\leq\|\Sigma\|_{\infty}\|u\|_{1}\|v\|_{1}$, we get
$\big|x(\gamma)^{\top}\Sigma x(\gamma)-x(1)^{\top}\Sigma x(1)\big|\leq\|\Sigma\|_{\infty}\,\|x(\gamma)-x(1)\|_{1}\,\|x(\gamma)+x(1)\|_{1}.$
In addition,
$x(\gamma)-x(1)=\begin{bmatrix}\omega_{\mathcal{E}^{c}}-\omega_{\mathcal{E}^{c}}\\[1.0pt] \omega_{\mathcal{E}}\circ(1-\gamma)\\[1.0pt] -\,\omega_{\mathcal{E}}\!\circ\!(1-\gamma)\end{bmatrix}\Rightarrow\|x(\gamma)-x(1)\|_{1}=2\,t(\gamma).$
Also,
$x(\gamma)+x(1)=\begin{bmatrix}2\,\omega_{\mathcal{E}^{c}}\\[1.0pt] \omega_{\mathcal{E}}\!\circ\!(1-\gamma)\\[1.0pt] \omega_{\mathcal{E}}\!\circ\!(1+\gamma)\end{bmatrix}\Rightarrow\|x(\gamma)+x(1)\|_{1}\leq 2\|\omega_{\mathcal{E}^{c}}\|_{1}+t(\gamma)+\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1+\gamma_{j}|.$
Using the triangular inequality for reals $u,v$, $|u|=|u-v+v|\leq|u-v|+|v|$, so that $|1+\gamma_{j}|\leq|1+\gamma_{j}-2|+2=|1-\gamma_{j}|+2$, we have $\sum_{j\in\mathcal{E}}|\omega_{j}||1+\gamma_{j}|\leq t(\gamma)+2\|\omega_{\mathcal{E}}\|_{1}.$ Therefore
$\|x(\gamma)+x(1)\|_{1}\leq 2\|\omega\|_{1}+2\,t(\gamma),$
and hence
$\big|x(\gamma)^{\top}\Sigma x(\gamma)-x(1)^{\top}\Sigma x(1)\big|\leq 4\,\|\Sigma\|_{\infty}\,t(\gamma)\,\big(\|\omega\|_{1}+t(\gamma)\big).$ (16)
(i) Case $A>0$. For any $\gamma$,
$R_{\bar{B}}(\gamma)-R_{\bar{B}}(\mathbf{1})$ $=\bar{B}^{2}\big[(A+t)^{2}-A^{2}\big]+\big[x(\gamma)^{\top}\Sigma x(\gamma)-x(1)^{\top}\Sigma x(1)\big]$
$\geq\bar{B}^{2}\big(2A\,t+t^{2}\big)-4\,\|\Sigma\|_{\infty}\,t\big(\|\omega\|_{1}+t\big)$
$=(\,\bar{B}^{2}-4\|\Sigma\|_{\infty}\,)t^{2}+(2A\bar{B}^{2}-4\|\Sigma\|_{\infty}\|\omega\|_{1}\,)t,$
where $t(\gamma)\geq 0$. If $\bar{B}^{2}\geq\max\{\,4\|\Sigma\|_{\infty},\ 2\,\|\Sigma\|_{\infty}\,\|\omega\|_{1}/A\,\},$ then both coefficients are nonnegative and the right-hand side is $\geq 0$, with equality iff $t=0$, i.e., $\gamma=\mathbf{1}$. Thus $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)=\mathbf{1}$ for all such $\bar{B}$.
(ii) Case $A=0$ (i.e., $\|\omega_{\mathcal{E}^{c}}\|_{1}=0$). Recall $t(\gamma)\equiv\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1-\gamma_{j}|\geq 0$ and
$R_{\bar{B}}(\gamma)-R_{\bar{B}}(\mathbf{1})~\geq~\bar{B}^{2}\,t(\gamma)^{2}\;-\;4\,\|\Sigma(\mathcal{E})\|_{\infty}\,t(\gamma)\big(\|\omega\|_{1}+t(\gamma)\big),$
which follows from (16) with $A=0$. Let $\gamma^{\star}=\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)$ be any minimizer, so $R_{\bar{B}}(\gamma^{\star})\leq R_{\bar{B}}(\mathbf{1})$ and hence
$0\;\geq\;R_{\bar{B}}(\gamma^{\star})-R_{\bar{B}}(\mathbf{1})\;\geq\;\bar{B}^{2}\,t_{\star}^{2}\;-\;4\,\|\Sigma(\mathcal{E})\|_{\infty}\,t_{\star}\big(\|\omega\|_{1}+t_{\star}\big),$
where $t_{\star}(\gamma^{\star})\geq 0$. Rearranging gives
$\big(\bar{B}^{2}-4\|\Sigma(\mathcal{E})\|_{\infty}\big)\,t_{\star}^{2}\;\leq\;4\,\|\Sigma(\mathcal{E})\|_{\infty}\,\|\omega\|_{1}\,t_{\star}.$
If $\bar{B}^{2}>4\|\Sigma(\mathcal{E})\|_{\infty}$, we can divide both sides by $B^{2}-4\|\Sigma(\mathcal{E})\|_{\infty}$ and obtain
$t_{\star}\;\leq\;\frac{\,4\,\|\Sigma(\mathcal{E})\|_{\infty}\,\|\omega\|_{1}\,}{\bar{B}^{2}-4\|\Sigma(\mathcal{E})\|_{\infty}}\;=\;O\!\left(\frac{1}{\bar{B}^{2}}\right).$
Hence $t(\gamma^{\star})\to 0$ as $\bar{B}\to\infty$.
Finally, since $t(\gamma^{\star})=\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1-\gamma^{\star}_{j}|$ and each summand is nonnegative, for any coordinate with $|\omega_{j}|>0$ we have
$|1-\gamma^{\star}_{j}|\;\leq\;\frac{t(\gamma^{\star})}{|\omega_{j}|}\;\longrightarrow\;0\quad\text{as }\bar{B}\to\infty.$
Thus $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)\to\mathbf{1}$ as $\bar{B}\rightarrow\infty$. ∎
The lemma below characterizes the behavior of the implicitly-defined functions of $\bar{B}$: $\bar{B}\mapsto\beta\left(\mathcal{E},\gamma^{\star}\left(\bar{B},\mathcal{E},\Sigma\right)\right)$ and $\bar{B}\mapsto\alpha\left(\mathcal{E},\gamma^{\star}\left(\bar{B},\mathcal{E},\Sigma\right)\right)$.
Lemma 3.
Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$. Let
$\alpha(\mathcal{E},\Sigma,\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\equiv\Big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\Big)^{2},$
and for $\bar{B}\geq 0$ define $R_{\bar{B}}(\gamma)\equiv\alpha(\mathcal{E},\Sigma,\gamma)+\bar{B}^{2}\,\beta(\mathcal{E},\gamma)$. Let $\gamma^{\star}(\bar{B})\in\arg\min_{\gamma}R_{\bar{B}}(\gamma)$. Then $\bar{B}\mapsto\beta\!\left(\mathcal{E},\gamma^{\star}(\bar{B})\right)$ is non-increasing and $\bar{B}\mapsto\alpha\!\left(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\right)$ is non-decreasing.
Proof.
Fix $0\leq\bar{B}_{1}<\bar{B}_{2}$ and choose minimizers $\gamma_{1}\in\arg\min_{\gamma}R_{\bar{B}_{1}}(\gamma)$ and $\gamma_{2}\in\arg\min_{\gamma}R_{\bar{B}_{2}}(\gamma)$. By optimality,
$\alpha(\gamma_{1})+\bar{B}_{1}^{2}\,\beta(\gamma_{1})$ $\leq\alpha(\gamma_{2})+\bar{B}_{1}^{2}\,\beta(\gamma_{2}),$ (16)
$\alpha(\gamma_{2})+\bar{B}_{2}^{2}\,\beta(\gamma_{2})$ $\leq\alpha(\gamma_{1})+\bar{B}_{2}^{2}\,\beta(\gamma_{1}).$ (17)
Subtracting $\alpha(\gamma_{2})$ from (16) and $\alpha(\gamma_{1})$ from (17) yields
$\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\leq\;\bar{B}_{1}^{2}\big(\beta(\gamma_{2})-\beta(\gamma_{1})\big),\qquad\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\geq\;\bar{B}_{2}^{2}\big(\beta(\gamma_{2})-\beta(\gamma_{1})\big).$
Let $\Delta_{\beta}\equiv\beta(\gamma_{2})-\beta(\gamma_{1})$. Then
$\bar{B}_{2}^{2}\,\Delta_{\beta}\;\leq\;\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\leq\;\bar{B}_{1}^{2}\,\Delta_{\beta}.$
If $\Delta_{\beta}>0$, the last display implies $\bar{B}_{2}^{2}\leq\bar{B}_{1}^{2}$, a contradiction. Hence $\Delta_{\beta}\leq 0$, i.e.
$\beta(\gamma_{2})\leq\beta(\gamma_{1}),$
so $\bar{B}\mapsto\beta(\gamma^{\star}(\bar{B}))$ is non-increasing. Plugging $\Delta_{\beta}\leq 0$ into the first inequality gives
$\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\leq\;\bar{B}_{1}^{2}\,\Delta_{\beta}\;\leq\;0,$
so $\alpha(\gamma_{1})\leq\alpha(\gamma_{2})$, i.e. $\bar{B}\mapsto\alpha(\gamma^{\star}(\bar{B}))$ is non-decreasing. Because the choice of minimizer at each $\bar{B}$ was arbitrary, the monotonicity holds for any $\bar{B}$. ∎
Proposition 1.
Let $\Sigma(\mathcal{E})\succ 0$ with uniformly bounded entries and fix $\mathcal{E}$. Let $|\omega_{j}|>0$ for all $j$. Let
$\alpha(\mathcal{E},\Sigma,\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\equiv\Big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\Big)^{2},$
and $\gamma^{\star}(\bar{B})\in\arg\min_{\gamma}\{\alpha(\mathcal{E},\Sigma,\gamma)+\bar{B}^{2}\beta(\mathcal{E},\gamma)\}$. Then, as $\bar{B}\to\infty$,
$\beta\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)\,\downarrow\,\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\big)^{2}\quad\text{and}\quad\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\,\uparrow\,\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1}),$
both monotonically.
Proof.
Write $A\equiv\|\omega_{\mathcal{E}^{c}}\|_{1}$ and $t(\gamma)\equiv\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1-\gamma_{j}|\geq 0$, so $\beta(\mathcal{E},\gamma)=(A+t(\gamma))^{2}$ and $\beta(\mathcal{E},\mathbf{1})=A^{2}=(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1})^{2}$.
Step 1 (monotonicity). By Lemma 3, $\bar{B}\mapsto\beta(\mathcal{E},\gamma^{\star}(\bar{B}))$ is non-increasing and $\bar{B}\mapsto\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing.
Step 2 (limit of $\beta$). Since $t(\gamma)\geq 0$, we have $\beta(\mathcal{E},\gamma)\geq A^{2}$ for all $\gamma$, hence the non-increasing sequence $\beta(\mathcal{E},\gamma^{\star}(\bar{B}))$ is bounded below by $A^{2}$ and thus converges to some limit $\geq A^{2}$. By Lemma 2, $\gamma^{\star}(\bar{B})\to\mathbf{1}$ as $\bar{B}\to\infty$, hence $t(\gamma^{\star}(\bar{B}))\to 0$ and therefore
$\lim_{\bar{B}\to\infty}\beta\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)=\lim_{\bar{B}\to\infty}\big(A+t(\gamma^{\star}(\bar{B}))\big)^{2}=A^{2}=(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1})^{2}.$
By Step 1 the convergence is monotone (decreasing).
Step 3 (limit of $\alpha$). Feasibility of $\gamma=\mathbf{1}$ implies, for all $\bar{B}$,
$\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)+\bar{B}^{2}\,\beta\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)~\leq~\alpha\big(\mathcal{E},\Sigma,\mathbf{1}\big)+\bar{B}^{2}\,\beta\big(\mathcal{E},\mathbf{1}\big)=\alpha(\mathcal{E},\Sigma,\mathbf{1})+\bar{B}^{2}A^{2},$
and since $\beta(\mathcal{E},\gamma^{\star}(\bar{B}))\geq A^{2}$, this yields
$\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\ \leq\ \alpha\big(\mathcal{E},\Sigma,\mathbf{1}\big)\qquad\forall\,\bar{B}.$
Thus $\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing and bounded above by $\alpha(\mathcal{E},\Sigma,\mathbf{1})$, so it converges to some limit $\leq\alpha(\mathcal{E},\Sigma,\mathbf{1})$. To identify the limit, we use the bound in Lemma 2
$\big|\alpha(\mathcal{E},\Sigma,\gamma)-\alpha(\mathcal{E},\Sigma,\mathbf{1})\big|=\big|x(\gamma)^{\top}\Sigma x(\gamma)-x(\mathbf{1})^{\top}\Sigma x(\mathbf{1})\big|\leq 4\,\|\Sigma(\mathcal{E})\|_{\infty}\,t(\gamma)\big(\|\omega\|_{1}+t(\gamma)\big),$
where $x(\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)$. Because $t(\gamma^{\star}(\bar{B}))\to 0$ from Step 2, the RHS tends to 0, so
$\lim_{\bar{B}\to\infty}\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)=\alpha\big(\mathcal{E},\Sigma,\mathbf{1}\big)=\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1}).$
By Step 1 the convergence is monotone (increasing). ∎
B.1.2 The researcher’s problem
To characterize the adaptation regret, it suffices to characterize
$\sup_{\bar{B}\geq 0}\,\sup_{\|b\|_{\infty}\leq\bar{B}}\frac{\text{MSE}_{b}\left[\,\widehat{\tau}\left(\mathcal{E},\gamma\right)\,\right]}{R^{\star}_{o}\left(\bar{B}\right)}=\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\sup_{\bar{B}\geq 0}\frac{R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)}{R_{\bar{B}}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma_{o}^{\star}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime}))}.$
Write
$\rho(\gamma,\bar{B},\mathcal{E},\mathcal{E}^{\prime},\Sigma,\Sigma^{\prime})\equiv\frac{R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)}{R_{\bar{B}}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma_{o}^{\star}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime}))}.$
Lemma 4.
Let $\Sigma(\mathcal{E}),\Sigma^{\prime}(\mathcal{E}^{\prime})\succ 0$ have uniformly bounded entries and assume $|\omega_{j}|>0$ for all $j$. For any fixed $(\mathcal{E},\Sigma,\gamma)$ and $(\mathcal{E}^{\prime},\Sigma^{\prime})$, define
$\alpha\equiv\alpha(\mathcal{E},\Sigma,\gamma),\quad\beta\equiv\beta(\mathcal{E},\gamma),\qquad\alpha^{\star}(\bar{B})\equiv\alpha(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})),\quad\beta^{\star}(\bar{B})\equiv\beta(\mathcal{E}^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})).$
Let
$\rho(\bar{B})\;\equiv\;\frac{\bar{B}^{2}\,\beta+\alpha}{\bar{B}^{2}\,\beta^{\star}(\bar{B})+\alpha^{\star}(\bar{B})}.$
[…]
Proof.
By Proposition 1, along the oracle path $\bar{B}\mapsto\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})$ we have
$\beta^{\star}(\bar{B})\ \downarrow\ \big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2},\qquad\alpha^{\star}(\bar{B})\ \uparrow\ \tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1})^{\top}\Sigma^{\prime}(\mathcal{E}^{\prime})\,\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1}),$
hence the ratio
$\psi(\bar{B})\;\equiv\;\frac{\beta^{\star}(\bar{B})}{\alpha^{\star}(\bar{B})}$
is non-increasing in $\bar{B}$ (numerator $\downarrow$, denominator $\uparrow$, both nonnegative).
Fix $0\leq\bar{B}_{1}<\bar{B}_{2}$. Using optimality of the oracle at each $\bar{B}$,
$\alpha^{\star}(\bar{B}_{1})+\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})~\leq~\alpha^{\star}(\bar{B}_{2})+\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{2}),\qquad\alpha^{\star}(\bar{B}_{2})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})~\leq~\alpha^{\star}(\bar{B}_{1})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{1}).$
Therefore
$\rho(\bar{B}_{2})-\rho(\bar{B}_{1})$ $=\frac{(\bar{B}_{2}^{2}\beta+\alpha)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)-(\bar{B}_{1}^{2}\beta+\alpha)\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}$
$\leq\frac{(\bar{B}_{2}^{2}\beta+\alpha)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)-(\bar{B}_{1}^{2}\beta+\alpha)\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}$
$=\frac{(\bar{B}_{2}^{2}-\bar{B}_{1}^{2})\big[\alpha^{\star}(\bar{B}_{2})\,\beta-\beta^{\star}(\bar{B}_{2})\,\alpha\big]}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}$
$=\frac{(\bar{B}_{2}^{2}-\bar{B}_{1}^{2})\,\alpha^{\star}(\bar{B}_{2})\,\alpha}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}\left(\frac{\beta}{\alpha}-\frac{\beta^{\star}(\bar{B}_{2})}{\alpha^{\star}(\bar{B}_{2})}\right).$
Since all denominators are positive, the last display shows
$\rho(\bar{B}_{2})-\rho(\bar{B}_{1})\ \leq\ 0\quad\text{whenever}\quad\frac{\beta}{\alpha}\ \leq\ \psi(\bar{B}_{2}).$
A symmetric argument (now using $\alpha^{\star}(\bar{B}_{2})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})\leq\alpha^{\star}(\bar{B}_{1})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{1})$) yields
$\rho(\bar{B}_{2})-\rho(\bar{B}_{1})\ \geq\ 0\quad\text{whenever}\quad\frac{\beta}{\alpha}\ \geq\ \psi(\bar{B}_{1}).$
Define the threshold
$\tilde{B}\;\equiv\;\sup\Big\{B\geq 0:\ \psi(B)\geq\tfrac{\beta}{\alpha}\Big\}\ \in\ [0,\infty],$
[…]
Completion of the proof.
Fix $(\mathcal{E},\Sigma,\gamma)$ and write, for any $(\mathcal{E}^{\prime},\Sigma^{\prime})$,
$\rho(\bar{B})\equiv\rho(\gamma,\bar{B},\mathcal{E},\mathcal{E}^{\prime},\Sigma,\Sigma^{\prime})=\frac{\bar{B}^{2}\,\beta(\mathcal{E},\gamma)+\alpha(\mathcal{E},\Sigma,\gamma)}{\bar{B}^{2}\,\beta^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})+\alpha^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})},$
[…]
$\sup_{\bar{B}\geq 0}\rho(\bar{B})=\max\!\Big\{\rho(0),\ \lim_{\bar{B}\to\infty}\rho(\bar{B})\Big\}.$
Endpoint $\bar{B}=0$. Since $\rho(0)=\alpha(\mathcal{E},\Sigma,\gamma)\big/\alpha^{\star}(0;\mathcal{E}^{\prime},\Sigma^{\prime})$,
$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\rho(0)=\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\inf_{\mathcal{E}^{\prime},\Sigma^{\prime}}\alpha^{\star}(0;\mathcal{E}^{\prime},\Sigma^{\prime})}.$
[…]
$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\rho(0)=\frac{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)}{\displaystyle\min_{\mathcal{E}^{\prime},\Sigma^{\prime}}\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{0})^{\top}\Sigma^{\prime}(\mathcal{E}^{\prime})\,\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{0})}.$
Endpoint $\bar{B}\to\infty$. By Proposition 1, $\beta^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})\downarrow\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2}$ and $\alpha^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})\uparrow\alpha(\mathcal{E}^{\prime},\Sigma^{\prime},\mathbf{1})$. Hence, after diving the numerators and denominator by $B^{2}$
$\lim_{\bar{B}\to\infty}\rho(\bar{B})=\frac{\beta(\mathcal{E},\gamma)}{\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2}}.$
Maximizing the ratio over $(\mathcal{E}^{\prime},\Sigma^{\prime})$ at this endpoint amounts to minimizing the denominator. This completes the proof. ∎
B.2 Two parameters examples: proof of Corollary 1
By Theorem 1, for fixed $(\mathcal{E},\Sigma)$ the adaptation regret as a function of the shrinkage vector equals
$\mathcal{R}(\gamma)\;=\;\max\!\Big\{\alpha(\mathcal{E},\Sigma,\gamma)/\alpha^{\star},\;\beta(\mathcal{E},\gamma)/\beta^{\star}\Big\},$
with $\alpha^{\star},\beta^{\star}$ constants.
Step 1. With independence and two parameters,
$\alpha(\mathcal{E},\Sigma,\gamma)=C+\omega_{j}^{2}\!\Big(\gamma_{j}^{2}v_{j}^{2}+(1-\gamma_{j})^{2}\sigma_{j}^{2}\Big),$
where $C$ does not depend on $\gamma_{j}$. Hence $\alpha(j,\gamma_{j})\equiv\alpha(\mathcal{E},\Sigma,\gamma)$ is a strictly convex quadratic in $\gamma_{j}$ with unique minimizer at
$\gamma_{j}^{\mathrm{var}}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}.$
Moreover, $\alpha(j,\gamma_{j})$ is strictly increasing on $\big(\gamma_{j}^{\mathrm{var}},1\big)$ and strictly decreasing on $\big[0,\gamma_{j}^{\mathrm{var}}\big)$.
Write the worst–case bias component holding the other coordinate fixed as
$\beta(j,\gamma_{j})\;=\;\Big(A_{j}+|\omega_{j}|\cdot|1-\gamma_{j}|\Big)^{2},$
where $A_{j}\geq 0$ collects all terms not involving $\gamma_{j}$ (including the contribution from the other coordinate). On $[0,1]$, $|1-\gamma_{j}|=1-\gamma_{j}$, so $\beta(j,\gamma_{j})$ is strictly decreasing in $\gamma_{j}$ and convex with minimum at $\gamma_{j}=1$. Therefore, any minimizer satisfies
$\gamma_{j}^{\star}\in\big[\gamma_{j}^{\mathrm{var}},\,1\big].$
Step 2. On $\big(\gamma_{j}^{\mathrm{var}},1\big)$ the map $\gamma_{j}\mapsto\alpha(j,\gamma_{j})/\alpha^{\star}$ is strictly increasing, while $\gamma_{j}\mapsto\beta(j,\gamma_{j})/\beta^{\star}$ is strictly decreasing. Hence the function
$f(\gamma_{j})\equiv\max\!\Big\{\alpha(j,\gamma_{j})/\alpha^{\star},\ \beta(j,\gamma_{j})/\beta^{\star}\Big\},\qquad\gamma_{j}\in\big[\gamma_{j}^{\mathrm{var}},1\big],$
is minimized either (i) at a boundary point, or (ii) at the unique interior point where the two arguments are equal (by strict monotonicity, at most one intersection exists). Therefore,
  • If $\alpha(j,\gamma_{j})/\alpha^{\star}<\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\beta(j,\gamma_{j})/\beta^{\star}$ on that interval. Since this term is strictly decreasing, the minimizer is the right boundary $\gamma_{j}^{\star}=1$.
  • If $\alpha(j,\gamma_{j})/\alpha^{\star}>\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\alpha(j,\gamma_{j})/\alpha^{\star}$ on that interval. Since this term is strictly increasing there, the minimizer is the left boundary $\gamma_{j}^{\star}=\gamma_{j}^{\mathrm{var}}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
  • Otherwise, by the intermediate value theorem and strict monotonicity of the two curves, there exists a unique $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$ such that
    $\frac{\alpha(j,\gamma_{j})}{\alpha^{\star}}\;=\;\frac{\beta(j,\gamma_{j})}{\beta^{\star}}.$
Appendix C Proofs of the extensions
C.1 Proof of Theorem 2
The proof of Theorem 2 follows similarly the one of Theorem 1 with minor modifications.
Let
$\bar{\alpha}(\mathcal{E},\Sigma,\gamma)\equiv\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},\qquad\bar{\beta}(\mathcal{E},\gamma)\equiv\|\tilde{\omega}(\mathcal{E},\gamma)\|_{1},$
so that
$|L_{B}(\mathcal{E},\Sigma,\gamma)|=2\{z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E},\Sigma,\gamma)+B\,\bar{\beta}(\mathcal{E},\gamma)\}.$
For any comparison design $(\mathcal{E}^{\prime},\Sigma^{\prime})$, let the oracle choose
$\gamma^{\star}(B,\mathcal{E}^{\prime},\Sigma^{\prime})\in\arg\min_{\gamma^{\prime}\in\mathbb{R}^{|\mathcal{E}^{\prime}|}}\Big\{z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})+B\,\bar{\beta}(\mathcal{E}^{\prime},\gamma^{\prime})\Big\},$
and denote the induced oracle path
$\bar{\alpha}^{\star}(B)\equiv\bar{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\star}(B,\mathcal{E}^{\prime},\Sigma^{\prime})),\qquad\bar{\beta}^{\star}(B)\equiv\bar{\beta}(\mathcal{E}^{\prime},\gamma^{\star}(B,\mathcal{E}^{\prime},\Sigma^{\prime})).$
Define the ratio (the factor $2$ cancels)
$\tilde{\rho}(B)\equiv\frac{z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E},\Sigma,\gamma)+B\,\bar{\beta}(\mathcal{E},\gamma)}{z_{1-\eta/2}\,\bar{\alpha}^{\star}(B)+B\,\bar{\beta}^{\star}(B)}.$
Step 1: Endpoint reduction (CI analogue of Lemma 4).
Write $a\equiv z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E},\Sigma,\gamma)$ and $b\equiv\bar{\beta}(\mathcal{E},\gamma)$, and $A(B)\equiv z_{1-\eta/2}\,\bar{\alpha}^{\star}(B)$, $B^{\star}(B)\equiv\bar{\beta}^{\star}(B)$. By Proposition 1 (applied with square roots), $\bar{\beta}^{\star}(B)$ is non-increasing in $B$, and $\bar{\alpha}^{\star}(B)$ is non-decreasing in $B$. Exactly as in Lemma 4 (replace $\alpha,\beta$ there with $\bar{\alpha},\bar{\beta}$ here, and use $x=B$ in place of $x=\bar{B}^{2}$), one verifies that
$\tilde{\rho}^{\prime}(B)\ \text{is non-decreasing in $B$}.$
Hence $\tilde{\rho}$ is first non-increasing and then non-decreasing, so its supremum is attained at the endpoints:
$\sup_{B\geq 0}\tilde{\rho}(B)=\max\Big\{\tilde{\rho}(0),\ \lim_{B\to\infty}\tilde{\rho}(B)\Big\}.$ (18)
Step 2: Evaluate the endpoints and optimize over comparison designs.
At $B=0$,
$\tilde{\rho}(0)=\frac{\bar{\alpha}(\mathcal{E},\Sigma,\gamma)}{\bar{\alpha}^{\star}(0)}.$
Since the oracle at $B=0$ minimizes standard error, $\bar{\alpha}^{\star}(0)=\min_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})}\bar{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})=\sqrt{\alpha^{\star}},$ where $\alpha^{\star}$ is as in (5). Thus
$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\tilde{\rho}(0)=\frac{\bar{\alpha}(\mathcal{E},\Sigma,\gamma)}{\sqrt{\alpha^{\star}}}=\left(\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}}\right)^{\!1/2}.$
As $B\to\infty$, the BB–term dominates. By Proposition 1 (again in square–root form), $\lim_{B\to\infty}\bar{\beta}^{\star}(B)=\min_{(\mathcal{E}^{\prime},\gamma^{\prime})}\bar{\beta}(\mathcal{E}^{\prime},\gamma^{\prime})=\sqrt{\beta^{\star}},$ with $\beta^{\star}$ from (5). Therefore
$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\ \lim_{B\to\infty}\tilde{\rho}(B)=\frac{\bar{\beta}(\mathcal{E},\gamma)}{\sqrt{\beta^{\star}}}=\left(\frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right)^{\!1/2}.$
Step 3: Combine.
Taking the supremum over comparison designs and using (18),
$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)=\sup_{B\geq 0}\ \sup_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}=\max\!\left\{\left(\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}}\right)^{\!1/2},\ \left(\frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right)^{\!1/2}\right\},$
which is exactly the statement of Theorem 2. ∎
C.2 Proof of Theorem 3
Define the recursive solution to the oracle’s problem in (9):
$\gamma^{\star}_{o}\left(B,\mathcal{E},\Sigma\right)$ $=$ $\operatorname*{arg\,min}_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)$
$\Sigma^{\star}_{o}\left(B,\mathcal{E}\right)$ $=$ $\operatorname*{arg\,min}_{\Sigma\in\mathcal{G}(\mathcal{E})}\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}\left(\mathcal{E},\Sigma,\gamma^{\star}_{o}\left(B,\mathcal{E},\Sigma\right)\right)$
$\mathcal{E}^{\star}_{o}(B)$ $=$ $\operatorname*{arg\,min}_{\mathcal{E}\in\mathcal{S}}\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}\left(\mathcal{E},\Sigma^{\star}_{o}\left(B,\mathcal{E}\right),\gamma^{\star}_{o}\left(B,\mathcal{E},\Sigma\right)\right),$
and let
$\check{R}_{B}\left(\mathcal{E},\Sigma,\gamma\right)\equiv\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\beta}(\mathcal{E},\gamma)B^{2}+\check{\alpha}(\mathcal{E},\Sigma,\gamma),$
with $\check{\beta},\check{\alpha}$ as defined in Equation (10). The proof mimics the proof of Theorem 1. We will refer to $\check{R}_{B}(\gamma)$ omitting its arguments $\mathcal{E},\Sigma$ whenever clear from the context.
Lemma 5 (Convexity and strict convexity).
Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$. Then the map
$\gamma~\longmapsto~\check{R}_{\bar{B}}(\mathcal{E},\Sigma,\gamma)~=~\check{\alpha}(\mathcal{E},\Sigma,\gamma)\;+\;\bar{B}^{2}\,\check{\beta}(\mathcal{E},\gamma)$
is convex in $\gamma$. Moreover, if for each $j\in\mathcal{E}$ there exists at least one $\ell\in\{1,\dots,L\}$ with $|\omega^{\ell}_{j}|>0$, then the map is strictly convex in $\gamma$.
Proof.
We show convexity of the two addends separately.
1) Convexity (and strict convexity) of $\check{\alpha}(\mathcal{E},\Sigma,\gamma)$.
For each $\ell$, define $D_{\ell}:=\mathrm{diag}(\omega^{\ell}_{\mathcal{E}})$ and
$d_{\ell}~:=~\begin{bmatrix}\omega_{\mathcal{E}^{c}}\\ \omega^{\ell}_{\mathcal{E}}\\ 0\end{bmatrix},\qquad A_{\ell}~:=~\begin{bmatrix}0\\ -D_{\ell}\\ D_{\ell}\end{bmatrix}\ \in\mathbb{R}^{(p+|\mathcal{E}|)\times|\mathcal{E}|}.$
Then, by the definition of $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)$,
$\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\;=\;d_{\ell}\;+\;A_{\ell}\,\gamma^{\ell},$
where $\gamma^{\ell}\in\mathbb{R}^{|\mathcal{E}|}$ is the row of $\gamma$ for target $\ell$. Using (10),
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)=\mathrm{tr}\!\left(\tilde{\omega}\,\Sigma\,\tilde{\omega}^{\top}\right)=\sum_{\ell=1}^{L}(d_{\ell}+A_{\ell}\gamma^{\ell})^{\top}\Sigma(\mathcal{E})\,(d_{\ell}+A_{\ell}\gamma^{\ell}).$
Hence $\check{\alpha}$ is a (separable across $\ell$) quadratic form in $\gamma$ with Hessian
$\nabla^{2}_{\gamma}\check{\alpha}(\mathcal{E},\Sigma,\gamma)=2\,\mathrm{blkdiag}\!\big(A_{1}^{\top}\Sigma(\mathcal{E})A_{1},\ldots,A_{L}^{\top}\Sigma(\mathcal{E})A_{L}\big)\ \succeq\ 0,$
where $\mathrm{blkdiag}$ is a block-diagonal matrix with entries as above. It follows that $\check{\alpha}$ is convex.
For strict convexity: if $\Sigma(\mathcal{E})\succ 0$ and $A_{\ell}$ has full column rank for at least one $\ell$ in every coordinate direction, then $\sum_{\ell}(A_{\ell}^{\top}\Sigma A_{\ell})\succ 0$ and thus $\check{\alpha}$ is strictly convex. Since $A_{\ell}=[0;-D_{\ell};D_{\ell}]$, $A_{\ell}u=0$ iff $D_{\ell}u=0$. Thus $A_{\ell}$ has full column rank iff $D_{\ell}$ is invertible, i.e. iff $|\omega^{\ell}_{j}|>0$ for all $j\in\mathcal{E}$. A sufficient condition ensuring strict convexity is: for each $j\in\mathcal{E}$ there exists at least one $\ell$ with $|\omega^{\ell}_{j}|>0$; then, for any nonzero $u\in\mathbb{R}^{|\mathcal{E}|}$ we can pick an $\ell$ with $u_{j}\neq 0$ and $|\omega^{\ell}_{j}|>0$, whence
$u^{\top}\Big(\sum_{\ell}A_{\ell}^{\top}\Sigma A_{\ell}\Big)u=\sum_{\ell}(A_{\ell}u)^{\top}\Sigma(A_{\ell}u)\ \geq\ (A_{\ell}u)^{\top}\Sigma(A_{\ell}u)\ >\ 0.$
Therefore $\check{\alpha}$ is strictly convex under the stated condition.
2) Convexity of $\check{\beta}(\mathcal{E},\gamma)$.
From (10) and the definition of $\bar{v}(\gamma)$,
$\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}\ \sum_{\ell=1}^{L}\Big((\omega^{\ell})^{\top}v-(\omega^{\ell}_{\mathcal{E}})^{\top}v_{\mathcal{E}}+(1-\gamma^{\ell})^{\top}(\,v_{\mathcal{E}}\circ\omega_{\mathcal{E}}\,)\Big)^{\!2}.$
For each fixed $v$, the inner expression is an affine function of $\gamma$:
$(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\mathcal{E}})^{\top}v_{\mathcal{E}}+\underbrace{\mathbf{1}^{\top}(\,v_{\mathcal{E}}\circ\omega_{\mathcal{E}}\,)}_{\text{constant in }\gamma}\;-\;(\gamma^{\ell})^{\top}(\,v_{\mathcal{E}}\circ\omega_{\mathcal{E}}\,).$
The square of an affine function is convex; a finite sum of convex functions is convex; and the pointwise supremum of convex functions is convex. Therefore $\check{\beta}(\mathcal{E},\gamma)$ is convex in $\gamma$. This concludes the proof. ∎
Lemma 6.
Suppose $\Sigma(\mathcal{E})\succ 0$ with uniformly bounded entries and $|\omega^{\ell}_{j}|>0$ for all $j\in\{1,\dots,p\}$ and $\ell\in\{1,\dots,L\}$. Then, for fixed $(\mathcal{E},\Sigma)$,
$\gamma^{\star}_{o}(B,\mathcal{E},\Sigma)\ \longrightarrow\ \mathbf{1}\qquad\text{as }B\to\infty,$
where $\mathbf{1}$ denotes the $L\times|\mathcal{E}|$ matrix of ones.
Proof.
Recall that for fixed $(\mathcal{E},\Sigma)$ the oracle criterion is
$\check{R}_{B}(\mathcal{E},\Sigma,\gamma)=\sup_{\|b\|_{\infty}\leq B}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\alpha}(\mathcal{E},\Sigma,\gamma)+B^{2}\,\check{\beta}(\mathcal{E},\gamma),$
with (see (10))
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)=\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\gamma)\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\big),\qquad\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}\ \sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2},$
where
$a_{\ell}(v):=\sum_{j\in\mathcal{E}^{c}}\omega^{\ell}_{j}v_{j},\qquad\delta_{\ell}(\gamma,v):=\sum_{j\in\mathcal{E}}(1-\gamma^{\ell}_{j})\,\omega^{\ell}_{j}v_{j}.$
Step 1: The bias term is uniquely minimized at $\gamma=\mathbf{1}$.
For each $j\in\mathcal{E}$, define the $L$-vector
$d_{j}(\gamma):=\big((1-\gamma_{j}^{1})\omega_{j}^{1},\ldots,(1-\gamma_{j}^{L})\omega_{j}^{L}\big)^{\top}\in\mathbb{R}^{L},\qquad S(\gamma):=\max_{j\in\mathcal{E}}\|d_{j}(\gamma)\|_{2}.$
Choose $v\in[-1,1]^{p}$ with $v_{\mathcal{E}^{c}}=0$, $v_{j}=1$ for some fixed $j\in\mathcal{E}$, and $v_{j^{\prime}}=0$ for $j^{\prime}\in\mathcal{E}\setminus\{j\}$. Then
$\sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2}=\sum_{\ell=1}^{L}\big(\delta_{\ell}(\gamma,v)\big)^{2}=\|d_{j}(\gamma)\|_{2}^{2}.$
Taking the supremum over $v$ yields
$\check{\beta}(\mathcal{E},\gamma)\ \geq\ \max_{j\in\mathcal{E}}\|d_{j}(\gamma)\|_{2}^{2}\;=\;S(\gamma)^{2}.$ (19)
In particular, $\check{\beta}(\mathcal{E},\gamma)>\check{\beta}(\mathcal{E},\mathbf{1})$ whenever $\gamma\neq\mathbf{1}$ (since then $S(\gamma)>0$ by $|\omega^{\ell}_{j}|>0$). Thus $\gamma\mapsto\check{\beta}$ is uniquely minimized at $\gamma=\mathbf{1}$, with a quadratic gap at least $S(\gamma)^{2}$.
Step 2: Variance bound.
[…]
$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma)-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\leq\|\Sigma(\mathcal{E})\|_{\infty}\sum_{\ell=1}^{L}\|x_{\ell}(\gamma)-y_{\ell}\|_{1}\,\|x_{\ell}(\gamma)+y_{\ell}\|_{1}.$
A direct calculation gives
$\|x_{\ell}(\gamma)-y_{\ell}\|_{1}=2\sum_{j\in\mathcal{E}}|(1-\gamma^{\ell}_{j})\omega^{\ell}_{j}|=2\,t_{\ell}(\gamma),$
and, using $|1+\gamma^{\ell}_{j}|\leq|1-\gamma^{\ell}_{j}|+2$,
$\|x_{\ell}(\gamma)+y_{\ell}\|_{1}\leq 2\|\omega^{\ell}\|_{1}+2\,t_{\ell}(\gamma).$
Hence
$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma)-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\leq 4\,\|\Sigma(\mathcal{E})\|_{\infty}\!\left(\sum_{\ell=1}^{L}t_{\ell}(\gamma)\,\|\omega^{\ell}\|_{1}+\sum_{\ell=1}^{L}t_{\ell}(\gamma)^{2}\right).$
Let $d(\gamma)\in\mathbb{R}^{L\times|\mathcal{E}|}$ stack the $d_{j}(\gamma)$ as columns; then $t_{\ell}(\gamma)=\|d(\gamma)_{\ell,\cdot}\|_{1}$ and
$\sum_{\ell=1}^{L}t_{\ell}(\gamma)=\sum_{j\in\mathcal{E}}\|d_{j}(\gamma)\|_{1}\leq|\mathcal{E}|\,\sqrt{L}\,S(\gamma),\qquad\sum_{\ell=1}^{L}t_{\ell}(\gamma)^{2}\leq\Big(\sum_{\ell=1}^{L}t_{\ell}(\gamma)\Big)^{2}\leq|\mathcal{E}|^{2}L\,S(\gamma)^{2}.$
Therefore there exist finite constants $C_{1},C_{2}$ (depending only on $\Sigma,\omega,L,|\mathcal{E}|$) such that
$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma)-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\ \leq\ C_{1}\,S(\gamma)\;+\;C_{2}\,S(\gamma)^{2}.$ (20)
Step 3: convergence of the minimizer.
Combine (19)–(20) to get, for any $\gamma$,
$\check{R}_{B}(\mathcal{E},\Sigma,\gamma)-\check{R}_{B}(\mathcal{E},\Sigma,\mathbf{1})=\big[\check{\alpha}(\gamma)-\check{\alpha}(\mathbf{1})\big]+B^{2}\big[\check{\beta}(\gamma)-\check{\beta}(\mathbf{1})\big]\ \geq\ -\,C_{1}S(\gamma)+\big(B^{2}-C_{2}\big)S(\gamma)^{2}.$
Fix $\varepsilon>0$ and choose $B$ such that $(B^{2}-C_{2})\varepsilon^{2}>C_{1}\varepsilon$. Then the RHS is >0>0 for all $\gamma$ with $S(\gamma)\geq\varepsilon$. Hence any minimizer $\gamma^{\star}_{o}(B)$ must satisfy $S(\gamma^{\star}_{o}(B))<\varepsilon$. Because $\varepsilon>0$ is arbitrary and $B\to\infty$, we obtain $S(\gamma^{\star}_{o}(B))\to 0$.
Finally, $S(\gamma^{\star}_{o}(B))\to 0$ means $(1-\gamma^{\star,\ell}_{j}(B))\omega^{\ell}_{j}\to 0$ for every $j\in\mathcal{E}$ and $\ell\in\{1,\dots,L\}$. Since $|\omega^{\ell}_{j}|>0$, this implies $\gamma^{\star,\ell}_{j}(B)\to 1$ for all $j,\ell$, i.e. $\gamma^{\star}_{o}(B,\mathcal{E},\Sigma)\to\mathbf{1}$ componentwise. ∎
Lemma 7.
Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$. Let
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)=\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\gamma)\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\big),\qquad\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}\ \sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2},$
where
$a_{\ell}(v):=\sum_{j\in\mathcal{E}^{c}}\omega^{\ell}_{j}v_{j},\qquad\delta_{\ell}(\gamma,v):=\sum_{j\in\mathcal{E}}(1-\gamma^{\ell}_{j})\,\omega^{\ell}_{j}v_{j},$
and for $\bar{B}\geq 0$ define $R_{\bar{B}}(\gamma)\equiv\alpha(\mathcal{E},\Sigma,\gamma)+\bar{B}^{2}\,\beta(\mathcal{E},\gamma)$. Let $\gamma^{\star}(\bar{B})\in\arg\min_{\gamma}R_{\bar{B}}(\gamma)$. Then $\bar{B}\mapsto\beta\!\left(\mathcal{E},\gamma^{\star}(\bar{B})\right)$ is non-increasing and $\bar{B}\mapsto\alpha\!\left(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\right)$ is non-decreasing.
Proof.
The proof is identical to that of Lemma 3, after replacing $\alpha$ and $\beta$ with $\check{\alpha}$ and $\check{\beta}$, respectively. ∎
Proposition 2.
Let $\Sigma(\mathcal{E})\succ 0$ with uniformly bounded entries and fix $\mathcal{E}$. Let $|\omega_{j}|>0$ for all $j$. Define $\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ and $\check{\beta}(\mathcal{E},\gamma)$ as in Lemma. Then, as $\bar{B}\to\infty$,
$\check{\beta}\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)\,\downarrow\,\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\big)^{2}\quad\text{and}\quad\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\,\uparrow\,\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\mathbf{1})\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\big),$
both monotonically.
Proof.
Write $A\equiv\|\omega_{\mathcal{E}^{c}}\|_{1}$ and, following the notation in Lemma 7, define $g(\mathcal{E},\gamma,v)=\sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2}$, so $\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}g(\mathcal{E},\gamma,v)$.
Step 1 (monotonicity). By Lemma 7, $\bar{B}\mapsto\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))$ is non-increasing and $\bar{B}\mapsto\check{\alpha}(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing.
Step 2 (limit of $\beta$). Letting $\mathbf{1}_{\mathcal{E}}$ denote the $p\times 1$ vector with $j^{th}$ entry $1\{j\notin\mathcal{E}\}$, we have $\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))\geq g(\mathcal{E},\gamma^{\star}(\bar{B}),\mathbf{1}_{\mathcal{E}})=A^{2}$, hence the non-increasing sequence $\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))$ is bounded below by $A^{2}$ and thus converges to some limit $\geq A^{2}$.
Because $v\mapsto g(\mathcal{E},\gamma,v)$ is convex, $\check{\beta}(\mathcal{E},\gamma)=\max_{v\in V}g(\mathcal{E},\gamma,v)$, where $V=\{-1,1\}^{p}$. Additionally, Lemma 6 implies $\lim_{\bar{B}\rightarrow\infty}g(\mathcal{E},\gamma^{\star}(\bar{B}),v)=g(\mathcal{E},\mathbf{1},v)$ $\forall v\in V$, since $\gamma\mapsto g(\mathcal{E},\gamma,v)$ is continuous. Therefore, since $V$ is finite,
$|\max_{v\in V}g(\mathcal{E},\gamma^{\star}(\bar{B}),v)-\max_{v\in V}g(\mathcal{E},\mathbf{1},v)|\leq\max_{v\in V}|g(\mathcal{E},\gamma^{\star}(\bar{B}),v)-g(\mathcal{E},\mathbf{1},v)|\rightarrow 0\text{ as $\bar{B}\rightarrow\infty$.}$
We conclude that $\lim_{\bar{B}\rightarrow\infty}\check{\beta}\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)=\max_{v\in V}g(\mathcal{E},\mathbf{1},v)=A^{2}$. By Step 1, the convergence is monotone (decreasing).
Step 3 (limit of $\alpha$). Feasibility of $\gamma=\mathbf{1}$ implies, for all $\bar{B}$,
$\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)+\bar{B}^{2}\,\check{\beta}\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)~\leq~\check{\alpha}\big(\mathcal{E},\Sigma,\mathbf{1}\big)+\bar{B}^{2}\,\check{\beta}\big(\mathcal{E},\mathbf{1}\big)=\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})+\bar{B}^{2}A^{2},$
and since $\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))\geq A^{2}$, this yields
$\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\ \leq\ \check{\alpha}\big(\mathcal{E},\Sigma,\mathbf{1}\big)\qquad\forall\,\bar{B}.$
Thus $\check{\alpha}(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing and bounded above by $\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})$, so it converges to some limit $\leq\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})$. To identify the limit, we use the bound in Lemma 6
$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\leq C_{1}\,S(\gamma^{\star}(\bar{B}))+C_{2}\,S(\gamma^{\star}(\bar{B}))^{2}\rightarrow C_{1}\,S(\mathbf{1})+C_{2}\,S(\mathbf{1})^{2}=0,$
as $\bar{B}\rightarrow\infty$. Therefore,
$\lim_{\bar{B}\to\infty}\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)=\check{\alpha}\big(\mathcal{E},\Sigma,\mathbf{1}\big)=\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\mathbf{1})\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\big).$
By Step 1, the convergence is monotone (increasing). ∎
Lemma 8.
Let $\Sigma(\mathcal{E}),\Sigma^{\prime}(\mathcal{E}^{\prime})\succ 0$ have uniformly bounded entries and assume $|\omega_{j}^{\ell}|>0$ for all $j,\ell$. For any fixed $(\mathcal{E},\Sigma,\gamma)$ and $(\mathcal{E}^{\prime},\Sigma^{\prime})$, define
$\check{\alpha}\equiv\check{\alpha}(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}\equiv\check{\beta}(\mathcal{E},\gamma),\qquad\check{\alpha}^{\star}(\bar{B})\equiv\check{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})),\quad\check{\beta}^{\star}(\bar{B})\equiv\check{\beta}(\mathcal{E}^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})).$
Let
$\check{\rho}(\bar{B})\;\equiv\;\frac{\bar{B}^{2}\,\check{\beta}+\check{\alpha}}{\bar{B}^{2}\,\check{\beta}^{\star}(\bar{B})+\check{\alpha}^{\star}(\bar{B})}.$
[…]
Proof.
By Proposition 2, along the oracle path $\bar{B}\mapsto\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})$ we have
$\check{\beta}^{\star}(\bar{B})\ \downarrow\ \big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2},\qquad\check{\alpha}^{\star}(\bar{B})\ \uparrow\ \mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1})\,\Sigma(\mathcal{E}^{\prime})\,\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1})^{\top}\big),$
hence the ratio
$\check{\psi}(\bar{B})\;\equiv\;\frac{\check{\beta}^{\star}(\bar{B})}{\check{\alpha}^{\star}(\bar{B})}$
is non-increasing in $\bar{B}$ (numerator $\downarrow$, denominator $\uparrow$, both nonnegative).
The rest of the proof is identical to that of Lemma 4, after replacing $\alpha$, $\beta$, $\alpha^{\star}$, $\beta^{\star}$, $\rho$, and $\psi$ with $\check{\alpha}$, $\check{\beta}$, $\check{\alpha}^{\star}$, $\check{\beta}^{\star}$, $\check{\rho}$, and $\check{\psi}$, respectively. ∎
The completion of the proof is a analogous to that of B.1, after replacing $\alpha$, $\beta$, $\alpha^{\star}$, $\beta^{\star}$, and $\rho$ with $\check{\alpha}$, $\check{\beta}$, $\check{\alpha}^{\star}$, $\check{\beta}^{\star}$, and $\check{\rho}$, respectively. ∎

### v2

@@
 When such estimates are insufficient to extrapolate effects for broader policy questions as for external validity and general-equilibrium (GE) effects, researchers combine trials with external evidence from reduced-form or structural observational estimates, or prior experiments.
-We develop a unified framework for designing experiments in this setting: the researcher selects which parameters to identify experimentally from a feasible set (which treatment arms and/or individuals to include in the experiment), allocates sample size, and specifies how to weight experimental and observational estimators.
+We develop a unified framework for designing experiments in this setting: the researcher selects which parameters (or moments) to identify experimentally from a feasible set of possible experiments, allocates sample size, and specifies how to weight experimental and observational estimators.
 Because observational inputs may be biased in ways unknown ex ante, we adopt a minimax proportional-regret objective that evaluates any candidate design relative to an oracle that knows the bias and jointly chooses the design and estimator.
@@
 1 Introduction
-Randomized controlled trials have transformed empirical economics by delivering credible, internally valid estimates.
+Randomized controlled trials (RCTs) have improved empirical economics by providing internally valid causal estimates.
 However, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism.
-These effects, although useful, are often not sufficient to answer broader questions about external validity, generalizability, or equilibrium effects.
-In response, a growing literature in development economics (e.g., Meghir et al., 2022; Gechter et al., 2024; Bassi et al., 2022; Attanasio et al., 2012), political economy (e.g., de Albuquerque et al., 2025), education (e.g., Allende et al., 2019; Larroucau et al., 2024), and labor economics (e.g., Chetty et al., 2016) complements localized experiments with external evidence from reduced-form or structural observational estimates, and/or trials in other contexts, with the goal of estimating complex counterfactuals that no single experiment can fully identify.
-For example, in our review of AEA journals (in 2015-2025), over 30% of experimental papers report, in addition to the experiment, at least one observational estimate – either reduced-form (e.g., via matching or instrumental variable regression), or structural (see Figure 12).
-This raises a design question: given constraints on what experiments can identify, which experiments should be run (and how) when they will be combined with external evidence to estimate those counterfactuals?
+These effects, although useful, are often not sufficient to answer broader questions about external validity and (general) equilibrium effects of at-scale interventions.
+For instance, the median evaluation of experiments in top economics journals is “representative of a population of 10,885 units, studied a treatment delivered to 5,340 units, and randomized in clusters of 26 units per cluster” (Muralidharan and Niehaus, 2017).
+In response, a growing literature in development economics (e.g., Meghir et al., 2022; Bassi et al., 2022; Attanasio et al., 2012; de Albuquerque et al., 2025), education (e.g., Allende et al., 2019; Larroucau et al., 2024), labor economics (e.g., Chetty et al., 2016), and industrial organization (e.g., Katz and Allcott, 2025; Allcott et al., 2025) combine experiments with external evidence—observational estimates and/or results from experiments in other contexts—to estimate complex counterfactuals that no single experiment can identify.
+In our review of AEA journals over the past decade, more than 30% of experimental papers complement RCT evidence with at least one observational input, either reduced-form or structural (Figure 12).
+This trend poses a design question: given constraints on feasible experiments, which experiments should be run (and how) to estimate such counterfactuals?
+Our main contribution is a framework for designing experiments combined with external evidence.
+By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments conducted in other contexts.
+We provide a procedure that selects which experiments to run subject to a budget (e.g., which treatment arms and/or sub-populations to include), sets their precision via sample allocation, and prescribes how to combine experimental and external evidence.
+External input may be biased in ways not known ex ante (e.g., due to confounding or lack of external validity).
 For an illustrative example, consider a government piloting a cash-transfer program in a small set of districts to increase children’s school attendance.
@@
 The design question is twofold: (i) which experiment to run (i.e., which parameters/effects are the most valuable to learn experimentally) given feasibility constraints and (ii) how to allocate sample across sites and arms so that, once combined with external inputs, we precisely estimate the effect at scale.
-To answer these questions, we develop a framework for designing experiments to be used alongside external evidence.
-By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments conducted in other times or places.
-Our main contribution is a joint procedure that selects which experiments to run subject to a budget (i.e., which treatment arms and/or sub-populations to include in the study), sets their precision via sample allocation, and prescribes how to combine experimental estimates with external evidence.
-A central consideration in our design is that external inputs may be biased in ways not known ex ante.
-We consider a setting where the object of interest is a known function $\tau(\theta)$ with $\theta$ a vector of unknown parameters; $\tau$ may be univariate or multivariate and typically encodes a policy-relevant target.
-Researchers have access to external estimates of $\theta$ that may be biased—henceforth, observational estimates.
-They then run one or more experiments in the population of interest to learn a subset of parameters without bias, choosing a sample allocation across sites and arms subject to a budget.
-We parameterize the design so that some, but not necessarily all, components of $\theta$ can be learned experimentally.
-After the experiment, each experimental estimate is combined with its observational counterpart via parameter-specific shrinkage weights (chosen jointly with the design), while parameters not learned experimentally rely on the observational estimates alone.
-Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in a region receive the subsidy.
-The parameter vector $\theta$ bundles (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium, and (iii) wage/price adjustments that shift the returns to schooling.
-A large-scale trial that identifies GE effects is often infeasible; instead, one can run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$.
-For instance, researchers may choose between two (small-scale) treatment arms such as a CCT to estimate the direct effect or an unconditional-cash arm to measure the income elasticity, and extrapolate the GE effects using price effects from other experimental or observational studies.
-When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ may target the average effect across sites.
-A natural starting point is to minimize the mean-squared error (MSE) for estimating $\tau(\theta)$, perhaps in a worst-case sense over bias in the observational estimates.
-In practice, however, the size and even the direction of bias are unknown (before the experiment is run).
+The first step is to specify the estimand and the underlying parameters.
+Define $\tau(\theta)$ the (policy-relevant) object of interest, where $\tau$ is a known smooth function and $\theta$ is a vector of unknown parameters; $\tau$ may be scalar or vector-valued.
+Researchers also observe external baseline estimates of $\theta$—henceforth observational estimates—which may be biased.
+Each feasible experiment can identify a subset of parameters without bias.
+This parameterization makes explicit which components of $\theta$ are learned experimentally and which may rely on biased observational evidence; not all components need (or can) be learned experimentally.
+Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in rural Kenya receive the subsidy.
+The parameter vector $\theta$ bundles (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium , and (iii) wage adjustments that shift the returns to schooling.
+Due to cost constraints, researchers may only run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$.
+They may choose between a CCT treatment arm to estimate the direct effect, or an unconditional-cash arm (UCT) to measure the income elasticity, and extrapolate the GE effects using evidence from a previous study in Mexico (Todd and Wolpin, 2006).
+(When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ targets the average effect across sites.)
+The second step is to design the experiment: (i) selecting which experiment to run; (ii) allocating sample sizes across arms; and (iii) choosing how to combine observational and experimental evidence.
+We allow flexibility in (iii): researchers may choose to either optimally combine both sources when available or rely solely on experimental estimates.
+Parameters not identified experimentally are based on observational estimates.
+A natural starting point is to design the experiment to minimize the mean-squared error (MSE) for estimating $\tau(\theta)$, perhaps worst-case over the observational bias.
+In practice, however, the size and even the direction of bias are unknown (before the experiment).
 A worst-case approach would therefore be overly-conservative and disregard information about the variance.
-We instead adapt the definition of proportional (adaptation) regret previously studied when combining estimators generated by a fixed design (Armstrong et al., 2024; Tsybakov, 1998) and generalize it to the experimental design problem.
-Here, the regret is the ratio of the worst-case researcher’s mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the (shrinkage) estimator.
-Taking the supremum of this ratio over any value of the bias yields a procedure that is robust to any level of misspecification without requiring priors about the bias.
-Our main result gives an explicit, easy-to-compute characterization of proportional regret.
-For any design and shrinkage weights, the regret is the maximum of two normalized components.
-The first is a variance component: the estimator’s sampling variance under the chosen design and weights, divided by the smallest variance attainable by any feasible design.
-The second is a bias component: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias.
-The variance component depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the shrinkage weights, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters.
+We instead adapt the definition of proportional (adaptation) regret previously studied when combining two estimators generated by a fixed design (Armstrong et al., 2024; Tsybakov, 1998) and generalize it to the experimental design problem (with multiple estimators).
+Here, the regret is the ratio of the worst-case researcher’s mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the estimator.
+Taking its supremum over any value of the bias yields a procedure that is robust to any level of misspecification without requiring priors about the bias.
+Our main result gives an explicit novel characterization of proportional regret.
+For any design and weights that combine the experimental and observational point estimates, the regret is the maximum of two normalized components.
+The first is a variance regret: the estimator’s sampling variance under the chosen design and weights, divided by the smallest attainable variance.
+The second is a bias regret: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias over the class of designs.
+The variance regret depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the shrinkage weights, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters.
 This gradient is evaluated, under mild conditions, at the observational estimates.
-[^4] For this result to hold, we assume that $\tau(\theta)$ is a smooth and differentiable function in $\theta$.
+[^4] For this result to hold, we assume that $\tau(\theta)$ is twice differentiable with bounded derivatives.
 For non-linear $\tau(\theta)$ in $\theta$, we also require that the observational estimates bias is local to zero in the spirit of Andrews et al.
-(2020); Bonhomme and Weidner (2022), but consider a more general framework where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error.
+(2020); Armstrong and Kolesár (2021); Bonhomme and Weidner (2022), where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error.
 Section 5.1 provides details.
-By contrast, the bias component depends only on the sensitivity vector and the shrinkage weights, both known ex-ante.
+The bias regret depends only on the sensitivity vector and the shrinkage weights, known ex-ante.
 This characterization makes the bias-variance trade-off transparent.
-At the optimum, the two normalized components tend to be equalized: when the bias component dominates, weight shifts toward experimental evidence and the design prioritizes learning high-sensitivity coordinates; when variance dominates, the procedure invests sample where it most reduces variance and relies more (via shrinkage) on the most precise observational inputs.
-This yields a nested procedure: we (i) find (regret-optimal) shrinkage weights for each candidate design to balance the two ratios; (ii) allocate sample, subject to the budget, to reduce the variance; and (iii) select the experiment set, trading off designs that improve precision the most against those targeting parameters to which the policy estimand is the most sensitive.
+At the optimum, the two normalized components tend to be equalized: when the bias component dominates, the design prioritizes high-sensitivity parameters; when variance dominates, the procedure invests sample where it most reduces variance and relies more on the most precise observational inputs.
+This yields a nested procedure that chooses jointly the estimator, the sample allocation and the experiment to run, trading off precision with sensitivity of the policy estimand.
 The resulting design and estimator only require knowledge of the expected variance-covariance matrix (standard in experimental design problems, Gerber and Green, 2012) and of the observational estimates.
 It can then be reported in a pre-analysis plan.
-We illustrate the framework in the cash‐transfer application.
-As an illustrative example, we consider a researcher evaluating a conditional cash transfer (CCT) to increase schooling in rural Kenya.
-Budget and feasibility constraints permit only small-scale (partial-equilibrium) randomization, while the goal is to predict general-equilibrium effects.
-As preliminary (external) inputs, the researcher uses evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006).
+That is, the inputs for the design are the same that would be required by standard power-analysis plans.
+From a theoretical perspective, our result leverages a key insight.
+Because here bias arise from multiple parameters, the choice of the uncertainty set is pivotal.
+While we do not require researchers to specify such a set, regret is with respect to an oracle that knows the worst-case bias within an $\ell_{\infty}$-ball.
+Compared with $\ell_{1}$ or $\ell_{2}$, the $\ell_{\infty}$-ball prevents that a large bias in one component is “offset” by smaller biases elsewhere.
+This is attractive when biases may be positively correlated across parameters.
+We show that the $\ell_{\infty}$ choice combined with our design problem (with decisions made before experimental data is realized), yields a quasi-convex regret in the bias radius as the oracle optimizes over the designs.
+This result, worst-case over the radius, delivers a transparent and easy-to-compute variance-bias trade-off.
+We then extend the model to general method-of-moments estimators that combine experimentally identified moments with observational moments.
+As in the baseline case, observational moments may incur unknown misspecification.
+We optimize (i) which moments to identify experimentally and (ii) how to combine experimental and observational moments to estimate the counterfactual of interest.
+Our main result continues to hold in this more general scenario provided that the moment estimators are (asymptotically) linear in the moments.
+Linearity occurs (under smoothness conditions) as long as the ratio of the bias to standard error is no faster than $n^{1/4}$, with $n$ denoting the observational sample size; standard local-asymptotics with the bias and standard error of the same order is a special case.
+We illustrate the framework in the cash‐transfer application, when the researcher uses evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006; Attanasio et al., 2012).
 We are concerned that these preliminary estimates may lack external validity in Kenya.
@@
 If only one arm can be run, the choice hinges on sample size: for very small $n$, the lower-variance of the UCT arm makes this preferred over a (too noisy) CCT arm; once $n\geq 500$, the more informative but noisier CCT yields lower regret.
-Relative to an oracle that knows the bias, the optimal design closely tracks oracle performance, and regret approaches one as sample size grows.
+Relative to an oracle that knows the bias, the optimal design closely tracks the oracle.
+We provides a step-by-step guide for practitioners applicable in other contexts.
 As a second application, we design where to run a microfinance experiment in rural India, integrating evidence from earlier nonrandomized microfinance introductions.
@@
 The optimal design concentrates sample in high-sensitivity, high-variance areas and then spreads as the budget grows; compared to a benchmark that randomly chooses areas and splits sample equally, it reduces MSE by roughly 20% under the calibrated bias.
-In summary, this paper links the experimental-design literature—which typically focuses on settings where all parameters are identified within the experiment and leaves aside questions of data combination—to recent work that integrates experimental and (reduced-form or structural) observational evidence to extrapolate effects in complex scenarios.
-Recent advances for experimental design include balancing and variance-minimizing allocations (Tabord-Meehan, 2018; Bai, 2019; Cytrynbaum, 2021; Kallus, 2018; Bertsimas et al., 2015), adaptive designs for policy choice (Kasy and Sautmann, 2019; Russo et al., 2018; Kato et al., 2024; Cesa-Bianchi et al., 2025), and experimental design under correct model specification (Silvey, 2013; Chaudhuri and Mykland, 1993; Chaloner and Verdinelli, 1995; Higbee, 2024; Kiefer and Wolfowitz, 1959; Reeves et al., 2024; Viviano, 2020; Kasy, 2016).
+Related literature
+This paper links the experimental-design literature—which mostly focuses on settings where all parameters are identified within the experiment and leaves aside questions of data combination—to recent work that integrates experimental and (reduced-form or structural) observational evidence to extrapolate effects in complex scenarios.
+Recent advances for experimental design include balancing and variance-minimizing allocations (Tabord-Meehan, 2018; Bai, 2019; Cytrynbaum, 2021; Kallus, 2018; Bertsimas et al., 2015), adaptive designs for policy choice (Kasy and Sautmann, 2019; Russo et al., 2018; Cesa-Bianchi et al., 2025), and experimental design under correct model specification (Silvey, 2013; Chaudhuri and Mykland, 1993; Chaloner and Verdinelli, 1995; Higbee, 2024; Kiefer and Wolfowitz, 1959; Reeves et al., 2024; Viviano, 2020; Kasy, 2016).
 Traditional work on experimental design for robust-model estimation either focused on testing competing models (Atkinson and Fedorov, 1975; López-Fidalgo et al., 2007), or on using a-priori knowledge of (worst-case) bias for the design of an experiment (Box and Draper, 1959; Wiens, 1998; Tsirpitzi et al., 2023; Sacks and Ylvisaker, 1984).
-We complement this literature by introducing and studying the question of which experiment to design when combined with observational evidence (when not all parameters can be learned experimentally).
+All these references leave aside questions about observational data combination.
+In the context of minimizing the variance of an experiment, Rosenman and Owen (2021) proposes to use observational studies to construct high-confidence bounds on the experimental variance, and then focus on variance-optimal stratification.
+This problem differs from our design problem and analysis, which instead studies the use of observational studies in combination with experiments for extrapolation.
+In summary, we complement this literature by studying which experiment to run (and how) when not all parameters relevant for the target estimand can be learned experimentally.
 Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design.
@@
 (2025) among others.
-These focus on settings where researchers have only access to experimental variation.
-Here instead we study the problem of choosing which experiment to run (and how) when combined with observational evidence.
-We also complement literature that leverages correctly-specified models for site selection (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
-A related line of work analyzes estimation under misspecification (Armstrong and Kolesár, 2018; Armstrong et al., 2024; Andrews et al., 2017, 2020; Bonhomme and Weidner, 2022; Christensen and Connault, 2023; James et al., 1961), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; Bhattacharya, 2013; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024).
+These references focus on settings where researchers have only access to experimental variation, instead of combining it with observational evidence, motivating different designs and objective functions.
+In the context of site-selection, we complement literature that leverages correctly-specified models (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
+A related line of work analyzes estimation under misspecification (Armstrong and Kolesár, 2018; Armstrong et al., 2024; Andrews et al., 2025; Andrews et al., 2020; Bonhomme and Weidner, 2022; Christensen and Connault, 2023; James et al., 1961), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; Bhattacharya, 2013; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024).
 We adapt the worst-case proportional regret idea from Armstrong et al.
 (2024) and Tsybakov (1998) here studied in the context of experimental design.
-Our definition of sensitivity of the estimand to each parameter directly links to sensitivity analysis in Andrews et al.
+Our definition of sensitivity of the estimand to each parameter builds on sensitivity analysis in Andrews et al.
 (2020).
 Unlike both strands of this literature where the design is fixed and researchers optimize over the choice of the estimator only, here we optimize the design itself.
-Consequently, the objective changes: regret is computed against the best design-estimator combination, and the resulting estimator is learned jointly with the design.
-Finally, we connect to a growing empirical literature that combines observational (or model-based) evidence with experimental variation to learn economic quantities – spanning general equilibrium effects (Egger et al., 2022; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023), external-validity counterfactuals (Gechter et al., 2024; Banerjee et al., 2024), large-scale information campaigns (de Albuquerque et al., 2025; Larroucau et al., 2024), and market-system experiments (Bergquist and Dinerstein, 2020; Allende et al., 2019).
-We provide a unifying framework to decide what to learn experimentally, how to split the sample across experiments, and how to integrate experimental and observational evidence.
+The objective and analysis change: regret is computed against the best design-estimator combination, and the resulting estimator is learned jointly with the design.
 2 Problem description
-Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$.
-The estimand may arise from a reduced-form (e.g., treatment effect in a given area) or a structural model.
+Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional unknown parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$.
 For simplicity, we assume that the estimand is univariate, while Section 5.3 shows how our framework directly extends to multivariate $\tau(\theta)\in\mathbb{R}^{q}$ for $q>1$.
@@
 Our question is how to design such experiments under feasibility constraints.
-2.1 Setup
+Basic setup
 We assume that researchers have access to estimators (and their variance) of $\theta$ denoted as $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$.
 We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model.
-However, $\tilde{\theta}^{\mathrm{obs}}$, that for exposition we define observational estimates, have unknown biases collected in a vector $b\in\mathbb{R}^{p}$.
-Examples of $\tilde{\theta}^{\mathrm{obs}}$ include a structural model estimate with confounding, an estimate from an instrumental variable regression that may fail the exclusion restriction, or an estimate from a country different from the one of interest that may lack external validity.
-Researchers may also collect evidence for a subset of parameters that they know is unbiased.
-For instance, researchers may generate exogenous variation via an experiment (or more generally acquiring more data) that identifies one or some of the parameters of interest.
-We refer to these parameters as experimental estimates, adopting here the definition of experimental estimates as parameters that the researcher knows are unbiased.
-[^5] In some applications researchers may be confident that a few observational estimates have no bias.
-It is possible to extend our framework to these settings with a change of notation where those observational estimates assumed to be unbiased are defined as experimental estimates, and such experimental estimates are always contained in the set $\mathcal{E}$ defined below.
-Setting 1 (Main setup).
-For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a known strictly positive-definite matrix $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})\subset\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$ with uniformly bounded entries, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
+However, $\tilde{\theta}^{\mathrm{obs}}$, defined as observational estimates, have unknown biases collected in a vector $b\in\mathbb{R}^{p}$.
+Examples include an estimate from an instrumental variable regression that may fail the exclusion restriction, or an estimate from a country different from the one of interest that may lack external validity.
+Researchers can also collect experimental evidence for a subset of parameters, producing estimates whose bias equals zero by design.
+Setting 1.
+For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a known positive-definite matrix $\Sigma(\mathcal{E})\in\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
 $\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{obs}}\right]-\theta=b,\qquad\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\right]-\theta_{\mathcal{E}}=0,$
-for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance as
-$\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E}).$ (1)
-Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators
+for an unknown bias vector $b\in\mathbb{R}^{p}$.
+Moreover, define its joint variance-covariance matrix as $\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E}).$ Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators
 $\hat{\tau}_{\gamma}\equiv\tau(\hat{\theta}(\gamma)),\qquad\hat{\theta}_{j}(\gamma)=\begin{cases}\gamma_{j}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
@@
 We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012) (in practice, any consistent estimator, e.g., from pilot studies would suffice).
-In addition, we assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite.
-In addition $\Sigma$ is strictly positive definite, therefore assuming that the variance is bounded away from zero.
-[^6] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance.
+The class of experiments and estimators is assumed to be in a user-specific set $\mathcal{D}$.
+Assumption 1 (Feasible experiments and estimators).
+We write compactly $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible design (i.e., choice of $\mathcal{E}$ and $\Sigma(\mathcal{E})$) and estimators (i.e., $\gamma$).
+The set $\mathcal{D}$ can be arbitrary as long as (i) each feasible $\Sigma$ has uniformly bounded entries and is strictly positive definite; (ii) $\gamma_{j}=1$ for all $j\in\mathcal{E}$ is one admissible choice (among potentially other values).
+We let $\mathcal{E},\Sigma(\mathcal{E})$ be within arbitrary constraint sets that may encode feasibility or budget constraints.
+In our framework, restrictions on the experiments researchers can run $\mathcal{E}$ and on their precision $\Sigma(\mathcal{E})$ can take any desired form.
+In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or constraints on their power (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011).
+We assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite, and strictly positive definite, therefore assuming that the variance is bounded away from zero.
+[^5] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined without loss as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance.
 See Section 5.1 for details.
-We let $\mathcal{E}\in\mathcal{S},\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$ for some constraint sets $(\mathcal{S},\mathcal{G}(\mathcal{E}))$ encoding feasibility or budget constraints.
-In our framework, restrictions on the experiments researchers can run $\mathcal{E}\in\mathcal{S}$ (and on their precision $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$) can take any desired form.
-In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or standard lower bound constraints on power for experimental estimates, which are common in empirical practice (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011)
-We write compactly $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible experiments (e.g., sample size, etc.), with $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ for a choice of experiments $\mathcal{E}$.
-Our goal is to optimize both over the design and the weights $\gamma$. Before introducing our main research question, we impose the following assumption.
-Assumption 1 (First-order estimation error).
-For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume
-$\tau(\theta)-\tau\!\big(\hat{\theta}(\gamma)\big)=\sum_{j=1}^{p}\omega_{j}\,\big(\theta_{j}-\hat{\theta}_{j}(\gamma)\big),$ (2)
-for known weights $\omega\in\mathbb{R}^{p}$, with $|\omega_{j}|>0$ for all $j$.
-Assumption 1 holds exactly for linear $\tau(\cdot)$ and serves as a first-order approximation for smooth nonlinear $\tau(\cdot)$.
-Remark 1 (Non-linear $\tau(\theta)$).
-Section 4.1 (and Example 3 below) extends our framework to non-linear estimands and shows how Assumption 1 serves as an approximation via a Taylor expansion within a local asymptotic framework, where the bias is local to zero, but at a possibly faster, equal or even slower rate than the estimator’s standard error.
-For non-linear $\tau$, $\omega$ is replaced by the gradient of $\tau$ evaluated at the baseline observational estimates (i.e., $\omega=\frac{\partial\tau(\theta)}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}+o_{p}(1)$), which is typically observed before the experiment is conducted.
-Remark 2 (When researchers can only identify linear combinations of the parameters).
-In some applications, researchers may be able to identify via an experiment a linear combination of the parameters of interest (up-to a first order approximation), defined $\phi\equiv W^{\top}\theta$, where $W\in\mathbb{R}^{p\times s}$ is a known matrix.
-Suppose that $\widetilde{\phi}$ is an unbiased estimator of $\phi$, and the researcher can access a subset of its entries via an experiment.
-Our framework extends to this setting as follows.
-Define the coefficient of the linear projection of $\omega$ onto the column space of $W$: $\pi=\left(W^{\top}W\right)^{-1}W^{\top}\omega$ (known ex-ante).
-Define the residual $v=\omega-W\pi$.
-Then,
-$\tau=\left[W\pi+v\right]^{\top}\theta=\pi^{\top}\phi+v^{\top}\theta\implies\tau=\check{\omega}^{\top}\check{\theta},$
-where $\check{\omega}\equiv(\beta^{\top},v^{\top})^{\top}$ and $\check{\theta}=(\phi^{\top},\theta^{\top})^{\top}$.
-This corresponds to our original setting where the researcher’s feasible set only allows access to unbiased estimates of a subset of the parameters, corresponding to the entries of $\phi$.
+The shrinkage estimator can be unrestricted (i.e., $\gamma$ can take any value on the real line) or restricted (when for example researchers aim to only use experimental evidence when available, setting $\gamma_{j}=1$ for $j\in\mathcal{E}$).
+We extend our framework to more general methods of moments estimators in Section 4.
+Remark 1 (Structure of $\Sigma$).
+Because observational and experimental moments are typically computed from disjoint samples, in practice, it is natural to model the covariance as block diagonal, $\Sigma=\operatorname{diag}(\Sigma_{\text{obs}},\Sigma_{\text{exp}})$.
+The first block corresponds to observational estimates (fixed and not subject to design choices), while the second corresponds to experimental estimates, and it is a function of the sample size allocated to treatment arms.
+This is a special case of our framework which can be encoded in the constraint set $\mathcal{D}$.
 ∎
-2.2 Research questions
+Research questions
 Our design problem can be described in three steps:
   1. 1.
-    Preliminary step: weights. For each $(\mathcal{E},\Sigma(\mathcal{E}))\in\mathcal{D}$, choose $\gamma$ to combine experimental and observational estimates on the selected coordinates.
+    Preliminary step: weights.
+For each feasible $\mathcal{E},\Sigma(\mathcal{E})$, choose $\gamma$ to combine experimental and observational estimates on the selected coordinates.
+(If researchers restrict their estimators to only use experimental evidence when available, set $\gamma=1$.)
   2. 2.
-    Middle step: precision.
-For each $\mathcal{E}$, choose $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$, where $\mathcal{G}(\mathcal{E})$ denotes an arbitrary set of positive definite covariances under feasibility constraints.
+    Middle step: precision. For each feasible $\mathcal{E}$, choose $\Sigma(\mathcal{E})$ (i.e., sample allocation).
   3. 3.
-    Outer step: experiment choice. Choose $\mathcal{E}\in\mathcal{S}$, the feasible set of parameters to learn experimentally for given arbitrary constraint set $\mathcal{S}$.
+    Outer step: experiment choice. Choose a feasible set of experiments $\mathcal{E}$.
 Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025).
-We study this question here in combination with the experimental design problem, step 2 and 3, which has not been studied by the references above; this, as we show, will change the underlying optimization problem (also for $\gamma$).
-The central challenge is that performance depends on the unknown bias $b$ in the observational estimates.
-2.3 Mapping to a simple two parameters model
-While we derive our results in general form, we will build intuition using a stylized description with a two-parameter model.
-Setting 2 (Illustration with two-parameter model).
-Consider two parameters $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates
-$\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\qquad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\qquad j=1,2.$
-We may run one experiment: $\mathcal{S}=\{\{1\},\{2\}\}$. If we pick $j$, we estimate
-$\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\qquad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$
-and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$
+We study this question here in combination with the experimental design problem, step 2 and 3; this, as we show, will change the underlying optimization problem (also for $\gamma$).
+The central challenge is that performance depends on the unknown bias $b$.
+Mapping $\tau$
+A key feature of our framework is that researchers explicitly parameterize $\theta$ and $\tau(\cdot)$, thereby pre-specifying and making transparent which biases drive the design problem; we see this as an advantage because it clarifies the role of the experiment in complementing existing evidence.
+The sensitivity of $\tau$ to $\theta$, captured by its gradient $\omega(\theta)=\partial\tau(\theta)/\partial\theta$, determines how the (first-order) bias propagates onto the final estimand.
+For simplicity, we impose exact linearity, $\tau(\theta)=\omega^{\top}\theta$; all results extend to smooth, nonlinear $\tau$ via a first-order expansion, as discussed below.
+Assumption 2 (First-order estimation error).
+For any $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume
+$\tau(\theta)-\tau\!\big(\hat{\theta}(\gamma)\big)=\sum_{j=1}^{p}\omega_{j}\,\big(\theta_{j}-\hat{\theta}_{j}(\gamma)\big),$ (1)
+for known weights $\omega\in\mathbb{R}^{p}$, with $|\omega_{j}|\in(0,\infty)$ for all $j$.
+Assumption 2 holds exactly when $\tau(\cdot)$ is linear and serves as a first-order approximation for smooth non-linear $\tau(\cdot)$.
+Section 5.1 (and Example 3) extend the framework to non-linear estimands encompassing as special cases the local-misspecification setups in Andrews et al.
+(2020); Armstrong and Kolesár (2021).
+For non-linear $\tau$, we replace $\omega$ with the gradient of $\tau$ evaluated at a preliminary observational estimate, i.e., $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}\;+\;o_{p}(1),$ which is typically available before the experiment is conducted.
+That is, building on references above, $\omega$ captures the first-order effect of the observational bias $b$ on the bias of $\hat{\tau}$.
+[^6] To see how this assumption holds under a first-order approximation, define $\theta$ the parameter rescaled by $\sqrt{n}$ (and similarly its bias $b_{n}$, indexed by $n$), where $n^{-1/2}$ is the rate of convergence of the observational estimators.
+Then second-order effects are negligible (and Assumption 2 holds asymptotically) provided that $b_{n}^{2}/\sqrt{n}\rightarrow 0$.
+This implies that Assumption 2 holds when the asymptotic bias $b_{n}$ is proportional to, smaller or even much larger than the asymptotic variance, provided that $b_{n}$ does not grow at rate faster than $n^{1/4}$.
+We conclude with three examples in simple two-parameter models.
 Example 1 (Choosing the site for an experiment for external validity).
@@
 Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al.
-(2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so $\mathcal{S}=\{\{1\},\{2\}\}$.
+(2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so that $\mathcal{E}\in\{\{1\},\{2\}\}$.
 If site $j$ is chosen, we obtain an unbiased $\tilde{\theta}^{\mathrm{exp}}_{j}$; the other site $k\neq j$ remains observational.
-Given $\mathcal{E}\in\mathcal{S}$,
+Given $\mathcal{E}$,
 $\widehat{\tau}_{\gamma}=\tfrac{1}{2}\!\left[\gamma_{j}\,\tilde{\theta}^{\mathrm{exp}}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j}+\tilde{\theta}^{\mathrm{obs}}_{k}\right],\qquad k\neq j.$
@@
 Example 3 (Supply or demand experiment with non linear target).
-(Bergquist and Dinerstein, 2020) conduct demand and supply experiments in food markets in Kenya.
+Bergquist and Dinerstein (2020) conduct demand and supply experiments in food markets in Kenya.
 Suppose here we are interested in similar applications in Uganda.
@@
 Let $(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top}$ denote baseline estimates from Kenya, that may lack external validity.
-Due to cost constraints, $\mathcal{S}=\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms).
+Due to cost constraints, $\mathcal{E}\in\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms).
 Let
-$\big(\tilde{\theta}^{\mathrm{obs}}-\theta\big)\ \sim\ \mathcal{N}\big(b,\Sigma\big),\quad\tilde{\theta}^{\mathrm{obs}}\equiv\sqrt{n}(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top},\quad\theta\equiv\sqrt{n}(\beta_{D},\beta_{S})^{\top}$
+$\big(\tilde{\theta}^{\mathrm{obs}}-\theta\big)\ \sim\ \mathcal{N}\big(b,\Sigma^{\mathrm{obs}}\big),\quad\tilde{\theta}^{\mathrm{obs}}\equiv\sqrt{n}(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top},\quad\theta\equiv\sqrt{n}(\beta_{D},\beta_{S})^{\top}$
 with $\theta$ and $\tilde{\theta}$ denoting the parameter and estimator here rescaled by the square-root of the sample size, and $b$ capturing bias.
 Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$.
-For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda, under local asymptotics and a first-order Taylor expansion described in Section 4.1, we can write [^7] Here, we are using a local asymptotic framework where the bias of the estimates in Kenya is non-negligible but small relative to $n$, i.e., $b/\sqrt{n}=o(1)$, so that $\omega$ can be consistently estimated using observational data.
-$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\check{\omega}(\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}}))^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{b}{n^{1/2}}\right)}_{\text{small higher order effects}},\qquad\check{\omega}(\beta)=\frac{t}{(\beta_{D}+\beta_{S})^{2}}\begin{pmatrix}-\beta_{S}\\ \beta_{D}\end{pmatrix}$,
+For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda under local asymptotics described in Section 5.1
+$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\check{\omega}(\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}}))^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{1}{n^{1/2}}\right)}_{\text{small higher order effects}},\qquad\check{\omega}(\beta)=\frac{t}{(\beta_{D}+\beta_{S})^{2}}\begin{pmatrix}-\beta_{S}\\ \beta_{D}\end{pmatrix}$,
 where we can replace $\check{\omega}$ with its consistent counterpart $\check{\omega}(\tilde{\beta}^{\mathrm{obs}})$.
@@
 3 Robust experimental design
-In this section we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$, a common measure of precision.
-In Section 4.2 we extend the framework to minimize the length of the confidence intervals.
-3.1 Adaptation regret for experimental design
-For a given design $(\mathcal{E},\Sigma(\mathcal{E}))$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
+Next, we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$, defined, for given bias level $b$ and design $(\mathcal{E},\Sigma(\mathcal{E}))$ as
 $\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[(\widehat{\tau}_{\gamma}-\tau)^{2}\right],$
 where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$.
-The MSE is a measure of precision of $\hat{\tau}$ and ideally, one would minimize $\mathrm{MSE}_{b}$ over $(\mathcal{E},\Sigma(\mathcal{E}),\gamma)\in\mathcal{D}$.
-However, because $b$ is unknown, we first consider an uncertainty set given by an $\ell_{\infty}$-ball,
+Section 5.2 shows that our results for MSE also extend when minimizing confidence intervals’ length.
+Ideally, one would minimize $\mathrm{MSE}_{b}$. However, because $b$ is unknown, we consider an uncertainty set given by an $\ell_{\infty}$-ball,
 $\mathcal{B}(B)\equiv\{b:\ \|b\|_{\infty}\leq B\},$
 with $B\geq 0$ an upper bound on the largest coordinate-wise bias.
-The $\ell_{\infty}$ choice has a desirable property relative to $\ell_{1}$ or $\ell_{2}$: it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters.
+In our framework, a key challenge is that biases may arise from multiple parameters and uncertainty sets defined by different norms may be possible.
+A key insight here is that the $\ell_{\infty}$ choice has a desirable property relative to $\ell_{1}$ or $\ell_{2}$: it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters.
 The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
@@
 However, having to specify $B$ can pose a large burden on the researchers and make the choice of the experiment sensitive to $B$.
-Nevertheless, we could seek designs that perform as close as possible to the oracle that knows $B$, uniformly over the values of $B$.
-This idea of performing as close as possible to the oracle is based on an extensive literature on regret minimization (Manski, 2004; Manski and Tetenov, 2007; Montiel Olea et al., 2023; Kitagawa and Tetenov, 2018; Manski and Tetenov, 2016).
-We minimize the worst-case proportional increase in MSE relative to the oracle across all bias radii:
+Therefore, we seek designs that perform as close as possible to the oracle that knows $B$, uniformly over the values of $B$, following a long-standing tradition in decision theory (e.g.
+Manski, 2004; Manski and Tetenov, 2007; Montiel Olea et al., 2023; Kitagawa and Tetenov, 2018).
+Specifically, we minimize the worst-case proportional regret
 $\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;\equiv\;\sup_{B\geq 0}\ \frac{\displaystyle\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\displaystyle\mathrm{MSE}^{*}(B)},$
 defined adaptation regret by Armstrong et al. (2024) (building in turn on Tsybakov, 1998).
-Whereas the above references consider the choice of an estimator for fixed design, here, the regret is relative to an oracle that chooses both the estimator and the design.
-Our key contribution – different from the references above – is the study of the optimal experimental design informed by biased (observational) estimates.
-This leads to a different definition of regret optimization, both because we optimize over the class of designs and $\Sigma$, but also in $\gamma$ due to the different definition of oracle.
-The optimal solution consists of a pair of design and estimator that can be pre-specified.
-3.2 Optimal experimental design: main result
-Before stating the result, we introduce compact notation. For a given $(\mathcal{E},\gamma)$, with $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ define
-$\tilde{\omega}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega_{j},&j\notin\mathcal{E},\\ \omega_{j}(1-\gamma_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}_{\text{exp}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ\gamma$
-where $\circ$ denotes the component-wise product and $\omega_{\mathcal{E}}$ is the subvector of $\omega$ corresponding to the entries indexed by $\mathcal{E}$. Let
-$\tilde{\omega}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$ (3)
-The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
-$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},$ (4)
-with $|1-\gamma|$ indicating the absolute value of each entry of the vector $1-\gamma$.
-Define the best achievable variance and worst-case bias divided by $B^{2}$ as
-$\alpha^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\alpha(\mathcal{E},\Sigma,\gamma),\qquad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\beta(\mathcal{E},\gamma)=\min_{\mathcal{E}\in\mathcal{S}}\Big(||\omega||_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2}.$ (5)
-Here $\alpha^{\star}$ is the variance of the most precise feasible design, while $\beta^{\star}$ is the squared bias of the least biased design.
-[^8] The expression for $\beta^{\star}$ corresponds to the following experiment choice: choose $\mathcal{E}$ with the largest $\omega_{\mathcal{E}}||_{1}$ and set $\gamma_{\mathcal{E}}\equiv 1$ at those coordinates.
-Both quantities depend on the class of designs $\mathcal{S},\mathcal{G}(\mathcal{E})$ and are easy to compute: $\alpha^{\star}$ is a constrained quadratic minimization; $\beta^{\star}$ requires enumerating $\mathcal{E}\in\mathcal{S}$.
-Ideally, we would like a design with variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$.
+There are two key distinctions relative to such references using the adaptation regret.
+First (and importantly) we optimize over the design itself rather than fixing it ex ante; consequently, regret is defined relative to an oracle that jointly chooses both the estimator and the design.
+Second, we allow biases to arise across multiple parameters (rather than a single one), which necessitates a careful specification of the uncertainty set; we adopt an $\ell_{\infty}$-ball, which, as we show yields transparent and computationally-attractive solutions.
+3.1 Optimal design
+A natural question is whether the adaptation regret provides natural trade-offs between the variance and the bias.
+To do so, denote $\mathbb{V}_{\mathcal{E},\Sigma}(\cdot)$ the variance operator for given $(\mathcal{E},\Sigma)$.
+Definition 1 (Variance regret).
+Let
+$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\mathbb{V}_{\mathcal{E},\Sigma}(\hat{\tau}_{\gamma}),\qquad\alpha^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\alpha(\mathcal{E},\Sigma,\gamma),$ (2)
+with $\alpha^{\star}$ denoting the smallest feasible variance. We will refer to $\alpha/\alpha^{\star}$ as the variance regret.
+The variance regret denotes the ratio between the variance of a given design and estimator relative to the smallest achievable variance.
+We provide a similar definition for the bias.
+Let $|1-\gamma|$ define the absolute value of each entry of the vector $1-\gamma$.
+Definition 2 (Bias regret).
+Define
+$\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},\quad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\Big(||\omega||_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2},$
+with $\beta/\beta^{\star}$ defined as the bias regret.
+[…]
+An ideal design, would set the variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$.
 Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other.
-Instead, in our main theorem below we show that the regret-optimal pair of design and estimator trades-off the bias and variance and depends on how far each of these two components are from their smallest attainable value.
+This raises the question of how to trade-off these two components.
+This trade-off is formally characterized in our main theorem below.
 Theorem 1.
-Consider Setting 1 and let Assumption 1 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
+Consider Setting 1 and let Assumptions  1, 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$,
 $\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}.$
 Proof.
-See Appendix B.1. ∎
-Theorem 1 provides an explicit expression of the regret (which does not require researchers to specify $B$).
-To our knowledge, it is the first result to study adaptation regret for experimental design.
-Its derivation requires us to find precise binding patterns as a function of $B$, which accounts for its effect on optimizing the oracle solution over the design and estimator.
-The adaptation regret balances the variance of a given design relative to the smallest possible variance $\alpha^{\star}$ against the worst-case bias $\beta$ relative to the smallest possible worst-case bias $\beta^{\star}$.
-To gain insight into the optimal solution, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), we may expect the minimizer $(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})\in\arg\min\mathcal{R}(\mathcal{E},\Sigma,\gamma)$ to equalize
-$\frac{\alpha(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})}{\alpha^{\star}}\;=\;\frac{\beta(\mathcal{E}^{\star},\gamma^{\star})}{\beta^{\star}}.$
-That is, at the optimum the design equalizes the estimator’s variance $\mathbb{V}(\hat{\tau}_{\gamma})$ (over the design choice) divided by the smallest feasible variance $\alpha^{\star}$ to the worst-case squared bias (equal to $\beta B^{2}$) divided by the smallest achievable worst-case squared bias $\beta^{\star}B^{2}$.
+See Appendix A.1. ∎
+Theorem 1 provides a simple expression of the regret which does not require researchers to specify $B$.
+The key insight is that the adaptation regret in our context, leads to a (strictly) quasi-convex objective function, whose worst-case solution is the maximum between the variance and the bias regret.
+This provides a simple and intuitive trade-off between the two.
+Specifically, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), it follows that at the optimum we equalize the variance and bias regret as in Figure 1
+$\frac{\alpha}{\alpha^{\star}}\;=\;\frac{\beta}{\beta^{\star}}.$
+In practice, this amount to allocate sample size to noisier treatment arms when the variance dominates, and select treatment arms with largest sensitivity when the bias dominates.
 […]
@@
 Level sets of the objective $\mathcal{R}=\max\{\alpha/\alpha^{\star},\ \beta/\beta^{\star}\}$ are axis–aligned squares (the dotted inverted “L” shows the smallest such square touching the frontier).
-The minimizer outside boundary solutions is where the frontier meets the $45^{\circ}$ line (red dot), i.e., where the two normalized components are equalized: $\alpha(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})/\alpha^{\star}=\beta(\mathcal{E}^{\star},\gamma^{\star})/\beta^{\star}$.
-Theorem 1 provides us an immediate and simple to compute solution to the optimal design problem via backward induction.
-Specifically, we compute the estimator, the optimal variance and the experimental choice as follows:
+The minimizer outside boundary solutions is where the frontier meets the $45^{\circ}$ line (red dot).
+Theorem 1 provides us an immediate solution to the optimal design problem. We compute the estimator, the optimal variance and the experimental choice via backward induction:
 $\gamma^{\star}(\mathcal{E},\Sigma)$ $\in\mathrm{arg}\min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}$
@@
 $\mathcal{E}^{\star}$ $\in\mathrm{arg}\min_{\mathcal{E}\in\mathcal{S}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma^{\star}(\mathcal{E}),\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\beta^{\star}}\right\}.$
-Optimization proceeds as follows:
-  • we first choose $\gamma^{\star}$ for each design $\mathcal{E},\Sigma(\mathcal{E})$.
-Importantly, the choice of $\gamma^{\star}$ depends on the class of designs available to researchers ($\mathcal{D}$), because $\mathcal{D}$ determines the oracle values $\alpha^{\star},\beta^{\star}$ in the denominators of the objective function (a larger class of designs improves the oracle solution);
-  • the choice of the experimental variance $\Sigma$ (how precise each experiment is under budget constraints) also depends on the bias component $\beta$ because $\beta$ affects how we select the shrinkage weights $\gamma^{\star}$ in the previous step;
-  • the optimal design $\mathcal{E}^{\star}$ then accounts for both choices of $\gamma^{\star}$ and $\Sigma^{\star}$.
-3.3 Example with two parameters
-Consider the two-parameters model of Setting 2 where for simplicity we fix the choice of the variances $\sigma^{2},v^{2}>0$. The smallest variance and worst–case squared bias (per unit $B^{2}$) are
-$\alpha^{\star}=\min_{k\in\{1,2\}}\underbrace{\left\{\ \omega_{-k}^{2}\sigma_{-k}^{2}\;+\;\omega_{k}^{2}\frac{\sigma_{k}^{2}\,v_{k}^{2}}{\sigma_{k}^{2}+v_{k}^{2}}\ \right\}}_{\text{Variance with variance-optimal }\gamma},\qquad\beta^{\star}=\min_{k\in\{1,2\}}\underbrace{\omega_{k}^{2}}_{\text{Bias}^{2}/B^{2}\text{ with bias-optimal }\gamma}.$
-Here, $\alpha^{\star}$ follows from the variance-only optimal shrinkage $\gamma_{k}=\sigma_{k}^{2}/(\sigma_{k}^{2}+v_{k}^{2})$ (i.e., the oracle choice when $B=0$); $\beta^{\star}$ follows from the fact that the oracle that minimizes the worst-case bias chooses the experiment $k$ with the largest $\omega_{k}$, sets $\gamma_{k}=1$ and incurs a bias proportional to $|\omega_{-k}|$, since no experiment is conducted for k-k.
-The variance and worst-case squared bias (divided by $B^{2}$) incurred by an analyst choosing experiment $j$ and $\gamma_{j}$ are, respectively
+That is, we first choose $\gamma^{\star}$ for each design $\mathcal{E},\Sigma(\mathcal{E})$.
+We then choose $\Sigma$, whose choice depends on $\beta$ through the shrinkage weights.
+We finally choose the set of experiments $\mathcal{E}$.
+3.2 Intuition with two parameters
+To build further intuition, it is useful to study the two-parameter model.
+Setting 2 (Illustration with two-parameter model).
+Let $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\quad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\quad j=1,2.$ For simplicity $\sigma_{j}^{2},v_{j}^{2}$ are fixed.
+We may run one experiment due to budget constraints: $\mathcal{E}\in\{\{1\},\{2\}\}$.
+If we pick $j$, we estimate $\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\quad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$ and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$
+How do the variance and bias regret look like in this easier scenario? We can write
+$\alpha^{\star}=\min_{k\in\{1,2\}}\left\{\ \omega_{-k}^{2}\sigma_{-k}^{2}\;+\;\omega_{k}^{2}\frac{\sigma_{k}^{2}\,v_{k}^{2}}{\sigma_{k}^{2}+v_{k}^{2}}\ \right\},\qquad\beta^{\star}=\min_{k\in\{1,2\}}\omega_{k}^{2}.$
+Here, $\alpha^{\star}$ follows from the variance-only optimal shrinkage $\gamma_{k}=\sigma_{k}^{2}/(\sigma_{k}^{2}+v_{k}^{2})$ (i.e., the oracle choice when $B=0$); $\beta^{\star}$ equals the smallest between the two sensitivity parameters $\omega$.
+Both depend on the choice of the estimator and on the class of experiments.
+We cannot achieve both $\alpha^{\star}$ and $\beta^{\star}$, since these require different choices of $\gamma$. Instead,
 $\alpha(j,\gamma_{j})$ $\equiv\underbrace{\omega_{-j}^{2}\sigma_{-j}^{2}}_{\text{var obs estimate}}+\underbrace{\omega_{j}^{2}\!\big[(1-\gamma_{j})^{2}\sigma_{j}^{2}+\gamma_{j}^{2}v_{j}^{2}\big]}_{\text{var exp and obs estimate}},\quad\beta(j,\gamma_{j})$ $\equiv\big(\underbrace{|\omega_{-j}|}_{\begin{subarray}{c}\text{bias obs}\\ \text{estimate}/B\end{subarray}}+\underbrace{(1-\gamma_{j})|\omega_{j}|}_{\begin{subarray}{c}\text{bias obs minus}\\ \text{ bias exp}/B\end{subarray}}\big)^{2}.$
-By Theorem 1, the worst-case adaptation regret for choosing experiment $\{j\}$ and $\gamma_{j}$, equals
-$\mathcal{R}(\{j\},\gamma_{j})=\max\Big\{\underbrace{\alpha(j,\gamma_{j})/\alpha^{\star}}_{\text{Variance/}\alpha^{\star}},\ \underbrace{\beta(j,\gamma_{j})/\beta^{\star}}_{\text{Bias}^{2}/(B^{2}\beta^{\star})}\Big\}.$
-Both maximand components are evaluated with the same choice of $\gamma$.
-3.3.1 Choosing $\gamma_{j}$ for given design $\{j\}$
-The first step is to choose the shrinkage weight $\gamma$, noting that regret is computed relative to an oracle that optimizes both $\gamma$ and the experiment design $\{j\}$.
-Minimizing $\mathcal{R}$ over $\gamma_{j}$ yields a rule that either equalizes the two components or selects a boundary solution.
-Corollary 1.
-Consider Setting 2 and let Assumption 1 hold. Then
+The very first step is optimizing over the estimator (assuming here for simplicity $\gamma$ can take any value). The following characterizes the regret optimal $\gamma^{\star}$ (see Appendix C.1).
+[…]
+At the boundaries, if the variance regret $\alpha/\alpha^{\star}$ is always smaller than the bias regret $\beta/\beta^{\star}$, the optimal weight is $\gamma_{j}^{\star}=1$ (i.e., use only the experimental estimate).
+If instead the variance regret always dominates, the optimal choice is the variance-optimal $\gamma_{j}^{\star}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
+When the solution is an interior solution for $\gamma$, the optimal design minimizes both the bias and variance regret (now equalized), so that
+$j^{\star}\in\arg\min_{j\in\{1,2\}}\underbrace{\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}}_{\text{overall variance at }\gamma_{j}^{\star}}=\arg\min_{j\in\{1,2\}}\underbrace{\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}}_{\text{worst--case bias}/|B|\text{ at }\gamma_{j}^{\star}}.$
+However, cases at the extreme boundaries can also occur.
+For instance, if the variance regret $\alpha/\alpha^{\star}$ or the bias regret $\beta/\beta^{\star}$ uniformly dominates the other for one of the experiments, the optimal design minimizes the dominating term.
+These are desiderable properties as the optimal design always prioritizes the dominating term.
+Appendix Table 5 provides a complete overview of all possible scenarios.
+More complex scenarios where researchers may allocate sample sizes to each experiment are possible, see Section 6.
+Figure 2: Example for $j=2$.
+Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$.
+Bottom row: the optimal weight $\gamma_{2}^{\star}$.
+Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
+Figure 3: Example for $j=2$ (regret).
+Different colors correspond to the regret for choosing either experiment.
+Figure reports $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$).
+The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret.
+Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
+Example 4 (Numerical illustration).
+For an illustrative example, consider choosing $\gamma_{2}^{\star}$ when experiment $j=2$ is twice as precise as $j=1$, with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, $v_{1}=1$, $v_{2}=0.5$, and $\omega_{1}=1$.
+The optimal $\gamma_{2}^{\star}$ depends not only on features of design $j$ but also on those of the alternative design j-j, $(v_{-j}^{2},\sigma_{-j}^{2},\omega_{-j})$, through the benchmarks $(\alpha^{\star},\beta^{\star})$.
+Figure 2 (top row) plots $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$.
+The bottom row reports the corresponding $\gamma_{2}^{\star}$.
+Varying $\omega_{2}$ yields two key effects: both very small and very large $\omega_{2}$ push $\gamma_{2}^{\star}$ toward one, but for different reasons—when $\omega_{2}$ is small the oracle bias decreases faster than the experimental variance (since $\beta^{\star}=\omega_{2}$), while when $\omega_{2}$ is large the oracle bias is $\beta^{\star}=\omega_{1}$ and $\omega_{2}$ has only on the feasible experiment’s bias.
+Consequently, both regret and the optimal estimator depend on the oracle’s choice of the experiment.
+In the interior region, the optimal weight first shifts toward the variance–optimal solution (as the oracle bias declines) and then rises again toward a bias–preferred solution.
+When we vary the experimental and observational variances, we observe the expected patterns: $\gamma_{2}^{\star}$ moves inversely with experimental variance and in the same direction as observational variance.
+These patterns translate into implications for the regret of the chosen experiment (Figure 3).
+For the second arm $j=2$, regret decreases sharply as $\omega_{2}$ increases and then decreases, with further decreases becoming gradual once the bias component is already small.
+By contrast, for the first arm, regret is at best slightly decreasing for small changes in $\omega_{2}$ (when variance effects dominate), but eventually increases with $\omega_{2}$ because the oracle increasingly favors arm $2$, so the opportunity cost of selecting arm $1$ increases.
+A summary appears in Table 4, with a complete discussion in Appendix C.2.
+∎
+4 Generalization to moments selection
+In many scenarios, we may expect that experiments only identify relevant moments in the data. We now extend the framework to generalized method of moments (GMM) estimators.
+Consider a vector of moment conditions $g(\theta)\in\mathbb{R}^{p}$ that stacks all moments (observational and experimental) potentially available for estimation.
+Let $\Lambda$ denote the Jacobian of the moment function with respect to $\theta$, evaluated at $\theta$ (i.e., $\Lambda:=\partial g(\theta)/\partial\theta^{\top}$).
+For simplicity, we will assume that $\Lambda$ is known (i.e., moments are linear in $\theta$).
+This simplifies exposition but it is not necessary in the local asymptotic framework we present in Section 5.1, where $\Lambda$ can be replaced by the Jacobian of $g(\theta)$ evaluated at the observational estimates.
+For a given covariance matrix $\Sigma$ of the sample moment vector, let $\bar{g}_{\Sigma}$ denote the sample analogue with
+$\mathbb{E}[\bar{g}_{\Sigma}]=b,\qquad\mathbb{V}(\bar{g}_{\Sigma})=\Sigma,$
+where the vector $b$ captures the biases in the moments. The researcher may decide to access only some of the entries of $\bar{g}$ by selecting a subset of experiments due to budget constraints.
+Misspecification set.
+Let $\mathcal{I}\subseteq\{1,\dots,p\}$ index all potential experimental moments researchers may select (unbiased by design) and $\mathcal{I}^{c}$ the observational ones (potentially misspecified).
+We posit bounded misspecification
+$\mathcal{B}(B)\;=\;\Big\{b\in\mathbb{R}^{p}:\ b_{j}=0\ \text{for }j\in\mathcal{I},\ \ |b_{j}|\leq B\ \text{for }j\in\mathcal{I}^{c}\Big\}.$ (4)
+This restriction states that experimental moments have no bias, whereas observational moments have bias bounded by $B$, potentially unknown to the researcher.
+Estimator and design.
+Given the empirical moments $\bar{g}_{\Sigma}$, we consider (asymptotically) linear estimators of the form
+$\hat{\theta}(W,\Sigma)-\theta\;=\;\Gamma_{\Lambda}(W)\,\bar{g}_{\Sigma},\quad\Gamma_{\Lambda}(W)\equiv-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W$ (5)
+where $W$ and $\Sigma$ are chosen by the researcher from a feasible set.
+The structure of the estimator follows a standard GMM asymptotic expansion (e.g.
+Andrews et al., 2020).
+[^7] The weighting matrix $W$ may be singular.
+The expansion $\hat{\theta}-\theta\;\dot{=}\;-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\bar{g}_{\Sigma}$ holds whenever $\Lambda^{\top}W\Lambda$ is nonsingular, which can be forced through the set of constraints.
+Here, $W$ may have columns of (effectively) zeros, which excludes the corresponding moments from the estimator.
+This occurs when researchers faces a trade-off in which experiment to run, leading to different moments $\bar{g}$ observed in the experiment.
+The matrix $\Sigma$ depends on the experimental design (e.g., sample–size allocations across experimental moments).
+The constraints on $W$ and $\Sigma$ effectively limits the class of experiments, sample size allocations and potentially estimators the researcher may consider.
+Assumption 3 (Constraint set).
+Let $(W,\Sigma)\in\mathcal{D}^{\prime}$ for a set $\mathcal{D}^{\prime}$ so that $\Lambda^{\top}W\Lambda$ is invertible, $\Gamma_{\Lambda}(W)\Sigma\Gamma_{\Lambda}(W)^{\top}$ has uniformly bounded entries and is strictly positive definite.
+Assumption 3 requires that the variance of the selected (reweighted) moments is uniformly bounded and strictly positive definite (even when $W$ does not select some moments).
+This ensures that any subset of moments the researcher chooses to use is nondegenerate.
+Example 5 (GMM moment selection).
+Let $g(\theta)=(g^{\text{obs}}(\theta),\,g^{\text{exp}}_{2}(\theta),\,g^{\text{exp}}_{3}(\theta))^{\top}\in\mathbb{R}^{3}$, where the first component is observational and the other two are experimental.
+Suppose at most one experimental moment can be used.
+Selection is encoded by the GMM weight $W\in\mathbb{R}^{3\times 3}$.
+A simple example of class weighting matrices is
+$W^{(2)}=\begin{pmatrix}w_{\mathrm{obs}}&\rho&0\\ \rho&w_{2}&0\\ 0&0&0\end{pmatrix}\qquad W^{(3)}=\begin{pmatrix}w_{\mathrm{obs}}&0&\rho\\ 0&0&0\\ \rho&0&w_{3}\end{pmatrix}$,
+provided that $\Lambda^{\top}W\Lambda$ is invertible. ∎
+Target and MSE.
+Let the estimand be a linear functional $\tau(\theta)=\omega^{\top}\theta$ with $\omega\in\mathbb{R}^{d}$. Then
+$\mathrm{MSE}_{b}(W,\Sigma)$ $=\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\;+\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega.$ (6)
+The key distinction now is that the variance and bias regret depend on $\omega$ and $\Lambda$.
+Definition 3 (Moment based variance and bias regret).
+With a slight abuse of notation, let
+$\alpha(W,\Sigma)\;\equiv\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega,\qquad\beta(W)\;\equiv\;\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}^{\,2}.$
+and $\alpha^{\star},\beta^{\star}$ their corresponding minimizers over $(W,\Sigma)\in\mathcal{D}^{\prime}$.
+The moment-based variance and bias regret are defined as $\alpha/\alpha^{\star},\beta/\beta^{\star}$ respectively.
+The bias now depends on the interaction between the sensitivity $\omega$, the moment Jacobian $\Lambda$ and the identity of which moments are observational and therefore potentially biased.
+The weighting matrix $W$ determines how these moments influence the downstream estimator $\hat{\theta}$.
+As before, (and with a slight abuse of notation) we consider the adaptation regret of the following form:
+$\mathcal{R}(W,\Sigma)\;\equiv\;\sup_{B\geq 0}\ \frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\ \mathrm{MSE}_{b}(W^{\prime},\Sigma^{\prime})}.$ (7)
+The next theorem generalizes Theorem 1 to this more general framework.
+Theorem 2 (Regret characterization).
+Let Assumptions 2, 3 hold, $\mathcal{R}$ be as defined in Equation (7),with $\alpha,\beta,\alpha^{\star},\beta^{\star}$ as in Definition 3. Consider an estimator as in Equation (5). Then
+$\mathcal{R}(W,\Sigma)\;=\;\max\!\left\{\,\frac{\alpha(W,\Sigma)}{\alpha^{\star}}\,,\ \frac{\beta(W)}{\beta^{\star}}\,\right\}.$ (8)
+Proof.
+See Appendix B.2. ∎
+Theorem 2 shows how our framework generalizes to two related problems (i) choosing the weighting matrix $W$ for estimation and (ii) moment selection.
+The key insight is to show that the regret is quasi-convex under (approximate) linearity in the moments.
+5 Extensions
+5.1 Local asymptotics in the presence of non-linearities
+This subsection extends our framework for (i) nonlinear moment functions $g(\theta)$ used in GMM in Section 4 and (ii) a smooth, potentially nonlinear estimand $\tau(\theta)$.
+Linearization of the moments
+Assume the sample moments obey root–$n$ scaling, i.e., $\mathbb{V}(\bar{g}_{\Sigma})=O(n^{-1})$.
+Let $b=b_{n}$ be a sequence where retain the same misspecification structure as in Section 4: the moment vector has mean $b=b_{n}$ with $(b_{n})_{j}=0$ for $j\in\mathcal{I}$ (experiment-based moments) and $(b_{n})_{\mathcal{I}^{c}}$ potentially nonzero.
+Our key assumption is a local condition on the moments bias of the form
+$\|b_{n}\|^{2}\sqrt{n}\;\to\;0.$ (9)
+This condition builds on the local asymptotic frameworks in Andrews et al.
+(2020) and Armstrong and Kolesár (2021).
+In their context, $b_{n}||^{2}=1/n$ matching the rate of convergence of the variance.
+In our case, $b_{n}$ can grow faster, slower or at the same rate of the standard error, encompassing $b_{n}||^{2}=1/n$ as a special case.
+Under standard smoothness conditions, a GMM estimator with weighting matrix $W$ satisfies
+$\sqrt{n}\big(\hat{\theta}(W,\Sigma)-\theta\big)=-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)+o_{p}(1),$
+so on the original (unscaled) scale the mean shift is $(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,b_{n}$ and the remainder is $o_{p}(n^{-1/2})$.
+The condition $\|b_{n}\|^{2}\sqrt{n}\to 0$ ensures that higher order terms are negligible under sufficient smoothness conditions, justifying the use of the linear expansion (5) for our design analysis.
+Because $b_{n}\to 0$, the Jacobian $\Lambda$ can be consistently estimated by evaluating the derivative of $g(\theta)$ at a preliminary observational estimator $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$, assuming such estimates are available from an observational study (Andrews et al., 2020).
+Linearization of the estimand.
+Let $\tau:\mathbb{R}^{p}\to\mathbb{R}$ be twice continuously differentiable near $\theta$, with gradient $\omega(\theta):=\partial\tau(\theta)/\partial\theta$ and bounded Hessian.
+A Taylor expansion around $\theta$ yields
+$\tau(\hat{\theta})-\tau(\theta)=\omega(\theta)^{\top}(\hat{\theta}-\theta)\;+\;\tfrac{1}{2}\,(\hat{\theta}-\theta)^{\top}H_{\tau}(\bar{\theta})\,(\hat{\theta}-\theta),$
+for some $\bar{\theta}$ between $\hat{\theta}$ and $\theta$, and $H_{\tau}$ is the Hessian. It follows that we can write
+$\sqrt{n}\big(\tau(\hat{\theta})-\tau(\theta)\big)=\omega(\theta)^{\top}\,\Gamma_{\Lambda}(W)\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)\;+\;o_{p}(1),$
+because $\|\hat{\theta}-\theta\|=O_{p}(n^{-1/2}+\|b_{n}\|)$ and the local condition $\sqrt{n}\,\|b_{n}\|^{2}\to 0$ renders the quadratic remainder $o_{p}(1)$ after $\sqrt{n}$ scaling.
+Consequently, the regret expressions derived for the linear case continue to apply up to $o_{p}(1)$, with $\omega$ interpreted as the gradient $\omega(\theta)$ (or any consistent plug-in $\omega(\tilde{\theta}^{\mathrm{obs}})$ from a preliminary estimator $\tilde{\theta}^{\mathrm{obs}}\to_{p}\theta$).
+Remark 2 (Rescaling and change of variables linking Example 3).
+An equivalent formulation uses $\sqrt{n}$–rescaled coordinates (as in Example 3).
+Define the rescaled moment vector $\tilde{g}:=\sqrt{n}\,\bar{g}$ and the rescaled estimation error $\tilde{\theta}\equiv\sqrt{n}\,(\hat{\theta}-\theta)$.
+Then
+$\mathbb{V}(\tilde{g})=\Sigma=O(1),\qquad\mathbb{E}[\tilde{g}]=b_{n}\ \ \text{with}\ \ b_{n}\equiv\sqrt{n}\,b,$
+so $b_{n}$ is the asymptotic bias and $\Sigma$ the asymptotic variance.
+The local misspecification condition in (9) is equivalent to $\frac{\|b_{n}\|^{2}}{\sqrt{n}}\;\to\;0,$ which coincides with Example 3.
+∎
+5.2 Confidence interval length as a regret criterion
+In this subsection we consider an alternative loss based on the length of confidence intervals (CIs), focusing on Setting 1 for simplicity.
+For a candidate design $(\mathcal{E},\Sigma(\mathcal{E}))$, shrinkage weights $\gamma$, and a bias vector $b\in\mathbb{R}^{p}$, define the two–sided, $(1-\eta)$ lower and upper bounds for $\tau(\theta)$ by
+$\ell_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;-\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
+$u_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;+\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
+where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile.
+An audience, indexed by a worst–case bias radius $B\geq 0$, forms the worst–case CI
+$L_{B}(\mathcal{E},\Sigma,\gamma)~\equiv~\left[\inf_{\|b\|_{\infty}\leq B}\ell_{b}(\mathcal{E},\Sigma,\gamma),\;\sup_{\|b\|_{\infty}\leq B}u_{b}(\mathcal{E},\Sigma,\gamma)\right],$
+where we write $|L_{B}(\mathcal{E},\Sigma,\gamma)|$ for its total length. The proportional regret of $(\mathcal{E},\Sigma,\gamma)$ is
+$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{B\geq 0}\;\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{\inf_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}\!.$
+Theorem 3.
+Consider Setting 1 and let Assumption 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
 […]
 Proof.
-See Appendix B.2. ∎
-At the boundaries, if the relative variance $\alpha/\alpha^{\star}$ is always smaller than the relative bias $\beta/\beta^{\star}$, the optimal weight is $\gamma_{j}^{\star}=1$ (i.e., use only the experimental estimate).
-If instead the variance term always dominates the bias term, the optimal choice is the variance–minimizing weight $\gamma_{j}^{\star}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
-In most applications, the solution is interior: we choose $\gamma_{j}^{\star}$ strictly between the boundaries to equalize the two contributions, $\frac{\alpha(j,\gamma_{j}^{\star})}{\alpha^{\star}}=\frac{\beta(j,\gamma_{j}^{\star})}{\beta^{\star}},$ so that the variance ratio matches the worst–case squared bias ratio.
-The optimal solution depends not only on the features of the experiment design $j$ but also on the features of the other feasible design j-j ($v_{-j}^{2},\sigma_{-j}^{2}$ and $\omega_{-j}$) through $\alpha^{\star}$ and $\beta^{\star}$.
-This differs from choosing the optimal $\gamma$ for a single fixed design, since the choice of $\gamma$ will then interact with the choice of the class of design the researcher (and oracle) can choose.
-Illustration in Figure 2
-Figure 2 illustrates three cases for $j=2$.
-In the top row, we plot $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$.
+See Appendix B.1
+Theorem 3 shows that the objective function is the same as the MSE-objective function up to a monotonic (square-root) transformation which does not affect the minimizer.
+5.3 Vector valued estimands
+Next, we extend our framework to vector value estimands of the form $\tau(\theta)\in\mathbb{R}^{q}$ for $q\geq 1$.
+Setting 3 (Vector-valued estimands).
+Consider a vector-valued target estimand with entries $\tau^{\ell}(\theta)\equiv(\omega^{\ell})^{\top}\theta$ for $\ell=1,\dots q$, and $\omega^{\ell},\theta\in\mathbb{R}^{k}$.
+Given $(\mathcal{E},\Sigma)$, and observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}},\tilde{\theta}^{\mathrm{exp}}$, consider a class of linear plug-in estimators to allow the optimal shrinkage weights to depend on the specific entry of $\tau$.
+$\hat{\tau}^{\ell}_{\gamma}\equiv\tau^{\ell}(\hat{\theta}^{\ell}(\gamma)),\qquad\hat{\theta}^{\ell}_{j}(\gamma)=\begin{cases}\gamma_{j}^{\ell}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j}^{\ell})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
+where $\gamma=(\gamma^{\ell}_{j})_{\ell\in\{1,\cdots,q\},j\in\mathcal{E}}\in\mathbb{R}^{q\times|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$.
+Let $\hat{\theta}(\gamma)\equiv(\hat{\theta}^{1}(\gamma),\dots,\hat{\theta}^{q}(\gamma))\in\mathbb{R}^{p\times q}$ and $\hat{\tau}_{\gamma}\equiv(\hat{\tau}^{1}_{\gamma}(\hat{\theta}^{1}(\gamma)),\dots,\hat{\tau}^{q}_{\gamma}(\hat{\theta}^{q}(\gamma)))^{\top}\in\mathbb{R}^{q}$.
+Setting 3 is a flexible generalization of Setting 1, allowing each entry of $\tau$ to have its own vector of shrinkage weights.
+Assumption 4 (First-order estimation error).
+For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume for all $\ell=1,\dots,q$
+$\tau^{\ell}(\theta)-\tau^{\ell}\!\big(\hat{\theta}^{\ell}(\gamma)\big)=(\omega^{\ell})^{\top}\,\big(\theta-\hat{\theta}^{\ell}(\gamma)\big),$ (10)
+for a known weighting vector $\omega^{\ell}\in\mathbb{R}^{p}$, with $|\omega^{\ell}_{i}|\in(0,\infty)$, for all $i,\ell$.
+We begin by generalizing the mean squared error to higher-dimensional parameters.
+For a given design $(\mathcal{E},\Sigma)$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
+$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[\|\widehat{\tau}_{\gamma}-\tau\|_{2}^{2}\right],$ (11)
+where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$, and $\|\cdot\|_{2}$ denotes the $L^{2}$ Euclidean norm.
+We introduce more compact notation before proceeding to the main result. For a given $(\mathcal{E},\gamma)$ define
+$\tilde{\omega}^{\ell}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega^{\ell}_{j},&j\notin\mathcal{E},\\ \omega^{\ell}_{j}(1-\gamma^{\ell}_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}^{\ell}_{\text{exp},j}(\mathcal{E},\gamma)=\begin{cases}0,&j\notin\mathcal{E},\\ \omega^{\ell}_{j}\gamma^{\ell}_{j},&j\in\mathcal{E},\end{cases}$
+and let $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}^{\ell}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}^{\ell}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$ We can write $\hat{\tau}^{\ell}_{\gamma}=\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}\,\hat{\theta}$, where $\hat{\theta}=(\tilde{\theta}^{\text{obs}\top},\tilde{\theta}^{\text{exp}\top})^{\top}\in\mathbb{R}^{p+|\mathcal{E}|}$.
+Stacking across the $\ell$ index gives the estimator for the full vector $\tau$: $\hat{\tau}_{\gamma}=\tilde{\omega}(\mathcal{E},\gamma)\,\hat{\theta}$, where the $\ell^{th}$ row of $\tilde{\omega}(\mathcal{E},\gamma)\in\mathbb{R}^{q\times(p+|\mathcal{E}|)}$ is $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}$.
+The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
+$\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ $\equiv\text{Trace}\left(\tilde{\omega}(\mathcal{E},\gamma)\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\right),$ (12)
+$\check{\beta}(\mathcal{E},\gamma)$ $\equiv\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}\bar{v}(\gamma)-(\omega^{\ell}_{\,\mathcal{E}})^{\top}\bar{v}(\gamma)_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(\bar{v}(\gamma)_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
+where $\omega^{\ell}_{\mathcal{E}}$ is the subvector of $\omega^{\ell}$ corresponding to the entries indexed by the experiment $\mathcal{E}$, and
+$\bar{v}(\gamma)\in\operatorname*{arg\,max}_{\{v\in\mathbb{R}^{p}:|v_{j}|\leq 1\forall j\}}\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\,\mathcal{E}})^{\top}v_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(v_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
+where the dependence of $\bar{v}(\gamma)$ on $\mathcal{E}$ is omitted for conciseness and $\circ$ denotes the component-wise multiplication. The smallest worst-case variance and bias are defined as
+$\check{\alpha}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\beta(\mathcal{E},\gamma).$
+As before, we are interesting in minimizing
+$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\equiv\sup_{B\geq 0}\frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}.$
+The result below shows how our results directly extend to this case.
+Theorem 4.
+Consider Setting 3 and let Assumption 4 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
+$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\check{\alpha}(\mathcal{E},\Sigma,\gamma)}{\check{\alpha}^{\star}},\ \ \frac{\check{\beta}(\mathcal{E},\gamma)}{\check{\beta}^{\star}}\right\}.$
+Proof.
+See Appendix B.2
+6 Empirical applications
+In this section we provide two empirical applications, one for measuring GE effect and the second one for externally valid site selection.
+6.1 An implementation guide for measuring GE effects
+Combining experimental or pre-program observational data with structural models is increasingly used for program evaluation (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Allcott et al., 2025).
+[^8] Across all AEA articles in 2015-2025, about $10\%$ of the experimental papers combine experimental results with structural models (i.e., one-third of the papers that use observational in combination with experimental estimates).
+Because large-scale experimentation are often infeasible due to cost constraints (Muralidharan and Niehaus, 2017) these models are meant to extrapolate effects from small-scale experiments to learn general-equilibrium (GE) effects.
+This extrapolation typically relies on other experiments or observational data.
+We present a (step-by-step) illustration below when these models are calibrated using previous studies conducted in a different country.
+Question
+Consider a researcher evaluating a cash transfers for sending children to school in rural Kenya where such a program is not in place.
+Due to budget constraints, the researcher can only randomize at a small scale (partial equilibrium), but ultimately wishes to predict GE effects.
+Researchers have access to preliminary estimates from the Mexican PROGRESA experiment (Todd and Wolpin, 2006), which we may be concerned, lack external validity.
+6.1.1 Step 1: model description
+The first step for a researcher is to describe the modeling choices. We present a stylized model of school choice from Bonhomme and Weidner (2022) and Todd and Wolpin (2006).
+Individual choices
+Let $S\in\{0,1\}$ denote school attendance, $C$ consumption, $Y$ (pre-transfer) household income, $W$ the child’s potential wage, and $t$ the stipend when enrolled.
+Abstracting from covariates (we will introduce covariates in the estimation), utility is
+$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{3}-\xi_{1}-\xi_{2})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (13)
+with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS.
+The parametrization $(\xi_{3}-\xi_{1}-\xi_{2})$ is without loss and simplifies expressions below.
+Enrollment satisfies $S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$ Letting $Z(Y,W,t)\equiv\xi_{1}W-\xi_{2}Y-\xi_{3}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$, with $\Phi(\cdot)$ denoting the standard Gaussian CDF.
+General equilibrium effects
+We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya.
+GE feedback is allowed through income and wages: for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, let
+$Y_{i}(t)=y(t)+\varepsilon_{Yi},\qquad W_{i}(t)=w(t)+\varepsilon_{Wi}.$
+Let $y_{0}\equiv\partial_{t}y(t)|_{t=0}$ and $w_{0}\equiv\partial_{t}w(t)|_{t=0}$.
+Define $\phi_{0}\equiv\mathbb{E}\!\left[\phi\!\big(-Z\big(Y(0),W(0),0\big)\big)\right]$.
+Our estimand of interest is the marginal effect of an increase in a small transfer $t$ to all eligible households
+$\left.\frac{\partial\,\mathbb{E}\!\left[\Pr(S{=}1\mid Y(t),W(t),t)\right]}{\partial t}\right|_{t=0}=\;\phi_{0}\cdot\big(\,\xi_{3}+\xi_{2}y_{0}-\xi_{1}w_{0}\,\big),$ (14)
+which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{2}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$.
+Under simple market clearing for labor supply and demand, we can show [^9] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$.
+Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working.
+From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{3}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{1}$ at $t=0$.
+Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{3}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{1}H$.
+Dividing the equation by $H$ we obtain the desired result.
+$w_{0}=\;\frac{\phi_{0}\,\xi_{3}}{\phi_{0}\,\xi_{1}-d},$ (15)
+where $d$ denotes the slope of the demand curve divided by total hours of work.
+6.1.2 Step 2: parametrization and transparency on biases
+Given the model, the following step for researchers is to define the parameters of interest. Here, we define the parameters of interest as
+$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{3},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{1},$ (16)
+with corresponding estimand
+$\tau(\theta)\;=\;\underbrace{\theta_{2}}_{\text{Direct effect}}\;+\;\underbrace{y_{0}}_{\text{Income multiplier}}\,\underbrace{\theta_{1}}_{\text{Income effect}}\;+\;\underbrace{w_{0}(\theta)\,\theta_{3}}_{\text{Wage effect}},\qquad w_{0}(\theta)\;=\;\frac{\theta_{2}}{-\theta_{3}-d},$ (17)
+where the explicit form of $w_{0}(\theta)$ follows from (15).
+We treat $d$ as known for simplicity, using meta-analyses on demand elasiticities, and $y_{0}=1.5$ because computed experimentally in Kenya by Egger et al.
+(2022).
+[^10] Egger et al.
+(2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}=2.5$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$.
+Specifically, $d=0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained from meta-studies in Espey and Thilmany (2000).
+[^11] One could also calibrate $S_{0},W_{0}$ to pre-experimental data in Kenya directly when available.
+In this example, the primary concern is bias in the observational estimator $\tilde{\theta}^{\mathrm{obs}}$ constructed as described below, abstracting from potential biases in $y_{0},d$ which, for simplicity, we assume to be second-order.
+When these biases are considered first-order, researchers should let $y_{0},d$ be additional parameters in $\theta$.
+This difference illustrates a feature of our framework: the chosen parameterization ensures transparency about which biases’ concerns drive the design; these considerations can be collected in a pre-analysis plan.
+6.1.3 Step 3: observational study estimates
+Using experimental data from Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ via probit with standard controls (age, distance to school, eligibility, year, highest grade attained).
+Estimation is conducted separately by gender, and we will be focusing here on the effects on female students.
+We pool treated and control observations to estimate the schooling effect $\xi_{4}$, controlling for the individual-level subsidy.
+In turn, we obtain observational estimates $\big(\tilde{\theta}_{1}^{\mathrm{obs}},\tilde{\theta}_{2}^{\mathrm{obs}},\tilde{\theta}_{3}^{\mathrm{obs}}\big)$ that map to the model parametrization.
+Standard errors are constructed using the Delta method with clustering at the village level.
+Following Section 5.1, we then construct $\omega=\frac{\partial\tau}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}$.
+Point estimates, estimated sensitivity $\omega$, and standard errors are reported below.
+Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sigma$
+[…]
+[…]
+[…]
+Observational estimates for female students, corresponding sensitivity parameter $\omega$, and standard errors.
+$\Sigma^{\mathrm{obs}}\;=\;\begin{bmatrix}4.31&-11.31&5.57\\ -11.31&973.11&-126.16\\ 5.57&-126.16&1038.56\end{bmatrix}\times e^{-9}$
+Variance–covariance matrix of observational estimates for female students.
+We observe heterogeneity in both the sensitivity weights $\omega$ the precision of the observational estimates.
+6.1.4 Step 4: describing the candidate experiments
+The following step is to list the feasible experiments, subject to budget constraints.
+In this example, we consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, and examine three possible experiments for a given sample size $n$:
+  • $j=\{1\}$: Unconditional transfer (income shock).
+The researcher randomizes a small income shock to a small fraction of households (implying no general-equilibrium effects).
+Under a first-order (Taylor) approximation, the experiment identifies
+    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,0)]}{\partial Y}\right|_{t=0}=\theta_{1},$
+    with precision $v_{1}^{2}/n$. The researcher optimizes over $\gamma_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
+  • $j=\{2\}$: Conditional cash transfer (stipend).
+The researcher randomizes a small stipend $t$, conditional on attending school, to a small fraction of households (with prices held fixed).
+Under a first-order approximation, this identifies
+    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,t)]}{\partial t}\right|_{t=0}=\theta_{2},$
+    with precision $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
+  • $j=\{1,2\}$: Two-arm design.
+The researcher runs (j=1j{=}1) and (j=2j{=}2) on independent samples, identifying $(\theta_{1},\theta_{2})$ with precisions $(v_{1}^{2}/n_{1},\ v_{2}^{2}/n_{2})$ where $n_{1}+n_{2}=n$.
+They optimize over both $\gamma_{1}$ and $\gamma_{2}$, while $\theta_{3}$ is calibrated to the PROGRESA study.
+We consider two scenarios.
+In the first, the researcher (and oracle) can choose one or both experiments.
+In the second, they can choose only one of the two (either $j=\{1\}$ or $j=\{2\}$).
+Regret is defined accordingly, conditional on the set of feasible designs.
+For simplicity, we calibrate $v_{1}^{2}$ as $\sigma_{1}^{2}\times n^{\mathrm{obs}}$, where $\sigma_{1}^{2}$ is the variance of $\tilde{\theta}_{1}^{\mathrm{obs}}$ and $n^{\mathrm{obs}}$ is the observational-sample size ($n^{\mathrm{obs}}=1089$ for the female sample using data from eligible students in Todd and Wolpin (2006)).
+We calibrate $v_{2}^{2}$ analogously as $\sigma_{2}^{2}\times n^{\mathrm{obs}}$.
+In practice, researchers may use pilot studies to calibrate the experimental variances when available.
+Experiment type Identified parameter Choice Variables Calibrated standard error Description
+$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $6.56e^{-5}/\sqrt{n}$ Unconditional transfer (income shock, UCT). Randomize an income shock to a small fraction of households.
+$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $9.86e^{-4}/\sqrt{n}$ Conditional cash transfer (education shock, CCT). Randomize a conditional transfer to a small fraction of households.
+$j=\{1,2\}$ $(\theta_{1},\,\theta_{2})$ $\gamma_{1},\ \gamma_{2},\ n_{1}$ $6.56e^{-5}/\sqrt{n_{1}}$ (arm 1), $9.86e^{-4}/\sqrt{n-n_{1}}$ (arm 2) Two-arm design.
+Run (j=1j{=}1) and (j=2j{=}2) on independent samples with sample sizes respectively $n_{1},n-n_{1}$.
+Table 1: Design options, identified parameters, choice variables, and per-unit variances.
+6.1.5 Step 5: experimental design
+Figure 4: Regret–optimal sample allocation across treatment arms when both the unconditional cash transfer (UCT; income shock) and the conditional cash transfer (CCT; stipend) arms are feasible.
+The optimal solution is interior: the vast majority of participants (over 90%) are assigned to CCT, with a small but nonzero fraction assigned to UCT to hedge misspecification.
+This pattern reflects the higher payoff to learning about CCT in this application relative to UCT.
+Figure 5: Optimal shrinkage weights and design choice as functions of the total sample size $n$.
+Left panel: When both arms are run, the weight on the UCT experimental estimator is essentially one across $n$, while the weight on the CCT experimental estimator, $\gamma^{\star}_{\text{CCT}}$, rises with precision (from about 0.5 at small $n$ to roughly 0.9 by $n=1000$), reflecting the increasing value of experimental evidence.
+Right panel: When restricted to a single arm, the optimal choice switches from UCT at small $n$ (variance dominates under CCT) to CCT once $n$ crosses a threshold (around $n\approx 500$), where bias considerations dominate and favor CCT.
+Notes: $\gamma_{j}^{\star}$ is the optimal weight on the experimental (vs.
+observational) estimator for parameter $j$; we set $\gamma_{j}^{\star}=0$ to indicate that arm $j$ is not chosen.
+Figure 6: Worst–case regret by design.
+Left panel: Allowing both arms with an interior allocation yields regret that approaches one as $n$ increases, indicating near–oracle performance under our normalization.
+In contrast, single–arm designs exhibit nonvanishing regret because bias remains first–order even as variance shrinks.
+Right panel: Focusing on the single–arm problem, UCT dominates only at small $n$; beyond the same threshold ($\approx 500$), CCT delivers uniformly lower regret as variance becomes second–order relative to bias.
+Notes: Regret is the worst–case objective normalized so that a value of one corresponds to the oracle benchmark; UCT = unconditional cash transfer, CCT = conditional cash transfer.
+The final step is to design the experiment and report it in a pre-analysis plan.
+Choice of the experiment when only one treatment arm is available
+Figure 5 (right panel) displays the selected arm as a function of $n$ (together with the corresponding $\gamma_{j}^{\star}$).
+We encode “not chosen” as $\gamma_{j}^{\star}=0$.
+For small samples ($n\lesssim 500$), the optimal choice is UCT: the experimental CCT estimator would be too noisy, so the resulting variance under CCT dominates, while the bias advantage of CCT over UCT is limited because $\omega_{1}$ and $\omega_{2}$ are of similar magnitude.
+Once $n$ crosses $n\approx 500$, CCT becomes dominant: by combining observational and experimental estimates for CCT, we can achieve comparable variance, and the worst-case bias is lower than under UCT.
+Choice of the estimator
+In addition to sample allocation, our method optimizes the shrinkage weights of the parameters of interest.
+Figure 5 (left panel) shows that when both arms can be run, we place essentially full weight on the experimental estimator for UCT ($\gamma^{\star}_{\text{UCT}}\approx 1$ across $n$), while for CCT we gradually shift toward the experimental estimator as precision improves: $\gamma^{\star}_{\text{CCT}}$ rises from about $0.5$ at small $n$ to approximately $0.9$ by $n=1000$.
+Intuitively, higher $n$ lowers experimental variance, making it optimal to rely more heavily on experimental evidence for the CCT parameter.
+Sample size allocation when both arms are feasible
+Figure 4 reports the sample allocation implied by the regret-optimal design when both treatment arms are feasible.
+The optimal allocation is interior: most participants (over roughly $90\%$) are assigned to the CCT arm (education), with a small but nonzero fraction assigned to the UCT arm (income shock).
+Two forces rationalize this pattern.
+First, the CCT parameter is more misspecification-sensitive ($\omega_{2}\approx 1.98$ vs.
+$\omega_{1}\approx 1.5$), so generating experimental evidence on CCT has a larger payoff in bias reduction.
+Second, the income effect $\theta_{1}$’s variance is about an order of magnitude smaller than the stipend effect $\theta_{2}$’s variance in our PROGRESA calibration, lowering the marginal returns to learn about UCT relative to CCT.
+Hence, while the designer keeps some allocation on UCT, most of the sample is placed on CCT; our framework makes this bias–variance trade-off explicit.
+Regret comparisons
+Figure 6 reports worst-case regret for each design choice.
+The left panel shows that the regret of our method (green line, running both arms with an interior allocation) declines toward one as $n$ increases, reflecting near-oracle performance under our normalization.
+In contrast, single-arm designs exhibit nonvanishing regret because bias remains first order even as variance shrinks.
+The right panel focuses on the single-arm problem: both the oracle and the researcher can only choose a single arm.
+In this case, UCT is preferable only for small $n$; beyond the same threshold ($\approx 500$), CCT yields uniformly lower regret as variance becomes second order relative to bias.
+The regret vanishes relative to the oracle that can only choose a single arm.
+Implications
+When only one arm is feasible, CCT is preferable only once the total sample size is sufficiently large; otherwise UCT is preferred.
+This pattern reflects the interaction between bias and variance: for small samples, CCT’s bias advantage is not large enough to compensate for its higher variance relative to UCT, whereas at larger samples the bias reduction becomes dominant.
+Existing designs based on variance considerations would fail to provide solutions that balance bias and variance comparisons.
+When feasible, it is instead optimal to run two arms with highly unbalanced sample, reflecting the relative contribution of the variance for each treatment arm.
+Following these steps, researchers can similarly report their design in a pre-analysis plan.
+6.2 Measuring the performance of our method for site selection
+Next, we show how observational evidence can inform whom to recruit into an experiment.
+We consider the setting in Banerjee et al.
+(2024), who study how the expansion of microfinance affects village social networks in Karnataka, India.
+We focus on the first outcome reported by Banerjee et al.
+(2024) corresponding to the density of the network, i.e., the percentage of connections a random household in a village has relative to the village size.
+Background
+In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance.
+Because program rollout was not randomized across villages, they estimate effects using Difference-in-Differences (DiD).
+While DiD is informative, the lack of randomized assignment may bias estimates in the presence of selection (e.g., Ghanem et al., 2022).
+Motivated by these limitations, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods.
+We only use observational variation for choosing the experimental design and estimator.
+We then use experimental variation generated by Banerjee et al.
+(2024) to validate our procedure.
+Research question
+In practice, implementation costs often depend on how geographically dispersed the study sites are due to coordination and survey costs.
+Using preliminary observational estimates from Banerjee et al.
+(2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled?
+Figure 7: Observational villages and area-level DiD effects in Karnataka.
+The background heat map partitions the state into the four contiguous areas used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value.
+Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey).
+Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
+Observational study
+We partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period.
+Figure 7 maps these four areas and reports the corresponding DiD point estimates.
+Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas.
+Table 2 summarizes, for each area, samples sizes and variances, clustered at the village level.
+These area-specific DiD estimates and their variances serve as our primary observational inputs for experimental design.
+[^12] We use DiD estimates to mimic the estimator used by the authors.
+An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity.
+Table 2: Area-level observational inputs for Karnataka villages.
+The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village).
+For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$.
+Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$
+11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3
+22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140
+33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187
+44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130
+Experimental design
+We consider a family of designs that (i) select $E\in\{1,\ldots,4\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$).
+Here, $E=1$ corresponds to recruiting from a single area, while $E=4$ recruits from all four areas.
+We examine a grid of sample sizes that varies the number of participants from $10$ up to $104$ ($52$ treated units), with the latter corresponding to the size of the experiment in Banerjee et al.
+(2024).
+For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 2.
+We assume that the variance of a single treated–control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
+Figure 8: Optimal area selection and village allocation for a large experiment ($n_{1}=52$) under constraints on the number of areas $E$.
+Each panel corresponds to a value of $E$; shading indicates the total number of recruited villages in each area.
+With $E=1$ the design concentrates in the area with the noisiest observational estimate; as $E$ increases, allocation spreads across areas, reflecting a trade-off between observational uncertainty and experimental variance.
+Figure 9: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for each constraint on the number of eligible areas $E$.
+Lines show the number of villages assigned to each area.
+Figure 10: Relative MSE across designs under an hypothetical scenario where the bias equal the difference between the experimental and observational ATE estimate from Banerjee et al.
+(2024) as $n_{1}$ varies (from ten to fifty-two villages).
+The figure reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$ for (i) the proposed design (area targeting and optimal $\gamma^{\star}$), (ii) a random benchmark (uniform area selection and equal allocation), where $\mathrm{MSE}^{*}$ denotes the MSE of the oracle that knows the bias, as $n_{1}$ varies and for each $E$.
+The proposed design closely tracks the oracle across $n_{1}$ and $E$, while the random benchmark remains substantially above, especially when $E$ is small.
+Experimental design
+We begin by studying which areas are selected and how sample is allocated when the total experimental sample is large ($n_{1}=52$) and the number of admissible areas may or may not be constrained.
+Figure 8 visualizes the resulting allocation across Karnataka’s four observational areas for $E\in\{1,2,3,4\}$.
+With $E=1$, the algorithm concentrates recruitment in the Tumakuru area—the location with the highest uncertainty in the observational estimates.
+As $E$ relaxes to $2$ and $3$, the total sample splits across them in roughly (but not exactly) equal shares, reflecting a trade-off between (i) exploiting heterogeneity in the observational variance and (ii) keeping experimental variance low.
+When $E=4$, all areas are eligible and the allocation smooths further across space.
+In the rest of our discussion, we discuss properties when not all areas may be selected ($E<4$).
+Figure 9 tracks how area-level allocations evolve with the total number of villages.
+When the experiment is small (e.g., $10$ villages) and $E=1$, the algorithm favors an area with relatively low experimental variance, even if its observational estimate is less uncertain, because, at small $n$, experimental noise is first-order.
+As the total sample increases, the weight shifts toward areas with more uncertain observational estimates.
+Consistently, the area with the largest experimental variance (Area 2) is typically excluded.
+Appendix Figure 13 reports the optimal shrinkage $\gamma_{j}^{\star}$ by area as total villages vary.
+The solution is interior for all $E<4$: $\gamma_{j}^{\star}$ rises with sample size, approaching one from below as experimental noise diminishes.
+For example, at a moderate sample (around 15 villages), $\gamma^{\star}$ is about $0.65$ when $E\in\{2,3\}$; with $E=1$, concentrating the sample within one area lowers experimental variance and therefore yields a larger $\gamma^{\star}$.
+Main MSE comparisons
+We compare three strategies: (i) our chosen design which chooses the area, sample size and $\gamma^{\star}$ as described in Section 3, (ii) a standard benchmark that selects areas uniformly at random, allocates villages evenly across the chosen areas, and sets $\gamma_{j}=1$ for experimental estimates, and (iii) an oracle that knows the bias vector $b$ and optimizes both design (areas and allocation) and $\gamma$.
+Because $b$ is unknown in practice, for this exercise, we calibrate it using the difference between the Hyderabad RCT ATE obtained from the follow-up experiment of Banerjee et al.
+(2024) and the Karnataka DiD average effect across sites, treating $b$ as common across areas.
+Figure 10 plots the ratio $\text{MSE}/\text{MSE}^{\star}$, where $\text{MSE}^{\star}$ is the oracle’s MSE.
+As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone.
+Across $E\in\{1,2,3\}$, the random benchmark remains well above the oracle—by roughly 3–5 percentage points (pp) for $E=1$, about 10 pp for $E=2$, and around 20 pp for E=3E=3—whereas our design tracks the oracle closely (within a few percentage points) even without knowledge of $b$.
+The gap is the largest when a larger number of areas $E$ can be included, because our procedure can better emulate the oracle by selecting (and differentially weighting) favorable areas.
+This result illustrates the merit of the method that improves MSE up to $20$ percentage points compared to a standard benchmark and leads to a near-oracle solution.
+7 Implications for practice
+This paper studies experimental design in the presence of observational evidence.
+A key challenge is that the bias of the observational estimators is unknown in practice.
+We adopt a minimax proportional regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias.
+This reveals a fundamental trade-off between precision and robustness.
+The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap).
+We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
+In practice, the workflow is:
+  • Define the estimand(s) of interest.
+Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$.
+For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations.
+Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify.
+When such parametrization is not readily available, Section 4 shows how our procedure applies to $\tau$ identified through arbitrary moment restrictions.
+  • Assemble informative observational evidence.
+Collect observational estimates $\tilde{\theta}^{\rm obs}$ and their covariance $\tilde{\Sigma}^{\rm obs}$.
+These serve as informative (but potentially biased) baselines.
+Such evidence may come from observational designs, structural estimates with pre-experimental data, or prior experiments conducted in different contexts.
+  • Compute the sensitivity parameters.
+Using the observational baseline, compute the sensitivity weights $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\big|_{\tilde{\theta}^{\rm obs}}$, which quantify how bias in each coordinate of $\theta$ propagates to $\tau(\theta)$.
+Large $|\omega_{j}|$ indicates greater payoff to learning the $j$th component.
+  • Specify feasibility constraints and calibrate experimental variance.
+Enumerate the admissible design set, and the budget on sample size allocation.
+Calibrate per-unit experimental variances using pilot studies or historical data to form the set of feasible designs $\mathcal{D}$.
+  • Run the method and decide.
+Optimize over the design and corresponding estimators.
+The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$).
+The applicability of our framework spans a wide range of settings in economics and beyond.
+Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024).
+In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis (Allcott et al., 2025).
+Beyond economics, medical applications include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025).
+Several open questions remain for future research.
+These include settings where researchers have well-specified priors about bias such as from Bayesian models (Gechter et al., 2024), or other modeling choices (Olea et al., 2024).
+This may be potentially incorporated through additional moment conditions as in Section 4.
+Additional future extensions include sequential or adaptive experimental choices or settings where the target estimand is only be partially identified (e.g.
+Cesa-Bianchi et al., 2025).
+Appendix A Proofs of main results
+A.1 Proof of Theorem 1 and Theorem 2
+Theorem 1 is a special case of Theorem 2 proved below.
+Step 1 (worst-case MSE). By Hölder’s inequality,
+$\sup_{b\in\mathcal{B}(B)}\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\leq\Big(B\,\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}\Big)^{2}=B^{2}\,\beta(W).$
+In addition, this inequality is attained by $b_{j}=B\,\mathrm{sign}\!\big((\Gamma_{\Lambda}(W)^{\top}\omega)_{j}\big)$ for $j\in\mathcal{I}^{c}$ (and $b_{j}=0$ for $j\in\mathcal{I}$).
+Hence $\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)=\alpha(W,\Sigma)+B^{2}\beta(W).$
+Step 2 (oracle envelope and basic properties).
+Let $\delta(t):=\inf_{W,\Sigma}\{\alpha(W,\Sigma)+t\beta(W)\}$.
+Recall that $\beta(W)$ and $\alpha(W,\Sigma)$ are bounded from above and $\alpha(W,\Sigma)>0$ by Assumption 3.
+Being the pointwise infimum of affine functions of $t$, $\delta$ is concave, nondecreasing, and finite for $t\geq 0$.
+Moreover,
+$\delta(0)=\alpha^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=\beta^{\star}.$
+Step 3 (quasi-convexity and boundary maximization). Fix $(\alpha,\beta)$. Consider
+$\phi(t):=\frac{\alpha+t\beta}{\delta(t)},\qquad t\geq 0.$
+Because the numerator is affine and the denominator is positive and concave in $t$, $\phi$ is quasi-convex on $0,\infty)$.
+Hence, on any compact interval $[0,T]$, $\max_{t\in[0,T]}\phi(t)$ is attained at the boundary $\{0,T\}$.
+Letting $T\to\infty$ and using the limits in Step 2,
+$\sup_{t\geq 0}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \lim_{t\to\infty}\frac{\alpha+t\beta}{\delta(t)}\right\}=\max\!\left\{\frac{\alpha}{\alpha^{\star}},\ \frac{\beta}{\beta^{\star}}\right\},$
+with the stated conventions $0/0=0$.
+Appendix B Proofs of the extensions
+B.1 Proof of Theorem 3
+The proof of Theorem 3 follows similarly the one of Theorem 2 with minor modifications. It is easy to show that
+$|L_{B}(\mathcal{E},\Sigma,\gamma)|=2\{z_{1-\eta/2}\,\alpha(\mathcal{E},\Sigma,\gamma)+B\,\sqrt{\beta(\mathcal{E},\gamma)}\}.$
+We can now follow verbatim the proof of Theorem 2, with $\delta(t):=\inf_{\mathcal{E},\Sigma,\gamma}2\{z_{1-\eta/2}\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)}+t\sqrt{\beta(\mathcal{E},\gamma)}\}$ with limits
+$\delta(0)=2z_{1-\eta/2}\sqrt{\alpha^{\star}},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=2\sqrt{\beta^{\star}}.$
+Taking the ratios and limits as in Step 3 of the proof of Theorem 2 the proof completes.
+B.2 Proof of Theorem 4
+It is easy to show that
+$\check{R}_{B}\left(\mathcal{E},\Sigma,\gamma\right)\equiv\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\beta}(\mathcal{E},\gamma)B^{2}+\check{\alpha}(\mathcal{E},\Sigma,\gamma),$
+with $\check{\beta},\check{\alpha}$ as defined in Equation (12), where $\check{\beta}\geq 0$ by construction.
+In addition $\check{\alpha}>0$ and bounded away from zero by Assumption 1.
+We can therefore follow verbatim the proof of Theorem 1.
+Appendix C Examples
+C.1 Two parameters examples in Equation (3)
+By Theorem 1, for fixed $(\mathcal{E},\Sigma)$ the adaptation regret as a function of the shrinkage vector equals
+$\mathcal{R}(\gamma)\;=\;\max\!\Big\{\alpha(\mathcal{E},\Sigma,\gamma)/\alpha^{\star},\;\beta(\mathcal{E},\gamma)/\beta^{\star}\Big\},$
+with $\alpha^{\star},\beta^{\star}$ constants.
+Step 1. With independence and two parameters,
+$\alpha(\mathcal{E},\Sigma,\gamma)=C+\omega_{j}^{2}\!\Big(\gamma_{j}^{2}v_{j}^{2}+(1-\gamma_{j})^{2}\sigma_{j}^{2}\Big),$
+where $C$ does not depend on $\gamma_{j}$. Hence $\alpha(j,\gamma_{j})\equiv\alpha(\mathcal{E},\Sigma,\gamma)$ is a strictly convex quadratic in $\gamma_{j}$ with unique minimizer at
+$\gamma_{j}^{\mathrm{var}}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}.$
+Moreover, $\alpha(j,\gamma_{j})$ is strictly increasing on $\big(\gamma_{j}^{\mathrm{var}},1\big)$ and strictly decreasing on $\big[0,\gamma_{j}^{\mathrm{var}}\big)$.
+Write the worst–case bias component holding the other coordinate fixed as
+$\beta(j,\gamma_{j})\;=\;\Big(A_{j}+|\omega_{j}|\cdot|1-\gamma_{j}|\Big)^{2},$
+where $A_{j}\geq 0$ collects all terms not involving $\gamma_{j}$ (including the contribution from the other coordinate).
+On $[0,1]$, $|1-\gamma_{j}|=1-\gamma_{j}$, so $\beta(j,\gamma_{j})$ is strictly decreasing in $\gamma_{j}$ and convex with minimum at $\gamma_{j}=1$.
+Therefore, any minimizer satisfies
+$\gamma_{j}^{\star}\in\big[\gamma_{j}^{\mathrm{var}},\,1\big].$
+Step 2.
+On $\big(\gamma_{j}^{\mathrm{var}},1\big)$ the map $\gamma_{j}\mapsto\alpha(j,\gamma_{j})/\alpha^{\star}$ is strictly increasing, while $\gamma_{j}\mapsto\beta(j,\gamma_{j})/\beta^{\star}$ is strictly decreasing.
+Hence the function
+$f(\gamma_{j})\equiv\max\!\Big\{\alpha(j,\gamma_{j})/\alpha^{\star},\ \beta(j,\gamma_{j})/\beta^{\star}\Big\},\qquad\gamma_{j}\in\big[\gamma_{j}^{\mathrm{var}},1\big],$
+is minimized either (i) at a boundary point, or (ii) at the unique interior point where the two arguments are equal (by strict monotonicity, at most one intersection exists). Therefore,
+  • If $\alpha(j,\gamma_{j})/\alpha^{\star}<\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\beta(j,\gamma_{j})/\beta^{\star}$ on that interval.
+Since this term is strictly decreasing, the minimizer is the right boundary $\gamma_{j}^{\star}=1$.
+  • If $\alpha(j,\gamma_{j})/\alpha^{\star}>\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\alpha(j,\gamma_{j})/\alpha^{\star}$ on that interval.
+Since this term is strictly increasing there, the minimizer is the left boundary $\gamma_{j}^{\star}=\gamma_{j}^{\mathrm{var}}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
+  • Otherwise, by the intermediate value theorem and strict monotonicity of the two curves, there exists a unique $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$ such that $\frac{\alpha(j,\gamma_{j})}{\alpha^{\star}}\;=\;\frac{\beta(j,\gamma_{j})}{\beta^{\star}}.$
+C.2 Numerical example
+Here, we present a numerical example in Setting 2.
+Choice of $\gamma^{\star}$
+Figure 2 (top row) reports $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$.
 In the bottom row, we plot the corresponding $\gamma_{2}^{\star}$.
@@
 This is a typical (and desired) behavior of shrinkage estimators.
-Table 1: Regimes for $\gamma_{2}^{\star}$ as a function of $\omega_{2}$ for example in Figure 2 with $\sigma_{1}^{2}=\sigma_{2}^{2}=1,v_{1}=1,v_{2}=0.5,\omega_{1}=1$.
-Each row summarizes each observation 1-4 for $\gamma_{2}^{\star}$.
+Table 3: Example: Regimes for $\gamma_{2}^{\star}$ as a function of $\omega_{2}$ for example in Appendix Figure 2 with $\sigma_{1}^{2}=\sigma_{2}^{2}=1,v_{1}=1,v_{2}=0.5,\omega_{1}=1$.
 Regime $\beta^{\star}$ $\alpha^{\star}$ Solution regime ($\gamma_{2}^{\star}$) Trend of $\gamma_{2}^{\star}$
@@
 $\boldsymbol{\omega_{2}\gg\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{1}^{2}\sigma_{1}^{2}\;+\;\omega_{2}^{2}\dfrac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ Interior Slow increase in $\omega_{2}$
-Figure 2: Illustration for experiment $j=2$.
-Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$.
-Bottom row: the optimal weight $\gamma_{2}^{\star}$.
-Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
-3.3.2 Choosing the design $\{j\}$
-After computing $\gamma_{j}^{\star}$, we then optimize over the design index $j$, by choosing
-$j^{\star}\in\arg\min_{j\in\{1,2\}}\max\Big\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\ \beta(j,\gamma_{j}^{\star})/\beta^{\star}\Big\}.$ (6)
-To build intuition, suppose $\gamma_{j}^{\star}$ is interior for both $j\in\{1,2\}$. Then $\alpha(j,\gamma_{j}^{\star})/\alpha^{\star}=\beta(j,\gamma_{j}^{\star})/\beta^{\star}$, which implies
-$j^{\star}\in\arg\min_{j\in\{1,2\}}\underbrace{\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}}_{\text{overall variance at }\gamma_{j}^{\star}}=\arg\min_{j\in\{1,2\}}\underbrace{\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}}_{\text{worst--case bias}/|B|\text{ at }\gamma_{j}^{\star}}.$
-In this case, optimal $\gamma_{j}^{\star}$ balances variance and worst–case bias.
-Because $\alpha^{\star}$ and $\beta^{\star}$ do not depend on $j$, $j^{\star}$ minimizes either (and therefore both) criteria $\alpha(j,\gamma_{j}^{\star})$ and $\beta(j,\gamma_{j}^{\star})$.
-A second case arises when one experiment (say $j=1$) is bias–dominant (i.e., $\gamma_{1}^{\star}=1$) or variance–dominant (i.e., $\gamma_{1}^{\star}=\sigma_{1}^{2}/(\sigma_{1}^{2}+v_{1}^{2})$), while the other admits an interior solution $\gamma_{2}^{\star}$ for which $\beta(2,\gamma_{2}^{\star})/\beta^{\star}=\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$.
-In this case, the optimal design selects the experiment that yields the smaller value of the dominating criterion (worst–case bias in the bias–dominant scenario, variance in the variance–dominant scenario).
-Although we view these as the leading scenarios, cases at the extreme boundaries can occur.
-The first is when the relative variance $\alpha/\alpha^{\star}$ or the worst–case bias $\beta/\beta^{\star}$ uniformly dominates the other for both experiments.
-Then the optimal design minimizes the dominating term; alternatively, if for one experiment $j$ the variance ratio $\alpha/\alpha^{\star}$ is dominant while for the other experiment j-j the worst–case bias ratio $\beta/\beta^{\star}$ is dominant, the minimax regret is the smaller of the two dominating values.
-These are desiderable properties as the optimal design always prioritizes the dominating term.
-Table 2 summarizes the cases.
-Table 2: Optimal shrinkage and experiment choice across different regimes.
-The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases).
-The other cases correspond to boundary solutions.
-Case (condition) Optimal $\gamma_{j}^{\star}$ Optimal $j^{\star}$ solves Solution $(j^{\star},\gamma_{j}^{\star})$
-Leading cases
-Case 1 no bias/ variance dominance $\gamma_{j}^{\star}$ in the interior $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star}$ minimizes bias and variance
-($\alpha(j,\gamma_{j})/\alpha^{\star}-\beta(j,\gamma_{j})/\beta^{\star}$ can flip sign) for $j\in\{1,2\}$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $\gamma_{j}^{\star}$ balances bias/variance
-Case 2 $j=1$ is variance dominant $\gamma_{1}^{\star}=\frac{\sigma_{1}^{2}}{v_{1}^{2}+\sigma_{1}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes variance
-$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
-Case 3 $j=1$ is bias dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes bias
-$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
-Other (boundary) solutions
-Case 4 Both $j\in\{1,2\}$ are variance-dominant $\gamma_{j}^{\star}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\frac{\sigma_{j}^{2}v_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}\right\}$ Minimizes variance
-($\alpha/\alpha^{\star}>\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
-Case 5 Both $j\in\{1,2\}$ are bias-dominant $\gamma_{j}^{\star}=1$ $\mathrm{arg}\max_{j}|\omega_{j}|$ Minimizes bias
-($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
-Case 6 bias dominant vs. variance dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}$ Minimizes bias $\beta/\beta^{\star}$ of $j=1$
-($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for $j=1$ and vice versa for $j=2$) $\gamma_{2}^{\star}=\frac{\sigma_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ $\left\{\frac{|\omega_{2}|^{2}}{|\omega_{1}|^{2}}1\{j=1\}+(\omega_{1}^{2}\sigma_{1}^{2}+\omega_{2}^{2}\frac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}})1\{j=2\}\right\}$ vs.
-variance $\alpha/\alpha^{\star}$ of $j=2$
-Illustration in Figure 3
-For illustration, Figure 3 compares the maximum regret of choosing experiment $j=1$ versus $j=2$ in three scenarios.
+Choice of the design
+Figure 11 compares the maximum regret of choosing experiment $j=1$ versus $j=2$ in three scenarios.
 The vertical dashed line marks the value where the two curves intersect and the designer is indifferent between $j=1$ and $j=2$.
@@
 Table 4 summarizes the discussion.
-Table 3: Qualitative behavior of maximum regret as $\omega_{2}$, $v_{2}$, or $\sigma_{2}$ increase (rows) in Figure 3.
+Table 4: Example: Qualitative behavior of maximum regret as $\omega_{2}$, $v_{2}$, or $\sigma_{2}$ increase (rows) in Figure 11.
 Entries summarize the direction and relative speed of change in the maximum regret for choosing experiment $j=1$ or $j=2$.
@@
 Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
-4 Extensions
-4.1 Nonlinear estimand $\tau(\theta)$
-This subsection extends the framework to smooth, potentially nonlinear $\tau(\theta)$, and consider a local asymptotic framework in the spirit of Andrews et al. (2020); Bonhomme and Weidner (2022).
-Setup.
-Take any given $(\mathcal{E},\Sigma)$ and $\gamma$.
-Let $\tilde{\theta}^{\mathrm{obs}}$ and $\tilde{\theta}^{\mathrm{exp}}$ denote the corresponding observational and experimental estimators of $\theta\in\mathbb{R}^{p}$.
-Assume a common $\sqrt{n}$ rate and the following local asymptotics:
-$\sqrt{n}\!\left(\tilde{\theta}^{\mathrm{exp}}-\theta\right)$ $\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,\Sigma^{\mathrm{exp}}),$
-$\sqrt{n}\!\left(\tilde{\theta}^{\mathrm{obs}}-\theta-b_{n}\right)$ $\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,\Sigma^{\mathrm{obs}}),\qquad b_{n}\to 0,$
-where $b_{n}\in\mathbb{R}^{p}$ is an unknown drift capturing local misspecification, and $\Sigma^{\mathrm{exp}},\Sigma^{\mathrm{obs}}$ are finite positive–definite matrices.
-We take the experimental and observational estimators to be asymptotically independent with common rates; this is not necessary but simplifies exposition.
-Thus $\tilde{\theta}^{\mathrm{obs}}\stackrel{{\scriptstyle p}}{{\to}}\theta$ (consistency), while $\sqrt{n}\,b_{n}$ may be bounded or even diverge.
-To justify first order linearization when $\sqrt{n}\,b_{n}$ may grow, we impose $\sqrt{n}\,\|b_{n}\|^{2}\to 0$ (e.g., $\|b_{n}\|=n^{-\alpha}$ with $\alpha\in(1/4,1/2]$).
-This implies that even if $b_{n}$ converges to zero, the asymptotic bias is possibly comparable to, smaller or larger than the asymptotic variance, therefore imposing weak restrictions on the magnitude of $\sqrt{n}b_{n}$ relative to the asymptotic variance.
-Linearization of $\tau$.
-Let $\tau:\mathbb{R}^{p}\to\mathbb{R}$ be twice continuously differentiable in a neighborhood of $\theta$, with Lipschitz gradient and bounded Hessian. A Taylor expansion around $\theta$ yields
-$\tau(\hat{\theta}_{\gamma})-\tau(\theta)=\omega(\theta)^{\top}(\hat{\theta}_{\gamma}-\theta)\;+\;\tfrac{1}{2}(\hat{\theta}_{\gamma}-\theta)^{\top}H_{\tau}(\bar{\theta})\,(\hat{\theta}_{\gamma}-\theta),$
-[…]
-$\sqrt{n}\bigl(\tau(\hat{\theta}_{\gamma})-\tau(\theta)\bigr)=\omega(\theta)^{\top}\sqrt{n}\,(\hat{\theta}_{\gamma}-\theta)\;+\;o_{p}(1).$ (7)
-Using the fact that $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$, we can consistently estimate the asymptotic MSE $n\mathbb{E}\left[(\tau(\hat{\theta}_{\gamma})-\tau(\theta))^{2}\right]$ by replacing $\omega(\theta)$ with $\omega(\tilde{\theta}^{\mathrm{obs}})\rightarrow_{p}\omega(\theta)$ and ignoring the $o_{p}(1)$ on the right-hand side.
-As a result all our results continue to hold under such linearization up-to a negligible $o_{p}(1)$.
-4.2 Confidence interval length as a regret criterion
-In this subsection we consider an alternative loss based on the length of confidence intervals (CIs).
-For a candidate design $(\mathcal{E},\Sigma(\mathcal{E}))$, shrinkage weights $\gamma$, and a bias vector $b\in\mathbb{R}^{p}$, let $\tilde{\omega}(\mathcal{E},\gamma)$ be as in (3).
-Define the two–sided, $(1-\eta)$ bias–aware lower and upper bounds for $\tau(\theta)$ by
-$\ell_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;-\;z_{1-\eta/2}\,\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},$
-$u_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;+\;z_{1-\eta/2}\,\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},$
-where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile.
-An audience, indexed by a worst–case bias radius $B\geq 0$, forms the worst–case CI
-$L_{B}(\mathcal{E},\Sigma,\gamma)~\equiv~\left[\inf_{\|b\|_{\infty}\leq B}\ell_{b}(\mathcal{E},\Sigma,\gamma),\;\sup_{\|b\|_{\infty}\leq B}u_{b}(\mathcal{E},\Sigma,\gamma)\right],$
-where we write $|L_{B}(\mathcal{E},\Sigma,\gamma)|$ for its total length. The proportional regret of $(\mathcal{E},\Sigma,\gamma)$ is
-$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{B\geq 0}\;\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{\inf_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}\!.$
-Theorem 2.
-Consider Setting 1 and let Assumption 1 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
-[…]
-with $\alpha$ and $\beta$ as defined in Equation (4) and $\alpha^{\star},\beta^{\star}$ as defined in Equation (5).
-Proof.
-See Appendix C.1
-Theorem 2 shows that the objective function is the same as the MSE-objective function up to a monotonic (square-root) transformation which does not affect the choice of the minimizer.
-4.3 Vector valued estimands
-Next, we extend our framework to vector value estimands of the form $\tau(\theta)\in\mathbb{R}^{q}$ for $q\geq 1$.
-Setting 3 (Vector-valued estimands).
-Consider a vector-valued target estimand with entries $\tau^{\ell}(\theta)\equiv(\omega^{\ell})^{\top}\theta$ for $\ell=1,\dots L$, and $\omega^{\ell},\theta\in\mathbb{R}^{k}$.
-For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a positive-definite matrix $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})\subset\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
-$\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{obs}}\right]-\theta=b,\qquad\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\right]-\theta_{\mathcal{E}}=0,$
-for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance as
-$\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E})$
-assumed to be known, with uniformly bounded entries and strictly positive definite.
-Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators to allow the optimal shrinkage weights to depend on the specific entry of $\tau$.
-$\hat{\tau}^{\ell}_{\gamma}\equiv\tau^{\ell}(\hat{\theta}^{\ell}(\gamma)),\qquad\hat{\theta}^{\ell}_{j}(\gamma)=\begin{cases}\gamma_{j}^{\ell}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j}^{\ell})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
-where $\gamma=(\gamma^{\ell}_{j})_{\ell\in[L],j\in\mathcal{E}}\in\mathbb{R}^{L\times|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$.
-Let $\hat{\theta}(\gamma)\equiv(\hat{\theta}^{1}(\gamma),\dots,\hat{\theta}^{L}(\gamma))\in\mathbb{R}^{p\times L}$ and $\hat{\tau}_{\gamma}\equiv(\hat{\tau}^{1}_{\gamma}(\hat{\theta}^{1}(\gamma)),\dots,\hat{\tau}^{L}_{\gamma}(\hat{\theta}^{L}(\gamma)))^{\top}\in\mathbb{R}^{L}$.
-Setting 3 is a flexible generalization of Setting 1, allowing each entry of $\tau$ to have its own vector of shrinkage weights.
-Similar to the scalar case, we make the following assumption regarding the first-order estimation error.
-Assumption 2 (First-order estimation error).
-For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume for all $\ell=1,\dots,L$
-$\tau^{\ell}(\theta)-\tau^{\ell}\!\big(\hat{\theta}^{\ell}(\gamma)\big)=(\omega^{\ell})^{\top}\,\big(\theta-\hat{\theta}^{\ell}(\gamma)\big),$ (8)
-for a known weighting vector $\omega^{\ell}\in\mathbb{R}^{p}$, with $|\omega^{\ell}_{i}|>0$, for all $i,\ell$.
-We begin by generalizing the mean squared error to higher-dimensional parameters.
-For a given design $(\mathcal{E},\Sigma)$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
-$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[\|\widehat{\tau}_{\gamma}-\tau\|_{2}^{2}\right],$ (9)
-where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$, and $\|\cdot\|_{2}$ denotes the $L^{2}$ Euclidean norm.
-As before, we introduce more compact notation before proceeding to the main result. For a given $(\mathcal{E},\gamma)$ define
-$\tilde{\omega}^{\ell}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega^{\ell}_{j},&j\notin\mathcal{E},\\ \omega^{\ell}_{j}(1-\gamma^{\ell}_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}^{\ell}_{\text{exp},j}(\mathcal{E},\gamma)=\begin{cases}0,&j\notin\mathcal{E},\\ \omega^{\ell}_{j}\gamma^{\ell}_{j},&j\in\mathcal{E},\end{cases}$
-and let
-$\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}^{\ell}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}^{\ell}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$
-Using this notation, we can write $\hat{\tau}^{\ell}_{\gamma}=\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}\,\hat{\theta}$, where $\hat{\theta}=(\tilde{\theta}^{\text{obs}\top},\tilde{\theta}^{\text{exp}\top})^{\top}\in\mathbb{R}^{p+|\mathcal{E}|}$.
-Stacking across the $\ell$ index gives the estimator for the full vector $\tau$: $\hat{\tau}_{\gamma}=\tilde{\omega}(\mathcal{E},\gamma)\,\hat{\theta}$, where the $\ell^{th}$ row of $\tilde{\omega}(\mathcal{E},\gamma)\in\mathbb{R}^{L\times(p+|\mathcal{E}|)}$ is $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}$.
-The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
-$\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ $\equiv\text{Trace}\left(\tilde{\omega}(\mathcal{E},\gamma)\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\right),$ (10)
-$\check{\beta}(\mathcal{E},\gamma)$ $\equiv\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}\bar{v}(\gamma)-(\omega^{\ell}_{\,\mathcal{E}})^{\top}\bar{v}(\gamma)_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(\bar{v}(\gamma)_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
-where $\omega^{\ell}_{\mathcal{E}}$ is the subvector of $\omega^{\ell}$ corresponding to the entries indexed by the experiment $\mathcal{E}$, and
-$\bar{v}(\gamma)\in\operatorname*{arg\,max}_{\{v\in\mathbb{R}^{p}:|v_{j}|\leq 1\forall j\}}\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\,\mathcal{E}})^{\top}v_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(v_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
-where the dependence of $\bar{v}(\gamma)$ on $\mathcal{E}$ is omitted for conciseness.
-The smallest worst-case variance and bias are defined as
-$\check{\alpha}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\beta(\mathcal{E},\gamma).$
-As before, we are interesting in minimizing
-$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\equiv\sup_{B\geq 0}\frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}.$
-The result below shows how our results directly extend to this case.
-Theorem 3.
-Consider Setting 3 and let Assumption 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
-$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\check{\alpha}(\mathcal{E},\Sigma,\gamma)}{\check{\alpha}^{\star}},\ \ \frac{\check{\beta}(\mathcal{E},\gamma)}{\check{\beta}^{\star}}\right\}.$
-Proof.
-See Appendix C.2
-5 Empirical applications
-In this section we provide two empirical applications.
-In the first application we study the problem of choosing where to conduct an experiment using observational data from Banerjee et al.
-(2024).
-In the second application we illustrate the method for choosing treatment arms when estimating a structural model, calibrating our estimation to the PROGRESA experiment (Todd and Wolpin, 2006; Attanasio et al., 2012).
-We solve the optimization program by solving for each choice of experiments, both over $\gamma$ and the allocation of sample size via numerical optimizationand then enumerate the solutions across all possible choices.
-5.1 Choosing experimental participants
-As a first exercise, we show how observational evidence can inform whom to recruit into an experiment.
-We consider the setting in Banerjee et al.
-(2024), who study how the expansion of microfinance affects village social networks in Karnataka, India.
-In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance.
-Because program rollout was not randomized across villages, they estimate effects using difference-in-differences (DiD).
-While DiD is informative, the lack of randomized assignment may bias estimates in the presence of selection and lack of parallel-trends (e.g., Marx et al., 2024; Ghanem et al., 2022).
-Motivated by these limitations, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods.
-We only use observational variation for choosing the experimental design and estimator.
-We then use experimental variation generated by Banerjee et al.
-(2024) to validate our procedure.
-Using preliminary observational estimates from Banerjee et al.
-(2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled?
-Concretely, we consider designs that randomize the introduction of microfinance in $n$ villages within one or more selected areas and study how the choice of areas and the allocation of $n$ vary with design constraints.
-Our objective is to design an experiment that yields an externally valid estimate of the effect of microfinance for rural Karnataka.
-We focus on the first outcome reported by Banerjee et al.
-(2024) corresponding to the density of the network, i.e., the percentage of connections a random household in a village has relative to the village size.
-In practice, implementation costs often depend on how geographically dispersed the study sites are.
-Coordinating with microfinance partners, training and supervising field teams, and conducting surveys become more expensive when sites are far apart or span multiple administrative units.
-We therefore consider operational constraints that limit the dispersion of enrolled villages and number of villages enrolled in the area.
-Figure 4: Observational villages and area-level DiD effects in Karnataka.
-The background heat map partitions the state into the four contiguous areas used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value.
-Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey).
-Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
-Observational study
-We use observational evidence from Banerjee et al.
-(2024) to construct area-level difference-in-differences (DiD) estimates for Karnataka.
-Specifically, we partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period.
-Figure 4 maps these four areas and reports the corresponding DiD point estimates.
-Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas.
-Table 4 summarizes, for each area, the number of treated and untreated villages, the pre-treatment sample variances, and the DiD point estimates together with their variances, computed assuming independence across villages.
-[^10] Alternative variance estimators (e.g., allowing within-area correlation) can be incorporated without changing the design logic.
-These area-specific DiD estimates and their variances serve as our primary observational inputs to the experimental design problem.
-[^11] We use DiD estimates to mimic the estimator used by the authors.
-An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity.
-Table 4: Area-level observational inputs (pre-prior) for Karnataka villages.
-The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village).
-For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$.
-Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$
-11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3
-22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140
-33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187
-44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130
-Experimental design
-We consider a family of designs that (i) select $E\in\{1,\ldots,4\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$).
-Here, $E=1$ corresponds to recruiting from a single area, while $E=4$ recruits from all four areas.
-We examine a grid of sample sizes that varies the number of participants from $10$ up to $104$ ($52$ treated units), with the latter corresponding to the size of the experiment in Banerjee et al.
-(2024).
-For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 4.
-We assume that the variance of a single treated–control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
-Results
-We begin by studying which areas are selected and how sample is allocated when the total experimental sample is large ($n_{1}=52$) and the number of admissible areas may or may not be constrained.
-Figure 5 visualizes the resulting allocation across Karnataka’s four observational areas for $E\in\{1,2,3,4\}$.
-With $E=1$, the algorithm concentrates recruitment in the Tumakuru area—the location with the highest uncertainty in the observational estimates.
-As $E$ relaxes to $2$ and $3$, additional areas enter and the total sample splits across them in roughly—but not exactly—equal shares, reflecting a trade-off between (i) exploiting heterogeneity in the observational variance and (ii) keeping experimental variance low.
-When $E=4$, all areas are eligible and the allocation smooths further across space.
-In the rest of our discussion, we discuss properties when not all areas may be selected (E<4E<4).
-[^12] The solution for $E=4$ always favors minimizing the bias first, since $\beta^{\star}=0$.
-Figure 6 tracks how area-level allocations evolve with the total number of villages.
-When the experiment is small (e.g., $10$ villages) and $E=1$, the algorithm favors an area with relatively low experimental variance—even if its observational estimate is less uncertain—because, at small $n$, experimental noise is first-order.
-As the total sample increases, the weight shifts toward areas with more uncertain observational estimates.
-Consistently, the area with the largest experimental variance (Area 2) is typically excluded unless all four areas can be chosen.
-Figure 7 reports the optimal shrinkage $\gamma_{j}^{\star}$ by area as total villages vary.
-The solution is interior for all $E<4$: $\gamma_{j}^{\star}$ rises with sample size, approaching one from below as experimental noise diminishes.
-For example, at a moderate sample (around 15 villages), $\gamma^{\star}$ is about $0.65$ when $E\in\{2,3\}$; with $E=1$, concentrating the sample within one area lowers experimental variance and therefore yields a larger $\gamma^{\star}$.
-Main MSE comparison
-We compare three strategies: (i) our chosen design which chooses the area, sample size and $\gamma^{\star}$ as described in Section 3, (ii) a standard benchmark that selects areas uniformly at random, allocates villages evenly across the chosen areas, and sets $\gamma_{j}=1$ for experimental estimates, and (iii) an oracle that knows the bias vector $b$ and optimizes both design (areas and allocation) and $\gamma$.
-Because $b$ is unknown in practice, for this exercise, we calibrate it using the difference between the Hyderabad RCT ATE obtained from the follow-up experiment of Banerjee et al.
-(2024) and the Karnataka DiD average effect across sites, treating $b$ as common across areas.
-Figure 8 plots the ratio $\text{MSE}/\text{MSE}^{\star}$, where $\text{MSE}^{\star}$ is the oracle’s MSE.
-As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone.
-Across $E\in\{1,2,3\}$, the random benchmark remains well above the oracle—by roughly 3–5 percentage points (pp) for $E=1$, about 10 pp for $E=2$, and around 20 pp for $E=3$ at the top of the grid—whereas our design tracks the oracle closely (within a few percentage points) even without knowledge of $b$.
-The gap is the largest when a larger number of areas $E$ can be included, because our procedure can better emulate the oracle by selecting (and differentially weighting) favorable areas.
-On the other hand, the random benchmark ignores information about observational and experimental variance heterogeneity.
-As sample size increases, the gap between the two stabilizes: with large $n_{1}$, optimal shrinkage approaches $\gamma^{\star}\!\approx\!1$ and an even allocation across the selected areas is nearly optimal once the area set is well chosen.
-This empirical calibration highlights the benefit of the procedure that is able to track the oracle even in the absence of knowledge of the bias $b$, while significantly outperforming standard alternative benchmarks such as a random allocation strategy.
-Figure 5: Optimal area selection and village allocation for a large experiment ($n_{1}=52$) under constraints on the number of areas $E$.
-Each panel corresponds to a value of $E$; shading indicates the total number of recruited villages in each area.
-With $E=1$ the design concentrates in the area with the noisiest observational estimate; as $E$ increases, allocation spreads across areas, reflecting a trade-off between observational uncertainty and experimental variance.
-Figure 6: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for each constraint on the number of eligible areas $E$.
-Lines show the number of villages assigned to each area.
+Appendix D Additional figures and tables
+Figure 12: Share of experimental papers published in AEA journals also presenting experimental results in combination with observational estimates (either from a reduced form, or from a structural model or from both).
+Table 5: Optimal shrinkage and experiment choice across different regimes in the presence of two parameters (Setting 2) The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases).
+The other cases correspond to boundary solutions.
+Case (condition) Optimal $\gamma_{j}^{\star}$ Optimal $j^{\star}$ solves Solution $(j^{\star},\gamma_{j}^{\star})$
+Leading cases
+Case 1 no bias/ variance dominance $\gamma_{j}^{\star}$ in the interior $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star}$ minimizes bias and variance
+($\alpha(j,\gamma_{j})/\alpha^{\star}-\beta(j,\gamma_{j})/\beta^{\star}$ can flip sign) for $j\in\{1,2\}$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $\gamma_{j}^{\star}$ balances bias/variance
+Case 2 $j=1$ is variance dominant $\gamma_{1}^{\star}=\frac{\sigma_{1}^{2}}{v_{1}^{2}+\sigma_{1}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes variance
+$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
+Case 3 $j=1$ is bias dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes bias
+$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
+Other (boundary) solutions
+Case 4 Both $j\in\{1,2\}$ are variance-dominant $\gamma_{j}^{\star}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\frac{\sigma_{j}^{2}v_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}\right\}$ Minimizes variance
+($\alpha/\alpha^{\star}>\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
+Case 5 Both $j\in\{1,2\}$ are bias-dominant $\gamma_{j}^{\star}=1$ $\mathrm{arg}\max_{j}|\omega_{j}|$ Minimizes bias
+($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
+Case 6 bias dominant vs. variance dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}$ Minimizes bias $\beta/\beta^{\star}$ of $j=1$
+($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for $j=1$ and vice versa for $j=2$) $\gamma_{2}^{\star}=\frac{\sigma_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ $\left\{\frac{|\omega_{2}|^{2}}{|\omega_{1}|^{2}}1\{j=1\}+(\omega_{1}^{2}\sigma_{1}^{2}+\omega_{2}^{2}\frac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}})1\{j=2\}\right\}$ vs.
+variance $\alpha/\alpha^{\star}$ of $j=2$
 Figure 13: Optimal shrinkage $\gamma_{j}^{\star}$ by area as a function of total treated villages $n_{1}$ (from ten to fifty-two villages), under different constraints $E$.
 For $E<4$, the solution is interior ($0<\gamma_{j}^{\star}<1$) and increases with $n_{1}$, approaching one as experimental noise falls.
-Figure 8: Relative MSE across designs under an hypothetical scenario where the bias equal the difference between the experimental and observational ATE estimate from Banerjee et al.
-(2024) as $n_{1}$ varies (from ten to fifty-two villages).
-The figure reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$ for (i) the proposed design (area targeting and optimal $\gamma^{\star}$), (ii) a random benchmark (uniform area selection and equal allocation), where $\mathrm{MSE}^{*}$ denotes the MSE of the oracle that knows the bias, as $n_{1}$ varies and for each $E$.
-The proposed design closely tracks the oracle across $n_{1}$ and $E$, while the random benchmark remains substantially above, especially when $E$ is small.
-5.2 Choosing treatment arms for model estimation
-Next, we illustrate how our framework selects among treatment arms when the goal is to estimate and deploy a structural model.
-Combining experimental or pre-program observational data with structural models is increasingly used for program evaluation (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022).
-Because large-scale experimentation may be infeasible due to cost or fairness constraints (Muralidharan and Niehaus, 2017), researchers are often interested in designing small-scale experiments whose information, combined with a model, is informative about general-equilibrium (GE) effects.
-We consider a researcher evaluating a conditional cash-transfer (CCT) for sending children to school in rural Kenya where such a program is not in place.
-Due to budget and feasibility constraints, the researcher can only randomize at small scale (partial equilibrium), but ultimately wishes to predict large-scale GE effects.
-As preliminary observational estimates, we use information from the Mexican PROGRESA experiment.
-5.2.1 Model
-Individual choice model
-Following Bonhomme and Weidner (2022) (and in turn Todd and Wolpin (2006)), let $S\in\{0,1\}$ denote school attendance, $C$ consumption, $Y$ (pre-transfer) household income, $W$ the child’s potential wage, and $t$ the stipend when enrolled.
-Abstracting from covariates for exposition (we will introduce covariates in the estimation), utility is
-$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{3}-\xi_{1}-\xi_{2})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (11)
-with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS. The parametrization $(\xi_{3}-\xi_{1}-\xi_{2})$ is without loss and simplifies expressions below. Enrollment satisfies
-$S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$
-Letting $Z(Y,W,t)\equiv\xi_{1}W-\xi_{2}Y-\xi_{3}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$.
-General equilibrium effects
-We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya.
-GE feedback is allowed through income and wages: for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, we assume
-$Y_{i}(t)=y(t)+\varepsilon_{Yi},\qquad W_{i}(t)=w(t)+\varepsilon_{Wi}.$
-Let $y_{0}\equiv\partial_{t}y(t)|_{t=0}$ and $w_{0}\equiv\partial_{t}w(t)|_{t=0}$.
-Define $\phi_{0}\equiv\mathbb{E}\!\left[\phi\!\big(-Z\big(Y(0),W(0),0\big)\big)\right]$.
-Our estimand of interest is the marginal effect of an increase in a small transfer $t$ to all eligible households
-$\left.\frac{\partial\,\mathbb{E}\!\left[\Pr(S{=}1\mid Y(t),W(t),t)\right]}{\partial t}\right|_{t=0}=\;\phi_{0}\cdot\big(\,\xi_{3}+\xi_{2}y_{0}-\xi_{1}w_{0}\,\big),$ (12)
-which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{2}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$.
-Income and wage effects
-We estimate $y_{0}=1.5$ using a local fiscal/income multiplier $M_{\text{tot}}\approx 2.5$ from Egger et al.
-(2022), which in turn implies $y_{0}=M_{\text{tot}}-1$.
-[^13] Egger et al.
-(2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$.
-We write $w_{0}$ as a function of the parameters: Under simple market clearing for labor supply and demand, we can show [^14] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$.
-Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working.
-From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{3}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{1}$ at $t=0$.
-Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{3}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{1}H$.
-Dividing the equation by $H$ we obtain the desired result.
-$w_{0}=\;\frac{\phi_{0}\,\xi_{3}}{\phi_{0}\,\xi_{1}-d},$ (13)
-where $d$ denotes the slope of the demand curve divided by total hours of work.
-We calibrate $d$ to $0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained form the labor demand elasticity in Espey and Thilmany (2000).
-For expositional convenience, we treat $y_{0},d$ as constant, while $\phi_{0}\xi_{3}$ and $\phi_{0}\xi_{1}$ are parameters of the model and $w_{0}$ is a (non-linear) function of such parameters (it is possible to treat also $y_{0},d$ as additional possibly misspecified parameters, omitted for brevity).
-[^15] We treat $d$ as a fixed constant given the extensive literature on estimating demand elasiticities, and the fact that $S_{0}$, $W_{0}$ can be calibrated to pre-experimental data (for simplicity here using the PROGRESA pre-experimental data), and $y_{0}$ because computed experimentally in Kenya.
-When we are concerned with misspecification of those, researchers can augment $\theta$ to also incorporate $y_{0},d$ without having to estimate those estimated in the experiment.
-Map to the design parameters
-We can now define the parameters of interest as
-$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{3},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{1},$ (14)
-with
-$\tau(\theta)\;=\;\theta_{2}\;+\;y_{0}\,\theta_{1}\;+\;w_{0}(\theta)\,\theta_{3},\qquad w_{0}(\theta)\;=\;\frac{\theta_{2}}{-\theta_{3}-d},$ (15)
-where the explicit form of $w_{0}(\theta)$ follows from (13).
-Given observational estimates $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2},\tilde{\theta}^{\mathrm{obs}}_{3})$ described below, and following Section 4.1, we use $\omega=\frac{\partial\tau}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}$.
-[^16] Because $w_{0}$ is a function of $\theta$, we have $\frac{\partial\tau}{\partial\theta_{1}}=y_{0},\frac{\partial\tau}{\partial\theta_{2}}=\frac{d}{\theta_{3}+d},\frac{\partial\tau}{\partial\theta_{3}}=-\,\frac{\theta_{2}\,d}{(-\theta_{3}-d)^{2}}.$
-5.2.2 Observational study and design of the experiment
-Using experimental data from Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ via probit with standard controls (age, distance to school, eligibility, year, highest grade attained).
-Estimation is conducted separately by gender, focusing here on the effects on female students.
-We pool treated and control observations to estimate the schooling effect $\xi_{4}$, controlling for the individual-level subsidy.
-In turn, we obtain observational estimates $\big(\tilde{\theta}_{1}^{\mathrm{obs}},\tilde{\theta}_{2}^{\mathrm{obs}},\tilde{\theta}_{3}^{\mathrm{obs}}\big)$ that map to the model parametrization.
-Standard errors are constructed using the Delta method with clustering at the village level.
-Point estimates, estimated sensitivity $\omega$, and standard errors are reported below.
-Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sigma$
-[…]
-[…]
-[…]
-Observational estimates for female students, corresponding sensitivity parameter $\omega$, and standard errors.
-$\Sigma^{\mathrm{obs}}\;=\;\begin{bmatrix}4.31&-11.31&5.57\\ -11.31&973.11&-126.16\\ 5.57&-126.16&1038.56\end{bmatrix}\times e^{-9}$
-Variance–covariance matrix of observational estimates for female students.
-We observe heterogeneity in both the sensitivity weights $\omega$ and (more significantly) the precision of the observational estimates.
-Our goal is to trade off these features, together with prospective experimental variation, to design the experiment in Kenya.
-Because the setting differs in time and country, PROGRESA-based estimates provide informative but possibly biased baselines.
-Candidate small-scale treatment-arms in Kenya (partial equilibrium)
-We consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, as they may not be able to randomize treatments across the full population of interest.
-We examine three possible experiments for a given sample size $n$:
-  • $j=\{1\}$: Unconditional transfer (income shock).
-The researcher randomizes a small income shock to a small fraction of households (implying no general-equilibrium effects).
-Under a first-order (Taylor) approximation, the experiment identifies
-    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,0)]}{\partial Y}\right|_{t=0}=\theta_{1},$
-    with precision $v_{1}^{2}/n$.
-The researcher optimizes over $\gamma_{1}$, the weight used to combine the experimental and observational estimates of $\theta_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
-  • $j=\{2\}$: Conditional cash transfer (stipend).
-The researcher randomizes a small stipend $t$, conditional on attending school, to a small fraction of households (with prices held fixed).
-Under a first-order approximation, this identifies
-    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,t)]}{\partial t}\right|_{t=0}=\theta_{2},$
-    with precision $v_{2}^{2}/n$.
-The researcher optimizes over $\gamma_{2}$, the weight used to combine the experimental and observational estimates of $\theta_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
-  • $j=\{1,2\}$: Two-arm design.
-The researcher runs (j=1j{=}1) and (j=2j{=}2) on independent samples, identifying $(\theta_{1},\theta_{2})$ with precisions $(v_{1}^{2}/n_{1},\ v_{2}^{2}/n_{2})$ where $n_{1}+n_{2}=n$.
-They optimize over both $\gamma_{1}$ and $\gamma_{2}$, while $\theta_{3}$ is calibrated to the PROGRESA study.
-We consider two scenarios.
-In the first, the researcher (and oracle) can choose one or both experiments.
-In the second, they can choose only one of the two (either $j=\{1\}$ or $j=\{2\}$).
-Regret is defined accordingly, conditional on the set of feasible designs.
-We calibrate $v_{1}^{2}$ as $\sigma_{1}^{2}\times n^{\mathrm{obs}}$, where $\sigma_{1}^{2}$ is the variance of $\tilde{\theta}_{1}^{\mathrm{obs}}$ and $n^{\mathrm{obs}}$ is the observational-sample size ($n^{\mathrm{obs}}=1089$ for the female sample using data from eligible students in Todd and Wolpin (2006)).
-We calibrate $v_{2}^{2}$ analogously as $\sigma_{2}^{2}\times n^{\mathrm{obs}}$.
-Experiment type Identified parameter Choice variables Experiment standard error Description
-$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $6.56e^{-5}/\sqrt{n}$ Unconditional transfer (income shock, UCT). Randomize an income shock to a small fraction of households.
-$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $9.86e^{-4}/\sqrt{n}$ Conditional cash transfer (education shock, CCT). Randomize a conditional transfer to a small fraction of households.
-$j=\{1,2\}$ $(\theta_{1},\,\theta_{2})$ $\gamma_{1},\ \gamma_{2},\ n_{1}$ $6.56e^{-5}/\sqrt{n_{1}}$ (arm 1), $9.86e^{-4}/\sqrt{n-n_{1}}$ (arm 2) Two-arm design.
-Run (j=1j{=}1) and (j=2j{=}2) on independent samples with sample sizes respectively $n_{1},n-n_{1}$.
-Table 5: Design options, identified parameters, choice variables, and per-unit variances.
-5.2.3 Results
-Figure 9: Regret–optimal sample allocation across treatment arms when both the unconditional cash transfer (UCT; income shock) and the conditional cash transfer (CCT; stipend) arms are available.
-The optimal solution is interior: the vast majority of participants (over 90%) are assigned to CCT, with a small but nonzero fraction assigned to UCT to hedge misspecification.
-This pattern reflects the higher payoff to learning about CCT in this application relative to UCT.
-Figure 10: Optimal shrinkage weights and design choice as functions of the total sample size $n$.
-Left panel: When both arms are run, the weight on the UCT experimental estimator is essentially one across $n$, while the weight on the CCT experimental estimator, $\gamma^{\star}_{\text{CCT}}$, rises with precision (from about 0.5 at small $n$ to roughly 0.9 by $n=1000$), reflecting the increasing value of experimental evidence.
-Right panel: When restricted to a single arm, the optimal choice switches from UCT at small $n$ (variance dominates under CCT) to CCT once $n$ crosses a threshold (around $n\approx 500$), where bias considerations dominate and favor CCT.
-Notes: $\gamma_{j}^{\star}$ is the optimal weight on the experimental (vs.
-observational) estimator for parameter $j$; we set $\gamma_{j}^{\star}=0$ to indicate that arm $j$ is not chosen.
-Figure 11: Worst–case regret by design.
-Left panel: Allowing both arms with an interior allocation yields regret that approaches one as $n$ increases, indicating near–oracle performance under our normalization.
-In contrast, single–arm designs exhibit nonvanishing regret because bias remains first–order even as variance shrinks.
-Right panel: Focusing on the single–arm problem, UCT dominates only at small $n$; beyond the same threshold ($\approx 500$), CCT delivers uniformly lower regret as variance becomes second–order relative to bias.
-Notes: Regret is the worst–case objective normalized so that a value of one corresponds to the oracle benchmark; UCT = unconditional cash transfer, CCT = conditional cash transfer.
-We now summarize results as we vary both the set of experiments the researcher can run and the total sample size $n$.
-Specifically, Figure 9 reports the sample allocation implied by the regret-optimal design when both treatment arms are available.
-The optimal allocation is interior: most participants (over roughly $90\%$) are assigned to the CCT arm (education), with a small but nonzero fraction assigned to the UCT arm (income shock).
-Two forces rationalize this pattern.
-First, the CCT parameter is more misspecification-sensitive ($\omega_{2}\approx 1.98$ vs.
-$\omega_{1}\approx 1.5$), so generating experimental evidence on CCT has a larger payoff in bias reduction.
-Second, the income effect $\theta_{1}$’s variance is about an order of magnitude smaller than the stipend effect $\theta_{2}$’s variance in our PROGRESA calibration, lowering the marginal returns to learn about UCT relative to CCT.
-Hence, while the designer keeps some allocation on UCT, most of the sample is placed on CCT; our framework makes this bias–variance trade-off explicit.
-Alongside the sample allocation, our method optimizes the shrinkage weights on the parameters of interest.
-Figure 10 (left panel) shows that when both arms can be run, we place essentially full weight on the experimental estimator for UCT ($\gamma^{\star}_{\text{UCT}}\approx 1$ across $n$), while for CCT we gradually shift toward the experimental estimator as precision improves: $\gamma^{\star}_{\text{CCT}}$ rises from about $0.5$ at small $n$ to approximately $0.9$ by $n=1000$.
-Intuitively, higher $n$ lowers experimental variance, making it optimal to rely more heavily on experimental evidence for the CCT parameter.
-It is also informative to compare optimal choices when the researcher is restricted to one arm.
-Figure 10 (right panel) displays the selected arm as a function of $n$ (together with the corresponding $\gamma_{j}^{\star}$).
-We encode “not chosen” as $\gamma_{j}^{\star}=0$.
-For small samples ($n\lesssim 500$), the optimal choice is UCT: the experimental CCT estimator would be too noisy, so the resulting variance under CCT dominates, while the bias advantage of CCT over UCT is limited because $\omega_{1}$ and $\omega_{2}$ are of similar magnitude.
-Once $n$ crosses a critical threshold (around $n\approx 500$), CCT becomes dominant: by combining observational and experimental estimates for CCT, we can achieve comparable variance, and the worst-case bias is lower than under UCT.
-Finally, Figure 11 reports worst-case regret for each design choice.
-The left panel shows that the regret of our method (green line, running both arms with an interior allocation) declines toward one as $n$ increases, reflecting near-oracle performance under our normalization.
-In contrast, single-arm designs exhibit nonvanishing regret because bias remains first order even as variance shrinks.
-The right panel focuses on the single-arm problem: both the oracle and the researcher can only choose a single arm.
-In this case, UCT is preferable only at small $n$; beyond the same threshold ($\approx 500$), CCT yields uniformly lower regret as variance becomes second order relative to bias.
-The regret vanishes relative to the oracle that can only choose a single arm.
-These results illustrate how our approach systematically balances precision and misspecification risk to balance experimental variation with structural models.
-6 Implications for practice
-This paper studies experimental design in the presence of complementary observational evidence.
-Practitioners often have access to observational inputs—e.g., estimates from models trained on existing data (which may be misspecified), results from prior experiments that may not transport to the target context, or estimates from observational designs that may suffer from confounding.
-Our goal is to combine such evidence with new experimental data, which deliver unbiased estimators for specific parameters of interest.
-We ask which experiment to run, and how to run it, under budget constraints that limit the number of arms and/or impose per-unit (variable) costs on effective sample size.
-A key challenge is that the bias of the observational estimators is unknown in practice.
-We adopt a minimax proportional regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias.
-This reveals a fundamental trade-off between precision and robustness.
-The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap).
-We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
-In practice, the workflow is:
-  • Define the estimand(s) of interest.
-Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$.
-For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations.
-Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify.
-  • Assemble informative observational evidence.
-Collect observational estimates $\tilde{\theta}^{\rm obs}$ and their covariance $\tilde{\Sigma}^{\rm obs}$.
-These serve as informative (but potentially biased) baselines.
-Such evidence may come from observational designs, structural estimates with pre-experimental data, or prior experiments conducted in different contexts or periods.
-  • Compute the sensitivity parameters.
-Using the observational baseline, compute the sensitivity weights $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\big|_{\tilde{\theta}^{\rm obs}}$, which quantify how bias in each coordinate of $\theta$ propagates to $\tau(\theta)$.
-Large $|\omega_{j}|$ indicates greater payoff to learning the $j$th component.
-  • Specify feasibility constraints and calibrate experimental variance.
-Enumerate the admissible design set $\mathcal{S}$.
-That is, $\mathcal{S}$ is the set containing combinations of parameters that can be learned jointly via experiments (which arms, sites, or mechanisms can be randomized).
-Define, for each combination of parameters $\mathcal{E}\in\mathcal{S}$, the corresponding precision set $\mathcal{G}(\mathcal{E})$ as the feasible set of experimental variances for a given experiment $\mathcal{E}$.
-This requires specifying how the sample size $n_{j}$ allocated to each arm maps to the variance of each experimental estimate.
-Calibrate per-unit experimental variances using pilot studies or historical data.
-  • Run the method and decide.
-For each feasible set of experiments $\mathcal{E}\in\mathcal{S}$ and precisions $\Sigma\in\mathcal{G}(\mathcal{E})$, we first optimize over the shrinkage vector $\gamma$ (how to weight observational vs.
-experimental estimates).
-We then optimize over the precision/allocation $\Sigma$ (e.g., the sample size $n_{j}$ assigned to each arm).
-Finally, we optimize over the experiment choice $\mathcal{E}\in\mathcal{S}$.
-In practice, this involves enumerating feasible $\mathcal{E}$ and simple numerical optimization for $(\gamma,\Sigma)$.
-The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$).
-  • Reporting and diagnostics.
-In a pre-analysis plan, we recommend reporting (i) the estimands of interest; (ii) the chosen arm(s) and the final sample size allocation, (iii) the shrinkage weights $\gamma^{\star}$ by parameter between the observational and experimental study, and (iv) the two normalized components $\alpha/\alpha^{\star}$ (variance-gap) and $\beta/\beta^{\star}$ (bias-gap), that can help assess the quality of the allocation.
-Before committing to a single experiment choice, researchers may also want to explore sensitivity to a range of values for the experimental variances.
-The applicability of our framework spans a wide range of settings in economics and beyond.
-Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024).
-In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis.
-Beyond economics, medical applications include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025).
-In all cases, the same recipe identifies which parameters to target, how much to learn about each, and how to blend experimental and observational evidence.
-Several open questions remain for future research.
-These include settings where researchers have well-specified priors about bias (e.g., from meta-analysis), sequential or adaptive experimental choices, and environments where the exogenous source of variation yields only partial identification of some parameters of interest.
-Appendix A Additional figures
-Figure 12: Share of experimental papers published in AEA journals also presenting experimental results in combination with observational estimates (either from a reduced form, or from a structural model or from both).
-Appendix B Proofs of main results
-B.1 Proof of Theorem 1
-Preliminary notation
-It suffices to prove the theorem for any given feasible design and weights $(\mathcal{E},\Sigma(\mathcal{E}),\gamma)$.
-Throughout, we denote by $\mathcal{E}^{c}\equiv\{1,\dots,p\}\setminus\mathcal{E}$ the complement of $\mathcal{E}$.
-Let $\tilde{\omega}(\mathcal{E},\gamma)=[\,\tilde{\omega}^{\mathrm{obs}}(\mathcal{E},\gamma)^{\top},\ \tilde{\omega}^{\mathrm{exp}}(\mathcal{E},\gamma)^{\top}\,]^{\top}$ as in (3), with
-$\tilde{\omega}^{\mathrm{obs}}_{\mathcal{E}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ(1-\gamma),\quad\tilde{\omega}^{\mathrm{obs}}_{\mathcal{E}^{c}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}^{c}},\quad\tilde{\omega}^{\mathrm{exp}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ\gamma,$
-so that $\tilde{\omega}^{\mathrm{obs}}\in\mathbb{R}^{p}$ and $\tilde{\omega}^{\mathrm{exp}}\in\mathbb{R}^{|\mathcal{E}|}$.
-We let $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ and denote $|1-\gamma|$ a vector of dimension $|\mathcal{E}|$ with entry $j$ equal to $|1-\gamma_{j}|$.
-For an observational bias $b\in\mathbb{R}^{p}$ (i.e., $E[\tilde{\theta}^{\mathrm{obs}}]-\theta=b$), the mean–squared error is
-$\mathrm{MSE}_{\mathcal{E},\Sigma,b}\!\left[\hat{\tau}_{\gamma}\right]~=~\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)\;+\;\bigl(\tilde{\omega}^{\mathrm{obs}}(\mathcal{E},\gamma)^{\top}b\bigr)^{2}.$
-For an oracle that knows $\|b\|_{\infty}\leq\bar{B}$,
-$R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{\|b\|_{\infty}\leq\bar{B}}\mathrm{MSE}_{\mathcal{E},\Sigma,b}\!\left[\hat{\tau}_{\gamma}\right]~=~\bar{B}^{2}\!\left(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|_{\mathcal{E}}^{\top}|\omega_{\mathcal{E}}|\right)^{\!2}+\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),$
-since $\sup_{\|b\|_{\infty}\leq\bar{B}}\tilde{\omega}^{\mathrm{obs}\,\top}b=\bar{B}\,\|\tilde{\omega}^{\mathrm{obs}}\|_{1}=\bar{B}\big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|_{\mathcal{E}}^{\top}|\omega_{\mathcal{E}}|\big)$.
-B.1.1 The oracle problem
-Given $(\mathcal{E},\Sigma)$, the oracle first chooses the shrinkage vector
-$\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)~\in~\arg\min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}R_{\bar{B}}(\mathcal{E},\Sigma,\gamma),$
-and then (by backward induction) the design
-$\big(\mathcal{E}^{\star}_{o}(\bar{B}),\ \Sigma^{\star}_{o}(\bar{B})\big)~\in~\arg\min_{\mathcal{E},\,\Sigma}R_{\bar{B}}\big(\mathcal{E},\Sigma,\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)\big).$
-Let $R^{\star}_{o}(\bar{B})\equiv\min_{\mathcal{E},\Sigma,\gamma}R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)$ denote the oracle’s minimized worst–case MSE.
-Lemma 1 (Convexity and strict convexity).
-For fixed $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$, the map $\gamma\mapsto R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)$ is convex on $\gamma$.
-Moreover, if $|\omega_{j}|>0$ for all $j\in\mathcal{E}$, then $\gamma\mapsto R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)$ is strictly convex on $\gamma$.
-Proof.
-Write $R_{\bar{B}}=\bar{B}^{2}g+h$ with
-$g(\gamma)=\Big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\Big)^{2},\qquad h(\gamma)=\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma).$
-(i) Convexity of $g$.
-The map $\gamma\mapsto|1-\gamma|^{\top}|\omega_{\mathcal{E}}|$ is convex (sum of absolute values of affine functions).
-Adding the constant $\|\omega_{\mathcal{E}^{c}}\|_{1}$ preserves convexity, and squaring preserves convexity.
-(ii) Strict convexity of $h$.
-Let $D\equiv\mathrm{diag}(\omega_{\mathcal{E}})$.
-Without loss, reorder the entries of $\theta$ so that we have $\theta=(\theta_{\mathcal{E}^{c}},\theta_{\mathcal{E}})$.
-Then
-$\tilde{\omega}(\mathcal{E},\gamma)=\underbrace{\begin{bmatrix}\omega_{\mathcal{E}^{c}}\\ D\mathbf{1}\\ 0\end{bmatrix}}_{=:d}\;+\;\underbrace{\begin{bmatrix}0\\ -D\\ D\end{bmatrix}}_{=:A}\gamma,\quad\text{so}\quad h(\gamma)=(A\gamma+d)^{\top}\Sigma(\mathcal{E})\,(A\gamma+d).$
-Hence $h(\gamma)=\gamma^{\top}(A^{\top}\Sigma A)\gamma+2\gamma^{\top}A^{\top}\Sigma d+d^{\top}\Sigma d$, and the Hessian is
-$\nabla^{2}h(\gamma)=2\,A^{\top}\Sigma(\mathcal{E})\,A.$
-Because $\Sigma(\mathcal{E})\succ 0$ and, when $|\omega_{j}|>0$ for all $j\in\mathcal{E}$, the columns of A=[ 0;D;D]A=[\,0;\,-D;\,D\,] are linearly independent, we have for any nonzero $u$,
-$u^{\top}A^{\top}\Sigma A\,u=(Au)^{\top}\Sigma(Au)>0.$
-Thus $A^{\top}\Sigma A\succ 0$ and $h$ is strictly convex (or weakly convex if $\omega_{j}=0$ for some $j$).
-(iii) Conclusion. A sum of a convex function ($\bar{B}^{2}g$) and a strictly convex function ($h$) is strictly convex; when some $\omega_{j}=0$, the sum remains convex. ∎
-Next, we show that as the worst-case bias becomes large enough, the oracle will put zero weight on the observational estimators.
-Lemma 2.
-Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$ and write $A\equiv\|\omega_{\mathcal{E}^{c}}\|_{1}$, with $|\omega_{j}|>0$ for all $j$. Let
-$R_{\bar{B}}(\gamma)=\bar{B}^{2}\Big(A+t(\gamma)\Big)^{2}+\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad t(\gamma)\equiv|1-\gamma|^{\top}|\omega_{\mathcal{E}}|,$
-for $\gamma\in\mathbb{R}^{|\mathcal{E}|}$. Then:
-(i) If $A>0$ (equivalently, $\|\omega_{\mathcal{E}}\|_{1}<\|\omega\|_{1}$), there exists
-$\bar{B}^{\max}(\mathcal{E},\Sigma,\omega)\;\leq\;\sqrt{\,2\,\|\Sigma(\mathcal{E})\|_{\infty}\cdot\max\!\Big\{2,\ \frac{\|\omega\|_{1}}{A}\Big\}\,}$
-such that every minimizer satisfies $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)=\mathbf{1}$ for all $\bar{B}\geq\bar{B}^{\max}$.
-(ii) If $A=0$ (equivalently, $\|\omega_{\mathcal{E}}\|_{1}=\|\omega\|_{1}$), then for any sequence $\bar{B}\to\infty$ every sequence of minimizers obeys $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)\to\mathbf{1}$.
-Proof.
-Let $x(\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)$ and $x(1)\equiv\tilde{\omega}(\mathcal{E},\mathbf{1})$.
-Using $a^{\top}\Sigma a-b^{\top}\Sigma b=(a-b)^{\top}\Sigma(a+b)$ and the Holder’s bound $|u^{\top}\Sigma v|\leq\|\Sigma\|_{\infty}\|u\|_{1}\|v\|_{1}$, we get
-$\big|x(\gamma)^{\top}\Sigma x(\gamma)-x(1)^{\top}\Sigma x(1)\big|\leq\|\Sigma\|_{\infty}\,\|x(\gamma)-x(1)\|_{1}\,\|x(\gamma)+x(1)\|_{1}.$
-In addition,
-$x(\gamma)-x(1)=\begin{bmatrix}\omega_{\mathcal{E}^{c}}-\omega_{\mathcal{E}^{c}}\\[1.0pt] \omega_{\mathcal{E}}\circ(1-\gamma)\\[1.0pt] -\,\omega_{\mathcal{E}}\!\circ\!(1-\gamma)\end{bmatrix}\Rightarrow\|x(\gamma)-x(1)\|_{1}=2\,t(\gamma).$
-Also,
-$x(\gamma)+x(1)=\begin{bmatrix}2\,\omega_{\mathcal{E}^{c}}\\[1.0pt] \omega_{\mathcal{E}}\!\circ\!(1-\gamma)\\[1.0pt] \omega_{\mathcal{E}}\!\circ\!(1+\gamma)\end{bmatrix}\Rightarrow\|x(\gamma)+x(1)\|_{1}\leq 2\|\omega_{\mathcal{E}^{c}}\|_{1}+t(\gamma)+\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1+\gamma_{j}|.$
-Using the triangular inequality for reals $u,v$, $|u|=|u-v+v|\leq|u-v|+|v|$, so that $|1+\gamma_{j}|\leq|1+\gamma_{j}-2|+2=|1-\gamma_{j}|+2$, we have $\sum_{j\in\mathcal{E}}|\omega_{j}||1+\gamma_{j}|\leq t(\gamma)+2\|\omega_{\mathcal{E}}\|_{1}.$ Therefore
-$\|x(\gamma)+x(1)\|_{1}\leq 2\|\omega\|_{1}+2\,t(\gamma),$
-and hence
-$\big|x(\gamma)^{\top}\Sigma x(\gamma)-x(1)^{\top}\Sigma x(1)\big|\leq 4\,\|\Sigma\|_{\infty}\,t(\gamma)\,\big(\|\omega\|_{1}+t(\gamma)\big).$ (16)
-(i) Case $A>0$. For any $\gamma$,
-$R_{\bar{B}}(\gamma)-R_{\bar{B}}(\mathbf{1})$ $=\bar{B}^{2}\big[(A+t)^{2}-A^{2}\big]+\big[x(\gamma)^{\top}\Sigma x(\gamma)-x(1)^{\top}\Sigma x(1)\big]$
-$\geq\bar{B}^{2}\big(2A\,t+t^{2}\big)-4\,\|\Sigma\|_{\infty}\,t\big(\|\omega\|_{1}+t\big)$
-$=(\,\bar{B}^{2}-4\|\Sigma\|_{\infty}\,)t^{2}+(2A\bar{B}^{2}-4\|\Sigma\|_{\infty}\|\omega\|_{1}\,)t,$
-where $t(\gamma)\geq 0$.
-If $\bar{B}^{2}\geq\max\{\,4\|\Sigma\|_{\infty},\ 2\,\|\Sigma\|_{\infty}\,\|\omega\|_{1}/A\,\},$ then both coefficients are nonnegative and the right-hand side is $\geq 0$, with equality iff $t=0$, i.e., $\gamma=\mathbf{1}$.
-Thus $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)=\mathbf{1}$ for all such $\bar{B}$.
-(ii) Case $A=0$ (i.e., $\|\omega_{\mathcal{E}^{c}}\|_{1}=0$). Recall $t(\gamma)\equiv\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1-\gamma_{j}|\geq 0$ and
-$R_{\bar{B}}(\gamma)-R_{\bar{B}}(\mathbf{1})~\geq~\bar{B}^{2}\,t(\gamma)^{2}\;-\;4\,\|\Sigma(\mathcal{E})\|_{\infty}\,t(\gamma)\big(\|\omega\|_{1}+t(\gamma)\big),$
-which follows from (16) with $A=0$. Let $\gamma^{\star}=\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)$ be any minimizer, so $R_{\bar{B}}(\gamma^{\star})\leq R_{\bar{B}}(\mathbf{1})$ and hence
-$0\;\geq\;R_{\bar{B}}(\gamma^{\star})-R_{\bar{B}}(\mathbf{1})\;\geq\;\bar{B}^{2}\,t_{\star}^{2}\;-\;4\,\|\Sigma(\mathcal{E})\|_{\infty}\,t_{\star}\big(\|\omega\|_{1}+t_{\star}\big),$
-where $t_{\star}(\gamma^{\star})\geq 0$. Rearranging gives
-$\big(\bar{B}^{2}-4\|\Sigma(\mathcal{E})\|_{\infty}\big)\,t_{\star}^{2}\;\leq\;4\,\|\Sigma(\mathcal{E})\|_{\infty}\,\|\omega\|_{1}\,t_{\star}.$
-If $\bar{B}^{2}>4\|\Sigma(\mathcal{E})\|_{\infty}$, we can divide both sides by $B^{2}-4\|\Sigma(\mathcal{E})\|_{\infty}$ and obtain
-$t_{\star}\;\leq\;\frac{\,4\,\|\Sigma(\mathcal{E})\|_{\infty}\,\|\omega\|_{1}\,}{\bar{B}^{2}-4\|\Sigma(\mathcal{E})\|_{\infty}}\;=\;O\!\left(\frac{1}{\bar{B}^{2}}\right).$
-Hence $t(\gamma^{\star})\to 0$ as $\bar{B}\to\infty$.
-Finally, since $t(\gamma^{\star})=\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1-\gamma^{\star}_{j}|$ and each summand is nonnegative, for any coordinate with $|\omega_{j}|>0$ we have
-$|1-\gamma^{\star}_{j}|\;\leq\;\frac{t(\gamma^{\star})}{|\omega_{j}|}\;\longrightarrow\;0\quad\text{as }\bar{B}\to\infty.$
-Thus $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)\to\mathbf{1}$ as $\bar{B}\rightarrow\infty$. ∎
-The lemma below characterizes the behavior of the implicitly-defined functions of $\bar{B}$: $\bar{B}\mapsto\beta\left(\mathcal{E},\gamma^{\star}\left(\bar{B},\mathcal{E},\Sigma\right)\right)$ and $\bar{B}\mapsto\alpha\left(\mathcal{E},\gamma^{\star}\left(\bar{B},\mathcal{E},\Sigma\right)\right)$.
-Lemma 3.
-Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$. Let
-$\alpha(\mathcal{E},\Sigma,\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\equiv\Big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\Big)^{2},$
-and for $\bar{B}\geq 0$ define $R_{\bar{B}}(\gamma)\equiv\alpha(\mathcal{E},\Sigma,\gamma)+\bar{B}^{2}\,\beta(\mathcal{E},\gamma)$.
-Let $\gamma^{\star}(\bar{B})\in\arg\min_{\gamma}R_{\bar{B}}(\gamma)$.
-Then $\bar{B}\mapsto\beta\!\left(\mathcal{E},\gamma^{\star}(\bar{B})\right)$ is non-increasing and $\bar{B}\mapsto\alpha\!\left(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\right)$ is non-decreasing.
-Proof.
-Fix $0\leq\bar{B}_{1}<\bar{B}_{2}$ and choose minimizers $\gamma_{1}\in\arg\min_{\gamma}R_{\bar{B}_{1}}(\gamma)$ and $\gamma_{2}\in\arg\min_{\gamma}R_{\bar{B}_{2}}(\gamma)$. By optimality,
-$\alpha(\gamma_{1})+\bar{B}_{1}^{2}\,\beta(\gamma_{1})$ $\leq\alpha(\gamma_{2})+\bar{B}_{1}^{2}\,\beta(\gamma_{2}),$ (16)
-$\alpha(\gamma_{2})+\bar{B}_{2}^{2}\,\beta(\gamma_{2})$ $\leq\alpha(\gamma_{1})+\bar{B}_{2}^{2}\,\beta(\gamma_{1}).$ (17)
-Subtracting $\alpha(\gamma_{2})$ from (16) and $\alpha(\gamma_{1})$ from (17) yields
-$\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\leq\;\bar{B}_{1}^{2}\big(\beta(\gamma_{2})-\beta(\gamma_{1})\big),\qquad\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\geq\;\bar{B}_{2}^{2}\big(\beta(\gamma_{2})-\beta(\gamma_{1})\big).$
-Let $\Delta_{\beta}\equiv\beta(\gamma_{2})-\beta(\gamma_{1})$. Then
-$\bar{B}_{2}^{2}\,\Delta_{\beta}\;\leq\;\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\leq\;\bar{B}_{1}^{2}\,\Delta_{\beta}.$
-If $\Delta_{\beta}>0$, the last display implies $\bar{B}_{2}^{2}\leq\bar{B}_{1}^{2}$, a contradiction. Hence $\Delta_{\beta}\leq 0$, i.e.
-$\beta(\gamma_{2})\leq\beta(\gamma_{1}),$
-so $\bar{B}\mapsto\beta(\gamma^{\star}(\bar{B}))$ is non-increasing. Plugging $\Delta_{\beta}\leq 0$ into the first inequality gives
-$\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\leq\;\bar{B}_{1}^{2}\,\Delta_{\beta}\;\leq\;0,$
-so $\alpha(\gamma_{1})\leq\alpha(\gamma_{2})$, i.e.
-$\bar{B}\mapsto\alpha(\gamma^{\star}(\bar{B}))$ is non-decreasing.
-Because the choice of minimizer at each $\bar{B}$ was arbitrary, the monotonicity holds for any $\bar{B}$.
-∎
-Proposition 1.
-Let $\Sigma(\mathcal{E})\succ 0$ with uniformly bounded entries and fix $\mathcal{E}$. Let $|\omega_{j}|>0$ for all $j$. Let
-$\alpha(\mathcal{E},\Sigma,\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\equiv\Big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\Big)^{2},$
-and $\gamma^{\star}(\bar{B})\in\arg\min_{\gamma}\{\alpha(\mathcal{E},\Sigma,\gamma)+\bar{B}^{2}\beta(\mathcal{E},\gamma)\}$. Then, as $\bar{B}\to\infty$,
-$\beta\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)\,\downarrow\,\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\big)^{2}\quad\text{and}\quad\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\,\uparrow\,\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1}),$
-both monotonically.
-Proof.
-Write $A\equiv\|\omega_{\mathcal{E}^{c}}\|_{1}$ and $t(\gamma)\equiv\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1-\gamma_{j}|\geq 0$, so $\beta(\mathcal{E},\gamma)=(A+t(\gamma))^{2}$ and $\beta(\mathcal{E},\mathbf{1})=A^{2}=(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1})^{2}$.
-Step 1 (monotonicity).
-By Lemma 3, $\bar{B}\mapsto\beta(\mathcal{E},\gamma^{\star}(\bar{B}))$ is non-increasing and $\bar{B}\mapsto\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing.
-Step 2 (limit of $\beta$).
-Since $t(\gamma)\geq 0$, we have $\beta(\mathcal{E},\gamma)\geq A^{2}$ for all $\gamma$, hence the non-increasing sequence $\beta(\mathcal{E},\gamma^{\star}(\bar{B}))$ is bounded below by $A^{2}$ and thus converges to some limit $\geq A^{2}$.
-By Lemma 2, $\gamma^{\star}(\bar{B})\to\mathbf{1}$ as $\bar{B}\to\infty$, hence $t(\gamma^{\star}(\bar{B}))\to 0$ and therefore
-$\lim_{\bar{B}\to\infty}\beta\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)=\lim_{\bar{B}\to\infty}\big(A+t(\gamma^{\star}(\bar{B}))\big)^{2}=A^{2}=(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1})^{2}.$
-By Step 1 the convergence is monotone (decreasing).
-Step 3 (limit of $\alpha$). Feasibility of $\gamma=\mathbf{1}$ implies, for all $\bar{B}$,
-$\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)+\bar{B}^{2}\,\beta\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)~\leq~\alpha\big(\mathcal{E},\Sigma,\mathbf{1}\big)+\bar{B}^{2}\,\beta\big(\mathcal{E},\mathbf{1}\big)=\alpha(\mathcal{E},\Sigma,\mathbf{1})+\bar{B}^{2}A^{2},$
-and since $\beta(\mathcal{E},\gamma^{\star}(\bar{B}))\geq A^{2}$, this yields
-$\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\ \leq\ \alpha\big(\mathcal{E},\Sigma,\mathbf{1}\big)\qquad\forall\,\bar{B}.$
-Thus $\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing and bounded above by $\alpha(\mathcal{E},\Sigma,\mathbf{1})$, so it converges to some limit $\leq\alpha(\mathcal{E},\Sigma,\mathbf{1})$.
-To identify the limit, we use the bound in Lemma 2
-$\big|\alpha(\mathcal{E},\Sigma,\gamma)-\alpha(\mathcal{E},\Sigma,\mathbf{1})\big|=\big|x(\gamma)^{\top}\Sigma x(\gamma)-x(\mathbf{1})^{\top}\Sigma x(\mathbf{1})\big|\leq 4\,\|\Sigma(\mathcal{E})\|_{\infty}\,t(\gamma)\big(\|\omega\|_{1}+t(\gamma)\big),$
-where $x(\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)$. Because $t(\gamma^{\star}(\bar{B}))\to 0$ from Step 2, the RHS tends to 0, so
-$\lim_{\bar{B}\to\infty}\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)=\alpha\big(\mathcal{E},\Sigma,\mathbf{1}\big)=\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1}).$
-By Step 1 the convergence is monotone (increasing). ∎
-B.1.2 The researcher’s problem
-To characterize the adaptation regret, it suffices to characterize
-$\sup_{\bar{B}\geq 0}\,\sup_{\|b\|_{\infty}\leq\bar{B}}\frac{\text{MSE}_{b}\left[\,\widehat{\tau}\left(\mathcal{E},\gamma\right)\,\right]}{R^{\star}_{o}\left(\bar{B}\right)}=\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\sup_{\bar{B}\geq 0}\frac{R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)}{R_{\bar{B}}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma_{o}^{\star}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime}))}.$
-Write
-$\rho(\gamma,\bar{B},\mathcal{E},\mathcal{E}^{\prime},\Sigma,\Sigma^{\prime})\equiv\frac{R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)}{R_{\bar{B}}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma_{o}^{\star}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime}))}.$
-Lemma 4.
-Let $\Sigma(\mathcal{E}),\Sigma^{\prime}(\mathcal{E}^{\prime})\succ 0$ have uniformly bounded entries and assume $|\omega_{j}|>0$ for all $j$.
-For any fixed $(\mathcal{E},\Sigma,\gamma)$ and $(\mathcal{E}^{\prime},\Sigma^{\prime})$, define
-$\alpha\equiv\alpha(\mathcal{E},\Sigma,\gamma),\quad\beta\equiv\beta(\mathcal{E},\gamma),\qquad\alpha^{\star}(\bar{B})\equiv\alpha(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})),\quad\beta^{\star}(\bar{B})\equiv\beta(\mathcal{E}^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})).$
-Let
-$\rho(\bar{B})\;\equiv\;\frac{\bar{B}^{2}\,\beta+\alpha}{\bar{B}^{2}\,\beta^{\star}(\bar{B})+\alpha^{\star}(\bar{B})}.$
-[…]
-Proof.
-By Proposition 1, along the oracle path $\bar{B}\mapsto\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})$ we have
-$\beta^{\star}(\bar{B})\ \downarrow\ \big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2},\qquad\alpha^{\star}(\bar{B})\ \uparrow\ \tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1})^{\top}\Sigma^{\prime}(\mathcal{E}^{\prime})\,\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1}),$
-hence the ratio
-$\psi(\bar{B})\;\equiv\;\frac{\beta^{\star}(\bar{B})}{\alpha^{\star}(\bar{B})}$
-is non-increasing in $\bar{B}$ (numerator $\downarrow$, denominator $\uparrow$, both nonnegative).
-Fix $0\leq\bar{B}_{1}<\bar{B}_{2}$. Using optimality of the oracle at each $\bar{B}$,
-$\alpha^{\star}(\bar{B}_{1})+\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})~\leq~\alpha^{\star}(\bar{B}_{2})+\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{2}),\qquad\alpha^{\star}(\bar{B}_{2})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})~\leq~\alpha^{\star}(\bar{B}_{1})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{1}).$
-Therefore
-$\rho(\bar{B}_{2})-\rho(\bar{B}_{1})$ $=\frac{(\bar{B}_{2}^{2}\beta+\alpha)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)-(\bar{B}_{1}^{2}\beta+\alpha)\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}$
-$\leq\frac{(\bar{B}_{2}^{2}\beta+\alpha)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)-(\bar{B}_{1}^{2}\beta+\alpha)\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}$
-$=\frac{(\bar{B}_{2}^{2}-\bar{B}_{1}^{2})\big[\alpha^{\star}(\bar{B}_{2})\,\beta-\beta^{\star}(\bar{B}_{2})\,\alpha\big]}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}$
-$=\frac{(\bar{B}_{2}^{2}-\bar{B}_{1}^{2})\,\alpha^{\star}(\bar{B}_{2})\,\alpha}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}\left(\frac{\beta}{\alpha}-\frac{\beta^{\star}(\bar{B}_{2})}{\alpha^{\star}(\bar{B}_{2})}\right).$
-Since all denominators are positive, the last display shows
-$\rho(\bar{B}_{2})-\rho(\bar{B}_{1})\ \leq\ 0\quad\text{whenever}\quad\frac{\beta}{\alpha}\ \leq\ \psi(\bar{B}_{2}).$
-A symmetric argument (now using $\alpha^{\star}(\bar{B}_{2})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})\leq\alpha^{\star}(\bar{B}_{1})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{1})$) yields
-$\rho(\bar{B}_{2})-\rho(\bar{B}_{1})\ \geq\ 0\quad\text{whenever}\quad\frac{\beta}{\alpha}\ \geq\ \psi(\bar{B}_{1}).$
-Define the threshold
-$\tilde{B}\;\equiv\;\sup\Big\{B\geq 0:\ \psi(B)\geq\tfrac{\beta}{\alpha}\Big\}\ \in\ [0,\infty],$
-[…]
-Completion of the proof.
-Fix $(\mathcal{E},\Sigma,\gamma)$ and write, for any $(\mathcal{E}^{\prime},\Sigma^{\prime})$,
-$\rho(\bar{B})\equiv\rho(\gamma,\bar{B},\mathcal{E},\mathcal{E}^{\prime},\Sigma,\Sigma^{\prime})=\frac{\bar{B}^{2}\,\beta(\mathcal{E},\gamma)+\alpha(\mathcal{E},\Sigma,\gamma)}{\bar{B}^{2}\,\beta^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})+\alpha^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})},$
-[…]
-$\sup_{\bar{B}\geq 0}\rho(\bar{B})=\max\!\Big\{\rho(0),\ \lim_{\bar{B}\to\infty}\rho(\bar{B})\Big\}.$
-Endpoint $\bar{B}=0$. Since $\rho(0)=\alpha(\mathcal{E},\Sigma,\gamma)\big/\alpha^{\star}(0;\mathcal{E}^{\prime},\Sigma^{\prime})$,
-$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\rho(0)=\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\inf_{\mathcal{E}^{\prime},\Sigma^{\prime}}\alpha^{\star}(0;\mathcal{E}^{\prime},\Sigma^{\prime})}.$
-[…]
-$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\rho(0)=\frac{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)}{\displaystyle\min_{\mathcal{E}^{\prime},\Sigma^{\prime}}\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{0})^{\top}\Sigma^{\prime}(\mathcal{E}^{\prime})\,\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{0})}.$
-Endpoint $\bar{B}\to\infty$.
-By Proposition 1, $\beta^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})\downarrow\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2}$ and $\alpha^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})\uparrow\alpha(\mathcal{E}^{\prime},\Sigma^{\prime},\mathbf{1})$.
-Hence, after diving the numerators and denominator by $B^{2}$
-$\lim_{\bar{B}\to\infty}\rho(\bar{B})=\frac{\beta(\mathcal{E},\gamma)}{\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2}}.$
-Maximizing the ratio over $(\mathcal{E}^{\prime},\Sigma^{\prime})$ at this endpoint amounts to minimizing the denominator. This completes the proof. ∎
-B.2 Two parameters examples: proof of Corollary 1
-By Theorem 1, for fixed $(\mathcal{E},\Sigma)$ the adaptation regret as a function of the shrinkage vector equals
-$\mathcal{R}(\gamma)\;=\;\max\!\Big\{\alpha(\mathcal{E},\Sigma,\gamma)/\alpha^{\star},\;\beta(\mathcal{E},\gamma)/\beta^{\star}\Big\},$
-with $\alpha^{\star},\beta^{\star}$ constants.
-Step 1. With independence and two parameters,
-$\alpha(\mathcal{E},\Sigma,\gamma)=C+\omega_{j}^{2}\!\Big(\gamma_{j}^{2}v_{j}^{2}+(1-\gamma_{j})^{2}\sigma_{j}^{2}\Big),$
-where $C$ does not depend on $\gamma_{j}$. Hence $\alpha(j,\gamma_{j})\equiv\alpha(\mathcal{E},\Sigma,\gamma)$ is a strictly convex quadratic in $\gamma_{j}$ with unique minimizer at
-$\gamma_{j}^{\mathrm{var}}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}.$
-Moreover, $\alpha(j,\gamma_{j})$ is strictly increasing on $\big(\gamma_{j}^{\mathrm{var}},1\big)$ and strictly decreasing on $\big[0,\gamma_{j}^{\mathrm{var}}\big)$.
-Write the worst–case bias component holding the other coordinate fixed as
-$\beta(j,\gamma_{j})\;=\;\Big(A_{j}+|\omega_{j}|\cdot|1-\gamma_{j}|\Big)^{2},$
-where $A_{j}\geq 0$ collects all terms not involving $\gamma_{j}$ (including the contribution from the other coordinate).
-On $[0,1]$, $|1-\gamma_{j}|=1-\gamma_{j}$, so $\beta(j,\gamma_{j})$ is strictly decreasing in $\gamma_{j}$ and convex with minimum at $\gamma_{j}=1$.
-Therefore, any minimizer satisfies
-$\gamma_{j}^{\star}\in\big[\gamma_{j}^{\mathrm{var}},\,1\big].$
-Step 2.
-On $\big(\gamma_{j}^{\mathrm{var}},1\big)$ the map $\gamma_{j}\mapsto\alpha(j,\gamma_{j})/\alpha^{\star}$ is strictly increasing, while $\gamma_{j}\mapsto\beta(j,\gamma_{j})/\beta^{\star}$ is strictly decreasing.
-Hence the function
-$f(\gamma_{j})\equiv\max\!\Big\{\alpha(j,\gamma_{j})/\alpha^{\star},\ \beta(j,\gamma_{j})/\beta^{\star}\Big\},\qquad\gamma_{j}\in\big[\gamma_{j}^{\mathrm{var}},1\big],$
-is minimized either (i) at a boundary point, or (ii) at the unique interior point where the two arguments are equal (by strict monotonicity, at most one intersection exists). Therefore,
-  • If $\alpha(j,\gamma_{j})/\alpha^{\star}<\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\beta(j,\gamma_{j})/\beta^{\star}$ on that interval.
-Since this term is strictly decreasing, the minimizer is the right boundary $\gamma_{j}^{\star}=1$.
-  • If $\alpha(j,\gamma_{j})/\alpha^{\star}>\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\alpha(j,\gamma_{j})/\alpha^{\star}$ on that interval.
-Since this term is strictly increasing there, the minimizer is the left boundary $\gamma_{j}^{\star}=\gamma_{j}^{\mathrm{var}}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
-  • Otherwise, by the intermediate value theorem and strict monotonicity of the two curves, there exists a unique $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$ such that
-    $\frac{\alpha(j,\gamma_{j})}{\alpha^{\star}}\;=\;\frac{\beta(j,\gamma_{j})}{\beta^{\star}}.$
-Appendix C Proofs of the extensions
-C.1 Proof of Theorem 2
-The proof of Theorem 2 follows similarly the one of Theorem 1 with minor modifications.
-Let
-$\bar{\alpha}(\mathcal{E},\Sigma,\gamma)\equiv\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},\qquad\bar{\beta}(\mathcal{E},\gamma)\equiv\|\tilde{\omega}(\mathcal{E},\gamma)\|_{1},$
-so that
-$|L_{B}(\mathcal{E},\Sigma,\gamma)|=2\{z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E},\Sigma,\gamma)+B\,\bar{\beta}(\mathcal{E},\gamma)\}.$
-For any comparison design $(\mathcal{E}^{\prime},\Sigma^{\prime})$, let the oracle choose
-$\gamma^{\star}(B,\mathcal{E}^{\prime},\Sigma^{\prime})\in\arg\min_{\gamma^{\prime}\in\mathbb{R}^{|\mathcal{E}^{\prime}|}}\Big\{z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})+B\,\bar{\beta}(\mathcal{E}^{\prime},\gamma^{\prime})\Big\},$
-and denote the induced oracle path
-$\bar{\alpha}^{\star}(B)\equiv\bar{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\star}(B,\mathcal{E}^{\prime},\Sigma^{\prime})),\qquad\bar{\beta}^{\star}(B)\equiv\bar{\beta}(\mathcal{E}^{\prime},\gamma^{\star}(B,\mathcal{E}^{\prime},\Sigma^{\prime})).$
-Define the ratio (the factor $2$ cancels)
-$\tilde{\rho}(B)\equiv\frac{z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E},\Sigma,\gamma)+B\,\bar{\beta}(\mathcal{E},\gamma)}{z_{1-\eta/2}\,\bar{\alpha}^{\star}(B)+B\,\bar{\beta}^{\star}(B)}.$
-Step 1: Endpoint reduction (CI analogue of Lemma 4).
-Write $a\equiv z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E},\Sigma,\gamma)$ and $b\equiv\bar{\beta}(\mathcal{E},\gamma)$, and $A(B)\equiv z_{1-\eta/2}\,\bar{\alpha}^{\star}(B)$, $B^{\star}(B)\equiv\bar{\beta}^{\star}(B)$.
-By Proposition 1 (applied with square roots), $\bar{\beta}^{\star}(B)$ is non-increasing in $B$, and $\bar{\alpha}^{\star}(B)$ is non-decreasing in $B$.
-Exactly as in Lemma 4 (replace $\alpha,\beta$ there with $\bar{\alpha},\bar{\beta}$ here, and use $x=B$ in place of $x=\bar{B}^{2}$), one verifies that
-$\tilde{\rho}^{\prime}(B)\ \text{is non-decreasing in $B$}.$
-Hence $\tilde{\rho}$ is first non-increasing and then non-decreasing, so its supremum is attained at the endpoints:
-$\sup_{B\geq 0}\tilde{\rho}(B)=\max\Big\{\tilde{\rho}(0),\ \lim_{B\to\infty}\tilde{\rho}(B)\Big\}.$ (18)
-Step 2: Evaluate the endpoints and optimize over comparison designs.
-At $B=0$,
-$\tilde{\rho}(0)=\frac{\bar{\alpha}(\mathcal{E},\Sigma,\gamma)}{\bar{\alpha}^{\star}(0)}.$
-Since the oracle at $B=0$ minimizes standard error, $\bar{\alpha}^{\star}(0)=\min_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})}\bar{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})=\sqrt{\alpha^{\star}},$ where $\alpha^{\star}$ is as in (5).
-Thus
-$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\tilde{\rho}(0)=\frac{\bar{\alpha}(\mathcal{E},\Sigma,\gamma)}{\sqrt{\alpha^{\star}}}=\left(\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}}\right)^{\!1/2}.$
-As $B\to\infty$, the BB–term dominates.
-By Proposition 1 (again in square–root form), $\lim_{B\to\infty}\bar{\beta}^{\star}(B)=\min_{(\mathcal{E}^{\prime},\gamma^{\prime})}\bar{\beta}(\mathcal{E}^{\prime},\gamma^{\prime})=\sqrt{\beta^{\star}},$ with $\beta^{\star}$ from (5).
-Therefore
-$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\ \lim_{B\to\infty}\tilde{\rho}(B)=\frac{\bar{\beta}(\mathcal{E},\gamma)}{\sqrt{\beta^{\star}}}=\left(\frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right)^{\!1/2}.$
-Step 3: Combine.
-Taking the supremum over comparison designs and using (18),
-$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)=\sup_{B\geq 0}\ \sup_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}=\max\!\left\{\left(\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}}\right)^{\!1/2},\ \left(\frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right)^{\!1/2}\right\},$
-which is exactly the statement of Theorem 2. ∎
-C.2 Proof of Theorem 3
-Define the recursive solution to the oracle’s problem in (9):
-$\gamma^{\star}_{o}\left(B,\mathcal{E},\Sigma\right)$ $=$ $\operatorname*{arg\,min}_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)$
-$\Sigma^{\star}_{o}\left(B,\mathcal{E}\right)$ $=$ $\operatorname*{arg\,min}_{\Sigma\in\mathcal{G}(\mathcal{E})}\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}\left(\mathcal{E},\Sigma,\gamma^{\star}_{o}\left(B,\mathcal{E},\Sigma\right)\right)$
-$\mathcal{E}^{\star}_{o}(B)$ $=$ $\operatorname*{arg\,min}_{\mathcal{E}\in\mathcal{S}}\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}\left(\mathcal{E},\Sigma^{\star}_{o}\left(B,\mathcal{E}\right),\gamma^{\star}_{o}\left(B,\mathcal{E},\Sigma\right)\right),$
-and let
-$\check{R}_{B}\left(\mathcal{E},\Sigma,\gamma\right)\equiv\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\beta}(\mathcal{E},\gamma)B^{2}+\check{\alpha}(\mathcal{E},\Sigma,\gamma),$
-with $\check{\beta},\check{\alpha}$ as defined in Equation (10).
-The proof mimics the proof of Theorem 1.
-We will refer to $\check{R}_{B}(\gamma)$ omitting its arguments $\mathcal{E},\Sigma$ whenever clear from the context.
-Lemma 5 (Convexity and strict convexity).
-Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$. Then the map
-$\gamma~\longmapsto~\check{R}_{\bar{B}}(\mathcal{E},\Sigma,\gamma)~=~\check{\alpha}(\mathcal{E},\Sigma,\gamma)\;+\;\bar{B}^{2}\,\check{\beta}(\mathcal{E},\gamma)$
-is convex in $\gamma$. Moreover, if for each $j\in\mathcal{E}$ there exists at least one $\ell\in\{1,\dots,L\}$ with $|\omega^{\ell}_{j}|>0$, then the map is strictly convex in $\gamma$.
-Proof.
-We show convexity of the two addends separately.
-1) Convexity (and strict convexity) of $\check{\alpha}(\mathcal{E},\Sigma,\gamma)$.
-For each $\ell$, define $D_{\ell}:=\mathrm{diag}(\omega^{\ell}_{\mathcal{E}})$ and
-$d_{\ell}~:=~\begin{bmatrix}\omega_{\mathcal{E}^{c}}\\ \omega^{\ell}_{\mathcal{E}}\\ 0\end{bmatrix},\qquad A_{\ell}~:=~\begin{bmatrix}0\\ -D_{\ell}\\ D_{\ell}\end{bmatrix}\ \in\mathbb{R}^{(p+|\mathcal{E}|)\times|\mathcal{E}|}.$
-Then, by the definition of $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)$,
-$\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\;=\;d_{\ell}\;+\;A_{\ell}\,\gamma^{\ell},$
-where $\gamma^{\ell}\in\mathbb{R}^{|\mathcal{E}|}$ is the row of $\gamma$ for target $\ell$. Using (10),
-$\check{\alpha}(\mathcal{E},\Sigma,\gamma)=\mathrm{tr}\!\left(\tilde{\omega}\,\Sigma\,\tilde{\omega}^{\top}\right)=\sum_{\ell=1}^{L}(d_{\ell}+A_{\ell}\gamma^{\ell})^{\top}\Sigma(\mathcal{E})\,(d_{\ell}+A_{\ell}\gamma^{\ell}).$
-Hence $\check{\alpha}$ is a (separable across $\ell$) quadratic form in $\gamma$ with Hessian
-$\nabla^{2}_{\gamma}\check{\alpha}(\mathcal{E},\Sigma,\gamma)=2\,\mathrm{blkdiag}\!\big(A_{1}^{\top}\Sigma(\mathcal{E})A_{1},\ldots,A_{L}^{\top}\Sigma(\mathcal{E})A_{L}\big)\ \succeq\ 0,$
-where $\mathrm{blkdiag}$ is a block-diagonal matrix with entries as above. It follows that $\check{\alpha}$ is convex.
-For strict convexity: if $\Sigma(\mathcal{E})\succ 0$ and $A_{\ell}$ has full column rank for at least one $\ell$ in every coordinate direction, then $\sum_{\ell}(A_{\ell}^{\top}\Sigma A_{\ell})\succ 0$ and thus $\check{\alpha}$ is strictly convex.
-Since $A_{\ell}=[0;-D_{\ell};D_{\ell}]$, $A_{\ell}u=0$ iff $D_{\ell}u=0$.
-Thus $A_{\ell}$ has full column rank iff $D_{\ell}$ is invertible, i.e.
-iff $|\omega^{\ell}_{j}|>0$ for all $j\in\mathcal{E}$.
-A sufficient condition ensuring strict convexity is: for each $j\in\mathcal{E}$ there exists at least one $\ell$ with $|\omega^{\ell}_{j}|>0$; then, for any nonzero $u\in\mathbb{R}^{|\mathcal{E}|}$ we can pick an $\ell$ with $u_{j}\neq 0$ and $|\omega^{\ell}_{j}|>0$, whence
-$u^{\top}\Big(\sum_{\ell}A_{\ell}^{\top}\Sigma A_{\ell}\Big)u=\sum_{\ell}(A_{\ell}u)^{\top}\Sigma(A_{\ell}u)\ \geq\ (A_{\ell}u)^{\top}\Sigma(A_{\ell}u)\ >\ 0.$
-Therefore $\check{\alpha}$ is strictly convex under the stated condition.
-2) Convexity of $\check{\beta}(\mathcal{E},\gamma)$.
-From (10) and the definition of $\bar{v}(\gamma)$,
-$\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}\ \sum_{\ell=1}^{L}\Big((\omega^{\ell})^{\top}v-(\omega^{\ell}_{\mathcal{E}})^{\top}v_{\mathcal{E}}+(1-\gamma^{\ell})^{\top}(\,v_{\mathcal{E}}\circ\omega_{\mathcal{E}}\,)\Big)^{\!2}.$
-For each fixed $v$, the inner expression is an affine function of $\gamma$:
-$(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\mathcal{E}})^{\top}v_{\mathcal{E}}+\underbrace{\mathbf{1}^{\top}(\,v_{\mathcal{E}}\circ\omega_{\mathcal{E}}\,)}_{\text{constant in }\gamma}\;-\;(\gamma^{\ell})^{\top}(\,v_{\mathcal{E}}\circ\omega_{\mathcal{E}}\,).$
-The square of an affine function is convex; a finite sum of convex functions is convex; and the pointwise supremum of convex functions is convex.
-Therefore $\check{\beta}(\mathcal{E},\gamma)$ is convex in $\gamma$.
-This concludes the proof.
-∎
-Lemma 6.
-Suppose $\Sigma(\mathcal{E})\succ 0$ with uniformly bounded entries and $|\omega^{\ell}_{j}|>0$ for all $j\in\{1,\dots,p\}$ and $\ell\in\{1,\dots,L\}$. Then, for fixed $(\mathcal{E},\Sigma)$,
-$\gamma^{\star}_{o}(B,\mathcal{E},\Sigma)\ \longrightarrow\ \mathbf{1}\qquad\text{as }B\to\infty,$
-where $\mathbf{1}$ denotes the $L\times|\mathcal{E}|$ matrix of ones.
-Proof.
-Recall that for fixed $(\mathcal{E},\Sigma)$ the oracle criterion is
-$\check{R}_{B}(\mathcal{E},\Sigma,\gamma)=\sup_{\|b\|_{\infty}\leq B}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\alpha}(\mathcal{E},\Sigma,\gamma)+B^{2}\,\check{\beta}(\mathcal{E},\gamma),$
-with (see (10))
-$\check{\alpha}(\mathcal{E},\Sigma,\gamma)=\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\gamma)\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\big),\qquad\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}\ \sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2},$
-where
-$a_{\ell}(v):=\sum_{j\in\mathcal{E}^{c}}\omega^{\ell}_{j}v_{j},\qquad\delta_{\ell}(\gamma,v):=\sum_{j\in\mathcal{E}}(1-\gamma^{\ell}_{j})\,\omega^{\ell}_{j}v_{j}.$
-Step 1: The bias term is uniquely minimized at $\gamma=\mathbf{1}$.
-For each $j\in\mathcal{E}$, define the $L$-vector
-$d_{j}(\gamma):=\big((1-\gamma_{j}^{1})\omega_{j}^{1},\ldots,(1-\gamma_{j}^{L})\omega_{j}^{L}\big)^{\top}\in\mathbb{R}^{L},\qquad S(\gamma):=\max_{j\in\mathcal{E}}\|d_{j}(\gamma)\|_{2}.$
-Choose $v\in[-1,1]^{p}$ with $v_{\mathcal{E}^{c}}=0$, $v_{j}=1$ for some fixed $j\in\mathcal{E}$, and $v_{j^{\prime}}=0$ for $j^{\prime}\in\mathcal{E}\setminus\{j\}$. Then
-$\sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2}=\sum_{\ell=1}^{L}\big(\delta_{\ell}(\gamma,v)\big)^{2}=\|d_{j}(\gamma)\|_{2}^{2}.$
-Taking the supremum over $v$ yields
-$\check{\beta}(\mathcal{E},\gamma)\ \geq\ \max_{j\in\mathcal{E}}\|d_{j}(\gamma)\|_{2}^{2}\;=\;S(\gamma)^{2}.$ (19)
-In particular, $\check{\beta}(\mathcal{E},\gamma)>\check{\beta}(\mathcal{E},\mathbf{1})$ whenever $\gamma\neq\mathbf{1}$ (since then $S(\gamma)>0$ by $|\omega^{\ell}_{j}|>0$).
-Thus $\gamma\mapsto\check{\beta}$ is uniquely minimized at $\gamma=\mathbf{1}$, with a quadratic gap at least $S(\gamma)^{2}$.
-Step 2: Variance bound.
-[…]
-$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma)-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\leq\|\Sigma(\mathcal{E})\|_{\infty}\sum_{\ell=1}^{L}\|x_{\ell}(\gamma)-y_{\ell}\|_{1}\,\|x_{\ell}(\gamma)+y_{\ell}\|_{1}.$
-A direct calculation gives
-$\|x_{\ell}(\gamma)-y_{\ell}\|_{1}=2\sum_{j\in\mathcal{E}}|(1-\gamma^{\ell}_{j})\omega^{\ell}_{j}|=2\,t_{\ell}(\gamma),$
-and, using $|1+\gamma^{\ell}_{j}|\leq|1-\gamma^{\ell}_{j}|+2$,
-$\|x_{\ell}(\gamma)+y_{\ell}\|_{1}\leq 2\|\omega^{\ell}\|_{1}+2\,t_{\ell}(\gamma).$
-Hence
-$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma)-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\leq 4\,\|\Sigma(\mathcal{E})\|_{\infty}\!\left(\sum_{\ell=1}^{L}t_{\ell}(\gamma)\,\|\omega^{\ell}\|_{1}+\sum_{\ell=1}^{L}t_{\ell}(\gamma)^{2}\right).$
-Let $d(\gamma)\in\mathbb{R}^{L\times|\mathcal{E}|}$ stack the $d_{j}(\gamma)$ as columns; then $t_{\ell}(\gamma)=\|d(\gamma)_{\ell,\cdot}\|_{1}$ and
-$\sum_{\ell=1}^{L}t_{\ell}(\gamma)=\sum_{j\in\mathcal{E}}\|d_{j}(\gamma)\|_{1}\leq|\mathcal{E}|\,\sqrt{L}\,S(\gamma),\qquad\sum_{\ell=1}^{L}t_{\ell}(\gamma)^{2}\leq\Big(\sum_{\ell=1}^{L}t_{\ell}(\gamma)\Big)^{2}\leq|\mathcal{E}|^{2}L\,S(\gamma)^{2}.$
-Therefore there exist finite constants $C_{1},C_{2}$ (depending only on $\Sigma,\omega,L,|\mathcal{E}|$) such that
-$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma)-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\ \leq\ C_{1}\,S(\gamma)\;+\;C_{2}\,S(\gamma)^{2}.$ (20)
-Step 3: convergence of the minimizer.
-Combine (19)–(20) to get, for any $\gamma$,
-$\check{R}_{B}(\mathcal{E},\Sigma,\gamma)-\check{R}_{B}(\mathcal{E},\Sigma,\mathbf{1})=\big[\check{\alpha}(\gamma)-\check{\alpha}(\mathbf{1})\big]+B^{2}\big[\check{\beta}(\gamma)-\check{\beta}(\mathbf{1})\big]\ \geq\ -\,C_{1}S(\gamma)+\big(B^{2}-C_{2}\big)S(\gamma)^{2}.$
-Fix $\varepsilon>0$ and choose $B$ such that $(B^{2}-C_{2})\varepsilon^{2}>C_{1}\varepsilon$.
-Then the RHS is >0>0 for all $\gamma$ with $S(\gamma)\geq\varepsilon$.
-Hence any minimizer $\gamma^{\star}_{o}(B)$ must satisfy $S(\gamma^{\star}_{o}(B))<\varepsilon$.
-Because $\varepsilon>0$ is arbitrary and $B\to\infty$, we obtain $S(\gamma^{\star}_{o}(B))\to 0$.
-Finally, $S(\gamma^{\star}_{o}(B))\to 0$ means $(1-\gamma^{\star,\ell}_{j}(B))\omega^{\ell}_{j}\to 0$ for every $j\in\mathcal{E}$ and $\ell\in\{1,\dots,L\}$.
-Since $|\omega^{\ell}_{j}|>0$, this implies $\gamma^{\star,\ell}_{j}(B)\to 1$ for all $j,\ell$, i.e.
-$\gamma^{\star}_{o}(B,\mathcal{E},\Sigma)\to\mathbf{1}$ componentwise.
-∎
-Lemma 7.
-Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$. Let
-$\check{\alpha}(\mathcal{E},\Sigma,\gamma)=\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\gamma)\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\big),\qquad\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}\ \sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2},$
-where
-$a_{\ell}(v):=\sum_{j\in\mathcal{E}^{c}}\omega^{\ell}_{j}v_{j},\qquad\delta_{\ell}(\gamma,v):=\sum_{j\in\mathcal{E}}(1-\gamma^{\ell}_{j})\,\omega^{\ell}_{j}v_{j},$
-and for $\bar{B}\geq 0$ define $R_{\bar{B}}(\gamma)\equiv\alpha(\mathcal{E},\Sigma,\gamma)+\bar{B}^{2}\,\beta(\mathcal{E},\gamma)$.
-Let $\gamma^{\star}(\bar{B})\in\arg\min_{\gamma}R_{\bar{B}}(\gamma)$.
-Then $\bar{B}\mapsto\beta\!\left(\mathcal{E},\gamma^{\star}(\bar{B})\right)$ is non-increasing and $\bar{B}\mapsto\alpha\!\left(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\right)$ is non-decreasing.
-Proof.
-The proof is identical to that of Lemma 3, after replacing $\alpha$ and $\beta$ with $\check{\alpha}$ and $\check{\beta}$, respectively. ∎
-Proposition 2.
-Let $\Sigma(\mathcal{E})\succ 0$ with uniformly bounded entries and fix $\mathcal{E}$.
-Let $|\omega_{j}|>0$ for all $j$.
-Define $\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ and $\check{\beta}(\mathcal{E},\gamma)$ as in Lemma.
-Then, as $\bar{B}\to\infty$,
-$\check{\beta}\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)\,\downarrow\,\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\big)^{2}\quad\text{and}\quad\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\,\uparrow\,\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\mathbf{1})\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\big),$
-both monotonically.
-Proof.
-Write $A\equiv\|\omega_{\mathcal{E}^{c}}\|_{1}$ and, following the notation in Lemma 7, define $g(\mathcal{E},\gamma,v)=\sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2}$, so $\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}g(\mathcal{E},\gamma,v)$.
-Step 1 (monotonicity).
-By Lemma 7, $\bar{B}\mapsto\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))$ is non-increasing and $\bar{B}\mapsto\check{\alpha}(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing.
-Step 2 (limit of $\beta$).
-Letting $\mathbf{1}_{\mathcal{E}}$ denote the $p\times 1$ vector with $j^{th}$ entry $1\{j\notin\mathcal{E}\}$, we have $\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))\geq g(\mathcal{E},\gamma^{\star}(\bar{B}),\mathbf{1}_{\mathcal{E}})=A^{2}$, hence the non-increasing sequence $\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))$ is bounded below by $A^{2}$ and thus converges to some limit $\geq A^{2}$.
-Because $v\mapsto g(\mathcal{E},\gamma,v)$ is convex, $\check{\beta}(\mathcal{E},\gamma)=\max_{v\in V}g(\mathcal{E},\gamma,v)$, where $V=\{-1,1\}^{p}$.
-Additionally, Lemma 6 implies $\lim_{\bar{B}\rightarrow\infty}g(\mathcal{E},\gamma^{\star}(\bar{B}),v)=g(\mathcal{E},\mathbf{1},v)$ $\forall v\in V$, since $\gamma\mapsto g(\mathcal{E},\gamma,v)$ is continuous.
-Therefore, since $V$ is finite,
-$|\max_{v\in V}g(\mathcal{E},\gamma^{\star}(\bar{B}),v)-\max_{v\in V}g(\mathcal{E},\mathbf{1},v)|\leq\max_{v\in V}|g(\mathcal{E},\gamma^{\star}(\bar{B}),v)-g(\mathcal{E},\mathbf{1},v)|\rightarrow 0\text{ as $\bar{B}\rightarrow\infty$.}$
-We conclude that $\lim_{\bar{B}\rightarrow\infty}\check{\beta}\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)=\max_{v\in V}g(\mathcal{E},\mathbf{1},v)=A^{2}$.
-By Step 1, the convergence is monotone (decreasing).
-Step 3 (limit of $\alpha$). Feasibility of $\gamma=\mathbf{1}$ implies, for all $\bar{B}$,
-$\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)+\bar{B}^{2}\,\check{\beta}\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)~\leq~\check{\alpha}\big(\mathcal{E},\Sigma,\mathbf{1}\big)+\bar{B}^{2}\,\check{\beta}\big(\mathcal{E},\mathbf{1}\big)=\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})+\bar{B}^{2}A^{2},$
-and since $\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))\geq A^{2}$, this yields
-$\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\ \leq\ \check{\alpha}\big(\mathcal{E},\Sigma,\mathbf{1}\big)\qquad\forall\,\bar{B}.$
-Thus $\check{\alpha}(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing and bounded above by $\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})$, so it converges to some limit $\leq\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})$.
-To identify the limit, we use the bound in Lemma 6
-$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\leq C_{1}\,S(\gamma^{\star}(\bar{B}))+C_{2}\,S(\gamma^{\star}(\bar{B}))^{2}\rightarrow C_{1}\,S(\mathbf{1})+C_{2}\,S(\mathbf{1})^{2}=0,$
-as $\bar{B}\rightarrow\infty$. Therefore,
-$\lim_{\bar{B}\to\infty}\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)=\check{\alpha}\big(\mathcal{E},\Sigma,\mathbf{1}\big)=\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\mathbf{1})\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\big).$
-By Step 1, the convergence is monotone (increasing). ∎
-Lemma 8.
-Let $\Sigma(\mathcal{E}),\Sigma^{\prime}(\mathcal{E}^{\prime})\succ 0$ have uniformly bounded entries and assume $|\omega_{j}^{\ell}|>0$ for all $j,\ell$.
-For any fixed $(\mathcal{E},\Sigma,\gamma)$ and $(\mathcal{E}^{\prime},\Sigma^{\prime})$, define
-$\check{\alpha}\equiv\check{\alpha}(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}\equiv\check{\beta}(\mathcal{E},\gamma),\qquad\check{\alpha}^{\star}(\bar{B})\equiv\check{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})),\quad\check{\beta}^{\star}(\bar{B})\equiv\check{\beta}(\mathcal{E}^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})).$
-Let
-$\check{\rho}(\bar{B})\;\equiv\;\frac{\bar{B}^{2}\,\check{\beta}+\check{\alpha}}{\bar{B}^{2}\,\check{\beta}^{\star}(\bar{B})+\check{\alpha}^{\star}(\bar{B})}.$
-[…]
-Proof.
-By Proposition 2, along the oracle path $\bar{B}\mapsto\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})$ we have
-$\check{\beta}^{\star}(\bar{B})\ \downarrow\ \big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2},\qquad\check{\alpha}^{\star}(\bar{B})\ \uparrow\ \mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1})\,\Sigma(\mathcal{E}^{\prime})\,\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1})^{\top}\big),$
-hence the ratio
-$\check{\psi}(\bar{B})\;\equiv\;\frac{\check{\beta}^{\star}(\bar{B})}{\check{\alpha}^{\star}(\bar{B})}$
-is non-increasing in $\bar{B}$ (numerator $\downarrow$, denominator $\uparrow$, both nonnegative).
-The rest of the proof is identical to that of Lemma 4, after replacing $\alpha$, $\beta$, $\alpha^{\star}$, $\beta^{\star}$, $\rho$, and $\psi$ with $\check{\alpha}$, $\check{\beta}$, $\check{\alpha}^{\star}$, $\check{\beta}^{\star}$, $\check{\rho}$, and $\check{\psi}$, respectively.
-∎
-The completion of the proof is a analogous to that of B.1, after replacing $\alpha$, $\beta$, $\alpha^{\star}$, $\beta^{\star}$, and $\rho$ with $\check{\alpha}$, $\check{\beta}$, $\check{\alpha}^{\star}$, $\check{\beta}^{\star}$, and $\check{\rho}$, respectively.
-∎

### v3

@@
 Abstract
-Experiments deliver credible treatment effects estimates, but are often localized to specific sites, populations, or mechanisms.
-When such estimates are insufficient to extrapolate effects for broader policy questions as for external validity and general-equilibrium (GE) effects, researchers combine trials with external evidence from reduced-form or structural observational estimates, or prior experiments.
-We develop a unified framework for designing experiments in this setting: the researcher selects which parameters (or moments) to identify experimentally from a feasible set of possible experiments, allocates sample size, and specifies how to weight experimental and observational estimators.
-Because observational inputs may be biased in ways unknown ex ante, we adopt a minimax proportional-regret objective that evaluates any candidate design relative to an oracle that knows the bias and jointly chooses the design and estimator.
-This yields a transparent bias-variance trade-off that requires no prespecified bias bound and depends only on information about the precision of the estimators and the estimand’s sensitivity to underlying parameters.
-We illustrate the framework by (i) designing small-scale cash-transfer experiments aimed at estimating GE effects and (ii) optimizing site selection for microfinance interventions.
+Experiments deliver credible treatment-effect estimates but, because they are costly, are often restricted to specific sites, small populations, or particular mechanisms.
+A common practice across several fields is therefore to combine experimental estimates with reduced-form or structural external (observational) evidence to answer broader policy questions such as those involving general equilibrium effects or external validity.
+We develop a unified framework for the design of experiments when combined with external evidence, i.e., choosing which experiment(s) to run and how to allocate sample size under arbitrary budget constraints.
+Because observational evidence may suffer bias unknown ex-ante, we evaluate designs using a minimax proportional-regret criterion that compares any candidate design to an oracle that knows the observational study bias and jointly chooses the design and estimator.
+This yields a transparent bias-variance trade-off that does not require the researcher to specify a bias bound and relies only on information already needed for conventional power calculations.
+We illustrate the framework by (i) designing cash-transfer experiments aimed at estimating general equilibrium effects and (ii) optimizing site selection for microfinance interventions.
 1 Introduction
-Randomized controlled trials (RCTs) have improved empirical economics by providing internally valid causal estimates.
-However, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism.
-These effects, although useful, are often not sufficient to answer broader questions about external validity and (general) equilibrium effects of at-scale interventions.
-For instance, the median evaluation of experiments in top economics journals is “representative of a population of 10,885 units, studied a treatment delivered to 5,340 units, and randomized in clusters of 26 units per cluster” (Muralidharan and Niehaus, 2017).
-In response, a growing literature in development economics (e.g., Meghir et al., 2022; Bassi et al., 2022; Attanasio et al., 2012; de Albuquerque et al., 2025), education (e.g., Allende et al., 2019; Larroucau et al., 2024), labor economics (e.g., Chetty et al., 2016), and industrial organization (e.g., Katz and Allcott, 2025; Allcott et al., 2025) combine experiments with external evidence—observational estimates and/or results from experiments in other contexts—to estimate complex counterfactuals that no single experiment can identify.
-In our review of AEA journals over the past decade, more than 30% of experimental papers complement RCT evidence with at least one observational input, either reduced-form or structural (Figure 12).
-This trend poses a design question: given constraints on feasible experiments, which experiments should be run (and how) to estimate such counterfactuals?
-Our main contribution is a framework for designing experiments combined with external evidence.
-By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments conducted in other contexts.
-We provide a procedure that selects which experiments to run subject to a budget (e.g., which treatment arms and/or sub-populations to include), sets their precision via sample allocation, and prescribes how to combine experimental and external evidence.
-External input may be biased in ways not known ex ante (e.g., due to confounding or lack of external validity).
+Randomized controlled trials (RCTs) have improved empirical economics by providing internally valid estimates.
+However, in practice, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism, as Muralidharan and Niehaus (2017) document.
+[^4] Experiments in top economics journals are in median “representative of a population of 10,885 units […] and randomized in clusters of 26 units per cluster” (Muralidharan and Niehaus, 2017).These effects, although useful, are often not sufficient to answer relevant policy questions about external validity and general equilibrium (GE) effects of at-scale interventions.
+In response, a large literature in development (e.g., Meghir et al., 2022; Todd and Wolpin, 2006; Attanasio et al., 2012; de Albuquerque et al., 2025), education and labor economics (e.g., Allende et al., 2019; Larroucau et al., 2024; Chetty et al., 2016), and more recently industrial organization (e.g., Katz and Allcott, 2025; Allcott et al., 2025) combines experiments with external evidence—observational and/or results from experiments in other settings—to extrapolate policy-relevant counterfactuals that no single experiment can identify.
+In our review of AEA journals over the past decade, about one-third of experimental papers complement experimental with observational inputs, either reduced-form or structural (Figure 11).
+Since budget constraints often bind, such common practice raises the question of how to choose the most informative experiment(s) and design for policy-relevant estimands from a (typical large) menu of possible experiments (e.g.
+Ravallion, 2012; Niederle, 2025).
+Our main contribution is a framework for choosing which experiment to run and how to design it in this common setting of combining experimental with external evidence.
+By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments in other contexts.
+External evidence may be biased in ways not known ex ante (e.g., due to confounding or lack of external validity).
+We provide a procedure that (1) helps researchers disentangle the role of each experiment in this setting, (2) selects which experiments to run subject to a budget (e.g., which treatment arms and/or sub-populations), (3) optimally allocates sample size, and (4) prescribes how to combine experimental and external evidence for estimation.
+We allow for arbitrary (user-specific) constraints in each step, for both the choice of the design and of the estimator.
+[^5] For (4), researchers may constrain the estimator to either only use experimental evidence when available or optimally combine experimental with (biased) observational evidence for estimating certain parameters.
 For an illustrative example, consider a government piloting a cash-transfer program in a small set of districts to increase children’s school attendance.
@@
 Researchers therefore combine experimental evidence with a supply-demand model that leverages observational data to map local impacts into economy-wide outcomes.
-The design question is twofold: (i) which experiment to run (i.e., which parameters/effects are the most valuable to learn experimentally) given feasibility constraints and (ii) how to allocate sample across sites and arms so that, once combined with external inputs, we precisely estimate the effect at scale.
+The design question is (i) which mechanisms are the most valuable to learn experimentally given constraints on the size of the experiment and (ii) how to allocate sample size across treatment arms and sites.
 The first step is to specify the estimand and the underlying parameters.
 Define $\tau(\theta)$ the (policy-relevant) object of interest, where $\tau$ is a known smooth function and $\theta$ is a vector of unknown parameters; $\tau$ may be scalar or vector-valued.
-Researchers also observe external baseline estimates of $\theta$—henceforth observational estimates—which may be biased.
+Researchers also observe external baseline estimates of $\theta$–henceforth observational estimates–which may be biased.
 Each feasible experiment can identify a subset of parameters without bias.
-This parameterization makes explicit which components of $\theta$ are learned experimentally and which may rely on biased observational evidence; not all components need (or can) be learned experimentally.
+This parameterization makes explicit which components of $\theta$ are learned experimentally and which may rely on (biased) observational evidence; not all components need or can be learned experimentally.
 Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in rural Kenya receive the subsidy.
-The parameter vector $\theta$ bundles (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium , and (iii) wage adjustments that shift the returns to schooling.
+The entries of $\theta$ consist of (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium, and (iii) wage adjustments that shift the returns to schooling.
 Due to cost constraints, researchers may only run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$.
-They may choose between a CCT treatment arm to estimate the direct effect, or an unconditional-cash arm (UCT) to measure the income elasticity, and extrapolate the GE effects using evidence from a previous study in Mexico (Todd and Wolpin, 2006).
+They may choose between a CCT arm to estimate the direct effect, or an unconditional-cash arm (UCT) to measure the income elasticity, and extrapolate the GE effects using evidence from a previous study in Mexico (Todd and Wolpin, 2006).
 (When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ targets the average effect across sites.)
-The second step is to design the experiment: (i) selecting which experiment to run; (ii) allocating sample sizes across arms; and (iii) choosing how to combine observational and experimental evidence.
-We allow flexibility in (iii): researchers may choose to either optimally combine both sources when available or rely solely on experimental estimates.
-Parameters not identified experimentally are based on observational estimates.
-A natural starting point is to design the experiment to minimize the mean-squared error (MSE) for estimating $\tau(\theta)$, perhaps worst-case over the observational bias.
-In practice, however, the size and even the direction of bias are unknown (before the experiment).
-A worst-case approach would therefore be overly-conservative and disregard information about the variance.
-We instead adapt the definition of proportional (adaptation) regret previously studied when combining two estimators generated by a fixed design (Armstrong et al., 2024; Tsybakov, 1998) and generalize it to the experimental design problem (with multiple estimators).
-Here, the regret is the ratio of the worst-case researcher’s mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the estimator.
-Taking its supremum over any value of the bias yields a procedure that is robust to any level of misspecification without requiring priors about the bias.
-Our main result gives an explicit novel characterization of proportional regret.
-For any design and weights that combine the experimental and observational point estimates, the regret is the maximum of two normalized components.
+The second step is to select the experiment(s), sample size and estimator.
+A natural starting point would be to design the experiment to minimize the mean-squared error (MSE) for $\tau(\theta)$.
+However, in practice, the bias is unknown prior to the experiment.
+A worst-case approach would be overly-conservative and disregard information about the variance.
+We instead adapt the definition of adaptation regret, previously studied for combining two estimators under a fixed design (Armstrong et al., 2024; Tsybakov, 1998), to our experimental design problem (with multiple parameters).
+Here, the regret is the ratio of the researcher’s worst-case mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the estimator.
+Taking its supremum over the worst-case bias yields a robust procedure without requiring a prior about the bias or on its bound.
+However, our task of also choosing the experimental design and the lack of an unbiased estimate for $\tau(\theta)$ when not all parameters can be learned experimentally make existing solutions not applicable.
+Our main result provides an explicit and novel characterization of adaptation regret for (asymptotically) linear estimators in experimental and observational moments.
+These include generalized methods of moments where we may also optimize over the weighting matrix.
+For any design and weighting matrix that combine the experimental and observational point estimates, the regret is the maximum of two normalized components.
 The first is a variance regret: the estimator’s sampling variance under the chosen design and weights, divided by the smallest attainable variance.
 The second is a bias regret: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias over the class of designs.
-The variance regret depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the shrinkage weights, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters.
-This gradient is evaluated, under mild conditions, at the observational estimates.
+The variance regret depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the weighting matrix, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters evaluated, under mild conditions, at the observational estimates.
 [^6] For this result to hold, we assume that $\tau(\theta)$ is twice differentiable with bounded derivatives.
@@
 Section 5.3 provides details.
-The bias regret depends only on the sensitivity vector and the shrinkage weights, known ex-ante.
-This characterization makes the bias-variance trade-off transparent.
-At the optimum, the two normalized components tend to be equalized: when the bias component dominates, the design prioritizes high-sensitivity parameters; when variance dominates, the procedure invests sample where it most reduces variance and relies more on the most precise observational inputs.
-This yields a nested procedure that chooses jointly the estimator, the sample allocation and the experiment to run, trading off precision with sensitivity of the policy estimand.
-The resulting design and estimator only require knowledge of the expected variance-covariance matrix (standard in experimental design problems, Gerber and Green, 2012) and of the observational estimates.
-It can then be reported in a pre-analysis plan.
-That is, the inputs for the design are the same that would be required by standard power-analysis plans.
-From a theoretical perspective, our result leverages a key insight.
-Because here bias arise from multiple parameters, the choice of the uncertainty set is pivotal.
-While we do not require researchers to specify such a set, regret is with respect to an oracle that knows the worst-case bias within an $\ell_{\infty}$-ball.
-Compared with $\ell_{1}$ or $\ell_{2}$, the $\ell_{\infty}$-ball prevents that a large bias in one component is “offset” by smaller biases elsewhere.
-This is attractive when biases may be positively correlated across parameters.
-We show that the $\ell_{\infty}$ choice combined with our design problem (with decisions made before experimental data is realized), yields a quasi-convex regret in the bias radius as the oracle optimizes over the designs.
-This result, worst-case over the radius, delivers a transparent and easy-to-compute variance-bias trade-off.
-We then extend the model to general method-of-moments estimators that combine experimentally identified moments with observational moments.
-As in the baseline case, observational moments may incur unknown misspecification.
-We optimize (i) which moments to identify experimentally and (ii) how to combine experimental and observational moments to estimate the counterfactual of interest.
-Our main result continues to hold in this more general scenario provided that the moment estimators are (asymptotically) linear in the moments.
-Linearity occurs (under smoothness conditions) as long as the ratio of the bias to standard error is no faster than $n^{1/4}$, with $n$ denoting the observational sample size; standard local-asymptotics with the bias and standard error of the same order is a special case.
-We illustrate the framework in the cash‐transfer application, when the researcher uses evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006; Attanasio et al., 2012).
-We are concerned that these preliminary estimates may lack external validity in Kenya.
-We compare three designs: (i) a CCT arm to estimate the direct schooling effect; (ii) an unconditional cash transfer (UCT) arm to estimate income effects; and (iii) a two-arm design that allocates sample across CCT and UCT under a common budget.
-Running a single arm captures settings with fixed costs per arm; allowing both arms with budgeted allocation captures cases with non-binding fixed costs but binding variable costs.
-When both arms are feasible, the regret-optimal allocation puts most participants in the CCT and a small fraction in the UCT, reflecting the target’s greater sensitivity to the direct effect despite the CCT’s higher variance.
-If only one arm can be run, the choice hinges on sample size: for very small $n$, the lower-variance of the UCT arm makes this preferred over a (too noisy) CCT arm; once $n\geq 500$, the more informative but noisier CCT yields lower regret.
-Relative to an oracle that knows the bias, the optimal design closely tracks the oracle.
-We provides a step-by-step guide for practitioners applicable in other contexts.
-As a second application, we design where to run a microfinance experiment in rural India, integrating evidence from earlier nonrandomized microfinance introductions.
-We start from the observational estimates in Banerjee et al.
-(2024) and design an experiment that selects one or more areas for randomization.
+The bias regret depends only on the sensitivity vector and the weighting matrix, known ex-ante.
+Our characterization makes the bias-variance trade-off intuitive.
+At the optimum, the two normalized components tend to be equalized: when the bias dominates, the design prioritizes high-sensitivity parameters; when variance dominates, the procedure invests sample where it most reduces variance.
+This yields a nested procedure that jointly selects the estimator, the sample allocation and the experiment(s) to run, trading off precision with bias sensitivity.
+The resulting design and estimator only require knowledge of the expected variance-covariance matrix and observational estimates.
+It can be reported in a pre-analysis plan.
+That is, for given observational estimates, the inputs are the same as those required by standard power-analysis plans for $\tau(\theta)$ (e.g.
+Gerber and Green, 2012).
+We then generalize our framework to general moment selection problems that combine experimentally identified moments with observational (misspecified) moments.
+With parameters defined by non-linear but smooth moments our results directly apply under standard local asymptotics, where the bias is local to zero at same or possibly different rate than the standard error (e.g.
+Andrews et al., 2017, 2020).
+We also generalize to partially identified $\tau$ and settings where researchers have priors (confidence sets) about the bias.
+From a theoretical (and computational) perspective, our key insight is that, as we restrict the class of estimators to be (asymptotically) linear in the moments, the worst-case adaptation regret is quasi-convex in the radius of the bias, for any norm characterizing the bias’ uncertainty set.
+This result yields a transparent variance-bias trade-off and differs from existing solutions under the fixed-design optimal estimators as in Armstrong et al.
+(2024), Tsybakov (1998); there, quasi-convexity fails due to non-linearity of the estimator, and optimization corresponds to a minimax program over worst-case priors.
+In our joint design and estimation problem, quasi-convexity substantially simplifies optimization which is quadratic for the choice of the estimator, with a closed form expression for sample allocation and a mixed-integer formulation for the experiment choice only (typically from a finite set).
+This program is solved via off-the-shelf mixed-integer quadratic programming.
+We illustrate the framework in a cash‐transfer application in Kenya, where the researcher combines new experimental evidence with external evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006; Attanasio et al., 2012), where these preliminary estimates may lack external validity.
+We compare three designs and their combinations: (i) a CCT arm to estimate direct schooling effects; (ii) a UCT arm to estimate income effects; and (iii) an employment program that identifies individual wage elasticities.
+Experimentation costs are binding, as cash transfers and employment programs require distinct infrastructures and implementation teams.
+We show that the wage experiment is the most informative for sufficiently large sample sizes, while for smaller samples a combination of CCT and UCT—allocating most of the sample to the CCT arm—is optimal.
+With one thousand participants, relative to the variance-minimizing (Neyman) design, our optimal two-arm design delivers more than a 400% reduction in bias (and adaptation regret) with variance no larger than $\sim$30% of the Neyman variance; for a single-arm design, our procedure yields a 140% reduction in bias at the cost of only about a 12% increase in variance.
+These comparisons illustrate the benefit of our method in balancing robustness and precision.
+As a second application, we study where to run a microfinance experiment in India, integrating evidence from earlier nonrandomized microfinance introductions.
+We target the average effect in the entire region, but constraint the researcher to choose one or two smaller areas for randomization due to coordination (fixed) costs constraints.
+We use observational estimates in Banerjee et al.
+(2024) to select one or more areas for randomization.
 Because Banerjee et al.
-(2024) also implement a separate randomized expansion, we can calibrate performance of our design (that is agnostic of the bias) by comparing each candidate design’s MSE to that of an oracle that knows the bias, using the experimental estimates to calibrate the bias.
-The optimal design concentrates sample in high-sensitivity, high-variance areas and then spreads as the budget grows; compared to a benchmark that randomly chooses areas and splits sample equally, it reduces MSE by roughly 20% under the calibrated bias.
+(2024) also implement a separate randomized expansion, we can calibrate performance of our design (agnostic of the bias) by comparing each candidate design’s MSE to that of an oracle that knows the bias, using the experimental estimates to calibrate the bias.
+Compared to the variance-optimal benchmark, we reduce MSE by more than 250%.
 Related literature
@@
 This problem differs from our design problem and analysis, which instead studies the use of observational studies in combination with experiments for extrapolation.
-In summary, we complement this literature by studying which experiment to run (and how) when not all parameters relevant for the target estimand can be learned experimentally.
+In summary, we complement this literature by studying which experiment to run (and how) when not all parameters are learned experimentally, a common problem in economics.
 Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design.
@@
 (2025) among others.
-These references focus on settings where researchers have only access to experimental variation, instead of combining it with observational evidence, motivating different designs and objective functions.
-In the context of site-selection, we complement literature that leverages correctly-specified models (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
-A related line of work analyzes estimation under misspecification (Armstrong and Kolesár, 2018; Armstrong et al., 2024; Andrews et al., 2025; Andrews et al., 2020; Bonhomme and Weidner, 2022; Christensen and Connault, 2023; James et al., 1961), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; Bhattacharya, 2013; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024).
-We adapt the worst-case proportional regret idea from Armstrong et al.
-(2024) and Tsybakov (1998) here studied in the context of experimental design.
-Our definition of sensitivity of the estimand to each parameter builds on sensitivity analysis in Andrews et al.
-(2020).
-Unlike both strands of this literature where the design is fixed and researchers optimize over the choice of the estimator only, here we optimize the design itself.
-The objective and analysis change: regret is computed against the best design-estimator combination, and the resulting estimator is learned jointly with the design.
-2 Problem description
+These references focus on settings where researchers only have access to experimental variation, instead of combining it with external (observational) evidence, motivating different designs and objective functions.
+We also complement literature that leverages correctly-specified models for site-selection (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
+A related line of work analyzes estimation under misspecification (Armstrong et al., 2024; Armstrong and Kolesár, 2018; Andrews et al., 2025; Andrews et al., 2020; Bonhomme and Weidner, 2022; Donoho, 1994; Kwon and Sun, 2025), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024).
+We provide a novel characterization of adaptation regret as a quasi-convex function in the bias bound (and therefore as the maximum of the variance and bias regret) by leveraging linearity of the estimator.
+Such characterization shows how, in our problem, misspecification is captured through the first-order sensitivity of the estimand to each parameter, an important measure for sensitivity analysis in Andrews et al.
+(2020); it provides a decision-theoretic foundation for simultaneously protecting against a correctly specified and misspecified model (e.g.
+Bickel, 1984; Kempthorne, 1988), here through the maximum of a normalized variance and bias components.
+Unlike the references above, where the design is fixed and researchers optimize over the choice of the estimator, our main contribution is to optimize the design itself.
+This changes the optimization and regret, computed against the best design-estimator.
+2 Baseline scenario: problem description
+We will focus first on a simple shrinkage estimator and a univariate estimand. Our results generalize directly to general methods of moments and multivariate targets in Section 4.
 Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional unknown parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$.
-For simplicity, we assume that the estimand is univariate, while Section 5.3 shows how our framework directly extends to multivariate $\tau(\theta)\in\mathbb{R}^{q}$ for $q>1$.
 The goal is to construct an estimator $\hat{\tau}$ that accurately approximates $\tau(\theta)$ by combining existing evidence (e.g., observational studies) with experimental variation designed by the researcher.
@@
 We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model.
-However, $\tilde{\theta}^{\mathrm{obs}}$, defined as observational estimates, have unknown biases collected in a vector $b\in\mathbb{R}^{p}$.
-Examples include an estimate from an instrumental variable regression that may fail the exclusion restriction, or an estimate from a country different from the one of interest that may lack external validity.
-Researchers can also collect experimental evidence for a subset of parameters, producing estimates whose bias equals zero by design.
+However, $\tilde{\theta}^{\mathrm{obs}}$, defined as observational estimates, have biases collected in a vector $b\in\mathbb{R}^{p}$ unknown at the experimental design stage (see Remark 2).
+Examples include an estimate from an instrumental variable regression that may fail the exclusion restriction, or an experimental estimate from a country different from the one of interest that may lack external validity.
+Researchers can then collect experimental evidence for a subset of parameters, producing estimates which are unbiased by design.
 Setting 1.
@@
 where $\gamma=(\gamma_{j})_{j\in\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$.
-Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design (which parameters are learned and with what precision).
+Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design, encoding which parameters are learned and with what precision.
 When clear, we omit the subscript $\Sigma$ on $\tilde{\theta}^{\mathrm{exp}}$.
-We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012) (in practice, any consistent estimator, e.g., from pilot studies would suffice).
+We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012); when $\Sigma$ is unknown, we can interpret $\Sigma$ as the expected variance-covariance matrix.
+[^7] In this case, we interpret the MSE in Equation (2) as the expected MSE under a common expectation over the experimental variance-covariance matrix.
+It is also possible to extend our framework by allowing $\Sigma$ to be a function of $\tilde{\theta}^{\mathrm{obs}}$ provided that $\Sigma(\tilde{\theta}^{\mathrm{obs}})\rightarrow_{p}\Sigma(\mathrm{plim}(\tilde{\theta}^{\mathrm{obs}}))$, with $\Sigma(\mathrm{plim}(\tilde{\theta}^{\mathrm{obs}}))$ denoting the common expectation of the variance-covariance experimental matrix.
+In this case our analysis should be interpreted in an asymptotic regime similar to what discussed in Section 5.3.
 The class of experiments and estimators is assumed to be in a user-specific set $\mathcal{D}$.
 Assumption 1 (Feasible experiments and estimators).
-We write compactly $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible design (i.e., choice of $\mathcal{E}$ and $\Sigma(\mathcal{E})$) and estimators (i.e., $\gamma$).
-The set $\mathcal{D}$ can be arbitrary as long as (i) each feasible $\Sigma$ has uniformly bounded entries and is strictly positive definite; (ii) $\gamma_{j}=1$ for all $j\in\mathcal{E}$ is one admissible choice (among potentially other values).
-We let $\mathcal{E},\Sigma(\mathcal{E})$ be within arbitrary constraint sets that may encode feasibility or budget constraints.
+We write $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible designs (i.e., choice of $\mathcal{E}$ and $\Sigma(\mathcal{E})$) and estimators (i.e., $\gamma$).
+The set $\mathcal{D}$ can be arbitrary as long as (i) each feasible $\Sigma$ has uniformly bounded entries and is strictly positive definite; (ii) $\gamma_{j}=1$ for all $j\in\mathcal{E}$ is one admissible choice (among potentially others).
+We let $(\mathcal{E},\Sigma(\mathcal{E}))$ be within arbitrary constraint sets that may encode feasibility or budget constraints.
 In our framework, restrictions on the experiments researchers can run $\mathcal{E}$ and on their precision $\Sigma(\mathcal{E})$ can take any desired form.
@@
 The shrinkage estimator can be unrestricted (i.e., $\gamma$ can take any value on the real line) or restricted (when for example researchers aim to only use experimental evidence when available, setting $\gamma_{j}=1$ for $j\in\mathcal{E}$).
-We extend our framework to more general methods of moments estimators in Section 4.
-Remark 1 (Structure of $\Sigma$).
+Example 1 (Structure of $\Sigma$).
 Because observational and experimental moments are typically computed from disjoint samples, in practice, it is natural to model the covariance as block diagonal, $\Sigma=\operatorname{diag}(\Sigma_{\text{obs}},\Sigma_{\text{exp}})$.
-The first block corresponds to observational estimates (fixed and not subject to design choices), while the second corresponds to experimental estimates, and it is a function of the sample size allocated to treatment arms.
+The first block corresponds to observational estimates (fixed and not subject to design choices), while the second corresponds to experimental estimates, and is a function of the sample size allocated to treatment arms (that we may optimize).
 This is a special case of our framework which can be encoded in the constraint set $\mathcal{D}$.
 ∎
-Research questions
+Remark 1 (Some unbiased observational estimates).
+In some applications, researchers may assume that some of the observational study estimates are not biased ($b_{j}=0$ for some $j$).
+This is possible to accommodate as a special case of our general framework in Section 4.
+∎
+Main design problem
 Our design problem can be described in three steps:
@@
 Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025).
-We study this question here in combination with the experimental design problem, step 2 and 3; this, as we show, will change the underlying optimization problem (also for $\gamma$).
-The central challenge is that performance depends on the unknown bias $b$.
-Mapping $\tau$
-A key feature of our framework is that researchers explicitly parameterize $\theta$ and $\tau(\cdot)$, thereby pre-specifying and making transparent which biases drive the design problem; we see this as an advantage because it clarifies the role of the experiment in complementing existing evidence.
-The sensitivity of $\tau$ to $\theta$, captured by its gradient $\omega(\theta)=\partial\tau(\theta)/\partial\theta$, determines how the (first-order) bias propagates onto the final estimand.
-For simplicity, we impose exact linearity, $\tau(\theta)=\omega^{\top}\theta$; all results extend to smooth, nonlinear $\tau$ via a first-order expansion, as discussed below.
+We study this question here in combination with the experimental design problem, step 2 and 3, which, as we show will change the decision problem.
+The challenge is that performance depends on the bias $b$, unknown when designing the experiment.
+Properties of $\tau$
+A key feature of our framework is that researchers explicitly parameterize $\tau(\cdot)$, thereby pre-specifying and making transparent which biases drive the design problem; this feature clarifies the role of the experiment in complementing existing evidence.
+The sensitivity of $\tau$ to $\theta$, captured by its gradient $\omega(\theta)=\partial\tau(\theta)/\partial\theta$, determines how (to first order) the bias propagates onto the final estimand.
+For simplicity, we impose exact linearity, $\tau(\theta)=\omega^{\top}\theta$; all results extend to smooth, nonlinear $\tau$ via a first-order expansion.
+When $\tau$ is unknown, Section 5.1 studies settings where researchers only partially identify $\tau$.
 Assumption 2 (First-order estimation error).
@@
 for known weights $\omega\in\mathbb{R}^{p}$, with $|\omega_{j}|\in(0,\infty)$ for all $j$.
-Assumption 2 holds exactly when $\tau(\cdot)$ is linear and serves as a first-order approximation for smooth non-linear $\tau(\cdot)$.
-Section 5.1 (and Example 3) extend the framework to non-linear estimands encompassing as special cases the local-misspecification setups in Andrews et al.
-(2020); Armstrong and Kolesár (2021).
-For non-linear $\tau$, we replace $\omega$ with the gradient of $\tau$ evaluated at a preliminary observational estimate, i.e., $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}\;+\;o_{p}(1),$ which is typically available before the experiment is conducted.
-That is, building on references above, $\omega$ captures the first-order effect of the observational bias $b$ on the bias of $\hat{\tau}$.
-[^6] To see how this assumption holds under a first-order approximation, define $\theta$ the parameter rescaled by $\sqrt{n}$ (and similarly its bias $b_{n}$, indexed by $n$), where $n^{-1/2}$ is the rate of convergence of the observational estimators.
-Then second-order effects are negligible (and Assumption 2 holds asymptotically) provided that $b_{n}^{2}/\sqrt{n}\rightarrow 0$.
-This implies that Assumption 2 holds when the asymptotic bias $b_{n}$ is proportional to, smaller or even much larger than the asymptotic variance, provided that $b_{n}$ does not grow at rate faster than $n^{1/4}$.
+[…]
 We conclude with three examples in simple two-parameter models.
 Example 2 (Choosing the site for an experiment for external validity).
-Gechter et al. (2024) study where to run an experiment. For illustration, consider two sites $j\in\{1,2\}$ with site-specific ATEs $\theta_{j}$ and target the cross-site average
-$\tau(\theta)=\tfrac{1}{2}(\theta_{1}+\theta_{2})=\omega^{\top}\theta,\qquad\omega=\tfrac{1}{2}(1,1)^{\top}.$
+Gechter et al.
+(2024) study where to run an experiment.
+For illustration, consider two sites $j\in\{1,2\}$ with site-specific ATEs $\theta_{j}$ and target the cross-site average $\tau(\theta)=\omega_{1}\theta_{1}+\omega_{2}\theta_{2},$ with $\omega$ denoting the population share in each site.
 Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al.
@@
 Given $\mathcal{E}$,
-$\widehat{\tau}_{\gamma}=\tfrac{1}{2}\!\left[\gamma_{j}\,\tilde{\theta}^{\mathrm{exp}}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j}+\tilde{\theta}^{\mathrm{obs}}_{k}\right],\qquad k\neq j.$
+$\widehat{\tau}_{\gamma}=\omega_{j}\gamma_{j}\,\tilde{\theta}^{\mathrm{exp}}_{j}+\omega_{j}(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j}+\omega_{k}\tilde{\theta}^{\mathrm{obs}}_{k},\qquad k\neq j.$
 Our question is how to choose the set $\{j\}$ where to conduct the experiment (jointly with $\gamma$).
@@
 Our goal is to study which survey design to implement. (More complex designs with additional time periods or mixed precision across rounds are also possible here.)
-Example 3 (Supply or demand experiment with non linear target).
+Example 4 (Experiment with non linear target and equilibrium model).
 Bergquist and Dinerstein (2020) conduct demand and supply experiments in food markets in Kenya.
 Suppose here we are interested in similar applications in Uganda.
-For exposition, consider a basic linear demand and supply
+For exposition, consider a stylized linear demand and supply model (similar reasoning applies to more complex models)
 $Q^{D}=a-\beta_{D}P+u_{D},\qquad Q^{S}=c+\beta_{S}P+u_{S},\qquad\theta=(\beta_{D},\beta_{S})^{\top},$
-with $\beta_{D}\neq-\beta_{S}$. Several estimands may be of interest; one of such estimands is the effect of a tariff $t$ on prices
-$\tau\equiv\frac{\beta_{S}}{\beta_{D}+\beta_{S}}\,t.$
-Let $(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top}$ denote baseline estimates from Kenya, that may lack external validity.
-Due to cost constraints, $\mathcal{E}\in\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms).
+with $\beta_{D}\neq-\beta_{S}$.
+Several estimands may be of interest; one of such estimands is the effect of a tariff $t$ on prices $\tau\equiv\frac{\beta_{S}}{\beta_{D}+\beta_{S}}\,t.$ Let $(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top}$ denote baseline estimates from Kenya, that may lack external validity.
+Due to fixed costs of each experiment, $\mathcal{E}\in\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms).
 Let
@@
 Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$.
-For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda under local asymptotics described in Section 5.1
-$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\check{\omega}(\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}}))^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{1}{n^{1/2}}\right)}_{\text{small higher order effects}},\qquad\check{\omega}(\beta)=\frac{t}{(\beta_{D}+\beta_{S})^{2}}\begin{pmatrix}-\beta_{S}\\ \beta_{D}\end{pmatrix}$,
-where we can replace $\check{\omega}$ with its consistent counterpart $\check{\omega}(\tilde{\beta}^{\mathrm{obs}})$.
-Our goal is to choose between a supply or demand experiment accounting for first-order bias (and variance) effect.
+For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda under local asymptotics described in Section 5.3 and Remark 6
+$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\omega^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{1}{n^{1/2}}\right)}_{\text{small higher order effects}},\hskip 18.49988pt\omega=\frac{t}{(\mathrm{plim}(\tilde{\beta}_{D}^{\mathrm{obs}}+\tilde{\beta}_{S}^{\mathrm{obs}}))^{2}}\begin{pmatrix}-\mathrm{plim}(\tilde{\beta}_{S}^{\mathrm{obs}})\\ \mathrm{plim}(\tilde{\beta}_{D}^{\mathrm{obs}})\end{pmatrix}$,
+where we can replace $\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}})$ with its consistent counterpart $\tilde{\beta}^{\mathrm{obs}}$.
 3 Robust experimental design
-Next, we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$, defined, for given bias level $b$ and design $(\mathcal{E},\Sigma(\mathcal{E}))$ as
-$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[(\widehat{\tau}_{\gamma}-\tau)^{2}\right],$
-where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$.
-Section 5.2 shows that our results for MSE also extend when minimizing confidence intervals’ length.
+Next, we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$.
+We consider confidence interval length in Section 5.1.
+The MSE is defined, for given bias level $b$, design $(\mathcal{E},\Sigma(\mathcal{E}))$, and choice of weights $\gamma$ as
+$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[(\widehat{\tau}_{\gamma}-\tau)^{2}\right],$ (2)
+where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the design $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$.
 Ideally, one would minimize $\mathrm{MSE}_{b}$. However, because $b$ is unknown, we consider an uncertainty set given by an $\ell_{\infty}$-ball,
-$\mathcal{B}(B)\equiv\{b:\ \|b\|_{\infty}\leq B\},$
+$\mathcal{B}(B)\equiv\{b:\ \|b\|_{\infty}\leq B\},$ (3)
 with $B\geq 0$ an upper bound on the largest coordinate-wise bias.
-In our framework, a key challenge is that biases may arise from multiple parameters and uncertainty sets defined by different norms may be possible.
-A key insight here is that the $\ell_{\infty}$ choice has a desirable property relative to $\ell_{1}$ or $\ell_{2}$: it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters.
+In our framework, a key challenge is that biases may arise from multiple parameters.
+Here, for exposition we first focus on the $\ell_{\infty}$-ball given its interpretability, while Section 4 generalizes the framework to arbitrary norms.
+The $\ell_{\infty}$ norm imposes symmetric restrictions on the biases and it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters.
 The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
@@
 Specifically, we minimize the worst-case proportional regret
-$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;\equiv\;\sup_{B\geq 0}\ \frac{\displaystyle\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\displaystyle\mathrm{MSE}^{*}(B)},$
-defined adaptation regret by Armstrong et al. (2024) (building in turn on Tsybakov, 1998).
-There are two key distinctions relative to such references using the adaptation regret.
-First (and importantly) we optimize over the design itself rather than fixing it ex ante; consequently, regret is defined relative to an oracle that jointly chooses both the estimator and the design.
-Second, we allow biases to arise across multiple parameters (rather than a single one), which necessitates a careful specification of the uncertainty set; we adopt an $\ell_{\infty}$-ball, which, as we show yields transparent and computationally-attractive solutions.
+$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;\equiv\;\sup_{B\geq 0}\ \frac{\displaystyle\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\displaystyle\mathrm{MSE}^{*}(B)},$ (4)
+defined adaptation regret by Armstrong et al. (2024) and Tsybakov (1998) for estimation under fixed designs (see Remark 3 for a comparison with our experimental design problem).
+Remark 2 (Adaptation to $B$).
+We consider worst-case adaptation regret over $B$ because researchers typically do not know $B$ before running an experiment and learning $B$ would require running pilots for each design in the menu of feasible experiments.
+Also, Armstrong and Kolesár (2018) show that pointwise adaptation to unknown misspecification levels is impossible in general, motivating the adaptation regret.
+Note however that if researchers know that the misspecification level $B$ lies in a bounded interval $B\in[0,\bar{B}]$, our characterization extends by simply restricting the worst-case over $B$ to this interval; see Section 5.2.∎
 3.1 Optimal design
-A natural question is whether the adaptation regret provides natural trade-offs between the variance and the bias.
+Our main results show that the adaptation regret provides natural trade-offs between the variance and the bias.
 To do so, denote $\mathbb{V}_{\mathcal{E},\Sigma}(\cdot)$ the variance operator for given $(\mathcal{E},\Sigma)$.
@@
 Let
-$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\mathbb{V}_{\mathcal{E},\Sigma}(\hat{\tau}_{\gamma}),\qquad\alpha^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\alpha(\mathcal{E},\Sigma,\gamma),$ (2)
+$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\mathbb{V}_{\mathcal{E},\Sigma}(\hat{\tau}_{\gamma}),\qquad\alpha^{\star}\;\equiv\;\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),$ (5)
 with $\alpha^{\star}$ denoting the smallest feasible variance. We will refer to $\alpha/\alpha^{\star}$ as the variance regret.
-The variance regret denotes the ratio between the variance of a given design and estimator relative to the smallest achievable variance.
-We provide a similar definition for the bias.
-Let $|1-\gamma|$ define the absolute value of each entry of the vector $1-\gamma$.
+The variance regret denotes the ratio between the variance of a given design and estimator relative to the smallest achievable variance. We provide a similar definition for the bias.
 Definition 2 (Bias regret).
 Define
-$\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},\quad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\Big(||\omega||_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2},$
+$\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},\quad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\Big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2},$
 with $\beta/\beta^{\star}$ defined as the bias regret.
-[…]
-An ideal design, would set the variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$.
+To gain insight, note that $B^{2}\beta=B^{2}\Big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2}$ denotes the worst-case bias for a given design, value $B^{2}$ and shrinkage estimator.
+$B^{2}\beta^{\star}$ denotes its minimizer.
+An ideal design would set the variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$.
 Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other.
 This raises the question of how to trade-off these two components.
-This trade-off is formally characterized in our main theorem below.
+This trade-off is formally characterized in our main theorem below. We will use the convention throughout that 0/0 = 1.
 Theorem 1.
@@
 See Appendix A.1. ∎
-Theorem 1 provides a simple expression of the regret which does not require researchers to specify $B$.
-The key insight is that the adaptation regret in our context, leads to a (strictly) quasi-convex objective function, whose worst-case solution is the maximum between the variance and the bias regret.
+Theorem 1 offers a simple and novel expression of the regret which does not require researchers to specify $B$.
+As we show below, Theorem 1 will significantly simplify the optimization program.
+The key insight is that the adaptation regret for linear estimators and arbitrary designs leads to a (strictly) quasi-convex objective function, whose worst-case solution is the maximum between the variance and the bias regret.
 This provides a simple and intuitive trade-off between the two.
-Specifically, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), it follows that at the optimum we equalize the variance and bias regret as in Figure 1
+Specifically, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), it follows that at the optimum we equalize
 $\frac{\alpha}{\alpha^{\star}}\;=\;\frac{\beta}{\beta^{\star}}.$
-In practice, this amount to allocate sample size to noisier treatment arms when the variance dominates, and select treatment arms with largest sensitivity when the bias dominates.
+That is, we allocate sample size to noisier treatment arms when the variance dominates, and select treatment arms with largest sensitivity when the bias dominates.
 […]
@@
 The minimizer outside boundary solutions is where the frontier meets the $45^{\circ}$ line (red dot).
-Theorem 1 provides us an immediate solution to the optimal design problem. We compute the estimator, the optimal variance and the experimental choice via backward induction:
-$\gamma^{\star}(\mathcal{E},\Sigma)$ $\in\mathrm{arg}\min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}$
-$\Sigma^{\star}(\mathcal{E})$ $\in\mathrm{arg}\min_{\Sigma\in\mathcal{G}(\mathcal{E})}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\mathcal{E},\Sigma))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma))}{\beta^{\star}}\right\}$
-$\mathcal{E}^{\star}$ $\in\mathrm{arg}\min_{\mathcal{E}\in\mathcal{S}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma^{\star}(\mathcal{E}),\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\beta^{\star}}\right\}.$
-That is, we first choose $\gamma^{\star}$ for each design $\mathcal{E},\Sigma(\mathcal{E})$.
-We then choose $\Sigma$, whose choice depends on $\beta$ through the shrinkage weights.
-We finally choose the set of experiments $\mathcal{E}$.
-3.2 Intuition with two parameters
+Remark 3 (Comparison with previous literature on adaptation regret).
+Previous literature has studied adaptation regret to combine an unbiased and a biased estimator for a single parameter $\theta\in\mathbb{R}$ under a fixed design (Armstrong et al., 2024; Tsybakov, 1998).
+There are three key distinctions from this literature.
+First (and foremost), we optimize over the design itself rather than taking it as given; regret is defined relative to an oracle that chooses both the estimator and the design ex ante.
+Second, we allow biases to arise across multiple parameters (rather than a single one), and do not require that researchers have access to an unbiased estimator for each component of $\theta$.
+Third, we study a linear class of estimators, instead of arbitrary nonlinear rules.
+[^10] In the univariate bounded-normal-mean problem, Donoho et al.
+(1990) show that linear estimators lose no more than 25% efficiency relative to the optimal nonlinear estimator.
+As a consequence, the characterization in Theorem 1 differs from Armstrong et al.
+(2024), where regret is not quasi-convex in the bias radius and the optimal estimator solves a minimax problem via a least-favourable prior; see Appendix D.
+Their minimax solution for combining two estimators is not applicable here once we optimize over the choice of experiments, sample allocation and multiple estimators, since searching over nonlinear shrinkage rules becomes computationally demanding.
+By focusing on linear estimators, we obtain a tractable solution that also extends to common method-of-moments, see Section 4.
+∎
+Example 5 (Intuition with two parameters).
 To build further intuition, it is useful to study the two-parameter model.
-Setting 2 (Illustration with two-parameter model).
 Let $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\quad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\quad j=1,2.$ For simplicity $\sigma_{j}^{2},v_{j}^{2}$ are fixed.
 We may run one experiment due to budget constraints: $\mathcal{E}\in\{\{1\},\{2\}\}$.
-If we pick $j$, we estimate $\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\quad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$ and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$
-How do the variance and bias regret look like in this easier scenario? We can write
+If we pick $j$, we estimate $\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\quad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$ and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$ We can write
 $\alpha^{\star}=\min_{k\in\{1,2\}}\left\{\ \omega_{-k}^{2}\sigma_{-k}^{2}\;+\;\omega_{k}^{2}\frac{\sigma_{k}^{2}\,v_{k}^{2}}{\sigma_{k}^{2}+v_{k}^{2}}\ \right\},\qquad\beta^{\star}=\min_{k\in\{1,2\}}\omega_{k}^{2}.$
@@
 $\alpha(j,\gamma_{j})$ $\equiv\underbrace{\omega_{-j}^{2}\sigma_{-j}^{2}}_{\text{var obs estimate}}+\underbrace{\omega_{j}^{2}\!\big[(1-\gamma_{j})^{2}\sigma_{j}^{2}+\gamma_{j}^{2}v_{j}^{2}\big]}_{\text{var exp and obs estimate}},\quad\beta(j,\gamma_{j})$ $\equiv\big(\underbrace{|\omega_{-j}|}_{\begin{subarray}{c}\text{bias obs}\\ \text{estimate}/B\end{subarray}}+\underbrace{(1-\gamma_{j})|\omega_{j}|}_{\begin{subarray}{c}\text{bias obs minus}\\ \text{ bias exp}/B\end{subarray}}\big)^{2}.$
-The very first step is optimizing over the estimator (assuming here for simplicity $\gamma$ can take any value). The following characterizes the regret optimal $\gamma^{\star}$ (see Appendix C.1).
+The very first step is optimizing over the estimator (assuming here for simplicity $\gamma$ can take any value). The following characterizes the regret optimal $\gamma^{\star}$ (see Appendix B.1).
 […]
@@
 However, cases at the extreme boundaries can also occur.
-For instance, if the variance regret $\alpha/\alpha^{\star}$ or the bias regret $\beta/\beta^{\star}$ uniformly dominates the other for one of the experiments, the optimal design minimizes the dominating term.
+If the variance regret $\alpha/\alpha^{\star}$ or the bias regret $\beta/\beta^{\star}$ uniformly dominates the other for one of the experiments, the optimal design minimizes the dominating term.
 These are desiderable properties as the optimal design always prioritizes the dominating term.
 Appendix Table 5 provides a complete overview of all possible scenarios.
-More complex scenarios where researchers may allocate sample sizes to each experiment are possible, see Section 6.
-Figure 2: Example for $j=2$.
-Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$.
-Bottom row: the optimal weight $\gamma_{2}^{\star}$.
-Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
-Figure 3: Example for $j=2$ (regret).
-Different colors correspond to the regret for choosing either experiment.
-Figure reports $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$).
-The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret.
-Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
-Example 4 (Numerical illustration).
-For an illustrative example, consider choosing $\gamma_{2}^{\star}$ when experiment $j=2$ is twice as precise as $j=1$, with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, $v_{1}=1$, $v_{2}=0.5$, and $\omega_{1}=1$.
-The optimal $\gamma_{2}^{\star}$ depends not only on features of design $j$ but also on those of the alternative design j-j, $(v_{-j}^{2},\sigma_{-j}^{2},\omega_{-j})$, through the benchmarks $(\alpha^{\star},\beta^{\star})$.
-Figure 2 (top row) plots $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$.
-The bottom row reports the corresponding $\gamma_{2}^{\star}$.
-Varying $\omega_{2}$ yields two key effects: both very small and very large $\omega_{2}$ push $\gamma_{2}^{\star}$ toward one, but for different reasons—when $\omega_{2}$ is small the oracle bias decreases faster than the experimental variance (since $\beta^{\star}=\omega_{2}$), while when $\omega_{2}$ is large the oracle bias is $\beta^{\star}=\omega_{1}$ and $\omega_{2}$ has only on the feasible experiment’s bias.
-Consequently, both regret and the optimal estimator depend on the oracle’s choice of the experiment.
-In the interior region, the optimal weight first shifts toward the variance–optimal solution (as the oracle bias declines) and then rises again toward a bias–preferred solution.
-When we vary the experimental and observational variances, we observe the expected patterns: $\gamma_{2}^{\star}$ moves inversely with experimental variance and in the same direction as observational variance.
-These patterns translate into implications for the regret of the chosen experiment (Figure 3).
-For the second arm $j=2$, regret decreases sharply as $\omega_{2}$ increases and then decreases, with further decreases becoming gradual once the bias component is already small.
-By contrast, for the first arm, regret is at best slightly decreasing for small changes in $\omega_{2}$ (when variance effects dominate), but eventually increases with $\omega_{2}$ because the oracle increasingly favors arm $2$, so the opportunity cost of selecting arm $1$ increases.
-A summary appears in Table 4, with a complete discussion in Appendix C.2.
+Appendix B.2 provides a numerical example and Section 6 presents a more complex example that include optimally allocating sample size.
 ∎
-4 Generalization to moments selection
-In many scenarios, we may expect that experiments only identify relevant moments in the data. We now extend the framework to generalized method of moments (GMM) estimators.
-Consider a vector of moment conditions $g(\theta)\in\mathbb{R}^{p}$ that stacks all moments (observational and experimental) potentially available for estimation.
+3.2 Optimization
+Next, we provide an explicit optimization routine that can be solved using off-the-shelf software.
+We consider observational estimates independent of the experimental ones (but possibly dependent with each other), and experimental estimates independent of each other.
+This occurs when conducting experiments on independent samples different from the observational sample as in our applications.
+More general dependences are possible but omitted.
+Algorithm 1 Regret-optimal experimental design
+1:Inputs: target functional weights $\omega$, variance parameters $(v_{j}^{2})$, observational covariance $\Sigma_{\mathrm{obs}}$, per-unit costs $(c_{j})$, total budget $n$, and feasible experiment set $\mathcal{X}$.
+2:Define variance, bias, and sample sizes. For any $(x,\gamma)$ with $x\in\mathcal{X}$ and $0\leq\gamma_{j}\leq 1$, set $s_{j}=x_{j}\gamma_{j}$ and compute:
+  • the sample allocation (8), and the definition of $\alpha(s)$ in (9);
+  • the bias contribution $\beta(s)$ as defined below (9).
+3:Compute oracle quantities. Obtain $\alpha^{\star}$ and $\beta^{\star}$ by solving the two optimization problems, i.e.,
+  • minimize $\alpha(s)$ over $(x,\gamma,s)$ subject to $x\in\mathcal{X}$, $x_{j}\in\{0,1\}$, $s_{j}=x_{j}\gamma_{j}$;
+  • minimize $\sum_{j}x_{j}|\omega_{j}|$ over $x\in\mathcal{X}$ with $x_{j}\in\{0,1\}$.
+4:Solve the minimax MIQP. Solve (10)–(14) with a off-the-shelf solver.
+5:Outputs. Let $(x^{\star},\gamma^{\star},s^{\star},t^{\star})$ denote an optimizer of (10)–(14). Report:
+  • the regret-optimal experiment set $\mathcal{E}^{\star}=\{j:x_{j}^{\star}=1\}$ and shrinkage weights $\gamma^{\star}$;
+  • the implied $s^{\star}$ and regret factor $t^{\star}=\max\{\alpha/\alpha^{\star},\beta/\beta^{\star}\}$;
+  • the optimal sample sizes $n_{j}^{\star}$ obtained by plugging $s^{\star}$ into (8).
+Notation and decision variables
+For given $(\mathcal{E},\Sigma,\gamma)$, denote by $\Sigma_{\mathrm{obs}}$ the submatrix of $\Sigma$ corresponding to the variance–covariance matrix of the observational estimates (possibly non-diagonal), and let $\mathbb{V}(\tilde{\theta}_{j}^{\mathrm{exp}})=v_{j}^{2}/n_{j}$ be the variance of the experimental estimator for component $j$, where $n_{j}$ denotes the sample size allocated to that experiment.
+Let $x_{j}=1\{j\in\mathcal{E}\}\in\{0,1\}$ indicate whether the experiment identifies component $j$, and let $x\in\mathcal{X}$, where $\mathcal{X}$ denotes the constraint set of feasible experiments.
+We optimize over $(x,\gamma,n_{j})$ under a total budget constraint $\sum_{j\in\mathcal{E}}c_{j}n_{j}=n,$ where $c_{j}>0$ denotes a per-unit cost for experiment.
+Variance, bias and optimal sample size
+Under independence of experimental and observational estimators, the variance of $\tau(\theta)$ can be written as
+$\sum_{j}\omega_{j}^{2}s_{j}^{2}\frac{v_{j}^{2}}{n_{j}}\;+\;\big(\omega\odot(1-s)\big)^{\top}\Sigma_{\mathrm{obs}}\big(\omega\odot(1-s)\big),\qquad s_{j}\equiv x_{j}\gamma_{j},$ (7)
+where $\odot$ denotes the elementwise product.
+The sample sizes $n_{j}$ enter only through the first (experimental) component.
+By the standard Neyman allocation with variable costs (Gerber and Green, 2012), the minimizer over $(n_{j})$ for fixed $s$ is given by
+$n_{j}^{\star}\;=\;n\,\frac{|\omega_{j}|v_{j}s_{j}/\sqrt{c_{j}}}{\sum_{k}|\omega_{k}|v_{k}s_{k}\sqrt{c_{k}}}.$ (8)
+Substituting (8) into the experimental variance and with a slight abuse of notation, let
+$\alpha(s)\;\equiv\;\frac{1}{n}\Big(\sum_{j}s_{j}|\omega_{j}|v_{j}\sqrt{c_{j}}\Big)^{2}+\big(\omega\odot(1-s)\big)^{\top}\Sigma_{\mathrm{obs}}\big(\omega\odot(1-s)\big),\quad\beta(s)\;\equiv\;\Big(\sum_{j}(1-s_{j})|\omega_{j}|\Big)^{2},$ (9)
+the variance and bias contributions.
+Oracle solutions
+The next step is to compute $\alpha^{\star}$ and $\beta^{\star}$. These solve
+$\alpha^{\star}=\min_{x,\gamma,s}\alpha(s),$ $\text{s.t. }0\leq\gamma_{j}\leq 1,\;\;x_{j}\in\{0,1\},\;\;x\in\mathcal{X},$
+$0\leq s_{j}\leq\gamma_{j},\;\;s_{j}\leq x_{j},\;\;s_{j}\geq\gamma_{j}+x_{j}-1,\quad\forall j,$
+$\sqrt{\beta^{\star}}=\min_{x^{\prime}}\sum_{j}(1-x_{j}^{\prime})|\omega_{j}|,$ $\text{s.t. }x^{\prime}\in\mathcal{X},\;\;x_{j}^{\prime}\in\{0,1\},\quad\forall j.$
+The constraints for $s$ are standard inequalities enforcing $s_{j}=\gamma_{j}x_{j}$ when $x_{j}\in\{0,1\}$ and $0\leq\gamma_{j}\leq 1$.
+Because $\alpha(s)$ is convex quadratic and $\mathcal{X}$ typically admits a linear (or quadratic) description, the first problem is a mixed-integer convex quadratic program, while the second is a mixed-integer linear program.
+Both can be solved using off-the-shelf solvers.
+Regret-optimal solution
+We are now ready to solve for the regret-optimal design. By Theorem 1, we want to minimize $\max\{\alpha/\alpha^{\star},\beta/\beta^{\star}\}$. The corresponding optimization program can be written as
+$\min_{x,\gamma,s,t}\quad$ $t$ (10)
+s.t. $\alpha(s)\;\leq\;t\,\alpha^{\star},$ (11)
+$\beta(s)\;\leq\;t\,\beta^{\star},$ (12)
+$0\leq\gamma_{j}\leq 1,\quad x_{j}\in\{0,1\},\quad j=1,\dots,p,$ (13)
+$0\leq s_{j}\leq\gamma_{j},\quad s_{j}\leq x_{j},\quad s_{j}\geq\gamma_{j}+x_{j}-1,\quad j=1,\dots,p,$ (14)
+$x\in\mathcal{X}\quad\text{(constraint on feasible experiments).}$ (15)
+The constraints (14) again enforce $s_{j}=x_{j}\gamma_{j}$ for $x_{j}\in\{0,1\}$ and $0\leq\gamma_{j}\leq 1$.
+The variable $t$ satisfies $t\geq\alpha/\alpha^{\star}$ and $t\geq\beta/\beta^{\star}$, so at the optimum we have $t=\max\{\alpha/\alpha^{\star},\beta/\beta^{\star}\}$ as desired.
+Because both $\alpha(s)$ and $\beta(s)$ are convex quadratic functions, the optimization problem (10)–(14) is a mixed-integer quadratic program (MIQP) with linear objective and convex quadratic constraints, and can be passed to standard MIQP solvers.
+Remark 4 (“Bias-aware” Neyman allocation).
+The optimal sample size solution in Equation (8) has the desirable feature that, conditional on the chosen experiment and shrinkage weights, it takes the form of a Neyman (variance-optimal) allocation.
+However, because both the experiment and the estimator are selected to trade off variance and worst-case bias, the resulting allocation generally differs from the standard Neyman allocation one would obtain by ignoring the sensitivity of the estimand to misspecification (and the fixed-cost constraints).
+It is therefore natural to refer to our sample size allocation as a (novel) “bias-aware” Neyman allocation, where the choice of the experiment and shrinkage weights trade-off bias and variance, and optimal sample size minimizes the variance conditional on these choices.
+∎
+4 General framework for moment selection
+In many scenarios, we may expect that experiments only identify relevant moments in the data, and observational estimates are often based on model-implied moment conditions.
+We now extend the framework to generalized method of moments (GMM) estimators.
+In addition, we allow for a more general class of ambiguity sets and multivalued estimands.
+Consider a vector of moment conditions $g(\theta)\in\mathbb{R}^{p}$ that stacks the “menu of moments” (observational and experimental), i.e.
+all moments potentially available for estimation.
 Let $\Lambda$ denote the Jacobian of the moment function with respect to $\theta$, evaluated at $\theta$ (i.e., $\Lambda:=\partial g(\theta)/\partial\theta^{\top}$).
@@
 This simplifies exposition but it is not necessary in the local asymptotic framework we present in Section 5.3, where $\Lambda$ can be replaced by the Jacobian of $g(\theta)$ evaluated at the observational estimates.
-For a given covariance matrix $\Sigma$ of the sample moment vector, let $\bar{g}_{\Sigma}$ denote the sample analogue with
+Let $\bar{g}_{\Sigma}$ denote the sample analogue with
 $\mathbb{E}[\bar{g}_{\Sigma}]=b,\qquad\mathbb{V}(\bar{g}_{\Sigma})=\Sigma,$
-where the vector $b$ captures the biases in the moments. The researcher may decide to access only some of the entries of $\bar{g}$ by selecting a subset of experiments due to budget constraints.
-Misspecification set.
+where the vector $b$ captures the biases in the moments. The researcher may access only some of the entries of $\bar{g}$ by selecting a subset of experiments due to budget constraints.
+General misspecification set.
 Let $\mathcal{I}\subseteq\{1,\dots,p\}$ index all potential experimental moments researchers may select (unbiased by design) and $\mathcal{I}^{c}$ the observational ones (potentially misspecified).
-We posit bounded misspecification
-$\mathcal{B}(B)\;=\;\Big\{b\in\mathbb{R}^{p}:\ b_{j}=0\ \text{for }j\in\mathcal{I},\ \ |b_{j}|\leq B\ \text{for }j\in\mathcal{I}^{c}\Big\}.$ (4)
-This restriction states that experimental moments have no bias, whereas observational moments have bias bounded by $B$, potentially unknown to the researcher.
+We posit bounded misspecification under arbitrary norms
+[…]
+where the ambiguity set is defined relative to an arbitrary norm $||\cdot||_{l}$.
+Examples include the $\ell_{1},\ell_{2},\ell_{\infty}$ norms as special cases, as well as norms that may weight observational biases differently, or may even assume no bias for some observational moments (see Example 7).
 Estimator and design.
@@
 The structure of the estimator follows a standard GMM asymptotic expansion (e.g.
-Andrews et al., 2020).
+Andrews et al., 2020, 2017).
 [^11] The weighting matrix $W$ may be singular.
-The expansion $\hat{\theta}-\theta\;\dot{=}\;-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\bar{g}_{\Sigma}$ holds whenever $\Lambda^{\top}W\Lambda$ is nonsingular, which can be forced through the set of constraints.
-Here, $W$ may have columns of (effectively) zeros, which excludes the corresponding moments from the estimator.
-This occurs when researchers faces a trade-off in which experiment to run, leading to different moments $\bar{g}$ observed in the experiment.
-The matrix $\Sigma$ depends on the experimental design (e.g., sample–size allocations across experimental moments).
-The constraints on $W$ and $\Sigma$ effectively limits the class of experiments, sample size allocations and potentially estimators the researcher may consider.
+The expansion $\hat{\theta}-\theta\;\dot{=}\;-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\bar{g}_{\Sigma}$ holds whenever $\Lambda^{\top}W\Lambda$ is nonsingular, which can be forced through the set of constraints in Assumption 3.
+Here, $W$ may have columns of zeros, which excludes the corresponding moments from the estimator.
+This occurs when researchers face a trade-off in which experiment to run, leading to different moments $\bar{g}$ observed in the experiment.
+The matrix $\Sigma$ depends on the experimental design (e.g., sample-size allocations across experimental moments).
+The constraints on $W$ and $\Sigma$ effectively limit the class of experiments, sample size allocations and estimators the researcher may consider.
 Assumption 3 (Constraint set).
@@
 This ensures that any subset of moments the researcher chooses to use is nondegenerate.
+Note that for now we are abstracting from settings with incomplete models studied in Section 5.1.
 Example 6 (GMM moment selection).
-Let $g(\theta)=(g^{\text{obs}}(\theta),\,g^{\text{exp}}_{2}(\theta),\,g^{\text{exp}}_{3}(\theta))^{\top}\in\mathbb{R}^{3}$, where the first component is observational and the other two are experimental.
+Let $\theta\in\mathbb{R}^{2}$ and
+$g(\theta)=\big(g_{1}^{\text{obs}}(\theta),\,g_{2}^{\text{obs}}(\theta),\,g^{\text{exp}}_{1}(\theta),\,g^{\text{exp}}_{2}(\theta)\big)^{\top}\in\mathbb{R}^{4},$
+where the first two components are observational and the last two are experimental.
 Suppose at most one experimental moment can be used.
-Selection is encoded by the GMM weight $W\in\mathbb{R}^{3\times 3}$.
-A simple example of class weighting matrices is
-$W^{(2)}=\begin{pmatrix}w_{\mathrm{obs}}&\rho&0\\ \rho&w_{2}&0\\ 0&0&0\end{pmatrix}\qquad W^{(3)}=\begin{pmatrix}w_{\mathrm{obs}}&0&\rho\\ 0&0&0\\ \rho&0&w_{3}\end{pmatrix}$,
-provided that $\Lambda^{\top}W\Lambda$ is invertible. ∎
+Selection is encoded by the GMM weight matrix $W\in\mathbb{R}^{4\times 4}$.
+A simple class of weighting matrices is
+$W^{(1)}=\begin{pmatrix}w_{\mathrm{obs},1}&\rho&0&0\\ \rho&w_{\mathrm{obs},2}&0&0\\ 0&0&w_{\mathrm{exp},1}&0\\ 0&0&0&0\end{pmatrix}\qquad W^{(2)}=\begin{pmatrix}w_{\mathrm{obs},1}&\rho&0&0\\ \rho&w_{\mathrm{obs},2}&0&0\\ 0&0&0&0\\ 0&0&0&w_{\mathrm{exp},2}\end{pmatrix}$,
+where $W^{(1)}$ selects the second experimental moment $g^{\text{exp}}_{1}$ and sets the weight on $g^{\text{exp}}_{2}$ to zero, while $W^{(2)}$ selects $g^{\text{exp}}_{2}$ and excludes $g^{\text{exp}}_{1}$.
+∎
 Target and MSE.
-Let the estimand be a linear functional $\tau(\theta)=\omega^{\top}\theta$ with $\omega\in\mathbb{R}^{d}$. Then
-$\mathrm{MSE}_{b}(W,\Sigma)$ $=\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\;+\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega.$ (6)
+Let the estimand be potentially multivalued $\tau(\theta)\in\mathbb{R}^{q}$, where $\tau(\theta)=\Omega\theta$ with $\Omega\in\mathbb{R}^{q\times d}$ with $\Omega$ different from the zero matrix.
+Our objective is the sum of MSE for each entry of $\tau$, that is equal to
+[…]
 The key distinction now is that the variance and bias regret depend on $\Omega$ and $\Lambda$.
 Definition 3 (Moment based variance and bias regret).
-With a slight abuse of notation, let
-$\alpha(W,\Sigma)\;\equiv\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega,\qquad\beta(W)\;\equiv\;\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}^{\,2}.$
-and $\alpha^{\star},\beta^{\star}$ their corresponding minimizers over $(W,\Sigma)\in\mathcal{D}^{\prime}$.
-The moment-based variance and bias regret are defined as $\alpha/\alpha^{\star},\beta/\beta^{\star}$ respectively.
+[…]
+$\alpha_{\Omega}(W,\Sigma)\;\equiv\;\mathrm{Trace}\left(\Omega\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\Omega^{\top}\right),\qquad\beta_{l,\Omega}(W)\;\equiv\;\big\|[\Omega\Gamma_{\Lambda}(W)]_{.,\mathcal{I}^{c}}\big\|_{l,*}^{\,2}.$
+Let $\alpha_{\Omega}^{\star},\beta_{l,\Omega}^{\star}$ their corresponding minimizers over $(W,\Sigma)\in\mathcal{D}^{\prime}$.
+The moment-based variance and bias regret are defined as $\alpha_{\Omega}/\alpha_{\Omega}^{\star},\beta_{l,\Omega}/\beta_{l,\Omega}^{\star}$ respectively.
 The bias now depends on the interaction between the sensitivity $\Omega$, the moment Jacobian $\Lambda$ and the identity of which moments are observational and therefore potentially biased.
-The weighting matrix $W$ determines how these moments influence the downstream estimator $\hat{\theta}$.
-As before, (and with a slight abuse of notation) we consider the adaptation regret of the following form:
-$\mathcal{R}(W,\Sigma)\;\equiv\;\sup_{B\geq 0}\ \frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\ \mathrm{MSE}_{b}(W^{\prime},\Sigma^{\prime})}.$ (7)
+For unidimensional $\tau(\theta)$ the bias corresponds to the squared dual norm of $\Omega\Gamma_{\Lambda}(W)$ with respect to $||\cdot||_{l}$.
+The weighting matrix $W$ determines how these moments influence $\hat{\theta}$.
+As before, we consider the adaptation regret with respect to an oracle that knows $B$:
+$\mathcal{R}_{l}(W,\Sigma)\;\equiv\;\sup_{B\geq 0}\ \frac{\sup_{b\in\mathcal{B}_{l}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\sup_{b^{\prime}\in\mathcal{B}_{l}(B)}\ \mathrm{MSE}_{b^{\prime}}(W^{\prime},\Sigma^{\prime})}.$ (19)
 The next theorem generalizes Theorem 1 to this more general framework.
@@
 Let Assumptions 2, 3 hold, $\mathcal{R}$ be as defined in Equation (19),with $\alpha,\beta,\alpha^{\star},\beta^{\star}$ as in Definition 3. Consider an estimator as in Equation (17). Then
-$\mathcal{R}(W,\Sigma)\;=\;\max\!\left\{\,\frac{\alpha(W,\Sigma)}{\alpha^{\star}}\,,\ \frac{\beta(W)}{\beta^{\star}}\,\right\}.$ (8)
+$\mathcal{R}_{l}(W,\Sigma)\;=\;\max\!\left\{\,\frac{\alpha_{\Omega}(W,\Sigma)}{\alpha_{\Omega}^{\star}}\,,\ \frac{\beta_{l,\Omega}(W)}{\beta_{l,\Omega}^{\star}}\,\right\}.$ (20)
 Proof.
-See Appendix B.2. ∎
-Theorem 2 shows how our framework generalizes to two related problems (i) choosing the weighting matrix $W$ for estimation and (ii) moment selection.
-The key insight is to show that the regret is quasi-convex under (approximate) linearity in the moments.
-5 Extensions
-5.1 Local asymptotics in the presence of non-linearities
-This subsection extends our framework for (i) nonlinear moment functions $g(\theta)$ used in GMM in Section 4 and (ii) a smooth, potentially nonlinear estimand $\tau(\theta)$.
+See Appendix A.1. ∎
+Theorem 2 shows how our framework generalizes to (i) choosing the moments and the estimator (via $W$); and (ii) arbitrary norms characterizing the ambiguity set.
+The key insight is that the regret is quasi-convex under (approximate) linearity in the moments relative to the dual norm of the ambiguity set.
+This generalization encompasses, in addition to choosing which experiment to run, problems that involve whether to collect additional data (e.g., through surveys or data acquisition) since these can be directly incorporated as part of the moment conditions.
+Example 7 (Weighted bias).
+Let $\tau(\theta)\in\mathbb{R}$ and consider an ambiguity set of the form $\tilde{\mathcal{B}}(B)\;\equiv\;\Big\{b\in\mathbb{R}^{p}:|b_{j}|\leq k_{j}B\;\;\text{for all }j=1,\dots,p\Big\},$ for known rescaling factors $k$.
+Then the corresponding bias regret is $\tilde{\beta}/\tilde{\beta}^{\star}$ with $\tilde{\beta}(\mathcal{E},\gamma)\;\equiv\;(\sum_{j\notin\mathcal{E}}k_{j}|\omega_{j}|\;+\;\sum_{j\in\mathcal{E}}k_{j}|\omega_{j}|\big|1-\gamma_{j}\big|)^{2},\quad\tilde{\beta}^{*}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}(\sum_{j\notin\mathcal{E}}k_{j}|\omega_{j}|)^{2},$ and the overall regret equals $\max\!\left\{\,\frac{\alpha(W,\Sigma)}{\alpha^{\star}}\,,\ \frac{\tilde{\beta}(W)}{\tilde{\beta}^{\star}}\,\right\}$.
+∎
+5 Confidence sets, incomplete models and extensions
+5.1 Confidence sets and partially identified $\tau$
+Next, we generalize our framework to minimizing confidence interval length and partially identified $\tau(\theta)$.
+For simplicity, we let $\tau(\theta)\in\mathbb{R}$.
+Researchers may not know $\tau(\cdot)$ exactly and only know functions $\bar{\tau},\underline{\tau}$ with
+$\bar{\tau}(\theta)=\bar{\omega}^{\top}\theta,\qquad\underline{\tau}(\theta)=\underline{\omega}^{\top}\theta,\quad\tau(\theta)\in[\underline{\tau}(\theta),\bar{\tau}(\theta)]$
+where linearity can be justified as a first-order linear approximation for $\bar{\tau}(\theta)$ and $\underline{\tau}(\theta)$ under the local asymptotics in Section 5.3 for smooth functions $\bar{\tau},\underline{\tau}$.
+[^12] When the envelope for $\tau(\theta)$ is non-smooth, we recommend replacing it with smooth upper and lower envelopes.
+In this case, following Section 5.3, $\bar{\omega}$ and $\underline{\omega}$ can be interpreted as the gradients of these smooth bounds evaluated at the observational estimates.
+We implicitly assume that at least one of $\bar{\omega}$ and $\underline{\omega}$ differ from the zero vector to avoid trivial solutions.
+For a generic weight vector $\omega$ let $\alpha_{\omega}(W,\Sigma),\beta_{l,\omega}(W)$ as in Definition 3.
+Confidence intervals under bias.
+For a candidate design $(W,\Sigma)$ and a bias vector $b\in\mathbb{R}^{p}$, define two-sided $(1-\eta)$ lower and upper bounds for $\tau(\theta)$ by
+$\ell_{b}(W,\Sigma)$ $\equiv\underline{\tau}(\hat{\theta}(W))\;-\;\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b\;-\;z_{1-\eta/2}\,\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)},$
+$u_{b}(W,\Sigma)$ $\equiv\bar{\tau}(\hat{\theta}(W))\;-\;\bar{\omega}^{\top}\Gamma_{\Lambda}(W)b\;+\;z_{1-\eta/2}\,\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)},$
+where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile and $\hat{\theta}(W)$ is the GMM estimator in (17).
+These bounds adjust for the bias $\omega^{\top}\Gamma_{\Lambda}(W)b$ in $\omega^{\top}\hat{\theta}(W)$ for each envelope separately.
+As before, the Jacobian $\Lambda$ is known; when estimated see Section 5.3 and Remark 5.
+An audience, indexed by a worst-case bias radius $B\geq 0$, forms a worst-case confidence interval (i.e., a bias-aware confidence interval a la Armstrong and Kolesár (2018)) by choosing the most conservative endpoints over $b$ in the ambiguity set $\mathcal{B}_{l}(B)$ in (16):
+$L_{l,B}(W,\Sigma)~\equiv~\left[\inf_{b\in\mathcal{B}_{l}(B)}\ell_{b}(W,\Sigma),\;\sup_{b\in\mathcal{B}_{l}(B)}u_{b}(W,\Sigma)\right],$
+and we denote by $|L_{l,B}(W,\Sigma)|$ its total length.
+The researcher does not need to know the audience’s specific choice of $B$ and can simply report the point estimates and associated standard errors; the audience (indexed by $B$) can then form their own bias-aware confidence set from these statistics.
+The researcher would like to minimize the expected length of the confidence interval formed by the audience.
+An oracle that knows $B$ minimizes the worst-case expected length of the audience’s confidence length relative to the size of the sharp identified set (worst-case under the audience’s prior $b\in\mathcal{B}_{l}(B)$).
+Its (ex-ante) expected loss is defined as
+$\mathcal{L}_{l,B}(W,\Sigma)~\equiv~\sup_{b_{0}\in\mathcal{B}_{l}(B)}\Big\{\mathbb{E}_{W,\Sigma,b_{0}}\big[|L_{l,B}(W,\Sigma)|\big]-\big(\bar{\tau}(\theta)-\underline{\tau}(\theta)\big)\Big\}.$ (21)
+Here $\mathcal{L}_{l,B}(W,\Sigma)$ measures how informative the expected confidence interval is, relative to the length of the identified set $\bar{\tau}(\theta)-\underline{\tau}(\theta)$, worst-case over $b_{0}$ in $\mathcal{B}_{l}(B)$.
+Here, $b_{0}$ determines the expectation of $\hat{\theta}(W)$, and ultimately the length of the bias-aware identified set.
+As in previous sections, we define the corresponding regret as
+$\tilde{\mathcal{R}}_{l}(W,\Sigma)~\equiv~\sup_{B\geq 0}\;\frac{\mathcal{L}_{l,B}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\mathcal{L}_{l,B}(W^{\prime},\Sigma^{\prime})}\!.$
+Example 8 (Confidence set in a complete model).
+Consider a complete model with $\bar{\tau}(\theta)=\underline{\tau}(\theta)=\tau(\theta)$ and $\bar{\omega}=\underline{\omega}=\omega$. Then we can show
+$\mathcal{L}_{l,B}(W,\Sigma)=2\,z_{1-\eta/2}\sqrt{\alpha_{\omega}(W,\Sigma)}+2\,B\,\sqrt{\beta_{l,\omega}(W)}.$
+In particular, when the parameter is point identified, the excess confidence interval length depends only on the variance index $\alpha_{\omega}$ and the bias index $\beta_{l,\omega}$, corresponding to standard bias-aware confidence intervals (Armstrong and Kolesár, 2018).
+Example 9 (Simple incomplete model).
+[…]
+t=0t=0t=1t=1t=2t=2t=3t=3 Design
+Choose $(W,\Sigma)$
+Reporting
+Compute $\hat{\theta}(W)$
+[…]
+Report $\alpha_{\underline{\omega}}$, $\alpha_{\bar{\omega}}$
+Audience
+Observe $(W,\Sigma)$
+Choose $B\in\mathbb{R}$
+Form $L_{l,B}(W,\Sigma)$
+Ex-post loss
+Excess length
+$|L_{l,B}(W,\Sigma)|-(\bar{\tau}(\theta)-\underline{\tau}(\theta))$
+Figure 2: Timeline of design, reporting, audience confidence set, and loss.
+Regret optimal design (and estimator).
+Define
+$A(W,\Sigma)\equiv z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big),\quad C_{l}(W)\equiv\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}+\sqrt{\beta_{l,\bar{\omega}-\underline{\omega}}(W)},$
+and the corresponding oracle solutions
+$A^{\star}\equiv\inf_{(W,\Sigma)\in\mathcal{D}^{\prime}}A(W,\Sigma),\qquad C_{l}^{\star}\equiv\inf_{(W,\Sigma)\in\mathcal{D}^{\prime}}C_{l}(W).$
+The first component $A$ depends on the variance of the estimated parameters (and envelopes) and the second component $C_{l}$ depends on the bias.
+We now show how our results directly extend to this scenario.
+Theorem 3 (Regret for confidence interval length).
+Let Assumptions 2 and 3 hold and consider an estimator as in (17). Then, for any $(W,\Sigma)\in\mathcal{D}^{\prime}$,
+$\tilde{\mathcal{R}}_{l}(W,\Sigma)=\max\Big\{\frac{A(W,\Sigma)}{A^{\star}},\;\frac{C_{l}(W)}{C_{l}^{\star}}\Big\}.$
+Proof.
+See Appendix A.2. ∎
+Theorem 3 shows that the confidence-interval-length regret can be decomposed into a variance component $A(W,\Sigma)$ and a bias component $C_{l}(W)$, each divided by its smallest feasible value.
+This result mimics Theorem 2 allowing for partially identified models.
+It provides a ready-to-use expression for experimental design (and moment selection).
+The following corollary shows that whenever the model is point identified the solutions that minimize the regret relative to confidence interval length coincides with the MSE optimal solution we derived in Section 4.
+Corollary 1 (Equivalence of MSE and confidence-interval length optimal solutions).
+Let Assumptions 2 and 3 hold.
+Let $\alpha,\beta_{l},\alpha^{\star},\beta_{l}^{\star}$ be as in Definition 3, and consider an estimator as in (17).
+Suppose $\bar{\omega}=\underline{\omega}=\omega$ (complete model).
+Then, for any $(W,\Sigma)$,
+$\arg\min_{(W,\Sigma)\in\mathcal{D}^{\prime}}\mathcal{R}_{l}(W,\Sigma)~=~\arg\min_{(W,\Sigma)\in\mathcal{D}^{\prime}}\tilde{\mathcal{R}}_{l}(W,\Sigma),$
+where $\mathcal{R}_{l}(W,\Sigma)$ denotes the MSE-based regret in Equation (19).
+Proof.
+See Appendix A.3. ∎
+In summary, when $\tau(\theta)$ is point identified, minimizing the regret of the confidence interval length leads to the same optimal design as minimizing MSE-based regret.
+In partially identified models, Theorem 3 shows that the confidence interval length maintains the same bias-variance decomposition, a desirable property of the MSE solution in Section 4.
+5.2 Optimization with partial knowledge on the bias bound
+In this section we extend our results to settings where researchers have prior information about the bias upper bound.
+Following the framework in Section 4, suppose researchers know that $B\leq\bar{B}$ for a known upper bound $\bar{B}$, and consider an arbitrary (possibly weighted) norm $\|\cdot\|_{l}$ on the bias.
+The theorem below establishes the resulting characterization of the regret.
+Theorem 4 (Regret characterization).
+Consider the setting in Section 4.
+Let Assumptions 2 and 3 hold, and let $\mathcal{R}$ be as defined in Equation (19), with $\alpha_{\Omega},\beta_{l,\Omega},\alpha_{\Omega}^{\star},\beta_{l,\Omega}^{\star}$ as in Definition 3.
+Consider an estimator as in Equation (17).
+Then
+$\sup_{B\in[0,\bar{B}]}\ \frac{\sup_{b\in\mathcal{B}_{l}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\sup_{b^{\prime}\in\mathcal{B}_{l}(B)}\ \mathrm{MSE}_{b^{\prime}}(W^{\prime},\Sigma^{\prime})}\;=\;\max\!\left\{\,\frac{\alpha_{\Omega}(W,\Sigma)}{\alpha_{\Omega}^{\star}}\,,\ \frac{\alpha_{\Omega}(W,\Sigma)+\bar{B}^{2}\beta_{l,\Omega}(W)}{\min_{W^{\prime},\Sigma^{\prime}}\big\{\alpha_{\Omega}(W^{\prime},\Sigma^{\prime})+\bar{B}^{2}\beta_{l,\Omega}(W^{\prime})\big\}}\,\right\}.$ (22)
+Proof.
+See Appendix A.1. ∎
+Theorem 4 shows that when an upper bound $\bar{B}$ on the bias is known, the worst-case proportional regret over all $B\in[0,\bar{B}]$ is again the maximum of two terms.
+The first term, $\alpha_{\Omega}(W,\Sigma)/\alpha_{\Omega}^{\star}$, is the familiar variance regret.
+The second term compares the MSE evaluated at the largest feasible bias level, $\bar{B}$, to the smallest achievable MSE at the same bias level across all feasible designs and weight matrices.
+The optimization problem is therefore analogous to our baseline case, but with the “bias” component of regret evaluated at $B=\bar{B}$ rather than in the limit as $B\to\infty$.
+From an interpretation standpoint, this extension interpolates between two polar cases.
+When $\bar{B}$ is small, the second term is close to the variance ratio, and designs that minimize variance dominate.
+As $\bar{B}$ increases, the second term places more weight on robustness to misspecification, and the objective approaches the fully robust case studied in our baseline results.
+In practice, the theorem implies that any prior information that narrows $B$ to $0,\bar{B}]$ can be incorporated simply by replacing the worst-case bias component in our original criterion with the MSE at $B=\bar{B}$, while keeping the same quasi-convex structure.
+5.3 Local asymptotics with non linear target
+Next, we extend our framework for (i) nonlinear moment functions $g(\theta)$ used in Section 4 and (ii) a smooth, potentially nonlinear estimand $\tau(\theta)$.
+Although we focus on known $\tau(\cdot)$, similar reasoning applies to partially identified sets in Section 5.1, omitted for brevity.
 Linearization of the moments
-Assume the sample moments obey root–$n$ scaling, i.e., $\mathbb{V}(\bar{g}_{\Sigma})=O(n^{-1})$.
-Let $b=b_{n}$ be a sequence where retain the same misspecification structure as in Section 4: the moment vector has mean $b=b_{n}$ with $(b_{n})_{j}=0$ for $j\in\mathcal{I}$ (experiment-based moments) and $(b_{n})_{\mathcal{I}^{c}}$ potentially nonzero.
+Assume the sample moments obey root-$n$ scaling, i.e., $\mathbb{V}(\bar{g}_{\Sigma})=O(n^{-1})$.
+Let $b=b_{n}$ be a sequence where we retain the same misspecification structure as in Section 4: the moment vector has mean $b=b_{n}$ with $(b_{n})_{j}=0$ for $j\in\mathcal{I}$ (experiment-based moments) and $(b_{n})_{\mathcal{I}^{c}}$ potentially nonzero.
 Our key assumption is a local condition on the moments bias of the form
@@
 This condition builds on the local asymptotic frameworks in Andrews et al.
-(2020) and Armstrong and Kolesár (2021).
-In their context, $b_{n}||^{2}=1/n$ matching the rate of convergence of the variance.
-In our case, $b_{n}$ can grow faster, slower or at the same rate of the standard error, encompassing $b_{n}||^{2}=1/n$ as a special case.
+(2020) and Armstrong and Kolesár (2021), where $b_{n}||^{2}=1/n$ matching the rate of convergence of the variance.
+In our case, $b_{n}$ can grow faster, slower or at the same rate of the standard error (i.e., $b_{n}\propto n^{-\alpha},\alpha>1/4$), encompassing $b_{n}||^{2}=1/n$ as a special case.
 Under standard smoothness conditions, a GMM estimator with weighting matrix $W$ satisfies
-$\sqrt{n}\big(\hat{\theta}(W,\Sigma)-\theta\big)=-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)+o_{p}(1),$
-so on the original (unscaled) scale the mean shift is $(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,b_{n}$ and the remainder is $o_{p}(n^{-1/2})$.
+$\sqrt{n}\big(\hat{\theta}(W,\Sigma)-\theta\big)=-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)+o_{p}(1).$
 The condition $\|b_{n}\|^{2}\sqrt{n}\to 0$ ensures that higher order terms are negligible under sufficient smoothness conditions, justifying the use of the linear expansion (17) for our design analysis.
-Because $b_{n}\to 0$, the Jacobian $\Lambda$ can be consistently estimated by evaluating the derivative of $g(\theta)$ at a preliminary observational estimator $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$, assuming such estimates are available from an observational study (Andrews et al., 2020).
+Because $b_{n}\to 0$, the Jacobian $\Lambda$ can be consistently estimated by evaluating the derivative of $g(\theta)$ at a preliminary observational estimate $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$ (Andrews et al., 2020).
 Linearization of the estimand.
@@
 A Taylor expansion around $\theta$ yields
-$\tau(\hat{\theta})-\tau(\theta)=\omega(\theta)^{\top}(\hat{\theta}-\theta)\;+\;\tfrac{1}{2}\,(\hat{\theta}-\theta)^{\top}H_{\tau}(\bar{\theta})\,(\hat{\theta}-\theta),$
-for some $\bar{\theta}$ between $\hat{\theta}$ and $\theta$, and $H_{\tau}$ is the Hessian. It follows that we can write
 $\sqrt{n}\big(\tau(\hat{\theta})-\tau(\theta)\big)=\omega(\theta)^{\top}\,\Gamma_{\Lambda}(W)\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)\;+\;o_{p}(1),$
@@
 Consequently, the regret expressions derived for the linear case continue to apply up to $o_{p}(1)$, with $\omega$ interpreted as the gradient $\omega(\theta)$ (or any consistent plug-in $\omega(\tilde{\theta}^{\mathrm{obs}})$ from a preliminary estimator $\tilde{\theta}^{\mathrm{obs}}\to_{p}\theta$).
-Remark 2 (Rescaling and change of variables linking Example 3).
+Remark 5 (Ex-post valid coverage with estimated $\Lambda$).
+When some components of $\hat{\theta}$ are based on observational data and $\Lambda$ is estimated from the same observational sample used to select the design, the selection of the experiment and the construction of $\hat{\theta}$ both depend on the same data.
+Unless the observational sample is (asymptotically) much larger than the experimental sample, this case can lead to selection-induced distortions in the distribution of $\hat{\theta}$ and invalidate standard Gaussian confidence intervals (see, e.g., Andrews et al., 2024).
+A simple way to avoid this problem is to split the observational data into two independent subsamples (independent of the experimental samples): one used to estimate $\Lambda$ and choose the experiment at the design stage, and a second used, together with the experimental data to estimate the observational and experimental components of $\hat{\theta}$.
+Under this sample-splitting scheme the experimental and observational data used for inference are independent of the design choice, so standard CLT-based confidence intervals retain their nominal coverage.
+∎
+Remark 6 (Equivalent rescaling).
 An equivalent formulation uses $\sqrt{n}$–rescaled coordinates (as in Example 4).
 Define the rescaled moment vector $\tilde{g}:=\sqrt{n}\,\bar{g}$ and the rescaled estimation error $\tilde{\theta}\equiv\sqrt{n}\,(\hat{\theta}-\theta)$.
-Then
-$\mathbb{V}(\tilde{g})=\Sigma=O(1),\qquad\mathbb{E}[\tilde{g}]=b_{n}\ \ \text{with}\ \ b_{n}\equiv\sqrt{n}\,b,$
-so $b_{n}$ is the asymptotic bias and $\Sigma$ the asymptotic variance.
-The local misspecification condition in (9) is equivalent to $\frac{\|b_{n}\|^{2}}{\sqrt{n}}\;\to\;0,$ which coincides with Example 3.
+Then $\mathbb{V}(\tilde{g})=\tilde{\Sigma}=O(1),\mathbb{E}[\tilde{g}]=\tilde{b}_{n}\ \ \text{with}\ \ \tilde{b}_{n}\equiv\sqrt{n}\,b,$ so $\tilde{b}_{n}$ is the asymptotic bias and $\tilde{\Sigma}$ the asymptotic variance.
+Local misspecification in (23) is equivalent to $\frac{\|\tilde{b}_{n}\|^{2}}{\sqrt{n}}\;\to\;0,$ which coincides with notation in Example 4.
 ∎
-5.2 Confidence interval length as a regret criterion
-In this subsection we consider an alternative loss based on the length of confidence intervals (CIs), focusing on Setting 1 for simplicity.
-For a candidate design $(\mathcal{E},\Sigma(\mathcal{E}))$, shrinkage weights $\gamma$, and a bias vector $b\in\mathbb{R}^{p}$, define the two–sided, $(1-\eta)$ lower and upper bounds for $\tau(\theta)$ by
-$\ell_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;-\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
-$u_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;+\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
-where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile.
-An audience, indexed by a worst–case bias radius $B\geq 0$, forms the worst–case CI
-$L_{B}(\mathcal{E},\Sigma,\gamma)~\equiv~\left[\inf_{\|b\|_{\infty}\leq B}\ell_{b}(\mathcal{E},\Sigma,\gamma),\;\sup_{\|b\|_{\infty}\leq B}u_{b}(\mathcal{E},\Sigma,\gamma)\right],$
-where we write $|L_{B}(\mathcal{E},\Sigma,\gamma)|$ for its total length. The proportional regret of $(\mathcal{E},\Sigma,\gamma)$ is
-$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{B\geq 0}\;\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{\inf_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}\!.$
-Theorem 3.
-Consider Setting 1 and let Assumption 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
-[…]
-Proof.
-See Appendix B.1
-Theorem 3 shows that the objective function is the same as the MSE-objective function up to a monotonic (square-root) transformation which does not affect the minimizer.
-5.3 Vector valued estimands
-Next, we extend our framework to vector value estimands of the form $\tau(\theta)\in\mathbb{R}^{q}$ for $q\geq 1$.
-Setting 3 (Vector-valued estimands).
-Consider a vector-valued target estimand with entries $\tau^{\ell}(\theta)\equiv(\omega^{\ell})^{\top}\theta$ for $\ell=1,\dots q$, and $\omega^{\ell},\theta\in\mathbb{R}^{k}$.
-Given $(\mathcal{E},\Sigma)$, and observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}},\tilde{\theta}^{\mathrm{exp}}$, consider a class of linear plug-in estimators to allow the optimal shrinkage weights to depend on the specific entry of $\tau$.
-$\hat{\tau}^{\ell}_{\gamma}\equiv\tau^{\ell}(\hat{\theta}^{\ell}(\gamma)),\qquad\hat{\theta}^{\ell}_{j}(\gamma)=\begin{cases}\gamma_{j}^{\ell}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j}^{\ell})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
-where $\gamma=(\gamma^{\ell}_{j})_{\ell\in\{1,\cdots,q\},j\in\mathcal{E}}\in\mathbb{R}^{q\times|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$.
-Let $\hat{\theta}(\gamma)\equiv(\hat{\theta}^{1}(\gamma),\dots,\hat{\theta}^{q}(\gamma))\in\mathbb{R}^{p\times q}$ and $\hat{\tau}_{\gamma}\equiv(\hat{\tau}^{1}_{\gamma}(\hat{\theta}^{1}(\gamma)),\dots,\hat{\tau}^{q}_{\gamma}(\hat{\theta}^{q}(\gamma)))^{\top}\in\mathbb{R}^{q}$.
-Setting 3 is a flexible generalization of Setting 1, allowing each entry of $\tau$ to have its own vector of shrinkage weights.
-Assumption 4 (First-order estimation error).
-For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume for all $\ell=1,\dots,q$
-$\tau^{\ell}(\theta)-\tau^{\ell}\!\big(\hat{\theta}^{\ell}(\gamma)\big)=(\omega^{\ell})^{\top}\,\big(\theta-\hat{\theta}^{\ell}(\gamma)\big),$ (10)
-for a known weighting vector $\omega^{\ell}\in\mathbb{R}^{p}$, with $|\omega^{\ell}_{i}|\in(0,\infty)$, for all $i,\ell$.
-We begin by generalizing the mean squared error to higher-dimensional parameters.
-For a given design $(\mathcal{E},\Sigma)$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
-$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[\|\widehat{\tau}_{\gamma}-\tau\|_{2}^{2}\right],$ (11)
-where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$, and $\|\cdot\|_{2}$ denotes the $L^{2}$ Euclidean norm.
-We introduce more compact notation before proceeding to the main result. For a given $(\mathcal{E},\gamma)$ define
-$\tilde{\omega}^{\ell}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega^{\ell}_{j},&j\notin\mathcal{E},\\ \omega^{\ell}_{j}(1-\gamma^{\ell}_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}^{\ell}_{\text{exp},j}(\mathcal{E},\gamma)=\begin{cases}0,&j\notin\mathcal{E},\\ \omega^{\ell}_{j}\gamma^{\ell}_{j},&j\in\mathcal{E},\end{cases}$
-and let $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}^{\ell}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}^{\ell}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$ We can write $\hat{\tau}^{\ell}_{\gamma}=\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}\,\hat{\theta}$, where $\hat{\theta}=(\tilde{\theta}^{\text{obs}\top},\tilde{\theta}^{\text{exp}\top})^{\top}\in\mathbb{R}^{p+|\mathcal{E}|}$.
-Stacking across the $\ell$ index gives the estimator for the full vector $\tau$: $\hat{\tau}_{\gamma}=\tilde{\omega}(\mathcal{E},\gamma)\,\hat{\theta}$, where the $\ell^{th}$ row of $\tilde{\omega}(\mathcal{E},\gamma)\in\mathbb{R}^{q\times(p+|\mathcal{E}|)}$ is $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}$.
-The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
-$\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ $\equiv\text{Trace}\left(\tilde{\omega}(\mathcal{E},\gamma)\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\right),$ (12)
-$\check{\beta}(\mathcal{E},\gamma)$ $\equiv\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}\bar{v}(\gamma)-(\omega^{\ell}_{\,\mathcal{E}})^{\top}\bar{v}(\gamma)_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(\bar{v}(\gamma)_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
-where $\omega^{\ell}_{\mathcal{E}}$ is the subvector of $\omega^{\ell}$ corresponding to the entries indexed by the experiment $\mathcal{E}$, and
-$\bar{v}(\gamma)\in\operatorname*{arg\,max}_{\{v\in\mathbb{R}^{p}:|v_{j}|\leq 1\forall j\}}\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\,\mathcal{E}})^{\top}v_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(v_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
-where the dependence of $\bar{v}(\gamma)$ on $\mathcal{E}$ is omitted for conciseness and $\circ$ denotes the component-wise multiplication. The smallest worst-case variance and bias are defined as
-$\check{\alpha}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\beta(\mathcal{E},\gamma).$
-As before, we are interesting in minimizing
-$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\equiv\sup_{B\geq 0}\frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}.$
-The result below shows how our results directly extend to this case.
-Theorem 4.
-Consider Setting 3 and let Assumption 4 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
-$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\check{\alpha}(\mathcal{E},\Sigma,\gamma)}{\check{\alpha}^{\star}},\ \ \frac{\check{\beta}(\mathcal{E},\gamma)}{\check{\beta}^{\star}}\right\}.$
-Proof.
-See Appendix B.2
 6 Empirical applications
-In this section we provide two empirical applications, one for measuring GE effect and the second one for externally valid site selection.
-6.1 An implementation guide for measuring GE effects
-Combining experimental or pre-program observational data with structural models is increasingly used for program evaluation (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Allcott et al., 2025).
-[^8] Across all AEA articles in 2015-2025, about $10\%$ of the experimental papers combine experimental results with structural models (i.e., one-third of the papers that use observational in combination with experimental estimates).
-Because large-scale experimentation are often infeasible due to cost constraints (Muralidharan and Niehaus, 2017) these models are meant to extrapolate effects from small-scale experiments to learn general-equilibrium (GE) effects.
-This extrapolation typically relies on other experiments or observational data.
-We present a (step-by-step) illustration below when these models are calibrated using previous studies conducted in a different country.
-Question
-Consider a researcher evaluating a cash transfers for sending children to school in rural Kenya where such a program is not in place.
+6.1 Experimental design for General Equilibrium effects
+Because large-scale experimentation is often infeasible due to cost constraints (Muralidharan and Niehaus, 2017), researchers often combine experimental and observational data to extrapolate effects from small-scale experiments to learn general-equilibrium effects (e.g.
+Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022).
+The goal of this subsection is to present a step-by-step practitioner’s guide with an example, and compare our regret-optimal design to a standard variance-minimizing (Neyman) benchmark.
+Consider a researcher evaluating conditional cash transfers for sending children to school in rural Kenya.
 Due to budget constraints, the researcher can only randomize at a small scale (partial equilibrium), but ultimately wishes to predict GE effects.
-Researchers have access to preliminary estimates from the Mexican PROGRESA experiment (Todd and Wolpin, 2006), which we may be concerned, lack external validity.
-6.1.1 Step 1: model description
-The first step for a researcher is to describe the modeling choices. We present a stylized model of school choice from Bonhomme and Weidner (2022) and Todd and Wolpin (2006).
+Researchers have access to preliminary (external) estimates from the Mexican PROGRESA experiment (Todd and Wolpin, 2006), although these preliminary estimates may lack external validity in Kenya.
+Step 1: model description
+The first step for a researcher is to describe how observational/external and experimental estimates in Kenya are jointly used for estimation.
+Here, we consider a stylized model of school choice following Bonhomme and Weidner (2022), Todd and Wolpin (2006).
 Individual choices
@@
 Abstracting from covariates (we will introduce covariates in the estimation), utility is
-$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{3}-\xi_{1}-\xi_{2})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (13)
+$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{2}-\xi_{3}-\xi_{1})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (24)
 with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS.
-The parametrization $(\xi_{3}-\xi_{1}-\xi_{2})$ is without loss and simplifies expressions below.
-Enrollment satisfies $S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$ Letting $Z(Y,W,t)\equiv\xi_{1}W-\xi_{2}Y-\xi_{3}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$, with $\Phi(\cdot)$ denoting the standard Gaussian CDF.
-General equilibrium effects
+The parametrization $(\xi_{2}-\xi_{3}-\xi_{1})$ is without loss and simplifies expressions below.
+Enrollment satisfies $S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$ Letting $Z(Y,W,t)\equiv\xi_{3}W-\xi_{1}Y-\xi_{2}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$, with $\Phi(\cdot)$ denoting the standard Gaussian CDF.
+Model for GE effects
 We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya.
-GE feedback is allowed through income and wages: for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, let
-$Y_{i}(t)=y(t)+\varepsilon_{Yi},\qquad W_{i}(t)=w(t)+\varepsilon_{Wi}.$
+GE feedback is allowed through income and wages which we write as functions of a common transfer $t$ to all eligible households in the population.
+Specifically for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, for a common transfer $t$, denote the potential individual income and wage [^13] Because we focus on the effect of a common transfer, we keep a scalar notation for $t$ for expositional convenience.
+Alternatively, here $y(t),w(t)$ can also be interpreted as exposure mappings (Aronow et al., 2017) and $t$ can be defined as a $n$ dimensional vector of transfers.
+$Y_{i}(t)=y(t)+\varepsilon_{Yi},\hskip 18.49988ptW_{i}(t)=w(t)+\varepsilon_{Wi}.$
 Let $y_{0}\equiv\partial_{t}y(t)|_{t=0}$ and $w_{0}\equiv\partial_{t}w(t)|_{t=0}$.
@@
 Our estimand of interest is the marginal effect of an increase in a small transfer $t$ to all eligible households
-$\left.\frac{\partial\,\mathbb{E}\!\left[\Pr(S{=}1\mid Y(t),W(t),t)\right]}{\partial t}\right|_{t=0}=\;\phi_{0}\cdot\big(\,\xi_{3}+\xi_{2}y_{0}-\xi_{1}w_{0}\,\big),$ (14)
-which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{2}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$.
-Under simple market clearing for labor supply and demand, we can show [^9] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$.
+$\mathbb{E}\!\left[\left.\frac{\partial\,\Pr(S{=}1\mid Y(t),W(t),t)}{\partial t}\right|_{t=0}\right]=\;\phi_{0}\cdot\big(\,\xi_{2}+\xi_{1}y_{0}-\xi_{3}w_{0}\,\big),$ (25)
+which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{1}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$.
+Under standard market clearing for labor supply and demand, we write [^14] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$.
 Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working.
-From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{3}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{1}$ at $t=0$.
-Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{3}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{1}H$.
+From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{2}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{3}$ at $t=0$.
+Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{2}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{3}H$.
 Dividing the equation by $H$ we obtain the desired result.
-$w_{0}=\;\frac{\phi_{0}\,\xi_{3}}{\phi_{0}\,\xi_{1}-d},$ (15)
+$w_{0}=\;\frac{\phi_{0}\,\xi_{2}}{\phi_{0}\,\xi_{3}-d},$ (26)
 where $d$ denotes the slope of the demand curve divided by total hours of work.
-6.1.2 Step 2: parametrization and transparency on biases
-Given the model, the following step for researchers is to define the parameters of interest. Here, we define the parameters of interest as
-$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{3},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{1},$ (16)
+Wages in equilibrium depend both on the demand’s slope and the supply effect, captured through $\phi_{0}\xi_{2}$ (assuming that all non-students are employed for a constant number of hours).
+Step 2: parametrization and transparency on biases
+The following step is to define the parameters of interest, here defined as
+$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{1},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{3},$ (27)
 with corresponding estimand
 $\tau(\theta)\;=\;\underbrace{\theta_{2}}_{\text{Direct effect}}\;+\;\underbrace{y_{0}}_{\text{Income multiplier}}\,\underbrace{\theta_{1}}_{\text{Income effect}}\;+\;\underbrace{w_{0}(\theta)\,\theta_{3}}_{\text{Wage effect}},\qquad w_{0}(\theta)\;=\;\frac{\theta_{2}}{-\theta_{3}-d},$ (28)
-where the explicit form of $w_{0}(\theta)$ follows from (15).
-We treat $d$ as known for simplicity, using meta-analyses on demand elasiticities, and $y_{0}=1.5$ because computed experimentally in Kenya by Egger et al.
+where the explicit form of $w_{0}(\theta)$ follows from (26), a smooth function of the parameters.
+We treat $d$ as known for simplicity, using meta-analyses on demand elasticities, and $y_{0}=1.5$ based on experimental estimates in Kenya from Egger et al.
 (2022).
@@
 Specifically, $d=0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained from meta-studies in Espey and Thilmany (2000).
-[^11] One could also calibrate $S_{0},W_{0}$ to pre-experimental data in Kenya directly when available.
-In this example, the primary concern is bias in the observational estimator $\tilde{\theta}^{\mathrm{obs}}$ constructed as described below, abstracting from potential biases in $y_{0},d$ which, for simplicity, we assume to be second-order.
-When these biases are considered first-order, researchers should let $y_{0},d$ be additional parameters in $\theta$.
-This difference illustrates a feature of our framework: the chosen parameterization ensures transparency about which biases’ concerns drive the design; these considerations can be collected in a pre-analysis plan.
-6.1.3 Step 3: observational study estimates
-Using experimental data from Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ via probit with standard controls (age, distance to school, eligibility, year, highest grade attained).
-Estimation is conducted separately by gender, and we will be focusing here on the effects on female students.
+[^16] One could also calibrate $S_{0},W_{0}$ to pre-experimental data in Kenya directly when this data is available.
+Here, the primary concern is bias in $\tilde{\theta}^{\mathrm{obs}}$, abstracting from potential biases in $y_{0},d$.
+When biases of $(y_{0},d)$ are also considered first-order, researchers should let $y_{0},d$ be additional parameters in $\theta$, which is allowed in our framework.
+This difference illustrates a feature of our framework: the chosen parameterization ensures transparency about which bias concerns drive the design; these considerations can be collected in a pre-analysis plan.
+Step 3: observational study estimates
+Using experimental data from the large-scale Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ through probit with standard controls (age, distance to school, eligibility, year, highest grade achieved) and use these estimates to obtain $\tilde{\theta}^{\mathrm{obs}}$.
+Estimation is conducted separately by gender; we will focus here on the effects on female students.
 We pool treated and control observations to estimate the schooling effect $\xi_{4}$, controlling for the individual-level subsidy.
@@
 Point estimates, estimated sensitivity $\omega$, and standard errors are reported below.
-Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sigma$
+Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sqrt{\Sigma_{jj}^{\mathrm{obs}}}$
 […]
@@
 Variance–covariance matrix of observational estimates for female students.
-We observe heterogeneity in both the sensitivity weights $\omega$ the precision of the observational estimates.
-6.1.4 Step 4: describing the candidate experiments
-The following step is to list the feasible experiments, subject to budget constraints.
-In this example, we consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, and examine three possible experiments for a given sample size $n$:
+Step 4: candidate experiments
+In this example, we consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, and examine three stylized experiments or combinations thereof:
   • $j=\{1\}$: Unconditional transfer (income shock).
 The researcher randomizes a small income shock to a small fraction of households (implying no general-equilibrium effects).
-Under a first-order (Taylor) approximation, the experiment identifies
-    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,0)]}{\partial Y}\right|_{t=0}=\theta_{1},$
-    with precision $v_{1}^{2}/n$. The researcher optimizes over $\gamma_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
+Under a first-order (Taylor) approximation, the experiment identifies the average marginal effect of income on school attendance, $\theta_{1}\;=\;\mathbb{E}\left[\left.\frac{\partial}{\partial Y}\Pr\big(S{=}1\mid Y,W,t\big)\right|_{t=0}\right].$ The researcher optimizes over $\gamma_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to PROGRESA.
   • $j=\{2\}$: Conditional cash transfer (stipend).
 The researcher randomizes a small stipend $t$, conditional on attending school, to a small fraction of households (with prices held fixed).
-Under a first-order approximation, this identifies
-    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,t)]}{\partial t}\right|_{t=0}=\theta_{2},$
-    with precision $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
-  • $j=\{1,2\}$: Two-arm design.
-The researcher runs (j=1j{=}1) and (j=2j{=}2) on independent samples, identifying $(\theta_{1},\theta_{2})$ with precisions $(v_{1}^{2}/n_{1},\ v_{2}^{2}/n_{2})$ where $n_{1}+n_{2}=n$.
-They optimize over both $\gamma_{1}$ and $\gamma_{2}$, while $\theta_{3}$ is calibrated to the PROGRESA study.
+Under a first-order approximation, this identifies the average marginal effect of the stipend on school attendance, $\theta_{2}\;=\;\mathbb{E}\left[\left.\frac{\partial}{\partial t}\Pr\big(S{=}1\mid Y,W,t\big)\right|_{t=0}\right],$ with variance $v_{2}^{2}/n$.
+The researcher optimizes over $\gamma_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
+  • $j=\{3\}$: Youth job program (wage shock).
+Last, consider a job program for teenagers, that hires teenagers with small randomized differential wages.
+The program identifies $\theta_{3}$, holding constant $\theta_{1}$ and $\theta_{2}$
+The per-unit variance in the experiment is calibrated to the corresponding entry of $\sigma^{2}$ in observational study, appropriately multiplied by the observational study sample size.
 We consider two scenarios.
-In the first, the researcher (and oracle) can choose one or both experiments.
-In the second, they can choose only one of the two (either $j=\{1\}$ or $j=\{2\}$).
-Regret is defined accordingly, conditional on the set of feasible designs.
-For simplicity, we calibrate $v_{1}^{2}$ as $\sigma_{1}^{2}\times n^{\mathrm{obs}}$, where $\sigma_{1}^{2}$ is the variance of $\tilde{\theta}_{1}^{\mathrm{obs}}$ and $n^{\mathrm{obs}}$ is the observational-sample size ($n^{\mathrm{obs}}=1089$ for the female sample using data from eligible students in Todd and Wolpin (2006)).
-We calibrate $v_{2}^{2}$ analogously as $\sigma_{2}^{2}\times n^{\mathrm{obs}}$.
-In practice, researchers may use pilot studies to calibrate the experimental variances when available.
-Experiment type Identified parameter Choice Variables Calibrated standard error Description
-$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $6.56e^{-5}/\sqrt{n}$ Unconditional transfer (income shock, UCT). Randomize an income shock to a small fraction of households.
-$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $9.86e^{-4}/\sqrt{n}$ Conditional cash transfer (education shock, CCT). Randomize a conditional transfer to a small fraction of households.
-$j=\{1,2\}$ $(\theta_{1},\,\theta_{2})$ $\gamma_{1},\ \gamma_{2},\ n_{1}$ $6.56e^{-5}/\sqrt{n_{1}}$ (arm 1), $9.86e^{-4}/\sqrt{n-n_{1}}$ (arm 2) Two-arm design.
-Run (j=1j{=}1) and (j=2j{=}2) on independent samples with sample sizes respectively $n_{1},n-n_{1}$.
+First, the researcher may run only one of the three programs.
+Second, the researcher may run combinations of two of the three programs allocating a fixed total sample size to each, but not all three simultaneously, reflecting higher fixed costs and institutional constraints.
+For instance, the job program requires a different implementation infrastructure than a cash-transfer program; at the same time, conducting conditional and unconditional cash transfers in the same experimental location is often infeasible due to fairness concerns (e.g.
+Wingfield et al., 2023), which raises costs for each cash-transfer arm.
+Remark 7 (Additional constraints).
+In some contexts additional constraints may apply.
+For instance, we may be able to only run either $\{1,2\}$ or $\{3\}$, or only run either $\{1\}$ or $\{2\}$ but not $\{3\}$.
+Researchers can directly include these constraints in our framework.
+∎
+Experiment type Identified parameters Choice Variables Calibrated standard errors Description
+$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $\frac{6.56e^{-5}}{\sqrt{n_{\mathrm{tot}}/n_{\mathrm{obs}}}}$ Unconditional transfer (income shock, UCT).
+Randomize an income shock to a small fraction of households.
+$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $\frac{9.86e^{-4}}{\sqrt{n_{\mathrm{tot}}/n_{\mathrm{obs}}}}$ Conditional cash transfer (education shock, CCT).
+Randomize a conditional transfer to a small fraction of households.
+$j=\{3\}$ $\theta_{3}$ $\gamma_{3}$ $\frac{1.01e^{-3}}{\sqrt{n_{\mathrm{tot}}/n_{\mathrm{obs}}}}$ Job program (Wage shock). Teenager jobs program with (small) randomized wages differentials.
+$\{j,j^{\prime}\}$ $(\theta_{j},\theta_{j^{\prime}})$ $\gamma_{j},\gamma_{j^{\prime}},n_{j}$ $\frac{\sqrt{\Sigma_{jj}}}{\sqrt{n_{j}/n_{\mathrm{obs}}}}$, $\frac{\sqrt{\Sigma_{j^{\prime}j^{\prime}}}}{\sqrt{(n_{\mathrm{tot}}-n_{j})/n_{\mathrm{obs}}}}$ Combination of two of the three possible experiments.
+Choose two treatment arms and allocate optimally sample size, with given total experimental sample size $n_{\mathrm{tot}}$.
 Table 1: Design options, identified parameters, choice variables, and per-unit variances.
-6.1.5 Step 5: experimental design
-Figure 4: Regret–optimal sample allocation across treatment arms when both the unconditional cash transfer (UCT; income shock) and the conditional cash transfer (CCT; stipend) arms are feasible.
-The optimal solution is interior: the vast majority of participants (over 90%) are assigned to CCT, with a small but nonzero fraction assigned to UCT to hedge misspecification.
-This pattern reflects the higher payoff to learning about CCT in this application relative to UCT.
-Figure 5: Optimal shrinkage weights and design choice as functions of the total sample size $n$.
-Left panel: When both arms are run, the weight on the UCT experimental estimator is essentially one across $n$, while the weight on the CCT experimental estimator, $\gamma^{\star}_{\text{CCT}}$, rises with precision (from about 0.5 at small $n$ to roughly 0.9 by $n=1000$), reflecting the increasing value of experimental evidence.
-Right panel: When restricted to a single arm, the optimal choice switches from UCT at small $n$ (variance dominates under CCT) to CCT once $n$ crosses a threshold (around $n\approx 500$), where bias considerations dominate and favor CCT.
-Notes: $\gamma_{j}^{\star}$ is the optimal weight on the experimental (vs.
-observational) estimator for parameter $j$; we set $\gamma_{j}^{\star}=0$ to indicate that arm $j$ is not chosen.
-Figure 6: Worst–case regret by design.
-Left panel: Allowing both arms with an interior allocation yields regret that approaches one as $n$ increases, indicating near–oracle performance under our normalization.
-In contrast, single–arm designs exhibit nonvanishing regret because bias remains first–order even as variance shrinks.
-Right panel: Focusing on the single–arm problem, UCT dominates only at small $n$; beyond the same threshold ($\approx 500$), CCT delivers uniformly lower regret as variance becomes second–order relative to bias.
-Notes: Regret is the worst–case objective normalized so that a value of one corresponds to the oracle benchmark; UCT = unconditional cash transfer, CCT = conditional cash transfer.
-The final step is to design the experiment and report it in a pre-analysis plan.
-Choice of the experiment when only one treatment arm is available
-Figure 5 (right panel) displays the selected arm as a function of $n$ (together with the corresponding $\gamma_{j}^{\star}$).
-We encode “not chosen” as $\gamma_{j}^{\star}=0$.
-For small samples ($n\lesssim 500$), the optimal choice is UCT: the experimental CCT estimator would be too noisy, so the resulting variance under CCT dominates, while the bias advantage of CCT over UCT is limited because $\omega_{1}$ and $\omega_{2}$ are of similar magnitude.
-Once $n$ crosses $n\approx 500$, CCT becomes dominant: by combining observational and experimental estimates for CCT, we can achieve comparable variance, and the worst-case bias is lower than under UCT.
-Choice of the estimator
+Step 5: experimental design
+Figure 3: Optimal experimental sample sizes $n_{j}$ allocated to each arm (UCT, CCT, and Job program) as a function of the total experimental sample size $n_{\text{tot}}$.
+The left panel considers designs with only one treatment arm, while the right panel allows two treatment arms.
+Figure 4: Variance, bias, and adaptive regret of the Neyman (variance–optimal) allocation relative to the proposed optimal design, as a function of the total experimental sample size $n_{\text{tot}}$.
+Left panel: one feasible experiment; right panel: two feasible experiments.
+Finally, we design the experiment, with total number of experimental observations $n_{\mathrm{tot}}$; let $n_{\mathrm{obs}}$ denote instead the number of units in the original Mexican PROGRESA experiment.
+Choice of the experiment and sample size
+In Figure 3 we show the optimal sample size allocation as a function of the total experimental sample size, under the constraints that either one (left panel) or two (right panel) treatment arms can be selected.
+When the total sample size is small and only one arm can be chosen, the optimal design selects the UCT arm.
+Intuitively, the UCT strikes a good balance between sensitivity and precision of the experimental estimate when $n_{\text{tot}}$ is small.
+When two treatment arms can be selected, the optimal design initially chooses both UCT and CCT for $n_{\mathrm{tot}}\leq 500$.
+In this region, almost all of the experimental sample is allocated to the CCT arm (which is more sensitive but noisier), with a smaller fraction allocated to the more precise UCT arm.
+For total sample sizes above five-hundred, the optimal design instead selects the job program as the main experiment, reflecting the high sensitivity of the estimand to the wage effect (Appendix Figure 12 reports the bias and variance regret of the design).
+With two experiments and large sample sizes, roughly 90% of the total sample is allocated to the job program and the remaining fraction to the UCT arm, using the precise UCT effect to help balance variance against bias.
 In addition to sample allocation, our method optimizes the shrinkage weights of the parameters of interest.
-Figure 5 (left panel) shows that when both arms can be run, we place essentially full weight on the experimental estimator for UCT ($\gamma^{\star}_{\text{UCT}}\approx 1$ across $n$), while for CCT we gradually shift toward the experimental estimator as precision improves: $\gamma^{\star}_{\text{CCT}}$ rises from about $0.5$ at small $n$ to approximately $0.9$ by $n=1000$.
-Intuitively, higher $n$ lowers experimental variance, making it optimal to rely more heavily on experimental evidence for the CCT parameter.
-Sample size allocation when both arms are feasible
-Figure 4 reports the sample allocation implied by the regret-optimal design when both treatment arms are feasible.
-The optimal allocation is interior: most participants (over roughly $90\%$) are assigned to the CCT arm (education), with a small but nonzero fraction assigned to the UCT arm (income shock).
-Two forces rationalize this pattern.
-First, the CCT parameter is more misspecification-sensitive ($\omega_{2}\approx 1.98$ vs.
-$\omega_{1}\approx 1.5$), so generating experimental evidence on CCT has a larger payoff in bias reduction.
-Second, the income effect $\theta_{1}$’s variance is about an order of magnitude smaller than the stipend effect $\theta_{2}$’s variance in our PROGRESA calibration, lowering the marginal returns to learn about UCT relative to CCT.
-Hence, while the designer keeps some allocation on UCT, most of the sample is placed on CCT; our framework makes this bias–variance trade-off explicit.
-Regret comparisons
-Figure 6 reports worst-case regret for each design choice.
-The left panel shows that the regret of our method (green line, running both arms with an interior allocation) declines toward one as $n$ increases, reflecting near-oracle performance under our normalization.
-In contrast, single-arm designs exhibit nonvanishing regret because bias remains first order even as variance shrinks.
-The right panel focuses on the single-arm problem: both the oracle and the researcher can only choose a single arm.
-In this case, UCT is preferable only for small $n$; beyond the same threshold ($\approx 500$), CCT yields uniformly lower regret as variance becomes second order relative to bias.
-The regret vanishes relative to the oracle that can only choose a single arm.
-Implications
-When only one arm is feasible, CCT is preferable only once the total sample size is sufficiently large; otherwise UCT is preferred.
-This pattern reflects the interaction between bias and variance: for small samples, CCT’s bias advantage is not large enough to compensate for its higher variance relative to UCT, whereas at larger samples the bias reduction becomes dominant.
-Existing designs based on variance considerations would fail to provide solutions that balance bias and variance comparisons.
-When feasible, it is instead optimal to run two arms with highly unbalanced sample, reflecting the relative contribution of the variance for each treatment arm.
-Following these steps, researchers can similarly report their design in a pre-analysis plan.
-6.2 Measuring the performance of our method for site selection
+Appendix Figure 13 collects the estimated shrinkage weight across designs.
+We find that the shrinkage weight is above 0.7, and closer to one for most cases.
+This shows how the experimental variation gets significantly higher weight.
+Main bias and variance regret comparisons
+In Figure 4 we compare the proposed optimal design to the variance-optimal (Neyman) design in terms of both variance and bias.
+This comparison clarifies what is gained by our procedure relative to a benchmark that chooses the experiment, shrinkage weights, and sample size solely to minimize the variance of the estimated effect $\hat{\tau}$ (and which therefore achieves the smallest possible variance).
+We find that the variance ratio between the Neyman and optimal designs (the inverse of the variance regret) is close to one for all sample sizes and converges to one as the sample size grows.
+For experimental sample $n_{\mathrm{tot}}=1000$, the variance of the Neyman and our proposed procedure is only about $14\%$ smaller when choosing one treatment arm and $35\%$ when choosing two treatment arms.
+In contrast, the bias of the Neyman procedure is about 1.4 times larger for one treatment arm and about 4.5 times larger with two.
+As a result, the adaptation regret of the Neyman allocation is similarly 1.4 and 4.5 times larger than our proposed design.
+These comparisons illustrate the benefit of our method: its variance is only slightly larger than the Neyman variance, while the bias is more than four times smaller for two treatment arms.
+6.2 Site selection for external validity
 Next, we show how observational evidence can inform whom to recruit into an experiment.
@@
 (2024) corresponding to the density of the network, i.e., the percentage of connections a random household in a village has relative to the village size.
-Background
+Background and research question
 In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance.
-Because program rollout was not randomized across villages, they estimate effects using Difference-in-Differences (DiD).
-While DiD is informative, the lack of randomized assignment may bias estimates in the presence of selection (e.g., Ghanem et al., 2022).
-Motivated by these limitations, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods.
+Because program rollout was not randomized, they estimate effects using Difference-in-Differences (DiD).
+Although DiD is informative, the lack of randomized assignment can bias estimates in the presence of selection (e.g., Ghanem et al., 2022).
+Therefore, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods.
 We only use observational variation for choosing the experimental design and estimator.
-We then use experimental variation generated by Banerjee et al.
-(2024) to validate our procedure.
-Research question
+We then use experimental variation from Banerjee et al.
+(2024) to validate our procedure by calibrating its worst-case bias.
 In practice, implementation costs often depend on how geographically dispersed the study sites are due to coordination and survey costs.
 Using preliminary observational estimates from Banerjee et al.
-(2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled?
-Figure 7: Observational villages and area-level DiD effects in Karnataka.
-The background heat map partitions the state into the four contiguous areas used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value.
-Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey).
-Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
+(2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled in each area?
+We focus on the experiment randomizing treatments in one or two areas.
 Observational study
 We partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period.
+Areas are heterogeneous in terms of their overall population size.
 Figure 5 maps these four areas and reports the corresponding DiD point estimates.
 Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas.
-Table 2 summarizes, for each area, samples sizes and variances, clustered at the village level.
+Table 2 summarizes, for each area, samples sizes, variances clustered at the village level, and population sizes ($\omega$).
 These area-specific DiD estimates and their variances serve as our primary observational inputs for experimental design.
@@
 An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity.
+Using the 2011 population census, matched with the geographic boundaries of the areas obtained from merging village positions with the SHRUG platform (Asher et al., 2019), we compute $\omega$ as the population share of a given area.
 Table 2: Area-level observational inputs for Karnataka villages.
@@
 For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$.
-Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$
-11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3
-22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140
-33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187
-44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130
+Here $\omega$ denotes the relative population share obtained using the 2011 population census in India.
+Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$ $\omega$
+11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3 $0.033$
+22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140 $0.766$
+33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187 $0.121$
+44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130 $0.078$
+Figure 5: Observational villages and area-level DiD effects in Karnataka.
+The background heat map partitions the area of interest into the four contiguous zones used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value.
+Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey).
+Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
 Experimental design
-We consider a family of designs that (i) select $E\in\{1,\ldots,4\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$).
-Here, $E=1$ corresponds to recruiting from a single area, while $E=4$ recruits from all four areas.
-We examine a grid of sample sizes that varies the number of participants from $10$ up to $104$ ($52$ treated units), with the latter corresponding to the size of the experiment in Banerjee et al.
+We consider a family of designs that (i) select $E\in\{1,2\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$); (iii) optimizes over shrinkage weights as in Algorithm 1.
+We examine a grid of values for $n_{1}$ ranging from 40 to 80 (with 52 corresponding to the size of the experiment in Banerjee et al.
+(2024)).
+For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 2.
+We assume that the variance of a single treated-control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
+Choice of the areas
+Figure 6 reports the optimal sample size allocation across areas when either one or two areas can be selected.
+When only one area can be chosen, the method selects the area with the largest weight $\omega$, which in this application corresponds to the area containing Bengaluru (about 76% of the population).
+When two areas can be selected, the procedure chooses the two areas with the largest $\omega$ (Area 2 and Area 3, the latter containing Tumakuru), where the second area has weight $\omega\approx 12\%$.
+The method also optimally allocates the sample size across these areas, assigning roughly 10 to 20 villages in total (half of which treated) to Tumakuru and the remaining villages to the larger area containing Bengaluru.
+Figure 14 shows the optimal shrinkage weights: the shrinkage weight for the Bengaluru area is about 93%, reflecting an interior solution, while it is essentially one for Tumakuru.
+The resulting pattern is simple and intuitive: assign most of the sample size to the area with the largest weight, but, when possible, include a second area with substantial weight, letting the allocation of villages across areas balance robustness and variance.
+Figure 6: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for one or two areas included in the experiment.
+Figure 7: Relative worst-case MSE across designs under a calibrated bias bound $B$, set to the largest absolute discrepancy between the experimental and observational ATE estimates in Banerjee et al.
 (2024).
-For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 2.
-We assume that the variance of a single treated–control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
-Figure 8: Optimal area selection and village allocation for a large experiment ($n_{1}=52$) under constraints on the number of areas $E$.
-Each panel corresponds to a value of $E$; shading indicates the total number of recruited villages in each area.
-With $E=1$ the design concentrates in the area with the noisiest observational estimate; as $E$ increases, allocation spreads across areas, reflecting a trade-off between observational uncertainty and experimental variance.
-Figure 9: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for each constraint on the number of eligible areas $E$.
-Lines show the number of villages assigned to each area.
-Figure 10: Relative MSE across designs under an hypothetical scenario where the bias equal the difference between the experimental and observational ATE estimate from Banerjee et al.
-(2024) as $n_{1}$ varies (from ten to fifty-two villages).
-The figure reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$ for (i) the proposed design (area targeting and optimal $\gamma^{\star}$), (ii) a random benchmark (uniform area selection and equal allocation), where $\mathrm{MSE}^{*}$ denotes the MSE of the oracle that knows the bias, as $n_{1}$ varies and for each $E$.
-The proposed design closely tracks the oracle across $n_{1}$ and $E$, while the random benchmark remains substantially above, especially when $E$ is small.
-Experimental design
-We begin by studying which areas are selected and how sample is allocated when the total experimental sample is large ($n_{1}=52$) and the number of admissible areas may or may not be constrained.
-Figure 8 visualizes the resulting allocation across Karnataka’s four observational areas for $E\in\{1,2,3,4\}$.
-With $E=1$, the algorithm concentrates recruitment in the Tumakuru area—the location with the highest uncertainty in the observational estimates.
-As $E$ relaxes to $2$ and $3$, the total sample splits across them in roughly (but not exactly) equal shares, reflecting a trade-off between (i) exploiting heterogeneity in the observational variance and (ii) keeping experimental variance low.
-When $E=4$, all areas are eligible and the allocation smooths further across space.
-In the rest of our discussion, we discuss properties when not all areas may be selected ($E<4$).
-Figure 9 tracks how area-level allocations evolve with the total number of villages.
-When the experiment is small (e.g., $10$ villages) and $E=1$, the algorithm favors an area with relatively low experimental variance, even if its observational estimate is less uncertain, because, at small $n$, experimental noise is first-order.
-As the total sample increases, the weight shifts toward areas with more uncertain observational estimates.
-Consistently, the area with the largest experimental variance (Area 2) is typically excluded.
-Appendix Figure 13 reports the optimal shrinkage $\gamma_{j}^{\star}$ by area as total villages vary.
-The solution is interior for all $E<4$: $\gamma_{j}^{\star}$ rises with sample size, approaching one from below as experimental noise diminishes.
-For example, at a moderate sample (around 15 villages), $\gamma^{\star}$ is about $0.65$ when $E\in\{2,3\}$; with $E=1$, concentrating the sample within one area lowers experimental variance and therefore yields a larger $\gamma^{\star}$.
-Main MSE comparisons
-We compare three strategies: (i) our chosen design which chooses the area, sample size and $\gamma^{\star}$ as described in Section 3, (ii) a standard benchmark that selects areas uniformly at random, allocates villages evenly across the chosen areas, and sets $\gamma_{j}=1$ for experimental estimates, and (iii) an oracle that knows the bias vector $b$ and optimizes both design (areas and allocation) and $\gamma$.
-Because $b$ is unknown in practice, for this exercise, we calibrate it using the difference between the Hyderabad RCT ATE obtained from the follow-up experiment of Banerjee et al.
-(2024) and the Karnataka DiD average effect across sites, treating $b$ as common across areas.
-Figure 10 plots the ratio $\text{MSE}/\text{MSE}^{\star}$, where $\text{MSE}^{\star}$ is the oracle’s MSE.
+The y-axis reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$, where $\mathrm{MSE}$ is the worst-case MSE over $\|b\|_{\infty}\leq B$ for a given design and $\mathrm{MSE}^{\star}$ is the (infeasible) oracle minimum that knows $B$ and optimally chooses the design.
+We compare our design (“regret optimal”) that does not use information about $B$ and minimizes the worst-case adaptation regret to the variance-minimizing Neyman allocation.
+MSE comparisons
+We compare two strategies: (i) our robust design, which jointly chooses the area(s), the sample allocation, and the shrinkage weights $\gamma^{\star}$ as described in Section 3; and (ii) the variance–optimal (Neyman) allocation.
+We evaluate each strategy using the worst-case MSE over bias vectors satisfying $b||_{\infty}\leq B$.
+Rather than treating $B$ as a free sensitivity parameter, we calibrate it using experimental evidence from Banerjee et al.
+(2024).
+Specifically, we set $B$ equal to the largest absolute discrepancy between the Hyderabad RCT ATE reported in the follow-up experiment of Banerjee et al.
+(2024) and the corresponding Karnataka DiD estimate in each area.
+Importantly, our design procedure does not use this calibrated bound $B$ as an input; it is used only for the ex-post MSE comparison.
+Figure 7 plots regret (the ratio $\text{MSE}/\text{MSE}^{\star}$, $\text{MSE}^{\star}$ is the smallest achievable worst-case MSE for a bias upper bound at most $B$, and MSE is the worst case MSE for $b||_{\infty}\leq B$).
 As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone.
-Across $E\in\{1,2,3\}$, the random benchmark remains well above the oracle—by roughly 3–5 percentage points (pp) for $E=1$, about 10 pp for $E=2$, and around 20 pp for E=3E=3—whereas our design tracks the oracle closely (within a few percentage points) even without knowledge of $b$.
-The gap is the largest when a larger number of areas $E$ can be included, because our procedure can better emulate the oracle by selecting (and differentially weighting) favorable areas.
-This result illustrates the merit of the method that improves MSE up to $20$ percentage points compared to a standard benchmark and leads to a near-oracle solution.
+Across $E\in\{1,2\}$, the Neyman benchmark remains well above the oracle— with regret around 2.5 or more.
+On the other hand, our design tracks the oracle more closely, with the ratio closer to one and no larger than 1.2.
+Appendix Figure 15 collects similar comparisons for the worst-case regret over all possible values of $B$.
 7 Implications for practice
@@
 A key challenge is that the bias of the observational estimators is unknown in practice.
-We adopt a minimax proportional regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias.
+We adopt a minimax adaptation regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias.
 This reveals a fundamental trade-off between precision and robustness.
@@
 Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify.
-When such parametrization is not readily available, Section 4 shows how our procedure applies to $\tau$ identified through arbitrary moment restrictions.
+When such a parametrization is not readily available, Section 4 shows how our procedure applies to $\tau$ identified through arbitrary moment restrictions and Section 5.1 when it is partially identified.
   • Assemble informative observational evidence.
@@
 Optimize over the design and corresponding estimators.
-The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$).
+The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$), and optimization is conducted via a mixed-integer quadratic program in Algorithm 1.
 The applicability of our framework spans a wide range of settings in economics and beyond.
@@
 In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis (Allcott et al., 2025).
-Beyond economics, medical applications include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025).
+Beyond economics, medical applications may include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025).
 Several open questions remain for future research.
 These include settings where researchers have well-specified priors about bias such as from Bayesian models (Gechter et al., 2024), or other modeling choices (Olea et al., 2024).
-This may be potentially incorporated through additional moment conditions as in Section 4.
-Additional future extensions include sequential or adaptive experimental choices or settings where the target estimand is only be partially identified (e.g.
+This may be potentially incorporated through additional moment restrictions as in Section 4.
+Additional future extensions include sequential or adaptive experimental choices (e.g.
 Cesa-Bianchi et al., 2025).
 Appendix A Proofs of main results
-A.1 Proof of Theorem 1 and Theorem 2
-Theorem 1 is a special case of Theorem 2 proved below.
-Step 1 (worst-case MSE). By Hölder’s inequality,
-$\sup_{b\in\mathcal{B}(B)}\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\leq\Big(B\,\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}\Big)^{2}=B^{2}\,\beta(W).$
-In addition, this inequality is attained by $b_{j}=B\,\mathrm{sign}\!\big((\Gamma_{\Lambda}(W)^{\top}\omega)_{j}\big)$ for $j\in\mathcal{I}^{c}$ (and $b_{j}=0$ for $j\in\mathcal{I}$).
-Hence $\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)=\alpha(W,\Sigma)+B^{2}\beta(W).$
+A.1 Proof of Theorems 1, 2, 4
+[…]
+Step 1 (worst-case MSE). By definition of dual norm,
+[…]
+Hence $\sup_{b\in\mathcal{B}_{l}(B)}\mathrm{MSE}_{b}(W,\Sigma)=\alpha(W,\Sigma)+B^{2}\beta_{l}(W).$
 Step 2 (oracle envelope and basic properties).
-Let $\delta(t):=\inf_{W,\Sigma}\{\alpha(W,\Sigma)+t\beta(W)\}$.
-Recall that $\beta(W)$ and $\alpha(W,\Sigma)$ are bounded from above and $\alpha(W,\Sigma)>0$ by Assumption 3.
+Let $\delta(t):=\inf_{W,\Sigma}\{\alpha(W,\Sigma)+t\beta_{l}(W)\}$.
+Recall that $\beta_{l}(W)$ and $\alpha(W,\Sigma)$ are bounded from above and $\alpha(W,\Sigma)>0$ by Assumption 3 (using the fact that $\Omega$ contains some non zero entry).
 Being the pointwise infimum of affine functions of $t$, $\delta$ is concave, nondecreasing, and finite for $t\geq 0$.
 Moreover,
-$\delta(0)=\alpha^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=\beta^{\star}.$
-Step 3 (quasi-convexity and boundary maximization). Fix $(\alpha,\beta)$. Consider
-$\phi(t):=\frac{\alpha+t\beta}{\delta(t)},\qquad t\geq 0.$
+$\delta(0)=\alpha^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=\beta_{l}^{\star}.$
+Step 3 (quasi-convexity and boundary maximization) Fix $(\alpha,\beta_{l})$. Consider
+$\phi(t):=\frac{\alpha+t\beta_{l}}{\delta(t)},\qquad t\geq 0.$
 Because the numerator is affine and the denominator is positive and concave in $t$, $\phi$ is quasi-convex on $0,\infty)$.
@@
 Letting $T\to\infty$ and using the limits in Step 2,
-$\sup_{t\geq 0}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \lim_{t\to\infty}\frac{\alpha+t\beta}{\delta(t)}\right\}=\max\!\left\{\frac{\alpha}{\alpha^{\star}},\ \frac{\beta}{\beta^{\star}}\right\},$
-with the stated conventions $0/0=0$.
-Appendix B Proofs of the extensions
-B.1 Proof of Theorem 3
-The proof of Theorem 3 follows similarly the one of Theorem 2 with minor modifications. It is easy to show that
-$|L_{B}(\mathcal{E},\Sigma,\gamma)|=2\{z_{1-\eta/2}\,\alpha(\mathcal{E},\Sigma,\gamma)+B\,\sqrt{\beta(\mathcal{E},\gamma)}\}.$
-We can now follow verbatim the proof of Theorem 2, with $\delta(t):=\inf_{\mathcal{E},\Sigma,\gamma}2\{z_{1-\eta/2}\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)}+t\sqrt{\beta(\mathcal{E},\gamma)}\}$ with limits
-$\delta(0)=2z_{1-\eta/2}\sqrt{\alpha^{\star}},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=2\sqrt{\beta^{\star}}.$
-Taking the ratios and limits as in Step 3 of the proof of Theorem 2 the proof completes.
-B.2 Proof of Theorem 4
-It is easy to show that
-$\check{R}_{B}\left(\mathcal{E},\Sigma,\gamma\right)\equiv\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\beta}(\mathcal{E},\gamma)B^{2}+\check{\alpha}(\mathcal{E},\Sigma,\gamma),$
-with $\check{\beta},\check{\alpha}$ as defined in Equation (12), where $\check{\beta}\geq 0$ by construction.
-In addition $\check{\alpha}>0$ and bounded away from zero by Assumption 1.
-We can therefore follow verbatim the proof of Theorem 1.
-Appendix C Examples
-C.1 Two parameters examples in Equation (3)
+$\sup_{t\geq 0}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \lim_{t\to\infty}\frac{\alpha+t\beta_{l}}{\delta(t)}\right\}=\max\!\left\{\frac{\alpha}{\alpha^{\star}},\ \frac{\beta_{l}}{\beta_{l}^{\star}}\right\},$
+with the stated conventions $0/0=1$. This proves Theorem 2.
+Step 4 (worst-case with bound on the bias) From quasi-convexity, whenever $t\leq\bar{t}$, it follows that
+$\sup_{t\in[0,\bar{t}]}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \frac{\alpha+\bar{t}\beta_{l}}{\delta(\bar{t})}\right\},$
+proving Theorem 4.
+A.2 Proof of Theorem 3
+Step 1 (closed form of $\mathcal{L}_{l,B}(W,\Sigma)$). Fix $(W,\Sigma)$ and $B\geq 0$. By definition,
+$\ell_{b}(W,\Sigma)=\underline{\tau}(\hat{\theta}(W))-\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b-z_{1-\eta/2}\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)},$
+so
+$\inf_{\|b\|_{l}\leq B}\ell_{b}(W,\Sigma)=\underline{\tau}(\hat{\theta}(W))-z_{1-\eta/2}\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}-\sup_{\|b\|_{l}\leq B}\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b.$
+Since $b_{j}=0$ for $j\in\mathcal{I}$ and $\|b_{\mathcal{I}^{c}}\|_{l}\leq B$, the dual norm gives
+$\sup_{\|b\|_{l}\leq B}\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b=B\,\sqrt{\beta_{l,\underline{\omega}}(W)}.$
+Hence
+$\inf_{\|b\|_{l}\leq B}\ell_{b}(W,\Sigma)=\underline{\tau}(\hat{\theta}(W))-z_{1-\eta/2}\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}-B\,\sqrt{\beta_{l,\underline{\omega}}(W)}.$
+Similarly, for the upper bound,
+$\sup_{\|b\|_{l}\leq B}u_{b}(W,\Sigma)=\bar{\tau}(\hat{\theta}(W))+z_{1-\eta/2}\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+B\,\sqrt{\beta_{l,\bar{\omega}}(W)}.$
+Therefore the worst-case interval has length
+$|L_{l,B}(W,\Sigma)|=\big(\bar{\tau}(\hat{\theta}(W))-\underline{\tau}(\hat{\theta}(W))\big)+z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big)+B\big(\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}\big).$
+Let $b_{0}$ denote the true bias, with $\|b_{0}\|_{l}\leq B$. Under (17),
+$\bar{\tau}(\hat{\theta}(W))-\underline{\tau}(\hat{\theta}(W))=(\bar{\omega}-\underline{\omega})^{\top}\theta+(\bar{\omega}-\underline{\omega})^{\top}\Gamma_{\Lambda}(W)\bar{g}_{\Sigma},$
+so
+$\mathbb{E}_{W,\Sigma,b_{0}}\big[\bar{\tau}(\hat{\theta}(W))-\underline{\tau}(\hat{\theta}(W))\big]=(\bar{\omega}-\underline{\omega})^{\top}\theta+(\bar{\omega}-\underline{\omega})^{\top}\Gamma_{\Lambda}(W)b_{0}.$
+Subtracting the identified set length $\bar{\tau}(\theta)-\underline{\tau}(\theta)=(\bar{\omega}-\underline{\omega})^{\top}\theta$, we obtain
+$\mathbb{E}_{W,\Sigma,b_{0}}\big[|L_{l,B}(W,\Sigma)|\big]-\big(\bar{\tau}(\theta)-\underline{\tau}(\theta)\big)=$
+$(\bar{\omega}-\underline{\omega})^{\top}\Gamma_{\Lambda}(W)b_{0}+z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big)+B\big(\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}\big).$
+Taking the worst case over $\|b_{0}\|_{l}\leq B$ implies
+$\mathcal{L}_{l,B}(W,\Sigma)=z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big)+B\big(\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}+\sqrt{\beta_{l,\bar{\omega}-\underline{\omega}}(W)}\big),$
+which yields the linear form $\mathcal{L}_{l,B}(W,\Sigma)=A(W,\Sigma)+B\,C_{l}(W)$.
+Step 2 (oracle envelope and basic properties). Define
+$\delta(t)~\equiv~\inf_{(W,\Sigma)\in\mathcal{D}^{\prime}}\{A(W,\Sigma)+t\,C_{l}(W)\},\qquad t\geq 0.$
+By Assumption 3 (because either or both $\bar{\omega},\underline{\omega}$ are not zero vectors), $A(W,\Sigma)$ and $C_{l}(W)$ are bounded from above and $A(W,\Sigma)>0$ for all $(W,\Sigma)\in\mathcal{D}^{\prime}$, so $\delta(t)$ is finite and strictly positive for all $t\geq 0$.
+Being the pointwise infimum of affine functions of $t$, $\delta$ is concave and nondecreasing on $0,\infty)$.
+Moreover,
+$\delta(0)=A^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=C_{l}^{\star}.$
+Step 3 (quasi-convexity and boundary maximization). Fix $(W,\Sigma)$ and set $A\equiv A(W,\Sigma)$ and $C\equiv C_{l}(W)$. For $t\geq 0$ define
+$\phi(t)~\equiv~\frac{A+t\,C}{\delta(t)}.$
+The numerator is affine in $t$ and the denominator is positive and concave, so $\phi$ is quasi-convex on $0,\infty)$.
+Hence, on any compact interval $[0,T]$, $\max_{t\in[0,T]}\phi(t)$ is attained at the boundary $\{0,T\}$.
+Letting $T\to\infty$ and using the limits in Step 2,
+$\sup_{t\geq 0}\phi(t)=\max\left\{\frac{A}{\delta(0)},\;\lim_{t\to\infty}\frac{A+t\,C}{\delta(t)}\right\}=\max\left\{\frac{A(W,\Sigma)}{A^{\star}},\;\frac{C_{l}(W)}{C_{l}^{\star}}\right\}.$
+By the definition of $\tilde{\mathcal{R}}_{l}(W,\Sigma)$ and the identity $\mathcal{L}_{l,B}(W,\Sigma)=A(W,\Sigma)+BC_{l}(W)$, this yields
+$\tilde{\mathcal{R}}_{l}(W,\Sigma)=\sup_{B\geq 0}\frac{\mathcal{L}_{l,B}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\mathcal{L}_{l,B}(W^{\prime},\Sigma^{\prime})}=\max\left\{\frac{A(W,\Sigma)}{A^{\star}},\;\frac{C_{l}(W)}{C_{l}^{\star}}\right\},$
+which proves the theorem. ∎
+A.3 Proof of Corollary 1
+Whenever $\bar{\omega}=\underline{\omega}=\omega$, we can write $A(W,\Sigma)=\alpha_{\omega}(W,\Sigma)$ and $C_{l}(W)=\beta_{l,\omega}(W)$. From Theorem 3
+[…]
+[…]
+Appendix B Examples
+B.1 Two parameters examples in Equation (6)
 By Theorem 1, for fixed $(\mathcal{E},\Sigma)$ the adaptation regret as a function of the shrinkage vector equals
@@
   • Otherwise, by the intermediate value theorem and strict monotonicity of the two curves, there exists a unique $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$ such that $\frac{\alpha(j,\gamma_{j})}{\alpha^{\star}}\;=\;\frac{\beta(j,\gamma_{j})}{\beta^{\star}}.$
-C.2 Numerical example
-Here, we present a numerical example in Setting 2.
+B.2 Numerical example
+Here, we expand our discussion around Example 5.
+Figure 8: Example for $j=2$.
+Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$.
+Bottom row: the optimal weight $\gamma_{2}^{\star}$.
+Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
+Figure 9: Example for $j=2$ (regret).
+Different colors correspond to the regret for choosing either experiment.
+Figure reports $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$).
+The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret.
+Columns vary, respectively, (a) $\omega_{2}$ with $v_{2}=0.5$, $v_{1}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
+Consider choosing $\gamma_{2}^{\star}$ when experiment $j=2$ is more precise as $j=1$, with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, $v_{1}=1$, $v_{2}=0.5$, and $\omega_{1}=1$.
+The optimal $\gamma_{2}^{\star}$ depends not only on the features of arm $2$ but also on those of the alternative arm 2-2, $(v_{-j}^{2},\sigma_{-j}^{2},\omega_{-j})$, through the oracle benchmarks $(\alpha^{\star},\beta^{\star})$.
+Figure 8 illustrates this by plotting $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ at the optimal weight (top row) and the corresponding $\gamma_{2}^{\star}$ (bottom row), as we vary $\omega_{2}$, $v_{2}$, and $\sigma_{2}$ column by column.
+Varying $\omega_{2}$ yields two key effects: both very small and very large $\omega_{2}$ push $\gamma_{2}^{\star}$ toward one, but for different reasons.
+When $\omega_{2}$ is small, the oracle bias declines faster than the experimental variance (since $\beta^{\star}=\omega_{2}$), while when $\omega_{2}$ is large, the oracle bias is $\beta^{\star}=\omega_{1}$ and $\omega_{2}$ affects only the feasible experiment’s bias.
+Consequently, both regret and the optimal estimator depend on the oracle’s choice of which experiment it would run.
+In the interior region for $\omega_{2}$, the optimal weight first moves toward the variance-optimal solution (as the oracle bias falls) and then rises again toward a bias-preferred solution.
+When we vary the experimental and observational variances, we obtain the expected comparative statics: $\gamma_{2}^{\star}$ decreases with the experimental variance and increases with the observational variance.
+These patterns translate into the regret of each arm (Figure 9).
+For arm $j=2$, regret initially falls sharply as $\omega_{2}$ rises and then declines more gradually once the bias component is already small.
+For arm $j=1$, regret is at best slightly decreasing for small changes in $\omega_{2}$ (when variance effects dominate) but eventually increases with $\omega_{2}$, because the oracle increasingly favors arm $2$ and the opportunity cost of selecting arm $1$ rises.
+Table 4 summarizes these patterns.
+We now provide further details.
 Choice of $\gamma^{\star}$
 Figure 8 (top row) reports $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$.
-In the bottom row, we plot the corresponding $\gamma_{2}^{\star}$.
+In the bottom row, we plot $\gamma_{2}^{\star}$.
 Observation 1: small $\omega_{2}$ pushes $\gamma_{2}^{\star}$ toward one.
@@
 $\boldsymbol{\omega_{2}<\omega_{1}}$ $|\omega_{2}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Decrease in $\omega_{2}$
-$\boldsymbol{\omega_{2}>\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Increase in $\omega_{2}$
+$\boldsymbol{\omega_{2}>\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{1}^{2}\sigma_{1}^{2}\;+\;\omega_{2}^{2}\dfrac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ Interior Increase in $\omega_{2}$
 $\boldsymbol{\omega_{2}\gg\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{1}^{2}\sigma_{1}^{2}\;+\;\omega_{2}^{2}\dfrac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ Interior Slow increase in $\omega_{2}$
@@
 In each column, we keep the same parameterizations as in Figure 8.
-Observation 1: The regret for $j=2$ decreases rapidly in $\omega_{2}$ and then decreases more slowly.
+Observation 1: The regret for $j=2$ decreases rapidly in $\omega_{2}$ and then stabilizes/slowly increases.
 In the first column, we vary $\omega_{2}$ with the first experiment $j=1$ having a smaller experimental variance, $v_{1}=v_{2}/2$.
@@
 Bottom row: optimal weights $\gamma_{j}^{\star}$ for $j=1,2$, which explain how sensitivity and precision interact to drive the design switch at $x^{\star}$.
-Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
-Appendix D Additional figures and tables
-Figure 12: Share of experimental papers published in AEA journals also presenting experimental results in combination with observational estimates (either from a reduced form, or from a structural model or from both).
-Table 5: Optimal shrinkage and experiment choice across different regimes in the presence of two parameters (Setting 2) The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases).
+Columns vary, respectively, (a) $\omega_{2}$ with $v_{2}=0.5$, $v_{1}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
+Table 5: Optimal shrinkage and experiment choice across different regimes in the presence of two parameters (Example 5) The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases).
 The other cases correspond to boundary solutions.
@@
 variance $\alpha/\alpha^{\star}$ of $j=2$
-Figure 13: Optimal shrinkage $\gamma_{j}^{\star}$ by area as a function of total treated villages $n_{1}$ (from ten to fifty-two villages), under different constraints $E$.
-For $E<4$, the solution is interior ($0<\gamma_{j}^{\star}<1$) and increases with $n_{1}$, approaching one as experimental noise falls.
+Appendix C Additional figures and tables
+Figure 11: Share of experimental papers published in AEA journals also presenting experimental results in combination with observational estimates (either from a reduced form, or from a structural model or from both).
+Figure 12: Experimental design for General Equilibrium Effects: Variance and bias regret of the proposed optimal design relative to the corresponding variance and bias oracles, as a function of total experimental sample size $n_{\text{tot}}$.
+The left panel considers the case with one feasible experiment and the right panel the case with two feasible experiments; values close to one indicate little loss relative to the oracle.
+Figure 13: Experimental Design for General Equilibrium Effects: Optimal shrinkage weights $\gamma_{j}$ for each arm (UCT, CCT, and Training) as a function of the total experimental sample size $n_{\text{tot}}$.
+The left panel considers designs with one feasible experiment, while the right panel allows two feasible experiments, illustrating how the procedure gradually shifts weight from UCT toward the job program as sample size increases.
+Figure 14: Site selection for external validity: Optimal shrinkage weight $\gamma_{j}^{\star}$ as a function of the treated villages in the experiment.
+Figure 15: Site selection for external validity: Variance, bias, and adaptive regret of the Neyman (variance–optimal) allocation relative to the proposed optimal design, as a function of the total experimental sample size $n_{\text{tot}}$.
+Left panel: one feasible experiment; right panel: two feasible experiments.
+Appendix D Review of adaptation regret with fixed design
+This section reviews the definition of adaptation regret with a fixed design in Armstrong et al. (2024) and Tsybakov (1998) and illustrates distinctions from our framework.
+Single parameter with a fixed design
+Different from this paper, Armstrong et al.
+(2024) focus on a single parameter with two estimators under a fixed design where one of the two estimators is unbiased for the target estimand.
+In our notation, this corresponds to fixing a design $(E,\Sigma(E))$ with $E=\{1\}$, so there is an observational estimate $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}$ and an experimental estimate $\tilde{\theta}^{\exp}\in\mathbb{R}$ satisfying
+$\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]=\theta_{1}+b,\qquad\mathbb{E}[\tilde{\theta}^{\exp}]=\theta_{1},\quad\Delta\tilde{\theta}=\tilde{\theta}^{\mathrm{obs}}-\tilde{\theta}^{\exp}$
+for an unknown scalar bias $b\in\mathbb{R}$.
+Without loss, standardize each estimator as well as the estimand $\theta_{1}$ by the variance of the difference $\Delta\tilde{\theta}$.
+With this normalization, the authors further assume
+$\begin{pmatrix}\tilde{\theta}^{\exp}\\ \Delta\tilde{\theta}\end{pmatrix}\sim\mathcal{N}\!\left(\begin{pmatrix}\theta_{1}\\ b\end{pmatrix},\,\Sigma\right),\quad\Sigma=\begin{pmatrix}\sigma_{\exp}^{2}&\rho\,\sigma_{\exp}\\[4.0pt] \rho\,\sigma_{\exp}&1\end{pmatrix}$,
+where $\sigma_{\exp}^{2}$ is the variance of $\tilde{\theta}^{\exp}$, and $\rho$ is the correlation between the experimental estimate and $\Delta\tilde{\theta}$.
+The question of interest in Armstrong et al.
+(2024) is how to optimally combine $\tilde{\theta}^{\mathrm{exp}}$ and $\Delta\tilde{\theta}$ in a single estimator.
+Non linear estimator and corresponding adaptation regret
+Using invariance arguments, Armstrong et al. (2024) study estimators of the form
+$\hat{\tau}_{\delta}=\tilde{\theta}^{\exp}+\rho\,\sigma_{\exp}\,\big\{\delta(\Delta\tilde{\theta})-\Delta\tilde{\theta}\big\}$,
+for some measurable function $\delta:\mathbb{R}\to\mathbb{R}$ to be optimized. The authors show that for this choice, the adaptation regret worst case for $|b|\leq B$, can be written as
+$A(B,\delta)=\frac{\displaystyle\mathrm{MSE}_{B}(\delta)}{R^{*}(B)},\qquad R^{*}(B)=\sigma_{\exp}^{2}\Big[(1-\rho^{2})+\rho^{2}\,r_{\mathrm{BNM}}(B)\Big]$
+where $r_{\mathrm{BNM}}(\cdot)$ denotes the B-minimax risk in the bounded normal-mean problem.
+Properties of the objective function and solution
+Because of the non-linearity of the estimator (and of the $r_{\mathrm{BNM}}$ function) $A(B,\delta)$ is not quasi-convex in $B^{2}$ for the chosen non-linear shrinkage rule.
+The optimal solution is obtained by solving a minimax problem by using a discrete approximation over the least-favorable prior over $B$ (Chamberlain, 2000).
+This differs from our framework where quasi-convexity arises as we restrict the class of estimators to be linear.
+A formal argument is provided in the proof of Theorem 1.
+Such quasi-convexity simplifies estimation here, since the choice is not only for the estimator, but also for the class of designs and sample size allocations.